A Complete Bibliography of Publications in Statistical Science

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

18 October 2024 Version 1.27

Title word cross-reference

1: 1 [SJL^{+}12]. 2 × 2 [How98, RLQ17]. 3 [QOS19]. 50 [BM16, SB16]. α [JW94]. D [WWZ16]. E [VW23]. ℓ_{0} [CTB20]. ℓ_{1} [CTB20]. G [VJ14]. k [Men05, SZ14]. M [NRWY12]. A [STLD14]. Q [STLD14]. P [AG05, BC05, BCD05, GM05a, GM05b, Tho05, And22, FPBT22, Gel05, SSVV11]. ρ [TOS22]. u [Gel05].

-estimation [VJ14]. -Estimators [NRWY12]. -Learning [STLD14]. -Regularization [CTB20]. -Stable [JW94]. -Triangle-Faithfulness [SZ14]. -Values [FPBT22, Gel05, SSVV11, BC05, BCD05, GM05a, GM05b, Tho05, VW23]. -versus [CTB20]. -Wave [WWZ16]. -Year [BM16]. -Year-Old [SB16].

14 [HMRV00]. 19 [CWHD22, JCT22, Lin22, WBR^{+}22, YS22]. 1930s [MP93]. 1980 [Dem86, Eri86, Fel86, FN86a, FN86b, Kad86, Mad86, Mos86, Nat86, Wol86].

1980s [TW10]. 1988 [Sti90]. 1996 [Efr98, Swi97]. 19th [BG05].

2001a [HMRV00]. 2004 [Car11, DS11, Mar11, PS11]. 2005 [Bri08a]. 2016 [Car19]. 2020s [KZZW22]. 2021 [LB22]. 21st [Efr98, MFR24a, RP24]. 227 [BN90].

4 [BN90]. 447 [Sto14].

A. [Fri07, Sto14]. A.-M. [Fri07]. Abdel [Est03]. Abraham [BG07]. absolute [Ell98, PK97]. absolute-error [PK97]. Abundance [SS99]. Academe [Nol92]. Academic [Hot88a]. Accelerated [EM06]. Access [Cox91, DP91a, DP91b, GKRS05, KM91, Nor91]. Accessibility [LZD^{+}20]. Account [Sti89, GST97]. Accuracy [BFW07, ET86a, ET86b, Har86]. Accurate [BD08]. Accuse [Mur05]. ACE [RR14a]. ACIC [Car19]. Acknowledgment [Ano87a, Ano89a, Ano92a, Ano93a, Ano94a, Ano95a, MdL00]. Across [BH23, CV08, PB14, RTW13]. Act [Mal02, SH14]. Action [Bre88a, Eub88, HT88, Ram88a, Ram88b, Wah88, XR15]. Active [HJJ^{+}21, SGZW10]. Activities [Bar88a]. Acyclic [POSD13]. Ad [BGLC93, Bil93, Car93, CGP^{+}93, Gen93, vZ93]. Adapt [Har89]. Adaptation [HK02]. Adaptive [BV23, Cai12, DMJ23, GV23, Gio23, IR23, Jen23, Mar21, RLLKV23b, RLLKV23a, Ros96]. Add [Hut16]. Additive [Bri86, HT86a, HT86b, Hof21, LD24, McC86, Mey18a, Nel86, RR23, Sto86a]. Addressing [Shp14]. Adjusted [IRÁ22, WM04]. Adjusting [Dem86, Eri86, Fel86, FN86a, FN86b, Kad86, Mad86, Mos86, Nat86, Wol86]. Adjustment [BR94a, BR94b, Bre94a, Bre94b, DS94, EFK94, FW94a, FW94b, FC11, LP11a, LL94, MT11, Ros02, Ste94]. Administrative [YSR20]. Administrator [BF91]. Admissibility [Joh19]. Adopters [HPV06]. Advanced [Lan14]. Advances [ADR04, Sti08, WGHR06, PSES98]. Advancing [TX23]. Advection [RTW13]. Adversaries [Fei93]. AF [Sto14, SKKS14]. affecting [DWZ97]. Affects [BT18]. Affirmative [XR15]. Africa [DGA^{+}20]. After [SY12]. Afterwards [GH16]. Again [And22]. Age [Fri08, SW16]. Age-Period-Cohort [SW16]. Agreed [Ber03b]. Agricultural [CEN^{+}19, Rub90, SNDS90]. Ahead [DHS^{+}19b]. AIDS [FJH^{+}89, Gas87a, Gas87b, Gei87b, Gla87, Gol87a, Kay87, KR87, Wit87]. Aim [KEHSY19]. Aims [KJED09]. Air [BGST04, Gin88, SKKS14]. Aitchison [GGBS^{+}23]. Akaike [FP95]. al [CTB20, Law19]. Alan [CH15]. Albert [Olk87a]. Aldous [Bha22]. Alex [Oga18, Sch18, Zhu18]. Algorithm [BD01, FYL10, HRR11, HJJ^{+}21, Lai10, VACC22]. Algorithmic [IJ23]. Algorithms [AS93, Ber03a, BPV20, BH07a, BH07b, BMW07, CRSW13, FL24, Fei93, FL93, Has07, Hil89, Jor89, McC89, OS22, PSW15, Ram93, Rob95, SGZW10, Tha10, dP89b, dP89a, RR01]. Allelic [SCtW^{+}09]. Allen [Olk91]. Alone [CR17]. Also [ZKS19]. Alternative [CDW95, KS07a, KS07b, RM07, Ros99, TD07, WL10, Joh97]. Alum [CFL13]. Always [SCN^{+}10]. Amber [Sam91]. Ambiguity [HM16]. American [LRdP^{+}04]. Among [MM22]. Amplitude [MRSS15]. Analyses [LS03, ZWS^{+}19]. Analysis [Akr04, AY01, And87, ADR04, Ano89b, Ano89f, Atk86, Ban20, BBSW10, Bar92a, BH87a, BH87b, BB91, BB04a, BCW87a, BCW87b, Bec07, BW07, Ber12, Bic88, BFWW09, BH90, CP14, Cha86, CCLC09, CLSW16, CJLS93, Cox90a, CW93a, Cri90, Daw91, DB92, DS11, Dem93, Dia91, DHS^{+}19a, Dra88a, Dra88b, DHM^{+}17, DuM12a, DS92, Eas89, Edd87, Eva12, Fai89, Fel87, Fis90, Fre87, Fri07, GM21a, Gei87c, GKLP99, Gel14, GT03, GS93, GK87, Gol87b, Gol16, GKRS05, Gow90, GGBS^{+}23, Gre91, GvdL19, GZZ20, HT00a, HTT20, HR11, Hed92, Her19, HA88, Hil93, Hoa03, Hod87a, Hod87b, Hub86, Hub87a, Hub87b, Hym91, IKY10, JS06a, Jim11, Joe89, JY89, Kad87b, KP88, KJED09, LLS11, LW06, LYL22, LZD^{+}20, Lin08, Lin06, LSS^{+}19, LBK21, Mad87, Mad93, Mar89, MRSS15]. Analysis [Mar21, Mar88b, Mar11, MTP^{+}23, MT12, McK04, MTNO15, Mor89b, Mor91e, Mos91, MC92, MQ04, NA94, NRWY12, Nor93, O’H89, OKS89, OW16, Pea93, Peñ06, Per87, Pic89, Pol03, PBSO24, Por87, Pre87, PC04, Raf89, RD06, RHGF14, RBGK14, SWMW89a, SWMW89b, SMJN17, Sch87b, Sch87a, Sen87, Sha93, Sin89, SJL^{+}12, SW17, Smi87, Smi89b, Smi89c, Sob93, SDLC93a, SDLC93b, Ste89, SDD06, Stu90, Thi86a, Thi86b, Tuf87, Tuk87, Utt91a, Utt91b, VTH14, Wai87, WS90a, WS90b, Wei89, Wel88, Whi93a, WTLW95, WGHR06, You90, Zha11a, vDK04, And96, GST97, Lin98, MdL98, RS97]. Analyzing [BBF^{+}22, BGGM06, BM99, GGC13, JBM^{+}06, ZZS21, GG02]. Anaysis [CST00]. Ancestral [GT94]. Ancillaries [Fra04]. Ancillarity [SB16]. and/or [vdGM12]. Anderson [DeG86e]. Animal [BCM00, KT88, Ros88, SS99]. Annals [DeG86f]. Annealing [BT93, WWZ16]. Anniversary [BM16, Cas03b]. Annotated [BG07, Fre86, FJH^{+}89, GZ86a, GZ86b, Hog86a, Mor86, Sha86a, Win86]. anomaly [PT11]. ANOVA [MSM24]. Answer [Eva12]. Antibodies [Gas87a, Gas87b, Gei87b, Gla87, Gol87a, Kay87, KR87, Wit87]. Antithetic [CCFR24]. Antoine [BG24]. Anytime [RGVS23]. Anytime-Valid [RGVS23]. Apology [MdL00]. Apparent [AOW17, LRR17]. appendix [Twe98]. Application [CK03, DJ24, Fai89, Gas87a, Gas87b, Gei87b, Gla87, Gol87a, GZZ20, HRR11, IKN09a, Joe89, Kay87, KR87, Pic89, QKM20, QOS19, Raf89, Rub90, Rub06a, Sin89, Smi89b, Smi89c, SNDS90, Wei89, Wit87, ZS09]. Applications [BL10, Böh16, Bur92, CJLV06a, CJLV06b, DLM06, DE86, FWZZ21, FWB94, Goo86a, Goo86b, Hey86, Ize21, Kro92, LY21, LMN06, LR92a, LR92b, McC94, MM06b, Moo90c, MSM24, Oga18, POSD13, Que24, Rei18a, Rei18b, Rob90a, Sch18, Sin86, Sol86, THNC03, Yu06, Zhu18, deL92, PS00, PT11, STW00, Box90a, Dem90a, Moo90c, Net90, Pra90, Rob90b]. Applied [AET09a, Ald17, BJM12, GKLP99, Gre10]. Applying [SS03, SDD06]. Appraisal [CK94, GR94a, GR94b, Hol94b, Mee11, Rao11a, SC94, Sed11, Sin94, Slu11, Sta94, Tho94c]. Approach [ALS^{+}01, AG05, Ban20, BBSW10, BM99, Bre18, BZ21, CCFR24, DB92, EH03, GT03, LZD^{+}20, Lin87c, Meh92, SAY22, SRR^{+}17a, SQH18, TSW21, Wol04a, ZWS21, Zub87, GW02, GST97]. Approaches [CEH^{+}17, JT90, LD18, MMT07, PC04, BBC^{+}02]. Approximate [Ald93, And22, BL10, BNPS13, HBG15, MVA^{+}18, WH16a, WH16b]. Approximated [BBSW10]. Approximating [ES95, MFR24a, Ano96]. Approximation [AGG90a, AGG90b, Bar90a, Bha87, Bri87b, Che90, Che10, Dah87, Han87a, Han87b, LY21, MG08, Ris87, Shi87, Sol87, Ste90, Wat90, WV94, But00]. Approximations [BBB^{+}19, BBK^{+}19, BKB^{+}19, DKK19, FJL^{+}19, GK19, KG21, Law19, Lit19, Tjø19, LRI23, WSC19, vdG19]. Area [CK94, DG12, GR94a, GR94b, Hol94b, Kor06, Lah03, PT03, Pfe13, SC94, Sin94, Sta94, Tho94c]. Argument [Bjø14, Daw14, Eva14, Han14b, May14a, May14b, Sei92, Zab92]. Arguments [Fra14]. ARIMA [Raf92]. armed [VBW15]. Art [BK17, RD06]. Arthur [Nau16]. Articles [Arr88, Bra88, Dem88a, Gup88, Hog88, Hot88c, Hot88b, Hot88a, Moo88, Olk88, Zid88]. Artificial [Lin87c, Sha87c]. Aspects [Atk86, CDW95, Ize01, Mar88a, NS07, Raf92, RSK22, Sha16]. Asserting [Gus23]. Assess [CLSW16]. Assessing [CG93, DB92, Dem93, Gle98, GŠP12, SSP^{+}19, XR15]. Assessment [ÁEdBCAM17, And88, BS13, Bre88b, DuM88a, FZ88a, FZ88b, GST18, Has88, Hat88, KT88, LL09b, MD88, Mor88, Nee88, Rya03, Sil09, WGHR06, Smi02]. assistance [Sam99]. Assisted [BO17]. Associated [PGP09]. Association [AB09, CCLC09, DLS14, KZI09, LL09a, MMT07, PGP09, SCtW^{+}09, TCC^{+}09, ZMG09, Fre99]. Assumption [LMR20a, STR^{+}17, SZ14, LMR20b]. Assumption-Free [LMR20a, LMR20b]. Assumptions [DMS10, KLDR09, RR88]. Astronomy [BD04]. Astrophysics [vDK04]. Astrostatistics [GWC04]. Asymptotic [Ama89, BN89, BN90, Ber89b, Kas89a, Kas89b, KK15, Rao89, RF89]. Asymptotics [BM24, Pol89a, Wal04, CH89, Dud89, GZ89, Pol89b, Pyk89]. Atmospheric [Sol03]. Attacks [GNR^{+}24]. Attrition [DHR^{+}13]. Auction [RD06]. Auctions [BBK06]. Auditing [Ano89f]. Augmentation [HRR11, Tan23, TW10, GST97]. Augmented [BG07]. August [Swi97]. Auguste [Kad09]. author [Bre01, Ell98, Ano94b]. authors [AG01, ACS01, Ber99, HK02, HT00b, HZ00, HMRV99, SWY00]. Autoencoders [WMB19]. Automated [DHS^{+}19a, Dun17, GvdL19]. Automatic [FM93, HL93a, HL93b, Mul93]. Autopsy [KL08]. Autoregressive [HP04a]. Auxiliary [DPL15, JD21, JSH07]. Average [Ban20, BCvdL20, Li20]. averaging [HMRV99, HMRV00]. Aversion [HM16]. Avoiding [Cha91a, And91, Bai91, Cha91b, CN91, Gli91, MP91, Zah91]. Axioms [Ber86, Fin86, Fis86a, Fis86b, Goo86c, Sei86, Sto86b, Sud86, Sup86].

B [EW17, LM14b, Nau00, WZG24, EM96]. B-splines [EM96]. Backcalculation [BSJ93a, BSJ93b, Bro93, CG93, DP93, GR93, KS93, SW93]. backfitting [HT00b]. Background [Cha15, GMN^{+}04, Mar04b]. Backprojection [CG93]. Bad [Bre94a, DH05]. Bailar [Ell97b]. Balance [HB08]. Balanced [CFL13]. Balancing [TW17]. Baltic [KVP^{+}14]. Bandit [VBW15]. Bandwidth [CCS04]. Banff [DS08]. Bao [CO12]. Bao-lu [CO12]. Bar [Kas99]. Bar-Hillel [Kas99]. Bar-Natan [Kas99]. Barlow [Blo01a]. Barnard [DeG88]. Barndorff [BN90]. BART [Car19]. Bartlett [Olk89a]. Base [FW94b, KR87]. Based [Bic88, Bre18, CDW95, Cha15, Che10, Chu17, DG95, Din17a, Din17b, Dra88a, Dra88b, God95, HTT20, HA88, HPV06, JT04, KP88, LL09a, LS23, LZ95a, LZ95b, LZD^{+}20, LL95, Mar88b, McC95, MVA^{+}18, Rei95a, Rya95, Rya03, SMJN17, Sev95, Tuk90, Wel88, XR15, ZJ08, CZL^{+}20, GW02, KM98]. Basis [Ern04, SGZW10]. Baye [Sei92]. Bayes [Bel04, Lar11, Rob11, Sch11, SX11, Zha11b, Arm90, BBSW10, Ben08, Bre90a, Bre90b, Büh14, Cai08, Daw04, DR19, DG13, Efr08a, Efr08b, Efr14, Efr19a, Efr19b, EW17, Flu90, Fra11a, FBW^{+}16, Gel09, GR19, Jen90, Jia19, Lai19, Lit11a, Lou19, MFR24a, MFR24b, MR14, Mor08, OS21, RS08a, RC04, SSVV11, SF90, Sti13, WMB19, Zel90, vdV19]. BayesBag [Büh14]. Bayesian [Ano96, Ber12, Han11a, HMRV00, Kal11, MT12, Mee11, Sed11, Slu11, ADR04, DPCA14, ATH16, BHJK21, Bar90b, BB04a, BC07a, BC07b, BFW07, Ber90b, BBW97, BJM12, Ber90c, Ber04, BGHM95a, BGHM95b, BNPS13, Boc16, BCM00, CC24, CV95, CR17, CFL13, Cox90b, CW93a, Dem93, DPL15, DuM12a, ES95, Eva07, FJ16, Fie11a, FPDE14, Fre90, Fri95, GC95, GW02, Gel05, Gel07a, Gel10, Gel11, Gel14, Gey95, GV23, GST97, GS93, GLSS23, HT00b, HBG15, HDD24, Hil93, HRR11, HMRV99, How98, IRÁ22, JHS05, JLZL04, Joh07, JAAH14, Kad90b, Kad09, KG89, KVP^{+}14, KN18, LLS11, LL07, Leh90b, LYL22, Lin90a, Lin90b, LD18, LD24, LGA^{+}15, Mad93, MVA^{+}18, Mor14, Mou90, MQ04, Nor93, PHYW22, Pea93, POSD13, RAG14, RFIS24, Rao11a, RBGK14, RR16, RP24, RSG95, RR23]. Bayesian [Sob93, SDLC93a, SDLC93b, Spi04, Sti18, SBCC^{+}24, SCtW^{+}09, TW10, Teh16, Tha10, Tit04, VGB14, Wal04, Whi93a, WCL^{+}24, Won95, Yu95, Zas11, ZSCW06]. Be [Kas10, Nic16, RBGK14, Hwa90]. Beach [Smi02]. Before [Kit14]. beginnings [Gre97b]. Behind [Dem93]. Being [AMV19, Muk22]. Beliefs [MZL10]. Bell [Gil14, Mal03]. Belt [LRdP^{+}04]. Benchmark [CR17]. Benefit [AC89, Beg89, Ber89c, Cor89, Har89, KG89, LW89, Roy89, War89a, War89b]. Benefits [MHdL23, VBW15]. Benford [PT11]. Benoît [Taq13]. Berger [Wol04b]. Bergström [Råd97]. Berk [Lit19]. Bernard [Sha18]. Bernoulli [BF15]. Berry [SII12]. Bertsimas [CTB20]. Best [CCS04, HTT20, Hol88]. Best-Effort [CCS04]. Better [vDM10]. Between [Ban88, Bar88b, Ber88, Goo88a, Goo88b, Mol92, PS11, Smi92b, Sup88, LRI23, BZ21, DKN18, Kat02, OS21]. Beyond [Bes05, SLR15, TOS22, TJL23, Whe21]. Bi [OW16]. Bi-Cross-Validation [OW16]. Bias [ATH16, DB92, Gre09, Pac10, GST97]. Biased [Atk14]. Biased-Coin [Atk14]. Biases [MG20a, Ros23a, MG20b]. Bible [Kas99, BHBNKM99]. Bibliography [Fre86, FJH^{+}89, GZ86a, GZ86b, Hog86a, Mor86, Sha86a, Win86]. Bibliometrics [Sil09]. Bickel [Rit11]. Bid [BBK06]. Bidding [BBK06]. Big [Büh14, DHM^{+}17, Goo11, Kas11b, LY21, McC11, MR14, Ste11]. Big-Data [LY21]. Bill [Bra07, Goo07, Mar07, BF09]. Billard [Muk17]. Binary [AWJ^{+}18, CR17, SSP^{+}19, SJL^{+}12]. Binding [JLZL04]. Bing [SWY00]. Binomial [And22, BCD01, RLQ17, TT23, TGM24]. Binormal [Wal14b]. Bioconductor [LM14a]. Bioequivalence [BH96, Lin98]. Biography [Ano91a, Ano91b, Bel04, BG07]. Biologic [Mar88a, Men88]. Biological [Hat88, Joh88]. Biology [FYL10, WLLH21]. Biometric [Zub87]. Biostatistics [Bre90a, EGG97, Arm90, Bre90b, Flu90, Jen90, SF90, Zel90]. Bipartite [ZP21]. Birnbaum [Bjø14, Daw14, Eva14, Han14b, May14b, Mar90, May14a]. Birth [Bel04]. Black [HR11]. Blackwell [DeG86b]. Blind [BGLC93, Bil93, Car93, CGP^{+}93, Gen93, vZ93]. Block [BZ21, CFL13, ZLK10]. Block-Conditional [ZLK10]. blood [RS97]. Blue [Gel14]. Blue-State [Gel14]. BLUP [Cam91, Har91b, Rob91a, Rob91b, Spa91, Spe91, SK91, Tho91]. Board [Ano16a, Ano23a, Ano23b, Ano23c, Ano24a, Ano24b, Ano24c]. Book [WRR94, vDM10, Leh97]. Boosting [BH07a, BH07b, BMW07, Has07, SCtW^{+}09]. Bootstrap [Ber03a, Boo03, Cas03b, DHY03, DE96, ET86a, Efr03, EH03, Hal03, Hor03, Lah03, Lel03, Pol03, Sha03, SS03, You94a, Ber94a, ET86b, Efr94, GCL94, Har86, Hin94, MM94, Nav94, Sch94b, You94b]. Bootstrapping [Hol03a]. Bootstraps [Büh02]. Borch [JL13]. Borderline [SRR^{+}17b]. Borel [Daw04]. Boris [SS92]. Boundary [LW06]. bounded [Man02]. Bounds [RR14a, RHGF14]. Bowker [Olk87a]. Box [ARC21, DeG87b]. Boxi [Gei86]. Bradbury [GB10]. Bradley [Hol01, Hol03b]. Brain [WLLH21]. Breakdown [HRV08]. Breast [BC20]. Breiman [Ols01]. Brendan [Kas99]. Bricks [RR16]. bridge [GM98]. Bridges [HDD24]. Brief [Che94, Fie92]. Brillinger [Pan11]. Brown [Lit19, BD19, Das05, Joh19]. Brownian [Taq13]. Bru [Sha18]. Bruce [Sam99]. Buchanan [Smi02]. Buja [Lit19, Law19]. Bunts [Hod17]. Burden [DPCA14]. Burn [BS97]. Burn-In [BS97]. Business [Box90a, Dem90a, Moo90c, Net90, Pra90, Rob90a, Rob90b].

C [CT98, CS12c, Ell97b, FC11, GTWT12, LP11a, PT10, SR12, SWY00, ZLG15]. C. [CT98, DeG87a, Kad09, SWY00]. Calculations [Bes05, MH22]. Calculus [PB14]. Calibrated [FPBT22, Lar11, Lit11a, Sch11]. Calibration [CY14b]. California [BGST04]. Caliński [AB15]. Cam [Yan99]. Can [BR94a, FHJ08, Gra93a, Hwa90, JW94, Men05, RBGK14, Ros23a, BR94b, Bre94b, Con93, DS94, EFK94, FW94a, Gas93, Gra93b, LL94, Rob93a, Ste94]. Canada [BG99, BG05]. Cancer [BL10, BC20, Bre88b, DJK^{+}20, DuM88a, FZ88a, FZ88b, Has88, KT88, MD88, SW16, TR10]. Cannot [Gio23].

Capacities [AS21]. Capture [BF16, FSB16, SB16]. Capturing [CBD^{+}15]. Carcinogenicity [LGF95]. Care [Lin91, Spi04]. Career [Ans03]. Careful [Fie06a]. Carlo [Mir01, CTSH92, FBPR18, FHS15, FHJ08, Gel92, GR92b, GT03, Gey92a, Gey92b, HLO05, HJ01, JHS05, JJN13, LCL13, Mad92, Owe19, PHYW22, Pol92, RP92, RL92, Rob95, RC11, Ros92, Sch92, Tho94a, Tie92, WWZ16]. Carpentry [FM93, HL93a, HL93b, Mul93]. Carroll [LC24, Por94, Sen93]. Cascades [McK21]. cascading [LS02]. Case [ATH16, BS90, Car11, CCLC09, Con93, DHR^{+}13, Gas93, Gra93a, Gra93b, Law16, LMJ12, Rob93a, SFR10, SGKR06, TGM24]. Case-Control [CCLC09]. Case-Specific [LMJ12]. Cases [Ber91a, Ber91b, Che91, Eve91, Kay91, Lan91]. Categorical [FK03]. Caught [Sam91]. Causal [Car19, CC24, CMC^{+}24, Din17a, Din17b, DL18a, DHS^{+}19a, Fie06a, GS93, GPR99, GvdL19, Her19, Hol88, IKY10, Jen19, JD21, KEHSY19, Kor06, LMR20a, LG17, MHdL23, OV14, RR14b, Rub90, Rub06a, RW06, SSM23, SSP^{+}19, Shp14, SZ14, Stu10, TYC^{+}24, VMC20, VBBG11, ZKS19, ZP21, KB98, LS02, LMR20b]. Causality [Büh20a, Büh20b, Daw04, Gil14, MMT07, Pea93, Sup88]. causation [Fre99]. Cause [MAK06]. Causes [BDN94, KL08, Lap86]. Celebrate [Bel04]. Celebrating [SII12]. Cell [AWZ20, LZD^{+}20]. Censored [LS23, SQH18]. Censoring [Fie06a, RR23, Rub06a]. Census [BR94a, BR94b, Bre94a, Bre94b, DS94, EFK94, FW94a, FW94b, LL94, Ste94, Dem86, Eri86, Fel86, FN86a, FN86b, Kad86, Mad86, Mos86, Nat86, Wol86]. Centenary [Sen93]. Central [Doo86, Le 86a, Le 86b, Pol86, Tro86]. Centroids [THNC03]. Century [BG05, IR23, MFR24a, RP24, Sti91, Efr98]. Cepheid [BJM12]. Certainty [TGM24]. Chain [CW93b, CTSH92, FHJ08, Gel92, GR92b, GT03, Gey92a, Gey92b, HJ01, JHS05, JJN13, Mad92, NA94, PHYW22, Pol92, RP92, RL92, Rob95, RC11, Ros92, Sch92, Tie92, WWZ16]. Chains [Ald93, BR02, Coh91, Mir01, Sen16]. Challenge [DS08, FHK12]. Challenges [BD04, BGGM06, CKLP22, CEN^{+}19, Fri07, GS06, JS06b, PHYW22, Pin13, SGKR03, SDD06, TWDM13, VBW15]. Chance [Kad90a, Seg07]. Change [Ber03c, NHZ16]. Change-Point [NHZ16]. Changes [CP94]. Changing [Dav88, Lil88]. Chaos [Ber92c, Ber92d, CY92a, CY92b, Cut92, Gew92, Gra92, Gri92, Law16, Smi92b, Tsa92]. Chapters [Fie92]. character [HS98]. Characterizing [SCtW^{+}09]. Charles [DeG86a, GS12]. Chatterjee [BM07]. Checking [BC07a, BC07b, Eva07, Gel07a, Joh07, LL07, NWEE20]. Checks [LL07]. Chemical [ALS^{+}01]. Chen [AGG90a, AGG90b, Bar90a, Che90, Ste90, Wat90]. Chen-Stein [AGG90a, AGG90b, Bar90a, Che90, Ste90, Wat90]. Chernoff [Bat96]. Chess [Seg07]. Chevalier [BG24]. Chiang [Li99]. Children [BGST04, QOS19, RS97]. Chin [Li99]. China [CO12]. Choice [DTZ17, Dun17, Ros99, TW17, YB17]. Choosing [CM91a, MM22, CM91b, GJSE91, GH91, Har91a, Jon91, Sil91]. Choquet [AS21]. Chow [YZ06]. Chris [GK06]. Chromatin [LZD^{+}20]. Churchill [Olk92a]. Citation [AET09a, AET09b, Hal09, LLJ09, SG09, Sti94]. Civic [Ans03]. Clarification [Ano94b]. Class [THNC03, XR15]. Class-Based [XR15]. Classes [Tan23]. Classic [Arr88, Bra88, Dem88a, Gup88, Hog88, Hot88c, Hot88b, Hot88a, Moo88, Olk88, Zid88]. Classical [DZ91]. Classification [Ano89b]. Classifier [Fri06, Gay06, Han06a, Han06b, Hol06, Sti06b]. Classifiers [GNR^{+}24]. Clay [RR16]. Clickstream [GL06]. Climate [Ber03c, DJ24, Ste20]. Climbing [vDM10]. Clinical [BC91a, Ber04, Ber12, Bya91, Che10, DuM12a, DMJ23, Dup91, Gio23, Har89, JBM^{+}06, LZ00, Lev91, Lin91, MT12, MA87, MC92, NS07, RLLKV23b, RLLKV23a, RS08b, Roy91a, Roy91b, Sim91, Tha10, TR10, TX23, VBW15, Zel91]. Clinician [KB98]. Close [BSA16, Han14a]. Close-Kin [BSA16]. Closed [DZ91]. Closer [Lan15]. Cluster [BJP13, IKN09a, IKN09b, MTP^{+}23, ZS09]. Cluster-Randomized [IKN09a, IKN09b, ZS09]. Clustered [AMV19, HB08]. Clustering [Ano89b, Cha15, ZJ08]. Clyde [HMRV99]. CODA [RBGK14]. Code [BHBNKM99, Kas99]. Codifying [Kad90a]. Cognitive [KBH20, McC94, MG20a, MG20b]. Cohen [Nau16]. Cohort [SJL^{+}12, SW16, GL96]. Coin [Atk14]. Cointegration [DKN18]. Cokriging [ZC15]. Collaborators [Leh93]. Collapsibility [GPR99]. Collecting [BGGM06]. Collection [WGHR06, Wol04a]. Collinearities [HV87]. Collinearity [Bel87, Ste87a, Bel87, HV87, Mar87b, Ste87b, Thi87]. Combination [WGHR06]. Combinatorial [BJP13, Ste93c]. Combining [CMC^{+}24, GZ86a, LR17, PGP09, Fre86, GZ86b, Hog86a, Mor86, Sha86a, Win86]. Come [Tuk90]. Comment [AG05, Aka87, Aka90, Ald90, Ama89, Ama94, And87, And91, AC89, Arm90, Arr88, Atk86, AS21, Bai91, BDN94, Ban88, Bar90a, BK87, Bar88b, Bar89, Bar90b, BN88, BN89, BN90, Bar94, Bar92b, BC91a, BJM06, Bas87, Bat86, BC91b, Bay87, Bay88, BB91, BT88, BH92, Beg89, BC07c, Bel87, Ben08, Ber94a, Ber86, Ber88, Ber90a, Ber90b, Ber99, BC05, Ber89b, Ber89a, Ber90c, Ber09, Ber89c, Ber94b, BV23, BCPR08, BHP15, Bha87, Bic88, BG94, Bil86, Bil93, Bla88, Boo89, BS06b, Box90a, Box90b, Bra88, Bra86, BO07, Bre88a, Bre94c, Bre87, Bre88b, Bri86, Bri87a, Bri87b, Bri94, Bro93, Bro90, BCD05, BFK86, BH90, BMW07, Bur92, Bur89, But90, Bya91, Cai08, Cam86, Cam91, CG93, Car19, Car93, CB87, CDW95, Cha86]. Comment [CP94, CCCH08, Che90, Che91, CZL^{+}20, Chr07, Cla90, CN91, Cli90, Cna93, Coh90, Coh91, Con88, Con93, Coo86, Cor87, Cor89, Cox87, Cox90b, Cox91, Cre86, CK94, Cri90, CG92, CH89, CTSH92, Cut92, Dah87, Dal89, Dan88, DKK19, Daw86a, Daw86b, DG95, Daw91, DP93, DeM92, Dem88a, Dem90a, Dem86, DK87, Dem90b, DH92, Dem93, DLM06, DE86, Dia91, DS94, Don91, Doo86, Dos08, Dra87, DM93, DuM88a, DuM88b, DMJ23, Dud89, Duf92, Duf93, Dup91, Eas89, Eat87, Edd87, Efr90, Efr94, Ell92, EF92, Eri86, EFK94, Eub88, Eva07, Eve91, Fai89, FM93, FC92, Fay94, Fel86, Fel87, FN06, Fie88, Fie06a, Fin86, Fis86c, Fis88, Fis90, Fle87, Flo89, Flu90, Fre87, Fre89, FJL^{+}19, Fre86, Fre90, Fri06, Fri95, GR93, GJSE91, Gas88, Gas93]. Comment [Gay06, Gei87a, Gei87b, Gei87c, Gel92, GC95, Gel05, Gel07a, Gen93, Geo90, Ger87, Gew88, Gew92, Gey95, Gil86, GZ89, GV88, Gio23, Gla87, Gli91, GS93, GK87, Gna94, God95, Goe86, Gol87a, Gol88, Gol87b, Goo86c, Goo86d, Gow90, GCL94, Gra92, Gre87b, Gre87a, Gre91, Gre10, GR19, Gri92, GvdL19, GH91, Gup88, Gut87, Hac90, HV87, Hah93, Hal09, Har89, Har91a, Har86, Har91b, Has88, HT88, HZ06, Has07, Hed88, Her19, HA88, Hey86, HJ19, Hil89, Hil93, HS09, HW88, Hin94, HK86, Hog86a, Hog88, Hog93, Hol87, Hol88, Hol92, Hol94a, Hol94b, Hot88a, Hou88, Hub86, Hub87a, Hub87b, Hwa90, Hym91, IR23, Ize87, Jen90, Jen23, Jen19, Jia19, Joe89, JY89, Joh07, Jon91, JJ08, Jor89, Kad86, Kad87a, Kad87b, Kad90a]. Comment [Kad90b, KBH20, KT88, KEHSY19, KS93, Kas88, KG89, Kas09, Kas10, Kay87, Kay91, KM91, Ken86, Ken89c, KR87, Koe06, KP88, Kor06, Kre92, Kro92, Kru90, Kün08, Lag92, LL87, LPT88, Lai19, Lan91, LL07, Law19, LMN06, Leh90b, LLJ09, Lem94, LTH90, Let08, Lev91, LC08, Lew94, Li07, LN07, Li20, LW89, LW92, Lin86, Lin87a, Lin91, Lin09, LL95, Lit07, Lit19, LM21, Loh07, Lou92, Lou19, Lub91, LL94, MS90, Mad86, Mad87, Mad93, Mad92, MP91, Man92, Mar89, Mar87a, Mar87b, Mar88b, McC94, McC86, McC89, McC95, Meh92, MM94, MA87, Mit88, Mit89, Mol92, MD88, Moo88, Moo90c, Moo90b, Mor86, Mor89b, Mor91e, Mor08, Mos86, Mos91, Mou90, Mul93, MZ05, NP93, Nat86, Nav94, Nel86, Net90, Nor93, Nor91, Not88, O’H89]. Comment [Oga18, Olk88, Orc93, Owe88, OKS89, Owe19, Par92, Pea93, Per87, Pfe07, PY05, Pic89, Pla89, Pol86, Pol92, Por87, Pra86, Pra90, PM93, Pre87, Pyk89, RP92, Raf89, RL92, Raf92, Rao88, Rao89, RF89, Ric86, RS08a, RM07, RTW19, Rip94, Ris87, Rob89, Rob88, Rob93a, RSG95, RSLGR07, Rob93b, Ros06, RR88, Ros92, Roy89, Rub90, Ruk90, Rya95, Sak89, Sam86, Sam89, Saw91, Sch94a, SC94, Sch94b, Sch92, Sch18, Sco86, Sei86, Sen87, Sen09, Sev95, Sha86a, Sha87a, Sha20, Sha21, Shi87, Sil91, Sil09, Sim91, Sin94, Sin86, Sin87, Sin89, Ski94, Sma89, Smi87, Smi89a, Smi92a, Smi92b, Sob93, Sol87, Sol86, SW93, Sol94, Sør05, Spa91, Spe91, Spi87b, SF90, SG09, Spu88, Sta94, Ste94, Ste90, SK91, Ste89, SH19]. Comment [Sti06b, Sto86b, Sto86a, Sto89, Str90, Stu90, Sud94, Sud86, Sui92, Sun89, Sup86, Sup88, Tan07a, Tan90, Thi87, Tho91, Tho94b, Tho05, Tho94c, Tib94, Tie88, Tie92, Tit86, TX23, Tro86, Tsa92, TD07, Tuf87, Tuk87, Tur87, Utt86, Vel86, Ver87, Wah86, Wah88, Wah06, Wai87, WB90, Wan06, WMB19, Wat90, Wat87, Wat89, Weg86a, Weg89, Wei91, Wei87, Wei94, Wei86, Wei89, Wel86, Wel88, Whe21, WSC19, Whi93a, Win86, Win90, Wit87, Wol90, Wol86, Won95, Woo93, Wu93, Wyn93, Yan08, You90, You91, Yu95, Yu06, Zah91, Zas94, ZLH93, Zel90, Zel91, Zel87, Zel09, ZS09, ZKS19, ZL08, Zhu18, Zid88, Zid89, Zub87, deL92, vZ93, vdV19, vS94, Ell98]. Commentary [JT90]. Comments [Ano96, LB22, SR14, SAA^{+}92, AG01, ACS01, Bre01, HK02, HT00b, HZ00, HMRV99, SWY00]. Commerce [BS06a, BGGM06, Fie06b, GS06, JS06a, JS06b]. Committee [BGLC93, Bil93, Car93, CGP^{+}93, Gen93, Mar07, SAA^{+}92, vZ93]. Common [ZT09]. Communicating [Win90]. Communication [Law16]. Community [GM21a, LCX21]. Comparative [BBSW10, BNPS13, BL20, HB08, Lan15, OS22, YK16]. Compare [KAX^{+}18]. Comparing [BT88, BB04b, GV88, HJM88a, HJM88b, Not88, Owe88, SSS00, Spu88]. Comparison [Cat12, CFL13, FPW16, GST18, KS07a, KS07b, RAP14, RM07, Sav88, SW16, TD07, WV94, ZZS21]. Comparisons [HTT20, PA22, Tuk91]. Compatibility [CV08]. Competition [Car19, DHS^{+}19a, GvdL19, Wal14a, ZKS19, QM02]. Competition-Winning [ZKS19]. Competitions [Her19, KEHSY19]. Compilation [Lan14]. Complementary [LSA15]. Complex [Con93, Fie06a, Gas93, Gra93a, Gra93b, Gri92, MVA^{+}18, RBGK14, Rob93a, STG13, SW17, Tha10, ZT09, BBC^{+}02]. Complexity [RR17, SRR^{+}17a, Sup88]. complicated [GG02]. Component [GZZ20, JJN13, MTNO15, SRR^{+}17a]. Component-Wise [JJN13]. Components [BFWW09, CF08, TJL23]. Composition [JJN13]. Compositional [GGBS^{+}23]. Compound [GV23, GR19]. Compounds [ALS^{+}01]. compression [PCT^{+}98]. Computability [Vov09, PK97]. Computation [BGHM95a, BGHM95b, BNPS13, BOG^{+}19a, BOG^{+}19b, CR17, Fri95, GC95, Gey95, HDD24, HJ19, JCD^{+}05, MVA^{+}18, RSG95, SH19, TW10, Wan12, WCL^{+}24, Won95, Yu95]. Computational [APSAS17, ABB^{+}23, ADR04, BD04, FYL10, JLZL04, Raf92, SVS11, QM02]. Computations [RP24]. Compute [Lin19, TNM22]. Computer [BBF^{+}22, CLSW16, DJK^{+}01, Eas89, JY89, Mor89b, O’H89, OKS89, SWMW89a, SWMW89b, Ste89, BBC^{+}02]. Computers [Edd86a, WWZ16, Bat86, Bil86, BFK86, Edd86b, Goe86, Sco86, Utt86, Weg86a]. Computing [KZZW22, MFR24b, Thi86a, Cha86, Hub86, Thi86b]. Concave [Sam18, ZZ12, Wal09]. Concept [IRÁ22, FT96]. Concepts [PW15]. Concern [BDN94]. Concerning [Fra14]. Concert [Sin16b]. Concluding [Ste93b]. Concordance [LGF95]. Conditional [CDW95, Fra04, GS93, KG05, LN04, ZLK10]. Conditionally [ACS01, DR19]. Conditioned [Bel87]. Conditioning [Rei95b]. Conditions [RR14a]. Conducting [SC17]. Confidence [AG05, And22, BC05, BCD05, CH12, DRRS22, ET86a, ET86b, FPBT22, Fra11a, GM05a, GM05b, Har86, LB22, LPE15, LMR20a, NNLL18, PL21, Rob11, SX11, Tho05, VW23, Zha11b, BH96, DE96, LMR20b, Man02, PL22]. Confidentiality [Cox91, DP91a, DP91b, Fie06b, KM91, Nor91, ROSS18]. Conflict [NWEE20, POSD13]. Confounding [GPR99, KC14, Pac10, Shp14, SRC^{+}16]. Confuse [KG05]. Conjugate [DR19, MVDV10]. Connected [AS10]. Connection [Kat02]. Connections [LRI23]. Consensus [BR94a, BR94b, Bre94b, DS94, EFK94, FW94a, LL94, Ste94]. Consequences [BBB^{+}19, DH92, Lit19, SR14]. Considerations [DMJ23, LZ00, RLLKV23b, SSS20, Eud96]. Consistency [Che17, MGC15, FT96]. Consistent [SZ14, vHB21]. Conspiracy [Seg07]. constants [GM98]. Constrained [BGGM06, GJ18, KM18a, MMTW01]. Constraints [JJ18, SS18]. Constructing [NS94, SRR^{+}17a]. Construction [HB17, Que24]. Constructions [Bes05]. Consulting [Twe98]. Consumer [HPV06]. Consumption [Car14a, Car14b, FS14]. Contemporary [CC14, RLQ17]. Contents [Ano06, Ano07, Ano08, Ano09, Ano13, Ano16b, Ano23f, Ano23d, Ano23e, Ano24d, Ano24e, Ano24f]. Context [Sil09]. Contexts [Lew94]. Contextuality [Grü16]. Contingency [Agr92b, Agr92a, BH92, Duf92, EF92, Kre92, LW92, Meh92, Sui92]. Continual [OC10]. Continuous [AWJ^{+}18, Bur89, FBPR18, Hei89a, Hei89b, LSS^{+}19, SAC19]. Continuous-Time [FBPR18]. Contraction [Gus05]. Contracts [Sha18]. Contrast [FSB16]. Contributing [ZT09]. Contribution [EW17, Gir14, CR96, Fel99]. Contributions [Bec07, BD19, Car19, DS92, LS86, Mar89, McC16]. Contributor [Ans03]. Contrition [GR92b]. Control [Ban93a, Ban93b, Bre87, CCLC09, Cna93, Cor87, DGA^{+}20, DP07, Duf93, Hah93, HLO05, Hog93, Hol87, KC14, Mar87a, NP93, Orc93, Rob95, Rob93b, Ros87a, Ros87b, Ros99, SRC^{+}16, TR10, Woo93, Wu93, Wyn93]. Controls [BT88, GV88, HJM88a, HJM88b, Not88, Owe88, Spu88]. Controversy [Ano94b, BDN94, BG05, Ber94b, Lem94, Lew94, PB10, Roe94a, Roe94b, SR14, Sud94, Tho94b, Wei94]. Convergence [CTSH92, HRR11, Rob95]. Conversation [Ald13, Ald18, Ans88, AB15, AM19, BN04, BM07, BS16, BS18, Bar16, Bar05, Bec09, Beh19, BK24, Bha22, BS10, BL90, BD09, Buc16, BK08, BLL^{+}21, CH15, CY14a, CJ16, CJ07, Das05, DeG86a, DeG86b, DeG86c, DeG86d, DeG86e, DeG87a, DeG87b, DeG88, De 20, DW16, DT04, Est03, FM00, FM03, Fie94, Fie08, FP95, Fis15, Flo93, Fol95, GB10, GS08, GR10, GK06, GLW07, HKL17, Hah95, HS22, Hed05, Hey95, HM95, Hol08, Hol03b, HK22, Jon19, Ket01, KNR19, KK16, Kva06, Lai89, LM14b, LP11b, LC24, LSY21, Mac93, Mar90, MSY16, MH12, MM24, Muk24, Muk02, Muk05, Muk06, Muk17, NRSV21, Nau16, New02a, NMB20, Olk87a, Olk87b, Olk89a, Olk89b, Olk91, Olk92a, Ols01, Pan11, PT10, Ran07, Rea04, Rei94, Ric15]. Conversation [Ros15, Ros23b, RB20, Rya15, Sam08, Sam07, SC92, Sen03, She92, SS92, Smi95, SO94, ST13, Swi92, Tuf88, Wal18, WM07, Wel09, Whi93b, WH19, Wit91, Wol04b, YZ06, Zab94, Zei18, ZLG15, Ale96, Ban96, Bat96, BF98, Bin96, Blo01a, BH99, CT98, Ell97b, Ell97a, FK96, Gai97, Gel97, Gre97a, Han97, Hol01, Li99, MM00, McD98, Muk97, Muk00, Nad02, Nau00, New02b, Råd97, Sam99, Sim97, Sne97, Tho02, Yan99]. Conversations [EGG97, GPW23]. Convex [BJMO12, Bru18, CHP22, LCX21]. Cook [BLL^{+}21]. Cooperation [AET09a]. Coordinate [Eat07, SMJN17]. Coordinate-Based [SMJN17]. Coordinate-Free [Eat07]. Coprime [LK20]. Copula [STG13]. Copy [ZT09]. Cormack [Buc16, BM16, Nic16]. Correction [Ano91b, Ano00, BN90, Blo01b, GS13, HMRV00, New02b, NRB17, Olk89b, WZ02]. Corrections [AWZ20]. Correlated [OWY01].

Correlation [Gal89, Sti89]. Correlations [Ald95]. Correspondence [Bar92a, LBK21]. Cosmic [GMN^{+}04, Mar04b]. Cost [APSAS17]. Costs [MHdL23]. Could [Ber03b]. Council [AET09a]. Counting [Ald93, SAL^{+}20]. Counts [CLSW16]. County [CEN^{+}19, Smi02]. County-Level [CEN^{+}19]. Coupling [ABT16, STG13]. Course [Sch04]. CoV [CKLP22]. Covariance [BHP15, BFWW09, EG22, GK15a, KMR^{+}19, Per87, Pou11, Ros02, SLR15]. Covariate [HB08, IRÁ22, MY97]. Covariate-Adjusted [IRÁ22]. Covariates [CZL^{+}20, LYL22, QKM20, RS08b]. Coverage [LMR20a, LMR20b]. COVID [CWHD22, JCT22, Lin22, WBR^{+}22, YS22]. COVID-19 [CWHD22, JCT22, Lin22, WBR^{+}22, YS22]. Cows [Her19]. Cox [ARC21, DMRT13, Rei94]. Cracked [Gel14]. Cramér [Weg86b]. Credit [Gay06]. Cressie [WH19]. Criterion [EG22]. Critical [PT03]. Criticism [DJ15]. Critique [Fre86, GZ86a, GZ86b, Hog86a, Mor86, Sha86a, Win86]. CRM [MAK06]. Cross [GK15a, IMS90, LL07, MTP^{+}23, OW16, vDM10]. Cross-Covariance [GK15a]. Cross-Fertilizing [vDM10]. Cross-Study [MTP^{+}23]. Cross-Validated [LL07]. Crossing [Lin19]. Crossover [PC04]. Cryptic [AB09]. Cultures [Bre01]. Curiosities [SB16]. Curve [Aka87, Bri87a, Dra87, Gei87a, IRÁ22, Ize87, LL87, Rao87a, Rao87b, Sch04]. Cusum [Yu95]. Cuthbert [Tuf88]. Cutting [YNS23]. cytometry [Eud96].

D [Rob11, SX11, Zha11b, QOS19]. Dangers [Tan90]. Daniel [Tuf88]. Daniels [Whi93b]. Dark [Ber94a, Efr94, GCL94, Hin94, MM94, Nav94, Sch94b, You94a, You94b]. Darrell [Ste05]. Data [ABHK00, AWZ20, AWJ^{+}18, BS06a, BGGM06, BCW87a, BCW87b, Ber92a, Ber92b, Boo86a, Boo86b, BL20, BCN^{+}23, Bre94a, Bur89, CMN^{+}19, Cam86, Car19, Cat12, Cha86, Cha22, CFL13, Cre86, DTL05, DH05, DJ24, DH92, DL18a, DHS^{+}19a, Dra87, DH24, DHM^{+}17, Edd87, Eva12, FHS15, Fie06b, FM18, Gas87a, Gas87b, Gei87b, GS06, GN17, Gla87, GŠP12, GWVV09, Gol87a, GL06, GKRS05, Goo86d, GGBS^{+}23, GvdL19, Grü16, GLSS23, GZZ20, GG09, HT00a, Hei89a, Hei89b, Her19, Hoa03, HRR11, Hub86, Hub87a, JS06a, JBM^{+}06, JLMD16, JR18, KS07a, KS07b, Kay87, Ken86, KAX^{+}18, KR87, KJED09, LY21, Lar11, LL04, LYL22, LMW07, LZD^{+}20, Lin08, LD18, Lit11a, LSS^{+}19, LR17, LS17, MRSS15, Mar11, MH16, MQ04, MSM24, NA94, NRB16, NWEE20, Par92, Par04, PGP09, Raf92]. Data [RD06, RM07, RC11, SJW11, SMJN17, SSP^{+}19, Sam86, Sch11, SV18, SCN^{+}10, SW17, Smi92a, SAC19, SDD06, SQH18, Tan23, TW10, Thi86a, Thi86b, TD07, Tuf87, Tuk87, Tuk90, VMB08, Wai87, WH16c, WGHR06, Wit87, Wol04a, YS22, ZLK10, ZZS21, ZT09, GST97, KM98, PT11, RS97, NRB17]. data-augmentation [GST97]. Data-Based [Tuk90]. Data-Constrained [BGGM06]. Data-Dependent [Grü16]. Data-Enabled [BGGM06]. Databases [ALS^{+}01, YSR20]. David [Muk24, Rei94, Swi97, BN04, Bha22, Bin96, DeG86b, HMRV99, Lai89, Mac93, MH12, Pan11]. Davison [CS12c, GTWT12, SR12]. Dawson [NRSV21]. Days [Cra86]. DC [CHP22].

Death [Fie06a, KL08, KJED09, Ros06, Rub06a]. debate [GST97]. Decades [Tuk90]. Decision [Ber90a, BD19, Bro90, Coh90, Geo90, Hwa90, IJ23, Lin98, MC90a, MC90b, MS90, RR16, Ruk90, Str90, Tha10]. Decision-Making [IJ23]. Decision-Theoretic [Ber90a, Bro90, Coh90, Geo90, Hwa90, MC90a, MC90b, MS90, Ruk90, Str90]. Decisions [GKLP99, GR19, WH16a, WH16b]. Decoding [Men09]. Decomposable [NRWY12]. Deconvolution [JCT22]. Decreasing [Raf89]. Dedication [Mor91d]. Deep [FMZ21, HKL22]. Defeating [Fei93]. Defense [ZWS21]. Defining [JSH07, LQMT13, Par23]. Definition [GN17, NRB16, NRB17]. DeGroot [Ano91a, Ano91b]. Deja [Wai87]. Delay [VACC22]. Deliberate [WSC19]. democratic [GG02]. Demography [Coh91, GGC13]. Demonstration [RSK22]. Dempster [MZL10]. Demystifying [KS07a, KS07b, RM07, TD07]. Dennis [BLL^{+}21, Smi95]. Denoising [AWZ20]. Density [Cha15, FC11, KM18a, LP11a, MT11, Sam18, She04]. Density-Based [Cha15]. Dependence [Ber92a, Ber92b, CBD^{+}15, DH92, Par92, PW15, Que24, Raf92, RN13, Seg12, Smi92a, TOS22]. Dependencies [CW93b]. Dependent [Cat12, DKK10, Grü16, QMJM22]. Depth [GN17, LL04, MM22, NRB16, NRB17, Zuo21]. Derandomization [Ste93b]. Derivation [Hil95]. Derived [YSR20]. descriptive [Ald01, MdL98]. Design [Atk14, BBI^{+}23, BL10, BQW06a, BQW06b, BGLC93, Bil93, CC14, Car93, CV95, DHM^{+}17, FN06, Gen93, KMSG06, Koe06, RFIS24, RS08b, SWMW89a, SW17, TW17, VBW15, Wan06, WM04, ZSCW06, vZ93, Che99, Fel99, MW96, Eas89, JY89, Mor89b, O’H89, OKS89, SWMW89b, Ste89]. Designed [BFWW09, Mar87a]. Designs [BT88, Che10, GV88, Har89, HJM88a, HJM88b, Jen23, LL09a, Not88, OC10, Owe88, Spu88, TW17, TX23, VMC20, ZJ24, Ros96]. Detect [Ros23a]. Detecting [DJK^{+}01, LRdP^{+}04, SCtW^{+}09]. Detection [BH02, DS08, Fai89, GM21a, Joe89, LCX21, NHZ16, PT03, Pic89, Raf89, Sin89, Smi89b, Smi89c, Wei89, PT11]. Detective [Kad09]. Determination [CC24, GW02]. Determinism [Gew92, Sup88]. Deterministic [FBPR18]. Develop [JSH07]. Developing [RN13]. Development [Eat07, Weg86b, Gre97b]. Developments [ABB^{+}23, BS12, CCL^{+}04, DHY03, GJ18, Leh90a, MC90a, Pfe13, Ber90a, Bro90, Coh90, Geo90, Hwa90, MC90b, MS90, Ruk90, Str90]. Deviance [HBG15]. deviation [Ell98]. Diaconis [Ald13, DeG86d]. Diagnosing [HV87]. Diagnostic [Atk86, VK21, Yu95]. Diagnostics [BRM11, CZL^{+}20, POSD13, RL92]. Diagrams [OV14]. Dialogue [SCN^{+}10]. Dice [BB18, Sha18]. Dick [KNR19]. did [Edw97b]. Dietary [Car14a, Car14b, FS14]. Differ [WM04]. Difference [CHP22, DTZ17, SRC^{+}16]. Difference-in-Differences [SRC^{+}16]. Difference-of-Convex [CHP22]. Differenced [DPP18, Raf92]. Differences [CS12b, OS21, SRC^{+}16, KB98]. different [MW96]. Differential [ROSS18]. Differentially [BL20]. Diffusion [BDR^{+}22, Fen16, HDD24]. Diggle [AM19].

Digit [Hil95]. Dilation [GM21b]. Dilemma [LTRG17, Pol24]. Dimension [APSAS17, BNPS13, Chr07, Coo07a, Coo07b, CF08, GŠP12, Li07, LN07, MGC15, PQW14]. Dimensional [CHP22, FWZZ21, GHV12, HBM12, JCD^{+}05, MHdL23, NRWY12, NNLL18, RBGK14, SB23, ZZ12, ZWS21, ZLS10]. Dimensions [BC22, Boo86a, Boo86b, Cam86, Cre86, Goo86d, Ken86, Sam86, vdGM12]. Ding [AOW17]. Directed [POSD13]. Directions [WCL^{+}24, Ros96]. Dirichlet [QMJM22]. Dirty [Fra11a, Rob11, SX11, Zha11b]. Disasters [Cha22]. Disciplinary [IMS90]. Disclosure [Fie06b, GKRS05]. Discover [PGP09]. Discoverer [Wai03]. Discoveries [VW23]. Discovering [GT03]. Discovery [CC24, JLZL04]. Discrepancies [PS11]. Discrepancy [LL07]. Discrete [CS12a, CFL13, DM93, FLR93a, FLR93b, JBM^{+}06, PM93, Sut03, ZLH93]. Discriminant [Ano89b]. Discrimination [Bla88, Con88, Dem88b, Dem88c, Fie88, Fis88, Gas88, Gew88, Gol88, Hol88, Rob88]. Discussed [BF15]. Discussion [Ber11, Ber12, Bjø14, Büh14, Car14b, CT11, CTB20, CS12c, Daw14, Eva14, Fie06a, FS14, Fra14, FC11, GTWT12, GK19, Goo11, Han11a, Han14b, Han11b, Hel11, Ion11, Kal11, KMSZ91, LP11a, Lar11, Lou09, ML14, McC11, MT12, Mee11, Mei11, MKV09, Rob11, SSS20, Sch11, Sed11, SR12, SX11, Slu11, Ste11, Sto14, Swe11, Tjø19, Wel14, Wes11, Yao11, Zha11b, vdG19, BG99, How98]. Discussions [DHS^{+}19b]. Disease [Bre18, DGA^{+}20, GT03, KM18b, PGP09]. Disease-Associated [PGP09]. Diseases [ZT09]. Disequilibrium [ZJZ^{+}09]. Dispersion [BB04b, RTW13]. Display [Tuk90]. Dissemination [GKRS05, ROSS18]. Distance [EG22]. Distant [Mor14]. Distributed [GLSS23]. Distribution [BR02, Car14a, Car14b, FS14, Lin19, MG08, MN11, TT23, Wal14b]. Distributionally [RB23]. Distributions [Ano89f, ACS01, BS12, DFN12, DZ91, Fre86, Gel09, GZ86a, GZ86b, HJ01, Hog86a, KMR^{+}19, Lin19, Mor86, Sha86a, Wal09, Win86, CI00]. Distributions-to [Lin19]. Divergence [KM18a]. Divergences [NWEE20]. Divergent [PQW14]. Diversity [Aka90, Ald90, Box90b, Dem90b, Efr90, Hac90, Moo90b, Sha90b, Sha90a]. Dixon [Flo93]. DNA [Ano94b, BDN94, Ber91a, Ber91b, Ber94b, BM98, Che91, Eud96, Eve91, Kay91, Lan91, Lem94, Lew94, Roe94a, Roe94b, SSS00, Sud94, Tho94b, THNC03, Wei94]. Do [Con93, DHS^{+}19a, Gas93, Gra93a, Gra93b, GvdL19, KMSZ91, MAK06, PB14, Rob93a]. Do-Calculus [PB14]. Do-It-Yourself [DHS^{+}19a, GvdL19]. Does [Fre08, Hut16, GG02]. Doesn’t [ZC15]. Don [NRSV21, SII12]. Donald [DT04, Hah95, LM14b]. Doob [Sne97]. Dorothy [WM07]. Dose [Che10, OC10, TR10, ZJ24]. Dose-Finding [Che10, OC10, ZJ24]. Double [BGLC93, Bil93, Car93, CGP^{+}93, Gen93, KS07a, KS07b, RM07, RSLGR07, SV18, SB23, TD07, Wal14b, vZ93]. Double-Blind [BGLC93, Bil93, Car93, CGP^{+}93, Gen93, vZ93].

Double-Estimation-Friendly [SB23]. Double-Robust [RSLGR07]. Doubly [DELL14, Li20, LGA^{+}15]. Doubly-Intractable [LGA^{+}15]. Doubly-Robust [Li20]. DR [Tan07a]. Draper [HMRV99]. Drawer [Bay88, DuM88b, Hed88, IG88a, IG88b, LPT88, Rao88, RR88]. Draws [vHB21]. Drift [Rit11]. Dror [Kas99]. Drug [HR16]. Drugs [Ize01]. Dual [HJJ^{+}21, LSA15]. Dudley [KNR19]. Due [Fie06a, Rub06a]. DuMouchel [Ber12, MT12]. Dupin [Kad09]. During [CWHD22, Muk22]. Dynamic [BL10, BCW87a, Bri08a, Bri08b, Kün08, Peñ06, STLD14, Sin95, WBR^{+}22, Yan08, MW96, BCW87b, Edd87, Hub87a, Tuf87, Tuk87, Wai87]. Dynamics [Bre18, BNF^{+}07, KM18b, RD06].

e-Commerce [BGGM06, Fie06b, GS06]. E. [HMRV99]. Early [Cra86, Dav98a, Fie08, Kad09, Leh93, McC16, Tha10, Edw97a]. Earthquake [BS13, Oga13]. eBay [BBK06]. ECMO [AC89, Beg89, Ber89c, Cor89, Har89, KG89, LW89, Roy89, War89a, War89b]. École [Maz15]. Ecology [FPW16]. Econometrician [Imb14b]. Econometrics [Fan05a, Fan05b, Hor03, MZ05, PY05, Sør05]. Economic [Box90a, Dem90a, Gew88, HM16, Moo90c, Net90, Pra90, Rob90a, Rob90b]. Economics [JJ18]. Ederer [Gre97a]. Edge [CCFR24]. Editor [Ano91c, Mor92, Olk92b, Swi97]. Editorial [Ano86a, Ano16a, Ano23a, Ano23b, Ano23c, Ano24a, Ano24b, Ano24c, CDR23, DSW13, Kas92a, RP24, SS18]. Editors [SCN^{+}10]. EEG [GZZ20]. Effect [BCvdL20, BCN^{+}23, DTL05, Hol88, Lan15, Li20, Men88, PC12, SSP^{+}19]. Effective [KK15]. Effectiveness [MAK06, Owe19]. Effects [Ban20, BGST04, Cam91, CZL^{+}20, Chu17, DJ15, FC11, GWVV09, Har91b, Hed92, Hof21, IKY10, JD21, JSH07, LP11a, LG17, MY97, MN11, MVDV10, MT11, QKM20, RHGF14, Rob91a, Rob91b, RW06, Spa91, Spe91, SZ14, SK91, Tho91, WM04, RS97]. Efficiency [Ber04, LMJ12, SGKR06]. Efficient [BCvdL20, GHSS20, LLV10, Tan23, Wol04a, Edw97a]. Effort [CCS04]. Efron [Hol03b]. Eisenhart [Olk92a]. El-Shaarawi [Est03]. Elaboration [Hol06]. Election [Car11, AP02]. elections [CG02]. Electoral [HR11, PT11, QM02]. Electromagnetic [Mor88]. Electronic [BS06a, JS06a, JS06b, Mar11]. Elfving [Che99, Fel99, Nor99]. Elicitation [BQW06a, BQW06b, FN06, Koe06, Wan06]. Eliminating [SWY00, BLW99]. Elizabeth [BF91, FM00]. Elliptical [FMF13]. Elo [Ald17]. Elston [ZLG15]. EM-Type [SGZW10]. Emanuel [New02b, New02a]. Embedding [BNF^{+}07, TJL23]. Emergence [TW10, Fel99]. Emerging [WCL^{+}24]. Emperor [Ber99, PW99]. Emphasis [BM24, Cat12, ES95, Ano96]. Empirical [BBSW10, Ben08, BPV20, BBK06, Cai08, CH89, Daw04, Dud89, Efr08a, Efr08b, Efr14, Efr19a, Efr19b, GZ89, GR19, Jen19, Jia19, Lai19, Lou19, Mor08, Pol89a, Pol89b, Pyk89, RS08a, VMC20, WMB19, WM04, vdV19]. Employment [Dem88b, Bla88, Con88, Dem88c, Fie88, Fis88, Gas88, Gew88, Gol88, Hol88, Rob88]. Enabled [BGGM06]. Endogenous [CZL^{+}20, QKM20]. Energy [Pin13, vDK04]. Engage [DN06]. Engineering [Law16, Sco17]. Enhancing [DP91a, Lan14, Cox91, DP91b, KM91, Nor91]. Ensemble [STG13]. Entropy [LSA15]. envelope [STW00]. Environment [Ano03a, CK03, KLDR09]. Environmental [Bri03, Dav88, Fai89, Joe89, Lil88, Pic89, Raf89, Rya03, Sin89, Smi89b, Smi89c, Wei89, WM04, PSES98]. Environments [Cha86, Hub86, KZZW22, Lub91, Sin95, Thi86a, Thi86b]. Ephedra [Mor05]. Epic [Sti07a]. Epidemic [BDP18, FJH^{+}89, GST18, KN18, MVA^{+}18]. epidemiologic [GL96]. Epidemiological [KAX^{+}18]. Epidemiologically [Rya03]. Epidemiologist [SH14]. Epidemiology [Dav88, DMS10, FPW16, Lil88, Sav88]. Epilogue [Leh93]. Equality [GGM09]. Equations [NNLL18]. Equidistant [WRR94]. Equilibrium [RR14a]. Equitability [RRSM20]. equivalence [BH96]. Ergodicity [JJN13]. Erich [DeG86c]. Error [CK03, DP07, HKL22, PK97]. Errors [ET86a, ET86b, Har86, OWY01, Sti08, vHB21]. Escalation [TR10]. Especially [DJ15]. Essay [Fie92, Rub90, SNDS90, Sti13]. Esscher [LRI23]. Essential [HS09, IKN09a, ZS09]. establishment [Dav98b]. Estate [KK16]. Esther [SC92]. Estimate [CBD^{+}15, MG08]. Estimated [LMR20a, LMR20b]. Estimates [CEH^{+}17, CEN^{+}19, LGA^{+}15, SGKR06]. Estimating [Car14a, Car14b, CDW95, CPW15, DG95, FS14, FC11, God95, GWVV09, HR16, JSH07, KS07a, KS07b, KAX^{+}18, LP11a, Lel03, LZ95a, LZ95b, LL95, McC95, MT11, NNLL18, Rei95a, RM07, RW06, Rya95, STLD14, SS99, Sev95, TD07]. Estimation [DPCA14, ATPP21, Ber90a, BS90, BS12, BO17, BCM00, Bro90, BCD01, Bru18, Cai12, Cai19, CP14, Cam91, CCS04, CHP22, CJLV06a, CJLV06b, CS12a, Coh90, CK94, DTZ17, DG12, DTL05, DLM06, DL18b, Efr14, FPDE14, FW12, GM21a, Geo90, GLX12, GS06, GR94a, GR94b, GG09, Har91b, Hol94b, Hwa90, JCT22, Jia19, KDS^{+}15, KMR^{+}19, KM18a, Lah03, LMN06, LG17, MC90a, MC90b, MS90, Mey18a, ML12, Pfe13, Pou11, RRSM20, RT12, Rob91a, Rob91b, Ruk90, Sam18, SC94, SB23, She04, Sin94, STR^{+}17, Spa91, Spe91, Sta94, SK91, Str90, SH14, Tho91, Tho94c, VBBG11, Yu06, ZZ12, vdGM12, Edw97a, KB98, MY97, SWY00, VJ14]. Estimator [BCvdL20, CM91a, CM91b, GJSE91, GH91, Har91a, IRÁ22, Jon91, Men05, Sil91, SZ14]. Estimators [DD21, FW12, GŠP12, Li20, NRWY12, Pac10, RSLGR07, Sch04, Sti90, PK97]. Ethernet [WTLW95]. Ethical [Har89]. Ethics [Ber89c, Ber04, Roy91a, BC91a, Bya91, Dup91, Lev91, Lin91, Roy91b, Sim91, Zel91]. Ethos [Zel91]. Etiology [RR14b]. Euclidean [Ste93c]. Eugenio [LP11b]. Evaluating [Fle92a, GS06, Joh97, CG92, DeM92, Ell92, FC92, Fle92b, Lag92, Lou92]. Evaluation [DELL14, IKN09a, Jen19, LTRG17, Lin20, Spi04, ZS09, BBC^{+}02]. Evans [Ano96]. Evening [BF09]. Event [LW06, Peñ06]. Events [Lap86, LTRG17, VACC22]. Everyone [Hod17]. Everything [SRS91]. Everywhere [Arn07]. Evidence [DPCA14, BDP18, Efr10a, Efr10b, GS06, HR11, Kas10, Kay91, POSD13, Ros17, WM04]. Evolution [BH87a, BH87b, CKLP22, Fel87, Por87, SDD06]. Evolutionary [DJK^{+}20, Kar92]. Ewens [ABT16, Cra16a, Cra16b, FJ16, Fen16, McC16, Teh16]. EWOC [TR10]. Exact [Agr92b, Agr92a, BH92, Duf92, EF92, EH03, Ern04, Kre92, LW92, Meh92, Sui92]. Exactly [GM18]. Examine [BCN^{+}23]. Example [Mor88]. Examples [DG13, Han14a, JT04]. Exceedance [Raf89]. Exchange [SY12]. Exchangeability [Dra87, GR19, SSM23]. exchangeable [LS02]. Exciting [Mey18b, Oga18, Rei18a, Rei18b, Sch18, Zhu18]. exemplified [GST97]. Exercise [Sil09]. Exit [PS11]. Expansion [Gus05, LLV10]. Expected [JL13]. Experience [CTSH92]. Experiences [BGGM06, CG92, DeM92, Ell92, FC92, Fle92a, Fle92b, KVP^{+}14, Lag92, Lou92]. Experiment [BGLC93, Bil93, Car93, Gen93, Lan15, vZ93]. Experimental [BBI^{+}23, CV95, DHM^{+}17, RFIS24, BCN^{+}23, Che99, DWZ97, Fel99]. Experiments [Bai17, BBSW10, BFWW09, CLSW16, DK20, DSB03, Eas89, Fie06a, IKN09a, IKN09b, JY89, JCD^{+}05, KT88, Mar87a, Mar21, Mor89b, O’H89, OKS89, Ros02, Rub90, SWMW89a, SWMW89b, SNDS90, Ste89, ZS09]. Expert [BQW06a, BQW06b, CW93a, DK87, Dem93, FN06, GS93, Hil93, Koe06, Lin87c, Lin87a, Lin87b, Mad93, Nor93, Pea93, Sha87a, Sha87b, Sha87c, Sob93, Spi87a, Spi87b, Spi87c, SDLC93a, SDLC93b, Wan06, Wat87, Whi93a]. Explain [PB10, Shm10]. Explained [LRR17]. Explanation [BZ12]. Exploiting [ABT16]. Exploration [HJ01, vDM10]. Exploratory [BH90, Cri90, Fis90, GS11a, Gow90, Hel11, Mei11, Stu90, WS90a, WS90b, Wes11, You90]. Exponential [BCPR08, DKSC08a, DKSC08b, JJ08, RT12, SKBS20]. Exponential-Family [SKBS20]. Exposition [Bic88, Dra88a, Dra88b, HA88, KP88, Mar88b, Wel88]. Exposure [Men88, Tei88]. Exposures [SJL^{+}12]. Expression [LZD^{+}20]. Expressions [Cla90, Cli90, Kad90a, Kru90, MY90a, MY90b, Tan90, WB90, Win90, Wol90]. Extended [HK02]. Extending [DMRT13]. Extensions [ARC21, CZL^{+}20, XYY19]. Extensive [HTT20]. Extent [BBK06]. External [PB14]. Extracting [Yu95]. Extremal [PW15]. Extreme [LTRG17, Raf89, Smi89b, Fai89, HT00a, Joe89, Pic89, Sin89, Smi89c, Wei89]. Extreme-Value [HT00a]. Extremes [CS12c, DPR12b, GTWT12, SR12]. Extropy [LSA15]. Eyes [Han14a].

F [Bar16]. F. [Lai89]. Factor [AY01, DG13, EW17, FWZZ21, OS21, OW16]. Factorial [Bai17]. Factors [Ros17, SSVV11]. failure [MY97]. failures [LS02]. Fairness [IJ23]. Faithfulness [SZ14]. False [LL09b, Per05]. Families [BCPR08, DKSC08a, DKSC08b, JJ08]. Family [LL09a, MVDV10, SKBS20]. Family-Based [LL09a]. Fano [GMS20]. Faster [CRSW13, vHB21]. Father [Gre97b]. Feasible [ÁEdBCAM17]. Feature [BJP13, CT11, Han11b, Ion11, XT11a, Yao11]. Features [Car19, Lin19].

Federer [Hed05]. Feedback [YNS23, But00]. Feller [ABT16]. Fence [Hod17, Mol92]. Fertilizing [vDM10]. Few [SSP^{+}19]. Fiducial [Sei92, Zab92]. Field [DR19, vDM10]. Fields [Mor88, SAY22]. Fienberg [Han11a, Kal11, ST13]. Fifty [Ste05]. Figure [FHJ08]. File [Bay88, DuM88b, Hed88, IG88a, IG88b, LPT88, Rao88, RR88]. Filter [Spa91]. Filtering [Sin16a]. Financial [Fan05a, Fan05b, MZ05, PY05, Sør05]. Finding [Che10, Mor14, OC10, TR10, ZJ24]. Findings [Mur18]. Findley [MH12]. Finetti [Daw04, Arj16, CR96]. Fingerprinting [Ano94b, BDN94, Ber94b, Lem94, Lew94, Roe94a, Roe94b, Sud94, Tho94b, Wei94]. Fingerprints [SSS00]. Finite [SKBS20, XYY19]. Finney [Mac93]. Fire [TWDM13]. First [Arj16, Ber89a, Dal89, DeG86f, Flo89, GH16, Mit89, Nic16, Sch04, Smi89a, Sti18, Sun89, Tru89a, Tru89b, Weg89, Zid89]. Fisher [Efr98, Ald97, Ald05, And96, Bar89, Ber03b, Box87, Chr07, Coo07a, Coo07b, Edw97b, Efr98, Fie97, Kar92, Leh90a, Li07, LN07, Men05, PB10, Pla89, Rao92, SR14, Sei92, Sti05, Zab89a, Zab89b, Zab92, Bar92a]. Fitted [CF08]. Fitting [BH07a, BH07b, BMW07, HT00a, Has07, LS17]. Fixed [PQW14]. Flexibility [BHP15]. Flexible [EM96, FM18]. Flight [SKKS14]. Florida [AP02]. Flournoy [Ros15]. flow [Eud96]. Flowers [Sam91]. Fluorescence [SAL^{+}20]. fMRI [Lin08]. Focusing [AS21]. Follow [KJED09]. Follow-up [KJED09]. Forecast [LTRG17]. Forecaster [LTRG17]. Forecasting [BS13, Pin13, Wal14a, YS22]. Forensic [Ber91a, Ber91b, Che91, Eve91, Ize01, Jim11, Kay91, Lan91, Lew94, OS21]. Forensics [Car11]. forestry [PS00]. Foreword [Bai88, Cas03a, MR14]. Form [DZ91]. Formal [VMB08]. Formation [VGB14]. formative [BG99]. Formula [ABT16, Cra16a, Cra16b, FJ16, Fen16, Teh16]. Formulation [Fie06a]. Formulations [YB17]. Forward [HTT20, Stu10]. Foundations [BD19, ML14, Sin16a]. Founder [Rao92]. Four [CC14]. Fourth [MTNO15]. Fractal [GŠP12]. Fractals [CY92a]. Fraction [CCCH08, NMK08a, NMK08b, ZL08]. Fractional [Taq13, YK16]. Fractionally [DPP18, Raf92]. Framework [CHP22, CD21, DTZ17, GKRS05, KG21, MMTW01, Mar21, Mey18a, NRWY12, PRS07, RN13, RAP14]. Frameworks [ZZS21, ZJ24]. France [Fri07, Sha18, SKKS14]. Francis [Sti89, Wal14a]. Francisco [RB20]. Frank [HM95]. Fraser [DT04, Rob11, SX11, Zha11b]. Frasian [Was11]. Fraud [BH02, HR11]. Fred [Gre97a]. Frederick [Ans88]. Free [BBK^{+}19, Eat07, LMR20a, LMR20b]. Frequency [Mor88, ROSS18, Ver21]. Frequentist [BB04a, BFW07, Mee11, Rao11a, RLQ17, Sed11, Slu11, BBW97, KM98]. Frequentists [Ber11, Gho11a, Swe11]. Friedman [Fis15]. Friend [Tan07b]. Friendly [KMR^{+}19, SB23]. Friends [Hol03b]. FRS [Bel04]. Full [Zas94]. Fun [ABHK00]. Function [Bha87, Bri87b, Cai12, Cai19, Dah87, DL18b, EH03, Han87a, Han87b, Ris87, Shi87, SLR15, Sol87]. Functional [Cha14, GN17, JS06a, Lub91, MRSS15, NRB16, NRB17, RD06, SGKR03, SDD06]. Functions [CDW95, CRSW13, DG95, GK15a, GV23, God95, LQMT13, Lel03, LZ95a, LZ95b, LL95, LQ03, McC95, Rei95a, Rya95, Sev95, vdV14]. fundamental [FT96]. Further [CTSH92]. Fusion [JCT22]. Future [Aka87, Bri87a, Cox91, Dra87, DP91a, DP91b, Efr10a, Efr10b, Gei87a, IKN09b, Ize87, KM91, LL87, LKS04, Nor91, Rao87a, Rao87b, SBCC^{+}24]. Fuzzy [BC05, BCD05, Gel05, GM05a, GM05b, Tho05].

G [Ano91g, BS10]. G. [Bar16]. Gábor [GPW23]. Galaxy [VGB14]. Galton [Sti89, Wal14a]. Galtonian [Sti90]. Gambler [DE22]. Game [RGVS23]. Game-Theoretic [RGVS23]. Games [BB18, Sha18]. Gani [Hey95]. Gauss [Cel98]. Gaussian [BR02, Cai19, DMRT13, DP07, GHSS20, KG21, Lin19, PK97, SVS11, Wal14b]. Gaussianity [Mar04b]. Gaussianization [Cai19]. Gegenbauer [DPP18]. Geisser [CJ07]. Gelfand [CH15]. Gelman [VK21]. Gene [JLZL04, KLDR09, LZD^{+}20, WLLH21]. General [GN17, KG21, Lar11, Lit11a, MMTW01, Mar21, PRS07, Que24, Ros17, Sch11, SQH18, ZZ12, ZSCW06, Zuo21]. Generalised [CPW15]. Generalizable [RB23]. Generalization [SRC^{+}16]. Generalizations [BBFT17]. Generalized [ATH16, Ban20, EG22, EMLB19, HT86a, Hil89, Jor89, KMSG06, McC89, Mey18a, MVDV10, TN13, ZSCW06, dP89b, dP89a, Bri86, HT86b, McC86, Nel86, Sto86a]. Genes [GT03, DWZ97]. Genesis [WRR94]. Genetic [AB09, CCCH08, Coh91, DLS14, KAX^{+}18, LS03, NA94, NMK08a, NMK08b, Tho94a, Zha11a, ZL08]. Genetics [GT94, Lai10]. GENIUS [TSW21]. Genoese [Bel91]. Genome [GWVV09, KZI09, LL09a, PGP09, RW09, SCtW^{+}09, TCC^{+}09, ZMG09, ZJZ^{+}09]. Genome-Wide [GWVV09, KZI09, LL09a, PGP09, RW09, SCtW^{+}09, TCC^{+}09, ZMG09, ZJZ^{+}09]. Genomes [DJK^{+}20]. Genomic [FHS15]. Genomics [Lai10, LM14a, SGKR03]. Genotoxic [Men88]. Genton [BHP15]. Genuine [Ald95]. Geoff [BF98]. Geographic [PT03]. Geographical [Saw91]. Geometric [BD01, JJN13]. Geometrical [BZ12]. Geometry [Ama89, BN89, BN90, Ber89b, FMF13, Kas89a, Kas89b, Rao89, RF89, SAC19]. George [Lit19, BS10, DeG87b, DeG88, HMRV99, PT10]. Geostatistical [DMRT13]. Geostatistics [GK15a]. Get [Ber89a, Dal89, Flo89, Mit89, Smi89a, Sun89, Tru89a, Tru89b, Weg89, Zid89, GG02]. Getting [MN11]. Ghosh [Ber11, Swe11]. Ghosts [Maz15]. Gibbs [BBFT17, BHJK21, BCPR08, CTSH92, DKSC08a, DKSC08b, JJ08, Let08, LC08]. Gibbs-Type [BBFT17]. Gifi [MdL98]. Gil [Kas99]. Gilbert [Kat02]. Gilford [WM07]. Girolami [Gir14]. given [GW02]. Gleaned [PW00]. GLM [Pou11]. Global [Cha22, JJG17, Muk22]. Globally [SH14]. GMM [SGKR06]. Gnanadesikan [Ket01]. Gnedenko [SS92]. Godambe [Tho02]. Goeman [Hel11, Mei11, Wes11]. Golden [Hot88c, Arr88, Bra88, Dem88a, Fri08, Gup88, Hog88, Hot88b, Hot88a, Moo88, Olk88, Zid88]. Gombaud [BG24]. Good [Cam91, Gio23, Har91b, Hol03b, Rob91a, Rob91b, Spa91, Spe91, SK91, Tho91, Ban96, Zab23]. Goodman [Bec09]. Gordon [Gel97]. Gosset [Box87]. Government [Fie11a, Han11a, Kal11, DS92]. GPA [Joh97]. Grace [NMB20]. Graduate [Sch04]. Grandparent [CO12]. Grant [Ber89a, Dal89, Flo89, Mit89, Smi89a, Sun89, Tru89b, Weg89, Zid89, Tru89a]. Graph [SAY22]. Graphical [BH90, Cri90, Dem93, DKK10, DP07, Fis90, Gow90, JCD^{+}05, Jor04, LLW12, LS03, Pea93, Stu90, Wai90, WS90a, WS90b, You90, vDM10]. Graphics [BCW87a, BCW87b, Edd87, Fri08, Hub87a, Tuf87, Tuk87, Tuk90, Wai87, XHC14, ZLS10]. Graphon [GM21a]. Graphs [ATPP21, CW93b, POSD13, SKBS20, Shp14]. Great [AC89, Beg89, Ber89c, Cor89, Har89, KG89, LW89, Roy89, War89a, War89b]. Grenander [Muk06]. Groeneboom [Jon19]. Ground [Fai89, Joe89, Pic89, Raf89, Sin89, Smi89b, Smi89c, Wei89]. Ground-Level [Fai89, Joe89, Pic89, Raf89, Sin89, Smi89b, Smi89c, Wei89]. Group [Bre87, Cai08, Cor87, Hol87, HBM12, Jen23, Mar87a, Per87, Ros87a, Ros87b, YB17]. Grouped [Bur89, Hei89a, Hei89b]. Groups [Ben08, Efr08a, Efr08b, JSH07, Mar87a, Mor08, RS08a]. Groves [HKL17]. Grows [MGC15]. Growth [Aka87, Bri87a, Coh91, Coo16, Don91, Dra87, Gei87a, Ize87, Jag91a, Jag91b, LL87, Rao87a, Rao87b, Saw91]. Grundbegriffe [SV06]. Guerry [Fri07]. Guide [FL24, vDM10]. Guidelines [Ano94c]. Guido [MM24]. Guinness [Box87]. Gupta [McD98]. Gustav [Che99, Fel99, Nor99]. Guy [Mur05]. GWAS [KLDR09, ZT09].

H [Ano91a, Ano91b, CT11, Eat07, Est03, Han11b, Ion11, Men09, Nau00, Olk87a, Sto14, Tan07b, Yao11]. H-likelihood [Men09]. H. [BN04]. Haenszel [Han97]. Haldane [EW17, WZG24]. Half [Fie92]. Hall [DW16]. Handling [DK87, DHR^{+}13, Lin87a, Lin87b, RS08b, Sha87a, Sha87b, Spi87a, Spi87b, Spi87c, Wat87]. handwritten [HS98]. Hanfelt [SWY00]. Hannan [GR10]. Hansen [Olk87b, Olk89b]. Haplotypes [CCLC09]. Harald [Råd97, Weg86b]. hare [PK97]. Harms [Eva12]. Harness [KLDR09]. Harold [Bra88, Dar88, Dem88a, RCR09a, RCR09b]. Harris [Ale96]. Harry [Kva06]. Hartigan [Bar05]. Hastie [CTB20]. Hastings [BD01, RR01]. Hats [Arj16]. hazard [AG01]. Hazardous [Bar88a]. Hazards [Ros88, Sav88]. Health [BG05, BGST04, FPDE14, IKN09a, Lai10, Lin20, Muk22, QKM20, Sav88, Spi04, ZS09, EGG97]. Health-Care [Spi04]. Hear [Con93, Gas93, Gra93a, Gra93b, Rob93a]. Heavy [KMR^{+}19]. Heavy-Tailed [KMR^{+}19]. Hedayat [MSY16]. Henry [Whi93b]. Her [BF15]. Herbert [LS86, Swi92]. Here [Arn07]. Herman [Bat96]. Heterogeneity [BCN^{+}23, FW94b, SCtW^{+}09]. Heterogeneous [LYL22, ZWS^{+}19]. Heuristics [MG20a, MG20b]. Heyde [GK06, SWY00]. Hidden [JLMD16]. Hierarchical [BBFT17, BC07a, BC07b, Eva07, GKLP99, Gel07a, Joh07, LL07, PRS07, WMHL13, YB17]. High [Atk86, BC22, BR02, Bra86, CHP22, CH86a, CH86b, Coo86, FWZZ21, GHV12, HK86, HBM12, HRV08, JCD^{+}05, KZZW22, MHdL23, NRWY12, NS94, NNLL18, RBGK14, SB23, Vel86, Wei86, Wel86, WTLW95, ZZ12, ZWS21, ZLS10, vDK04, vdGM12]. High-Breakdown [HRV08]. High-Dimensional [CHP22, FWZZ21, GHV12, HBM12, JCD^{+}05, MHdL23, NRWY12, NNLL18, RBGK14, SB23, ZZ12, ZWS21, ZLS10]. High-Energy [vDK04]. High-Order [BR02]. High-Performance [KZZW22]. High-Speed [WTLW95]. Higher [BFW07, DJ15, vdV14]. Highlighting [OS21]. Highly [FPW16, RSLGR07, vDK04]. Hilbert [EG22]. Hillel [Kas99]. Him [Ken91]. Hirotugu [FP95]. Historical [KC14]. Histories [FSB16]. History [Bar89, Bel93, DJK^{+}20, Fie92, Han87c, Pla89, RC11, Sha21, VGB14, Zab89a, Zab89b, BG99, Fre99, Hal99, Sti96]. HIV [DPCA14, BSJ93a, BSJ93b, Bro93, CG93, DP93, GR93, KS93, SW93]. Hoc [BGLC93, Bil93, Car93, CGP^{+}93, Gen93, vZ93]. Hoeffding [FvZ08]. Hogg [Ran07]. Holds [Gus23]. Hollander [Sam08]. Holmes [Kad09]. Home [GKLP99]. Hominoid [BH87a, BH87b, Fel87, Por87]. Honest [HJ01]. Horse [BZ21]. Horseshoe [BDPW19]. Hospital [NS07]. Hotelling [Bra88, Dar88, Dem88a]. Household [GL06]. Household-Specific [GL06]. Howard [Fie08]. Howell [CY14a]. Hsu [CO12]. Huber [BK08]. Huff [Ste05]. Human [IJ23, Ros88, Sav88]. Hunter [De 20]. Hybrid [BHJK21]. Hypotheses [Bay87, BD87a, BD87b, CB87, Cox87, Eat87, Kad87a, Zel87, Leh97]. Hypothesis [CCCH08, DSB03, EW17, GM21a, Lan15, NMK08a, NMK08b, RWR23, RW09, WH16c, ZL08, BBW97].

I. [Ban96, Sam99, Zab23]. Ibragimov [BL90]. ICIAM [AET09a]. ICM [DE22]. Ideal [ÁEdBCAM17]. Ideas [LG17, Spi04]. Identifiability [Gus05]. Identification [Ber91a, Ber91b, Che91, Eve91, Gus23, HP04a, IKY10, JD21, Kay91, Lan91]. Identify [ZT09]. Identifying [ALS^{+}01, HPV06, LD18]. If [Men05]. Ignorability [Shp14]. II [BBK^{+}19, Law19, Tjø19, vdG19]. III [Ell97b, SS99]. Ildar [BL90]. Ill [Gil86, O’S86b, O’S86a, Ric86, Tit86, Wah86]. Ill-Posed [Gil86, O’S86b, O’S86a, Ric86, Tit86, Wah86]. Illicit [Ize01]. Illnesses [Zha11a]. Illusion [Fri06, Gay06, Han06a, Han06b, Hol06, Sti06b]. Illusions [KBH20, MG20a, Sha20, MG20b]. Illustrated [BBB^{+}19, Lit19]. Illustrative [Gus05]. image [PCT^{+}98]. Imbens [MM24]. Impact [Bar92b, Ber03a, Clo92a, Clo92b, GGM09, Hol92, Lah03, Lel03, Man92, Mee11, Mol92, Pol03, Rao11a, Sed11, Sha03, Slu11, XR15, Che99]. Impersonal [Lin91]. Implementation [RL92]. Implementations [Mey18b, TX23]. Implementing [Mal02]. Implications [Nee88]. Importance [APSAS17, AMV19, EMLB19, IRÁ22, Kas09, MH16, Pac10, GM98]. Important [Hol92, Pfe13]. Improved [FW12]. Improvement [ÁEdBCAM17]. Improvements [Hwa90]. Improving [CEH^{+}17, HRR11, Mir01]. Imputation [AWJ^{+}18, CCLC09, Fay94, Men94a, Men94b, Mur18, Sch94a, Ski94, YK16, Zas94, vHB21]. Imputations [vHB21]. IMS [AET09a]. IMU [AET09a]. Incidence [SW16]. Incompatibility [And88]. Incomplete [JBM^{+}06, KS07a, KS07b, RM07, RC11, SV18, TD07, VMB08, ZJZ^{+}09]. Incorporating [Spi04]. Incremental [CJLV06a, CJLV06b, DLM06, LMN06, Yu06]. Indefensible [ZWS21]. Independence [EG22, GS93]. Independent [GZZ20, MTNO15]. Indeterminism [Bri08a, Bri08b, Kün08, Yan08]. Index [Ano02, Ano03b, Ano05, Ano10, Ano11]. Indexed [Pem95]. Indians [CR17]. Indices [Bel87]. Indirect [DPL15, Efr10a, Efr10b, JT04, Kas10]. Indoor [Gin88]. Inductive [TGM24]. Industrial [AET09a, Ban93a, Ban93b, Cna93, Duf93, Hah93, Hog93, NP93, Orc93, Rob93b, Woo93, Wu93, Wyn93, You90]. Inequality [GMS20]. Infection [BSJ93a, BSJ93b, Bro93, CG93, DP93, GR93, KS93, SW93]. Infections [JCT22, Mey18b]. Infectious [Bre18, KM18b, SSS00]. Inference [Agr92b, Agr92a, Ama89, ATPP21, BT18, Bai17, Bar92a, BN88, BN89, BN90, BC22, BFW07, BH92, BD19, Ber89b, Boc16, BD08, Bre18, BNF^{+}07, Cai12, Car19, CDW95, CMC^{+}24, Dan88, DKK10, Din17a, Din17b, DL18a, DJK^{+}20, DJ15, DHS^{+}19a, DPL15, Duf92, EV17, EF92, Ern04, Fra04, GR92a, GMN^{+}04, Gew92, GS93, Goo11, GK02, GPR99, GT94, GJ18, GvdL19, GLSS23, Hei89a, Her19, HW88, Hou88, IKY10, Jen19, JT04, KEHSY19, Kas88, Kas89a, Kas89b, Kas11b, Kor06, Kre92, KVP^{+}14, KM18b, LN09a, LN09b, LLV10, LZ95a, LW92, LQ03, Lou09, LGA^{+}15, MH93, MZL10, ML14, McC11, MVA^{+}18, Meh92, Mey18a, Mit88, MKV09, MHdL23, RGVS23, Rao89, Rei88a, Rei88b, RF89, Rei95b, RTW19, RB23, Rub90, Rub06a, SSM23, SS18, SB16, SKBS20, Seg12]. Inference [SB23, Ste11, SBCC^{+}24, Stu10, Sui92, TN13, TSW21, TYC^{+}24, Tie88, TGM24, VMC20, Wal09, Was11, YNS23, ZKS19, ZWS21, ZP21, HZ00, KM98, Bur89, CDW95, DG95, God95, Hei89b, LZ95b, LL95, McC95, Rei95a, Rya95, Sev95]. Inferences [Ber91a, Fay94, Men94a, Men94b, Sch94a, Ski94, Zas94, Ber91b, Che91, Eve91, Kay91, Lan91]. Inferential [FPW16, IRÁ22, LM21]. Infinite [SKBS20]. Inflated [LSS^{+}19]. Influence [vdV14]. Influential [CH86a, Atk86, Bra86, CH86b, Coo86, HK86, Vel86, Wei86, Wel86]. Inform [TWDM13]. Informal [VMB08]. Information [CCCH08, Dos08, ESS10, EJ11, FM18, Gus05, MA87, NMK08a, NMK08b, SQH18, Wan12, Yu95, ZL08]. Informativity [EJ11]. Ingram [Bec07, Geo07, Sam07]. Ingredient [CST00]. Input [Fay94, Men94a, Men94b, Sch94a, Ski94, Zas94]. Inquiry [KK15]. Insights [FMF13]. Inspired [Sti08]. Instability [BT18, Ell98]. Institute [EGG97, AET09a, Hog86b]. Instrumental [CPW15, Imb14b, Kit14, STR^{+}17, SH14, VBBG11, ZWS^{+}19]. Instruments [Shp14]. Insurability [Bar88a]. Insurance [IKN09a, ZS09]. Integers [LK20]. Integral [GGC13]. Integrals [ES95, Lin19, Ano96]. Integrated [BLW99, QM02]. Integrating [BCN^{+}23, MvdKD19]. Integration [BOG^{+}19a, BOG^{+}19b, ES95, HJ19, SH19, Ano96]. Intelligence [Lin87c, Sha87c]. Intention [PA22]. Intention-to-Treat [PA22]. Interactions [KLDR09, Twe98]. Interactive [XHC14]. Intercensal [FW94b]. Interdisciplinary [Meh92]. Interface [AS93, Ban88, Bar88b, Ber88, Goo88a, Goo88b, Sup88]. Interfaces [RTW13]. Interference [OV14, VTH14, ZP21]. Interim [JT90]. Intermediate [JT04]. International [AET09a]. Internet [CCS04, Duf04]. Interoperability [NBB^{+}22]. Interplay [BB04a, Hol92]. Interpretable [RN13]. Interpretation [BD01]. Interpreting [FPBT22]. intersection [BH96]. intersection-union [BH96]. Interspecies [LGF95]. Interval [BCD01, CS12b, GG09, Jia19, LMR20a, RRSM20, LMR20b]. Intervals [AG05, And22, BC05, BCD05, DRRS22, ET86a, ET86b, FPBT22, GM05a, GM05b, Har86, LPE15, Tho05, TNM22, DE96, Man02]. Intervention [JSH07, Pea93]. Interventions [CG92, DeM92, Ell92, FC92, Fle92a, Fle92b, Lag92, Lin20, Lou92, RW06, SSP^{+}19]. Interview [Rit11, SII12, Swi97]. interviews [EGG97]. Intractable [HJ01, LGA^{+}15]. Intrinsic [APSAS17]. Introduction [AS93, Ano91c, Ano03a, Ano20, ACS01, Ber11, Boo03, Cas03b, Fie97, GWC04, Gho11a, JR18, Kas99, KM18b, LS11, Olk92b, RHC04, RC04, SV18, SW17, Swe11, TYC^{+}24, WZ12, ZMG09]. Introductory [Spe90]. Intrusion [DJK^{+}01]. Invariance [Büh20a, Büh20b, BW23]. Invention [Sti89]. Inventor [Wai03]. Inverse [Bar89, Gil86, O’S86b, O’S86a, Per05, Pla89, Ric86, RSLGR07, Sti86, Tit86, Wah86, Zab89a, Zab89b, Edw97b, Fie97, Hal99]. Investigating [War89a, AC89, Beg89, Ber89c, Cor89, Har89, KG89, LW89, Roy89, War89b]. Investigation [BBK06]. Invited [Efr98]. Involving [Gus05]. Iowa [Dav98b]. Isaac [Sti06a]. Isolation [Tan90]. Isotropy [WH16c]. Issue [Ano86b, Ano86c, Ano86d, Ano87b, Ano87c, Ano87d, Ano87e, Ano88a, Ano88b, Ano88c, Ano88d, Ano89c, Ano89d, Ano90a, Ano91d, Ano92b, Ano92c, Ano94d, Ano94e, Ano94f, Ano95b, Ano95c, Ano20, CDR23, DeG86f, JS06b, Kas92b, Kas92c, Kas93a, Kas93b, Kas93c, Kas94, Mor89a, Mor90a, Mor90b, Mor90c, Mor91a, Mor91b, Mor91c, RHC04, RC04, SS18, Ste93a, WZ12, ZMG09, Eat01, Swi97]. Issues [BGST04, Ber12, CG92, DeM92, DK87, DuM12a, Ell92, FC92, Fle92a, Fle92b, Gol16, KMSG06, Lag92, Lin87a, Lin87b, Lou92, MT12, NS94, Sha87a, Sha87b, Spi87a, Spi87b, Spi87c, Ste20, TCC^{+}09, Wat87, DWZ97, GST97]. Iterative [GR92a, Hil89, Jor89, McC89, dP89b, dP89a]. IV [DMS10]. Ivan [Ell98].

J [Ban96, Bha22, Flo93, GPW23, Hel11, Lar11, LC24, Mac93, Mee11, Mei11, Rit11, RB20, Sch11, Sed11, Slu11, SWY00, Wes11, Zab23]. J. [De 20, DZ23, EW17, WZG24]. James [GR10, RR14b, Wol04b]. Janet [Fie94]. Jayaram [Hol08]. Jeff [CJ16]. Jeffreys [Ber03b, Ber89b, Gel09, Kas09, RCR09a, RCR09b]. Jerry [Fis15]. Jim [Ald18]. Joe [Hey95, Sne97]. Johannes [Nau00]. John [Ans88, Ans03, Bar05, Cas03a, Ell97b, FM00, FM03, Fer03, Har03, Hoa03, Kaf03, Mal03, Mor03, Sen03, SWY00, Wai90, Wai03]. Johnson [Rea04]. Joint [DFN12, LZD^{+}20]. Jolly [Nic16, BM16]. Jon [BS18]. Joseph [MM00]. Journals [Sti94]. Journey [PBSO24]. Jr. [Ano91g]. Judgment [GM21b, Sha87c]. Judgments [Dra87]. Judicious [GM21b]. Jugalbandi [Sin16b]. Just [SRR^{+}17b, Fra11a, Rob11, SX11, Zha11b]. Justify [Fre08].

Kalai [Kas99]. Kalman [Spa91]. Kanti [Muk02]. Karl [Sti08]. Kass [BN90, Beh19, Goo11, McC11, Ste11]. Keep [SRR^{+}17b]. Kemperman [Nau00]. Kempthorne [Fol95, Hin01]. Kendall [Bin96, Swi97]. Kernel [DD21, CZL^{+}20, CM91a, CM91b, FM93, GJSE91, GH91, Har91a, HL93a, HL93b, Jon91, Mul93, QOS19, Sch04, Sil91]. Kernel-based [CZL^{+}20]. Khinchin [Ver21]. Khmaladze [KK16]. Kin [BSA16]. Kinship [Gal89]. Kish [FK96]. Kishore [BM07]. KL [Boc16]. Kleiber [BHP15]. Kolmogorov [Ken91, SV06, SRS91]. Kotz [Nad02]. Kramer [Ell97a]. Kriging [ZC15]. Kruskal [Bra07, Eat07, FST07, Fie07, Goo07, Mar07, Sti07b, Tan07b, Zab94]. Kuiper [LRdP^{+}04]. Kumar [GS08, Muk97].

L [DeG86c, Fie94, Lit19, Rea04]. Label [JHS05]. Laboratory [Dav98b]. Labs [Mal03]. Lai [LSY21]. Laird [Rya15]. Lancaster [Let08]. Land [Kro92]. Landmark [Boo86a, Boo86b, Cam86, Cre86, Goo86d, Ken86, Sam86]. Laplace [BBSW10, KBH20, MG20a, MG20b, Sti86]. Laplacian [PK97]. Large [ALS^{+}01, BT18, BGGM06, BFW07, DJ15, YSR20]. Large-Sample [BT18]. Large-Scale [BGGM06, DJ15]. Larry [BD19, Das05, Joh19]. Lasso [BDPW19, FS05, FHS17, HTT20, YB17]. Last [Nic16]. Late [ZJ24]. Late-Onset [ZJ24]. Latent [ATPP21, BHJK21, GHSS20, SAC19]. Later [Leh90a, Nic16, Kit14]. Law [Hil95, PT11]. Laws [Per05]. Layouts [EHX04]. Lead [ALS^{+}01]. League [Hod17]. Learn [JAAH14]. Learned [DHS^{+}19a, GvdL19, Lin22, YS22]. Learning [CJLP10, FWZZ21, FMZ21, FL24, LMR20a, PSW15, PBSO24, STLD14, SGZW10, WBR^{+}22, LMR20b]. Least [Bel87, HV87, Hil89, Jor89, Mar87b, McC89, Nyb23, Ste87a, Ste87b, Thi87, dP89b, dP89a, Ell98, Hal99]. Leave [CR17]. Lecoq [Kad09]. Lecture [Bri08b, Chr07, Coo07a, Coo07b, Kün08, Li07, LN07, Yan08, Bri08a, Efr98, Sti90]. Lectures [Bar90b, Ber90b, Ber90c, Cox90b, Fre90, Kad90b, Leh90b, Lin90a, Lin90b, Mou90]. Lee [LB22]. Legacy [FST07, Kas09]. Legal [Ize01, MMT07]. Legendre [Nyb23]. Lehmann [DeG86c]. Length [AS10]. Leo [Bec09, Ols01]. Leopold [Wit91]. Leslie [FK96]. Lessons [DHS^{+}19a, GvdL19, Lin22, YS22]. Let [Han14a]. Letter [WRR94]. Leung [LSY21]. Level [CEN^{+}19, Fai89, FPDE14, Joe89, LL07, Pic89, Raf89, Sin89, Smi89b, Smi89c, Wei89]. Levels [BC07a, BC07b, CP94, Eva07, Gel07a, Joh07, RW09]. Levene [GGM09]. Leverage [Atk86, Bra86, CH86a, CH86b, Coo86, HK86, Vel86, Wei86, Wel86]. Lewis [Por94, Sen93]. LGM [GHSS20]. Li [CT98, SWY00]. Lie [Bes05, DH05, Ste05]. Lies [Bes05]. Life [Ros06, Zab23, Nor99]. Light [MO00]. Likelihood [BFW07, Bjø90a, Bjø90b, Bjø14, BM24, Bre18, But90, CP14, Daw14, Eva14, Han14b, KM98, LB22, LN09a, LN09b, LTH90, Lou09, May14a, May14b, MKV09, OS21, PL21, PL22, Sti07a, Tho94a, vdGM12, vHB21, Ald97, BLW99, Fie97, Hal99, HZ00, Men09]. Likelihood-Based [Bre18]. Likelihoods [LGA^{+}15, WM04]. Likely [HPV06]. Limit [Doo86, DL18b, Le 86a, Le 86b, Pol86, Tro86]. Limitation [Fie06b, GKRS05]. Limitations [Sha21]. Lincoln [BH99]. Lindley [Smi95]. Line [RD06]. Linear [ATH16, Atk86, Bic88, Bra86, BBB^{+}19, CH86a, CH86b, CZL^{+}20, CV08, Coo86, CW93b, Dra88a, Dra88b, HK02, HA88, HK86, KMSG06, KP88, Lit19, Mar88b, McK04, MVDV10, MSW13, QKM20, Sha93, TN13, Vel86, Wei86, Wel86, Wel88, ZSCW06, Ell98]. Linkage [GT03, ZJZ^{+}09]. Linking [SFR10]. Lisp [Lub91]. Lisp-Stat [Lub91]. Litigation [Joh88, Wei88]. Little [Tei88, Lar11, Sch11]. Living [CCFR24]. Local [Ban20, BW23, GKLP99, HL93a, FM93, HL93b, Mul93]. Location [CK03]. Locations [LL04]. Log [DMRT13, Sam18, Wal09]. Log-Concave [Sam18, Wal09]. Log-Gaussian [DMRT13]. Logicist [Dem98]. Logistic [Ber12, BK17, DuM12a, DR19, Fre08, MT12]. Logrank [LS23]. Long [BT18, Ber92a, Ber92b, BGST04, DH92, DPP18, Li99, Par92, RL92, Raf92, Smi92a]. Long- [Raf92]. Long-Range [Ber92a, Ber92b, DH92, Par92, Raf92, Smi92a]. Long-Term [BGST04]. Longitudinal [DHR^{+}13, DM93, FLR93a, FLR93b, JBM^{+}06, KJED09, PM93, Sut03, VMC20, ZLH93, ZZS21]. Look [CTB20, DHS^{+}19b, Lan15, LS23, Stu10, MW96]. Lookahead [LCL13]. Loss [Dos08, FW12, GM21b]. Losses [Grü16]. Lottery [Bel91]. Louis [Wal18]. Love [SRR^{+}17b]. Low [FM18, Mor88]. Low-Rank [FM18]. Lu [CO12]. Lucien [Yan99]. Lying [Mon05, Mur05]. Lyngbya [JAAH14]. Lynne [Muk17].

M [Ano96, BK24, Ber11, Buc16, CS12c, GTWT12, Han97, RR14b, SR12, Swe11]. M. [Fri07, Gir14, HMRV99]. Machine [FWZZ21, FL24, LMR20a, PSW15, PBSO24, LMR20b]. Machines [Cai19, MM06b]. Magic [FS05]. Majorization [Arn07]. Make [CRSW13, SCN^{+}10]. Making [GWVV09, IJ23, Ald97]. Man [Bre88b, DuM88a, FZ88a, FZ88b, Has88, KT88, MD88]. Management [JAAH14, KVP^{+}14, Pin13, TWDM13]. Mandelbrot [Taq13]. Mantel [Gai97]. Mapping [Tho94a]. Mappings [Que24]. Maps [Mon05, NS94]. Mardia [Muk02]. Margaret [SO94]. Margin [ZJ08]. Margin-Based [ZJ08]. Marginal [CP14, LN04]. Marginalized [HZ00]. Margins [Gre91]. Marie [Sha18]. Marie-France [Sha18]. Mark [HR16, JLMD16]. Marker [GWVV09]. Markers [Hat88, Joh88, Mar88a, Men88]. Marketing [HPV06, MAK06, RW06]. Marketplaces [BBI^{+}23]. Markov [Ald93, BR02, Coh91, CTSH92, FBPR18, FHJ08, Gel92, GR92b, GT03, Gey92a, Gey92b, HJ01, JHS05, JJN13, JLMD16, Mad92, Mir01, PHYW22, Pol92, RP92, RL92, Rob95, RC11, Ros92, Sch92, Sen16, Tie92, WWZ16]. Marquardt [Hah95]. Martin [GB10, SO94]. Martingales [SSVV11]. Martz [Kva06]. Marvin [Sim97]. Mary [Ros23b]. Masquerades [DJK^{+}01].

Massive [GLSS23]. Match [KJED09]. Matched [IKN09b, SJL^{+}12]. Matching [CT11, Fie06b, Han11b, HS09, IKN09a, Ion11, Stu10, VGB14, XT11a, Yao11, YSR20, ZS09]. Materials [Bar88a]. Matérn [PBSO24, SAY22]. Mathematical [AET09a, DeG86f, Hog86b, LS86]. Mathematics [AET09a, Sco17, GKT02]. Matrix [Ize21]. Matter [MN11, Nol92, Ros23a]. Maty [BG07]. Maurice [Olk89a]. Max [OS22, Seg12]. Max-Stable [OS22, Seg12]. Maxima [BZ21, Lin19]. Maximization [FC11, LP11a, MT11]. Maximum [GZZ20, Sti07a, vHB21, Ald97, Hal99]. May [MN11]. Maya [Kas99]. McKay [Kas99]. MCMC [CRSW13, FT13, GHSS20, MH16, TW10]. Mean [CPW15, DR19, JL13, KS07a, KS07b, PQW14, RM07, TD07, Edw97b]. Mean-Field [DR19]. Means [MP93]. Meant [SGJC13]. Measure [BLNBP13, DN06, FT13]. Measurement [Duf04, GKLP99, HKL22, Pos23, Gle98]. Measurements [WTLW95]. Measures [BB04b, ET86a, ET86b, FPDE14, Har86, LL07, MVDV10, RN13, Dav98a]. Measuring [BCK^{+}07]. Mechanics [Sen16]. Mechanism [Hol88]. median [Ell98]. Mediation [IKY10]. Medical [Böh16, Gas87a, Gas87b, Gei87b, Gla87, Gol87a, JT90, Kay87, KR87, Kor06, PCT^{+}98, Wit87]. Medicine [DK87, Lin87a, Lin87b, Sha87a, Sha87b, Spi87a, Spi87b, Spi87c, Wat87]. Meeting [ABC^{+}91]. Meets [ABB^{+}23, BDPW19, GM21b]. Memoir [Lap86, Sti86]. Memorial [Bar90b, Ber90b, Ber90c, Cox90b, Fre90, Kad90b, Leh90b, Lin90a, Lin90b, Mou90, Sti90]. Memory [BT18, DPP18, Har03, Raf92]. Mendel [PB10]. Mendelian [TSW21]. Mental [Zha11a]. Mentor [Tan07b]. Mentors [Leh93, Nol92]. Menu [Zas11]. Méré [BG24]. Messy [GG09]. Meta [BB91, Bec07, BW07, Daw91, DB92, Dia91, Gre91, Hed92, Hym91, Mor91e, Mos91, MC92, SMJN17, Utt91a, Utt91b, GST97]. Meta-Analysis [BB91, Bec07, BW07, Daw91, DB92, Dia91, Gre91, Hed92, Hym91, Mor91e, Mos91, MC92, SMJN17, Utt91a, Utt91b, GST97]. Method [AY01, ABB^{+}23, AGG90a, AGG90b, Bar90a, BZ21, Che90, CPW15, JT04, KJED09, MO00, Nyb23, Ste90, SCtW^{+}09, Wat90, Cel98]. Methodological [IR23, TCC^{+}09]. Methodologies [SSS20]. Methodologists [CC14]. Methodology [Bar92b, Clo92a, Clo92b, DHY03, DK20, FJH^{+}89, Hol92, Man92, Mol92, RW06]. Methods [Ano96, BN88, Ber92a, Ber92b, BNPS13, BL20, BCN^{+}23, Bru18, CLSW16, CMC^{+}24, CRSW13, Dan88, DH92, DMS10, DHS^{+}19a, Eat07, ET86a, ET86b, Ern04, ES95, Fan05a, Fan05b, FWB94, FPW16, Fie11a, FMF13, GvdL19, Han11a, Har86, HR16, HW88, HP04b, Hol03a, Hou88, HKL22, HRV08, IRÁ22, JHS05, Jen19, Kal11, KDS^{+}15, Kas88, KL08, Lan14, LCX21, LRdP^{+}04, Lil88, Mee11, Mit88, MH16, Mor14, MZ05, Par92, PY05, Pol03, Raf92, Rao11a, Rei88a, Rei88b, RAP14, Rob95, STLD14, SV18, Sed11, Sil91, Slu11, Smi92a, Sør05, Stu10, TNM22, Tie88, Tit04, WLLH21, WZ12, WMHL13, YSR20, ZKS19, BLW99, BM98, PS00]. Metric [KAX^{+}18, MSM24]. Metrics [HS98]. Metropolis [BD01, RR01, SFR10]. Mexican [IKN09a, ZS09]. Michael [BS16]. Michigan [Cra86]. Microarray [BBSW10, DSB03]. Microarrays [Ben08, Cai08, Efr08a, Efr08b, Mor08, RS08a, THNC03]. Microdata [Cox91, DP91a, DP91b, GKRS05, KM91, Nor91]. Microscopy [SAL^{+}20]. Microwave [GMN^{+}04, Mar04b]. Mid [AG05]. Mid- [AG05]. Miller [Ano91g]. Milton [Muk00]. Minimax [Cai12, GM21a, GLX12]. Mining [ABHK00, BS06a, Fie06b]. Miscellaneous [Moo90a]. Mises [Ver21]. Mismeasured [Gus05]. Missing [CY14b, CCCH08, DTL05, DL18a, FM18, JR18, Lar11, LD18, Lit11a, NMK08a, NMK08b, Sch11, SGJC13, Ste93b, SQH18, ZL08, ZLK10, KM98]. Misspecification [Grü16, HM16, RR23]. Misspecified [SB23, YNS23]. Misspecifying [MN11]. Misuse [KG05]. Misvotes [AP02]. Mitigating [ATH16]. Mitra [Muk97]. Mixed [ATH16, CZL^{+}20, MSW13, QKM20, TN13, ZSCW06]. Mixing [JJN13, MG08]. Mixture [BLNBP13, BR02, Che17, EHX04, FT13, HRR11, JHS05, XYY19]. Mixtures [Ano89f]. Mizera [Ell98]. MLE [Che17]. MM [WL10]. Mobile [Lin20, QKM20]. Model [Ald17, Ben08, BR02, BFWW09, BO17, BH07a, BH07b, BMW07, BBK^{+}19, Cai08, Car19, Che10, CG04, CBD^{+}15, DFN12, DPL15, DP07, DG13, Efr08a, Efr08b, Fie06a, Fie07, GH16, Gus05, HM16, Has07, HBG15, LPE15, Leh90a, LZD^{+}20, LL09b, MGC15, Mor08, MSW13, PB10, PBSO24, RS08a, SRR^{+}17a, VMB08, GW02, HMRV99, HMRV00, QM02]. Model-Assisted [BO17]. Model-Based [Che10, LZD^{+}20]. Model-Free [BBK^{+}19]. Modeling [BLNBP13, BC07c, BO07, Bre01, Bre18, BJP13, CST00, CT11, CWHD22, CS12c, DPR12b, Dra87, Efr14, FM18, GTWT12, GKLP99, Gel07b, Gel07c, Gre09, Han11b, Hed92, HRR11, Ion11, JHS05, KK15, LW06, LMW07, Lin06, Lit07, Loh07, Par04, Peñ06, Pfe07, QKM20, Raf92, RD06, SJW11, SR12, Teh16, VACC22, Wal09, WLLH21, WTLW95, XT11a, YB17, Yao11, Dem98]. Modelling [Dem93, DGA^{+}20]. Models [Aka87, ÁEdBCAM17, Ano89f, ATH16, BHJK21, BRM11, BBF^{+}22, BLNBP13, Bay88, BC07a, BC07b, BHP15, Bic88, BDP18, BF16, BS13, BC20, Bri86, Bri87a, BNF^{+}07, BBB^{+}19, BBK^{+}19, BKB^{+}19, Bur92, Cai19, Cat12, Che17, CZL^{+}20, CS12a, CPW15, CV08, CFL13, Cox90a, DKK19, Dem86, DKK10, Dra87, Dra88a, Dra88b, DM93, DuM88b, DR19, DG13, EHX04, Eri86, EM06, Eva07, FWZZ21, FPW16, FT13, Fel86, FHS15, FSB16, Fie11a, FLR93a, FLR93b, FN86a, FN86b, FJL^{+}19, GHSS20, Gei87a, GGC13, Gel07a, GK19, GST18, HT00a, Han11a, HK02, HT86a, HT86b, Hed88, HA88, Hof21, HKL22, HBM12, IG88a, IG88b, Ize87, JJG17, Joh07, JLMD16, JCD^{+}05, Jor04, Kad86, Kal11, KDS^{+}15, KMSG06, KP88, Kro92, KN18, LLW12, LL87, LPT88, LL07, LS03, Law19]. Models [LN04, LN09a, LN09b, LQMT13, LR92a, LR92b, Lit19, LM21, Lou09, LS17, Mad86, Mar88b, McC86, McK04, MVA^{+}18, Mey18a, MKV09, MVDV10, ML12, Mos86, MSW13, Nat86, Nel86, Nic16, NBB^{+}22, PRS07, Pea93, Per87, PM93, QKM20, QMJM22, Rao87a, Rao87b, Rao88, RBGK14, RR88, SVS11, STG13, SB16, SKBS20, Sen16, SB23, SGZW10, SW16, SAC19, Sto86a, Sut03, TN13, Tan23, Tha10, Tit04, Tjø19, VJ14, VBW15, WBR^{+}22, WH16a, WH16b, Wel88, WSC19, Wol86, XYY19, YNS23, ZLH93, ZSCW06, ZLK10, deL92, vDK04, vdG19, Ald01, BBC^{+}02, CI00, GW02, HS98, HZ00, MY97, Dem98]. Modern [BO17, Che10, CD21, Kad09, RFIS24, Rao92, Wal04, WMHL13]. Modifying [CRSW13]. Moivre [BG07, DZ91]. Molecular [BH87a, BH87b, Fel87, Por87]. Molecule [DK20, SAL^{+}20]. Moment [Men05]. Moments [CPW15, MTNO15]. Monitoring [CJLV06a, CJLV06b, CTSH92, DLM06, JT90, LMN06, WBR^{+}22, Yu06]. Monotone [DL18b, Ram88a, Bre88a, Eub88, HT88, Ram88b, Wah88]. Monotonicity [STR^{+}17]. Monroe [GLW07]. Monsieur [Kad09]. Monte [CTSH92, FBPR18, FHS15, FHJ08, Gel92, GR92b, GT03, Gey92a, Gey92b, HLO05, HJ01, JHS05, JJN13, LCL13, Mad92, Mir01, Owe19, PHYW22, Pol92, RP92, RL92, Rob95, RC11, Ros92, Sch92, Tho94a, Tie92, WWZ16]. Montmort [BF15]. Moral [Fri07]. Morris [FC11, LP11a, Ano91a, Ano91b, Olk87b, Olk89b]. Morton [Ell97a]. Moses [BH99]. Most [Mor14]. Mosteller [Ans88]. Motifs [JLZL04]. Motion [Taq13]. Motivations [McC94]. Mountain [vDM10]. Mouse [Bre88b, DuM88a, FZ88a, FZ88b, Has88, KT88, MD88]. Mouse-to-Man [Bre88b, DuM88a, FZ88a, FZ88b, Has88, KT88, MD88]. Mouth [DN06]. Moving [BGGM06, Lin20, Whe21]. MR [HMRV00]. Much [Hwa90]. Multi [VBW15]. Multi-armed [VBW15]. Multidimensional [Kre92]. Multidisciplinary [LY21]. Multilevel [AWJ^{+}18, ML12, HZ00]. Multilinear [LR92a, Bur92, Kro92, LR92b, deL92]. multiparty [QM02]. Multiple [AWJ^{+}18, BH23, BBK06, CY14b, CPW15, CS12b, DP07, DSB03, EMLB19, Eva12, GR92a, GS11a, Hel11, KL08, LG17, Mar87a, Mei11, Men94a, Mur18, NHZ16, RWR23, Tuk91, Wes11, vHB21, SWY00, Fay94, Men94b, Sch94a, Ski94, Zas94]. Multiple-Imputation [Men94a, Fay94, Men94b, Sch94a, Ski94, Zas94]. Multiplicative [Hof21, McK21]. Multipoint [GT03]. Multipurpose [CY14b]. Multistage [Mag93, TCC^{+}09]. Multivariable [Fri07]. Multivariate [And87, Bec07, Ber89b, Ber12, BHP15, BFWW09, BH90, CBD^{+}15, Cri90, DuM12a, Fis90, GK15a, GK87, Gol87b, Gow90, HP04a, HRV08, JLMD16, LL04, MT12, MM22, OR04, Per87, Pre87, Que24, Sch87b, Sch87a, Sen87, Stu90, WS90a, WS90b, You90, And96, MdL98]. Multiway [ZJ08]. Murray [BD09]. Musical [BM99]. My [Fie07, SRR^{+}17b]. Myles [Sam08]. Myths [DMJ23, RLLKV23b].

N [Lai89, Mee11, Sed11, Slu11]. Naïve [ZWS21]. Nan [Rya15]. Nancy [Ros15]. Nanoscopy [SAL^{+}20]. Natan [Kas99]. Nathan [Gai97]. National [EGG97, Mar07]. Nations [RAG14]. Nationwide [XR15]. Near [HV87]. Nearest [THNC03]. Nearly [LMR20a, LMR20b]. Need [Gre10]. Needs [ABC^{+}91, CEN^{+}19]. Negative [SRC^{+}16]. Neglected [Ald17].

Neighbourhood [Boc16]. Nelder [Sen03]. Nested [CFL13, VJ14]. Netflix [FHK12]. Network [CCL^{+}04, CD21, GM21a, HPV06, Hof21, KK15, PT03, SAC19, WLLH21]. Network-Based [HPV06]. Networked [CJLV06a, CJLV06b, DLM06, LMN06, Yu06]. Networks [AS10, Ama94, Bar94, BG94, Bre94c, CT94a, CT94b, HPV06, HKL22, JAAH14, Mag93, McC94, PHYW22, Rip94, Tib94, Tit04, WLLH21]. Neural [CT94a, HKL22, McC94, PHYW22, Tit04, Ama94, Bar94, BG94, Bre94c, CT94b, McC94, Rip94, Tib94]. Neuroimaging [SMJN17]. Newcomb [PT11]. Newton [MG08, Sti06a]. Next [Mad93]. Neyman [Bri08a, Sti90, Ber03b, Bri08b, Kün08, Leh90a, Rub90, SR14, Spe90, Yan08]. Nielsen [BN90]. NIH [EGG97, Gre97b]. Nineteenth [Sti91]. Niño [Kat02]. No [Chu17, Lan15, HMRV00]. No-Treatment-Effect [Lan15]. Noel [WH19]. Noisy [DS08]. Nominal [LMR20a, LMR20b]. Non [BR02, BCN^{+}23, Cai19, LBK21, Mar04b]. Non-experimental [BCN^{+}23]. Non-Gaussian [BR02, Cai19]. Non-Gaussianity [Mar04b]. Non-symmetrical [LBK21]. Noncentered [TN13]. Noncommutative [McK21]. Nonconvex [HJJ^{+}21]. Nonidentified [Gre09]. Nonlinear [AY01, FPW16]. Nonlinearity [Tsa92]. Nonmonotone [Que24]. Nonnegative [LSS^{+}19]. Nonparametric [Akr04, BCvdL20, BJM12, Cai12, Cha15, DTZ17, DJ24, Fan05a, Fan05b, GMN^{+}04, GS18, HT00a, HP04b, HKL22, LLW12, LL04, MQ04, MZ05, OR04, OWY01, PY05, PC04, RHC04, RHGF14, RR16, SS18, SVS11, Sch04, Seg12, Sør05, Teh16, WH16c]. Nonparametrics [ATH16, FJ16, KN18]. Nonprobability [EV17]. Nonregular [BM24]. Nonstandard [Ano89f]. Normal [MVDV10, ML12, Wal14b]. Normale [Maz15]. Normalized [BLNBP13, FT13]. normalizing [GM98]. Norman [Rea04]. Norwood [Fie94]. Note [Ano91b, BN90, GS13, Nyb23, Olk89b, Swi97]. Notions [MM22, Zuo21]. Novel [TX23]. November [LB22]. Nuisance [CDW95, DG95, God95, LZ95a, LZ95b, LL95, McC95, MHdL23, Rei95a, Rya95, Sev95, BLW99]. Null [FPBT22, FHS15]. Number [FWB94, FHS17, ZT09]. Number-Theoretic [FWB94]. Numbers [Lag93]. Numerical [PBSO24].

O [Wol04b]. O. [BN90]. Object [Cha14]. Object-Oriented [Cha14]. Objective [Ber11, CFL13, Gho11a, Sha16, Swe11]. Objects [LRdP^{+}04]. Observable [VGB14]. Observational [Bre87, CMC^{+}24, Cor87, DMS10, GG09, Hol87, KC14, Mar87a, Ros87a, Ros87b, Ros99, Ros02, Ros17, Ros23a, Rub90, SSP^{+}19, YSR20]. Observations [Aka87, Atk86, Bra86, Bri87a, CH86a, CH86b, Coo86, DJ24, Dra87, Gei87a, HK86, Ize87, LL87, Rao87a, Rao87b, Vel86, Wei86, Wel86]. Observed [Coh91, MY97]. Obtain [EH03]. Obtuse [Por94]. Occultation [LRdP^{+}04]. Occultations [LRdP^{+}04]. Occurrence [VACC22]. Oceanography [Bri94, CP94, Che94, Gna94, Hol94a, Pan94, Sol94, WMHL13, vS94]. Odds [VBBG11]. Official [CEN^{+}19, DS11, PS11]. Old [CRSW13, Pol24, SB16, TT23]. Oldies [Arr88, Bra88, Dem88a, Gup88, Hog88, Hot88c, Hot88b, Hot88a, Moo88, Olk88, Zid88]. Olkin [Bec07, Geo07, Sam07]. Olshen [Ric15]. OMEGA [BH90, Cri90, Fis90, Gow90, Stu90, WS90a, WS90b, You90]. On-Line [RD06]. Once [And22]. Oncology [ZJ24]. One [EHX04, EJ11, Fie92, JW94, MAK06, Nic16, RL92]. One-Half [Fie92]. One-to-One [MAK06]. One-Way [EHX04]. Online [BBK06, BH90, Cri90, DN06, Fis90, Gow90, QOS19, RWR23, Stu90, Vov21, WS90a, WS90b, You90]. Onset [ZJ24]. Onward [Goo07]. Open [SDD06]. Opera [Gei86]. Operational [Pin13]. Operations [JJ18]. Opportunities [BD04, BBF^{+}22, JS06b, SDD06]. Optimal [HJM88a, MvdKD19, RR01, STLD14, LRI23, VBW15, Fel99, BT88, GV88, HJM88b, Not88, Owe88, Spu88]. Optimization [BJMO12, DELL14, Ste93c, ZLS10]. Options [SC17]. Oracle [CY14b, Efr19a, Efr19b, Lai19, Lou19, vdV19]. Order [BR02, HP04a, Mey18a, vdV14]. Ordered [LBK21, MH16]. Ordering [Mir01]. Organisms [SSS00]. Oriented [Cha14]. Origin [LY21, Wal14b]. Origins [And88, Hog86b]. Orthogonal [BCPR08, DKSC08a, DKSC08b, JJ08]. Oscar [Fol95, Hin01]. Oscillations [BJM12]. Other [BF15, BB04b, DG13, ET86a, ET86b, Har86, Lin19, LR17, Mur05]. Our [Han14a]. Outcome [DKK10, JSH07, LD18, SRC^{+}16, VMC20]. Outcome-Dependent [DKK10]. Outcome-Wide [VMC20]. Outcomes [LS23, NS07, Rub06a, SJL^{+}12, Twe98]. Outgoing [Mor92]. Outliers [Atk86, Bra86, CH86a, CH86b, Coo86, HK86, Vel86, Wei86, Wel86]. Outperform [ZC15]. Overdose [TR10]. Overview [DD21, Che94, Fan05a, Fan05b, FMZ21, MZ05, NHZ16, PY05, Sch04, Sør05, Sut03, XYY19]. overvotes [AP02]. Ozone [Fai89, Joe89, Pic89, Raf89, Sin89, Smi89b, Smi89c, Wei89].

P [Tho02]. Packet [WTLW95]. Packing [CJLS93]. Padoan [CS12c, GTWT12, SR12]. Pair [HS09, IKN09a, ZS09]. Paired [Cat12]. Pairs [IKN09b]. Pairwise [CS12b]. Palm [Smi02]. Pandemic [CWHD22, Lin22, Muk22, NBB^{+}22]. Panel [Ano89b, Ano89f, KMSZ91, SSP^{+}19]. Pao [CO12]. Pao-Lu [CO12]. Paper [BHP15, Sto14, Ver21, Efr98]. Papers [BM16, CC14, CTB20, Edw97a]. Paradigm [DMRT13]. Paradox [AOW17, Din17a, Din17b, LRR17, AH99, GM21b, JL13]. PARAFAC [Kro92]. Parallel [Ram93]. Parameter [CS12a, DFN12, ESS10, GM18, Gus23, KDS^{+}15, LLV10, MW96]. Parameters [CDW95, DG95, God95, GK02, LMJ12, LZ95a, LZ95b, LL95, LMR20a, McC95, MHdL23, Rei95a, Rya95, Sev95, BLW99, LMR20b, Man02].

Parametric [BD19, BD08, BBK^{+}19, DPL15, HT00a, RAP14, Edw97a]. Parametrization [PRS07]. Parametrizations [TN13]. Parapsychology [BB91, Daw91, Dia91, Gre91, Hym91, Mor91e, Mos91, Utt91a, Utt91b]. Parasites [DGA^{+}20]. Parity [Mal02]. Parrondo [AH99]. Part [Ber92d, CY92b]. Partially [Coh91, MH16, TN13, VJ14, MY97]. participation [GG02]. Particle [CJLP10, KDS^{+}15, LLS11]. Particular [Lar11, Lit11a, Sch11]. Parting [Mor92]. Partitioning [CJLS93]. Parts [Law19]. Parzen [New02b, New02a]. Passive [Duf04, GST97]. Past [SBCC^{+}24]. Paternity [Ber91a, Ber91b, Che91, Eve91, Kay91, Lan91]. Path [DG13, Yu95, GM98]. Pathways [CP14]. Patient [WBR^{+}22]. Patterns [BDR^{+}22, Car14a, Car14b, FS14, Sti94]. Pawitan [LB22]. Peak [BZ21]. Peak-over-Threshold [BZ21]. Pearson [Ald95, Sti08, TOS22]. Penalised [RR17]. Penalising [SRR^{+}17a]. Penalized [KM18a]. Penalties [Gre09, EM96]. Performance [BM99, BPV20, Car19, KZZW22, Mir01, RSLGR07, GW02, Joh97]. Performing [DB92]. Perhaps [Gus23, Sin16b]. Period [SW16]. Perlman [Ber99]. Permutation [Ern04]. Persi [Ald13, DeG86d]. Personal [Box90a, Dem90a, Lin91, Moo90c, Net90, Pra90, Rob90a, Rob90b, Weg86b]. Perspective [Ama94, Bar94, BG94, Bre94c, CT94a, CT94b, DL18a, EG22, Gil86, HM16, Imb14b, JLZL04, KG89, KC14, LZ00, LM21, McC94, Moo90c, O’S86b, O’S86a, Ric86, RTW19, Rip94, Sti90, SH14, Tib94, Tit86, LRI23, Wah86, WMHL13, ZJ08]. Perspectives [BC22, Pou11]. Peter [AM19, BK08, DW16, Rit11]. Petition [DS11]. Phase [DKN18, MRSS15, Tha10, BL10]. Philosophy [Ban88, Bar88b, Ber88, Gel09, Goo88a, Goo88b, Sup88, Tuk91]. Phylogenetic [Hol03a, Ald01]. Phylogeny [SS03]. Physical [Bri94, CP94, Che94, Gna94, Hol94a, NS94, Pan94, Sol94, vS94]. Physics [FL24, LLS11]. Picture [Goo11, Kas11b, McC11, Ste11]. Piece [Wal14b]. Pieces [Ste93b]. Piecewise [FBPR18]. Piet [Jon19]. Pillow [Por94, Sen93]. Pima [CR17]. pioneers [EGG97]. Pitfalls [And91, Bai91, Cha91a, Cha91b, CN91, Gli91, MP91, Zah91]. Pitkin [Lit19]. Pitman [Ald18]. Place [Hot88b, Ros06]. Placement [BBK06]. Plate [CKLR00]. Plausible [Gre09]. Playfair [CEMR90, Wai90]. Plot [Böh16, Yu95]. Plots [Mar04a]. Point [BRM11, BDR^{+}22, BS13, Mar11, Mey18b, NHZ16, Oga18, Rei18a, Rei18b, Sch18, Vov09, Zhu18, AG01, PS00]. Points [AS10, Atk86, Bra86, CH86a, CH86b, Coo86, HK86, Hol06, Vel86, Wei86, Wel86]. Poisson [AGG90a, AGG90b, Bar90a, Che90, DE86, Goo86a, Goo86b, Hey86, Sin86, Sol86, Ste90, TT23, Wat90, WV94]. Policy [Bre88b, Dav88, DELL14, Fie11a, Fre87, Gei87c, Han11a, Hod87a, Hod87b, Hub87b, Kad87b, Kal11, Mad87, Nee88, Smi87]. Politics [Hot88a]. Polls [PS11]. Pollutant [Gin88]. Pollution [BGST04]. Polygraph [FS05, Gas87a, Gas87b, Gei87b, Gla87, Gol87a, Kay87, KR87, Wit87]. Polymerase [NA94]. Polynomials [BCPR08, DKSC08a, DKSC08b, JJ08]. Pooled [CWHD22, DLS14]. Popper [Daw04, Arj16]. Population [AB09, BNF^{+}07, Cha15, FPDE14, GT94, KS07a, KS07b, RAG14, RM07, SKBS20, TD07, DN06]. Population-Level [FPDE14]. Populations [Coh91, Don91, Jag91a, Jag91b, PB14, Saw91, DWZ97]. Portnoy [Ell98, HS22]. Posed [Gil86, O’S86b, O’S86a, Ric86, Tit86, Wah86]. Position [Bar90b, Ber90b, Ber90c, Cox90b, Fre90, Kad90b, Leh90b, Lin90a, Lin90b, Mou90]. Positions [Mar04a]. Positivity [Li20]. Possible [Tan90]. Post [AWZ20, DN06, LPE15]. Post-Denoising [AWZ20]. Post-Model-Selection [LPE15]. Post-Purchase [DN06]. Posterior [Fra11a, LL07, MGC15, NWEE20, Rob11, SX11, Zha11b, vHB21]. Posttest [DTL05]. Postulates [BBFT17]. Potential [Rub06a, XR15]. Potentially [AC89, Beg89, Ber89c, Cor89, Har89, KG89, LW89, Roy89, War89a, War89b]. Power [Ban20, CI00, MH22, Per05, RRSM20, GKT02]. Powering [CMN^{+}19]. Practical [CWHD22, DK87, DMJ23, Gey92a, Lin87a, Lin87b, Mur18, RR17, RLLKV23b, SSS20, Sha87a, Sha87b, SRR^{+}17a, Spi87a, Spi87b, Spi87c, Wat87, CTSH92, Gel92, GR92b, Gey92b, Mad92, Pol92, RP92, RL92, Ros92, Sch92, Tie92]. Practice [BV23, GGM09, Lem94, Mee11, PW00, Rao11a, Sed11, SFR10, Slu11, ZKS19]. Practices [FHH23]. Practices-a [FHH23]. Pragmatism [Gel11]. Praise [DZ23, Hac90]. Pranab [GS08]. Precise [Bay87, BD87a, BD87b, CB87, Cox87, Eat87, Kad87a, Zel87, BBW97]. Precision [Gas87a, Gas87b, Gei87b, Gla87, Gol87a, Kay87, KR87, Pac10, Wit87]. Predict [CBD^{+}15, Shm10]. Predicting [WBR^{+}22]. Prediction [BO17, BH07a, BH07b, BMW07, ESS10, GLX12, Gew92, Has07, Oga13, Par23, Rao87a, TWDM13, TNM22, THNC03, Aka87, Bri87a, Dra87, Gei87a, Ize87, LL87, Rao87b]. Predictions [GWVV09]. Predictive [Bjø90a, Dra87, LL07, LQMT13, LMW07, RTW19, Bjø90b, But90, LTH90]. Predictor [FHS17]. Preference [FPBT22]. preferences [KB98]. Prehistory [Hal03]. Prentice [HK22]. Preparedness [NBB^{+}22]. Presence [CDW95, CK03, DG95, Gew92, God95, LZ95a, LZ95b, LL95, McC95, Rei95a, Rya95, Sev95]. Present [Bar90b, Ber90b, Ber90c, Cox90b, Fre90, Kad90b, Leh90b, Lin90a, Lin90b, Mou90, SBCC^{+}24]. presented [Efr98]. Presidential [Car11, Mar11, PS11, AP02]. pressure [RS97]. Pretest [DTL05]. Prevalence [HR16]. Price [Sti18]. Pricing [GS06]. Primal [HJJ^{+}21]. Principal [CF08, IJ23, JD21, Rub06a, TJL23]. Principle [Bjø14, Daw14, Eva14, Han14b, May14b, SQH18, May14a]. Principled [RR17, SRR^{+}17a]. Principles [DHM^{+}17, Fra14, NBB^{+}22, Rub90, SNDS90, TW17, YS22]. Prior [CP14, CV08, DB92, DFN12, Dun17, EJ11, Gel09, Gus05, NWEE20, Sco17, CI00]. Prior-Data [NWEE20]. Prior-Sensitivity [CP14]. Prior-to-Posterior [NWEE20]. Priors [BBFT17, Ber89b, Ber11, FPBT22, Gho11a, Gre09, RR17, SVS11, SRR^{+}17a, Swe11]. Privacy [Fie06b, ROSS18]. Private [BL20]. Probabilist [Sti06a]. Probabilistic [Bri03, BOG^{+}19a, BOG^{+}19b, Cla90, Cli90, CJLS93, Fei93, FL93, HJ19, Kad90a, Kru90, MY90a, MY90b, Sha93, Spi87a, SH19, Tan90, WB90, Win90, Wol90, DK87, Lin87a, Lin87b, Sha87a, Sha87b, Spi87b, Spi87c, Wat87]. Probabilities [Hac90, KG05, Let08, Whe21]. Probability [Aka90, Ald90, AS93, Ald17, Arr88, Bar89, Bel93, BF15, Ber86, Ber92c, Ber92d, Box90b, Bra88, CY92b, CO12, Cut92, Daw04, Dem88a, Dem90b, DS92, Efr90, Fin86, Fis86a, Fis86b, Fre86, GZ86a, GZ86b, Gew92, Goo86c, Gra92, Gri92, Gup88, Hac90, HJ01, Hog86a, Hog88, Hot88c, Hot88b, Hot88a, Lap86, LQMT13, LK20, Lin87c, McK21, Moo88, Moo90b, Mor86, Olk88, Par04, Pla89, Por94, RCR09a, RCR09b, RSLGR07, Rub90, Sei86, Sha86a, Sha87c, Sha90b, Sha90a, Sha16, Sha21, Smi92b, SNDS90, Ste93c, Sti86, Sti94, Sto86b, Sud86, Sup86, TW17, Tsa92, Ver21, Weg86b, Win86, Zab89a, Zab89b, Zid88, CR96, Edw97b, Fie97, Hal99]. Problem [Bay88, CY14b, DuM88b, Gel14, Hed88, IG88a, IG88b, JHS05, LPT88, LLS11, OS21, Por94, Rao88, RR88, Shp14, Fer89a, Fer89b, Fre89, Rob89, Sak89, Sam89]. Problems [BF15, Bri03, CJLS93, ES95, Eva12, Gil86, MC92, O’S86b, O’S86a, Ric86, Sen93, Ste93c, SQH18, Tit86, Wah86, ZZ12, Ano96, SWY00, Twe98]. Procedure [QOS19]. Procedures [CH12, Gas87a, Gas87b, Gei87b, Gla87, Gol87a, Kay87, KR87, RR88, Wit87]. Process [Ash88, BRM11, BS13, JJG17, LW06, Oga18, QMJM22, SVS11, Zhu18, AG01, PS00]. Processes [BJP13, CH89, DMRT13, DPP18, Dud89, FBPR18, Fen16, GZ89, JW94, KG21, Lin19, Mey18b, OS22, Pem95, Pol89a, Pol89b, Pyk89, Rei18a, Rei18b, Sch18, MY97, PT11]. Processing [ZLS10]. Producing [CEN^{+}19]. Professional [PW00]. Professor [AB15, Muk24, Swi97]. Profile [Kay91]. Profiles [BDN94, Lew94]. Profiling [Ber91a, Ber91b, Che91, Eve91, Kay91, Lan91, NS07]. Profit [ABHK00]. Programmes [DGA^{+}20]. Programming [BL10, Cha14, Lub91, Sha93, XHC14]. Progress [Fri06, Gay06, Han06a, Han06b, Har89, Hol06, MC92, Sam18, Sti06b]. Projection [GGC13, MMTW01]. Projections [RAG14]. Prokhorov [She92]. Promise [VJ14]. Propensity [DN06, Hut16, RW06, Sin16a]. proper [BW23]. Properties [HRR11, WH16c]. Property [CS12b]. Proportion [BCD01]. Proposal [Sun89]. Proschan [HM95]. Prospect [Oga13]. Prospects [Cox91, DP91a, DP91b, KM91, Nor91]. Protecting [Cox91, DP91a, DP91b, GNR^{+}24, KM91, Nor91]. Proximal [PSW15, TYC^{+}24]. PRR [Mar11]. PS [Tan07a]. Pseudo [BRM11]. Pseudo-Score [BRM11]. Pseudorandom [Lag93]. Psychiatric [MA87]. Psychiatry [Eve87a, Eve87b, Fle87, Gre87b, Gre87a, Gut87, MA87, Tur87, Ver87, Zub87]. Public [BG05, Fie11a, Han11a, Kal11, Lai10, Muk22, SCN^{+}10]. Publication [DB92, GST97, Hed92]. Publications [Ano86e, Ano87f, Ano88e, Ano89e, Ano90b, Ano91e, Ano92d, Ano93b, Ano94g, Ano95d]. Puerto [PT11]. Purchase [DN06]. Purposes [Har89]. Pushing [SRR^{+}17b]. Puzzle [BHBNKM99, Kas99]. PVP [Kay87].

QQ [Mar04a]. Quadratic [LQ03]. Quality [Ros06, WM04]. Quantification [STG13]. Quantifying [CCCH08, Dos08, MY90a, NMK08a, NMK08b, RN13, ZL08, Cla90, Cli90, Kad90a, Kru90, MY90b, Tan90, WB90, Win90, Wol90]. Quantile [CJLV06a, CJLV06b, DJ24, DLM06, EH03, LMN06, Par04, Yu06]. Quantitative [Bre88b, DuM88a, FZ88a, FZ88b, FJH^{+}89, Has88, KT88, MD88, Mor88, SAL^{+}20, DWZ97]. quantization [PCT^{+}98]. Quantum [MH93, RSK22, Wan12, WWZ16, WS20]. Quarter [IR23]. Quasars [Mor14]. Quasi [HLO05, vdGM12]. Quasi-Likelihood [vdGM12]. Quasi-Monte [HLO05]. Quest [FBW^{+}16]. Question [KK15]. Questionable [FHH23]. Questions [Fie06a]. Quick [Fra11a, PT11, Rob11, SX11, Zha11b]. Quiet [Ans03].

R [Bar92a, BN90, DeG87a, Efr98, FC11, Goo11, LP11a, Lar11, Lit19, McC11, Pan11, Sch11, Ste11, Wel09, Zei18, Cha14, Lan14, LM14a, SGKR06]. R-Estimates [SGKR06]. R. [Ald97, And96, Bar89, Efr98, Fie97, Kar92, Pla89, Rao92, Sei92, Zab89a, Zab89b, Zab92]. Race [BZ21]. Radiation [Mar04b]. Radon [GKLP99, Gin88, Tei88]. Raiffa [Fie08]. Raised [Hol06]. Ralph [Hol01]. Ramanathan [Ket01]. Random [AS10, ATPP21, BLNBP13, Cam91, FT13, FC11, GMS20, Har91b, Ize21, LP11a, LK20, LC08, LD18, MN11, MVDV10, MT11, RS97, Rit11, Rob91a, Rob91b, SKBS20, SGJC13, SFR10, Spa91, Spe91, SK91, Tho91, ZLK10, KM98]. Randomisation [Jen23]. Randomization [BV23, Chu17, Din17a, Din17b, DMJ23, Fre08, GV23, Gio23, IR23, LS23, Ram93, RLLKV23b, RLLKV23a, TSW21]. Randomization-Based [Chu17, Din17a, Din17b, LS23]. Randomized [AG05, Bai17, BC91a, BC05, BCD05, Bya91, CMC^{+}24, Dup91, GM05a, GM05b, IKN09a, IKN09b, Lev91, Lin91, PA22, Ros02, Roy91a, Roy91b, Sim91, Tho05, Zel91, ZS09]. Randomly [Mag93]. Randomness [Gri92, Vov21]. Range [Ber92a, Ber92b, DH92, Par92, Raf92, Smi92a]. Rank [Dra88a, FM18, HP04a, LYL22, MH16, Bic88, Dra88b, HA88, KP88, Mar88b, Wel88]. Rank-Based [Dra88a, Bic88, Dra88b, HA88, KP88, Mar88b, Wel88]. Ranked [Wol04a]. Rankers [LYL22]. Ranks [BT18]. Rao [DeG87a, Mee11, Sed11, Slu11]. Rapid [YS22]. Rapid-Response [YS22]. Rare [DLS14, DJ15]. rate [AG01]. Rates [BSJ93a, BSJ93b, Bro93, CG93, DP93, GR93, GM21a, KS93, KR87, Raf89, SW93]. Ratings [Ald17]. Ratio [Böh16, BM24, OS21]. Rational [Han87a, Bha87, Bri87b, Dah87, Han87b, Ris87, Shi87, Sol87]. Ratios [VBBG11]. Raymond [LC24]. Reach [BR94a, BR94b, Bre94b, DS94, EFK94, FW94a, LL94, Ste94]. Reaching [LW06]. Reaction [NA94]. Reactive [XHC14]. Reader [SAA^{+}92]. Real [JCT22, Twe98]. Real-Time [JCT22]. Reality [NBB^{+}22]. Realized [Hwa90, VJ14]. Reappraisal [GGBS^{+}23]. Reassessment [OC10]. Recall [Car11, Jim11, Mar11, PS11]. Recapture [BF16, BSA16, FSB16, HR16, JLMD16, SB16]. Received [Ano86e, Ano87f, Ano88e, Ano89e, Ano90b, Ano91e, Ano92d, Ano93b, Ano94g, Ano95d]. recognition [HS98]. Recollections [Dem88a, RC11, Weg86b]. Recommendations [HTT20]. Reconciling [Sha16]. Reconsidering [Sin16a]. Reconstruction [SS03]. Recovery [HJJ^{+}21]. Recursive [CP14]. Red [Gel14]. Red-State [Gel14]. Rediscoveries [Wal14b]. Reduction [BNPS13, Chr07, Coo07a, Coo07b, CF08, Li07, LN07]. Referee [Ano87a, Ano89a, Ano92a, Ano93a, Ano94a, Ano95a]. Refereeing [BGLC93, Bil93, Car93, CGP^{+}93, Gen93, vZ93]. Referendum [Mar11, Car11, DS11, Jim11, PS11]. Reflection [Lin22]. reflections [Gre97b]. Refreshment [DHR^{+}13]. regarding [Swi97]. Regazzini [LP11b]. Regimes [STLD14]. Regions [NNLL18]. Regressin [OWY01]. Regression [Ald05, Atk86, BW07, BC07c, Bel87, Ber12, BK17, BPV20, Böh16, Bra86, BO07, Bre88a, BBB^{+}19, BBK^{+}19, CKLR00, CH86a, CH86b, CR17, Chr07, CM91a, CM91b, Coo86, Coo07a, Coo07b, CF08, DTZ17, DJ24, DuM12a, EG22, Eub88, FM93, FLR93a, FK03, FN86a, Fre08, GJSE91, Gel07b, Gel07c, GHV12, GH91, GS18, HV87, Har91a, HT88, HL93a, HL93b, HK86, Jon91, LW06, Li07, LN07, Lit07, Lit19, Loh07, LS17, LD24, Mar87b, MT12, MvdKD19, Mul93, Pac10, Pfe07, QOS19, Ram88a, Ram88b, RAP14, SSS20, Sil91, Ste87a, Ste87b, Sut03, Thi87, Vel86, Wah88, Wei86, Wel86, Zuo21, CI00, Ell98, Dem86, DM93, Eri86, Fel86, FLR93b, FN86b, Kad86, Mad86, Mos86, Nat86, PM93, Wol86, ZLH93]. Regressions [GL06]. Regularization [BH07a, BH07b, BMW07, CTB20, Has07, LMJ12, MvdKD19, Pou11, WZ12, ZZ12]. Regularizations [CHP22]. Regularizers [NRWY12]. Regulatory [Ash88, JLZL04]. Reinhart [Oga18, Sch18, Zhu18]. Rejoinder [AET09b, Agr92a, Ano91f, AGG90b, BSJ93b, Ban93b, BH87b, BC07b, BCW87b, BQW06b, BR94b, Ber92b, BD87b, Ber92d, Ber91b, BGHM95b, Bjø90b, Boo86b, Bre94b, Bre90b, Bri08b, BOG^{+}19b, BH07b, Büh20b, BKB^{+}19, CJLV06b, Cha91b, CH86b, CY92b, CT94b, CM91b, Clo92b, Coo07b, CW93a, Cra16a, DPR12a, Dem88c, DKSC08b, Din17b, DHS^{+}19b, Dra88b, DuM12b, DP91b, Edd86b, ET86b, Efr08b, Efr10b, Efr19b, Eve87b, Fan05a, Fer89b, Fie11b, Fis86b, FLR93b, Fle92b, Fra11b, FN86b, FZ88b, FW94a, Gas87b, GR92b, Gel07b, GZ86b, GK15b, Gey92b, GM05b, GR94b, Gho11b, GS11b, Goo86b, Goo88b, Gra93b, Han06b, Han87b, HT86b, HL93b, HJM88b, Hei89b, Hod87b, IKN09b, Imb14a, IG88b, Jag91b, KS07b, Kas89b, Kas11a, Ken89a, Le 86b, LN09b, LR92b, LZ95b, Lin87b, Lin90b, Lit11b, LMR20b, MC90b, May14b, Men94b]. Rejoinder [MG20b, MM06a, Mor11, MY90b, NMK08b, O’S86a, PL22, Pol89b, Ram88b, Rao87b, Rao11b, Rei88b, Rei95a, Rei18a, RCR09b, Rob90b, RLLKV23a, Rob91b, Roe94b, Ros87b, Roy91b, Rub06b, RH87b, SWMW89b, Sch87a, Sha86c, Sha87b, Sha90a, SRR^{+}17b, Sin16b, Smi89c, Spi87c, SDLC93b, Ste87b, Thi86b, Tru89b, Utt91b, War89b, WH16b, WS90b, XT11b, You94b, Zab89b, dP89a, AG01, ACS01, Ber99, Bre01, Ell98, HK02, HT00b, HZ00, HMRV99, SWY00]. Related [CJLS93, OC10, QMJM22, Tit04]. Relatedness [AB09]. Relating [FJH^{+}89]. Relation [Smi92b, Hal99]. Relationship [DKN18]. Relative [EJ11]. Relatives [FJ16]. Relaxation [Gre09, LCX21]. Reliabilities [But00]. Reliability [BK87, Bas87, Coo16, Ger87, HP04b, KMWAC06, RH87a, RH87b, Sin87, Sin16a, Wei87, WGHR06]. Reliable [BQW06a, BQW06b, FN06, Koe06, Wan06]. Reliably [Ros23a]. Remarks [Mor92, Spe90, Ste93b, Fre99]. Remediation [GKLP99]. Remember [Ken91]. Remembered [Sti07b]. Remembering [Fer03, FvZ08, Hin01]. Reminiscences [Han87c, Leh93, Moo90a]. Remote [CK03, GKRS05]. Rényi [KM18a]. Repairable [Lin06]. Repeated [MVDV10]. Replicability [BH23, CDR23, MTP^{+}23, Par23]. Replicated [DJ24]. Replication [FHH23, GR92b, KZI09, MH22, Utt91a, BB91, Daw91, Dia91, Gre91, Hym91, Mor91e, Mos91, Utt91b]. Reply [Car14b]. Report [AET09a, BGLC93, CGP^{+}93, LKS04, Pan94, SAA^{+}92, Bil93, Bri94, Car93, CP94, Gen93, Gna94, Hol94a, Sol94, vZ93, vS94]. Reporting [VACC22]. Representations [SAY22]. Represented [CW93b]. Representing [Win90]. Reproducibility [CDR23, FBW^{+}16]. Require [Fie06a]. Research [BGGM06, Bat86, Ber89a, Bil86, BFK86, Dal89, Dra87, Edd86a, Edd86b, Flo89, FHH23, Goe86, GS11a, Hel11, IR23, JS06a, JS06b, JJ18, KMSZ91, KJED09, Lil88, Mei11, Mit89, Mol92, Oga13, RR14b, Sco86, SCN^{+}10, Sil09, Smi89a, Sun89, Tru89a, Tru89b, Utt86, Weg86a, Weg89, Wes11, Zid89]. Researcher [SAA^{+}92]. Researchers [ABC^{+}91]. Residual [BRM11]. Resolution [NS94, SAL^{+}20]. resources [Twe98]. Response [BV23, CY14b, DHS^{+}19b, DMJ23, Gio23, IR23, Jen23, RLLKV23b, RLLKV23a, Sto14, YS22, MW96]. Response-Adaptive [DMJ23, Gio23, IR23, Jen23, RLLKV23b, RLLKV23a]. Responses [DM93, FLR93a, FLR93b, PM93, Sut03, ZLH93]. Restricted [GS18]. Restriction [Gus23]. Restrictions [Gus23, LD18, Mey18a]. Results [DS11, DLS14, PS11, TNM22]. Reverend [Bel04]. Reversal [Nic16]. Reversing [PC12]. Review [Ama94, ABB^{+}23, And87, Ano94b, ARC21, BBF^{+}22, BDN94, Bar92a, Bar94, Ber94b, Bic88, BG94, Bjø90a, Bjø90b, BNPS13, BH02, BM24, Bre94c, Bur89, But90, Cat12, CV95, CT94a, CT94b, CMC^{+}24, DLS14, DPP18, Dra88a, Dra88b, Duf04, EM06, Fie92, GK87, Gol87b, HB17, Hei89a, Hei89b, HA88, HBM12, IRÁ22, JT90, KMSG06, KP88, Lem94, LTH90, Lew94, LSS^{+}19, LG17, Mar04b, Mar88b, McC94, Mur18, Oga18, Per87, Pre87, Rei18a, Rei18b, Rip94, Roe94a, Roe94b, Sch87b, Sch87a, Sch18, SS99, Sen87, SW16, Stu10, Sud94, Tho94b, TNM22, Tib94, Wei94, WH16c, Wel88, YK16, ZJ24, Zhu18, GG02, MY97]. Revision [AS21]. Revisited [Daw86a, Daw86b, Fis86c, Lin86, ML14, Pra86, RCR09a, RCR09b, Sha86b, Sha86c]. Revisiting [VK21, Wal14a]. Revolutionary [Wai03]. Reweighting [DD21]. Ribatet [CS12c, GTWT12, SR12]. Rice [Lin19]. Richard [Sam99, Blo01a, Buc16, Ric15, Sti18]. Richardson [Muk24]. Rico [PT11]. Rigorous [Lin20]. Rise [FYL10]. Risk [And88, BC20, Bre88b, Bri03, GL96, GKRS05, Hat88, KT88, Mor88, Nee88, Rya03]. Risks [Bre88b, DuM88a, FZ88a, FZ88b, Has88, KT88, MD88]. RNA [AWZ20, SJW11]. RNA-Seq [SJW11]. Robbins [LS86]. Robert [Beh19, HKL17, Ran07, ZLG15]. Robins [RR14b]. Robust [Bic88, Boc16, CY14b, Dra88a, Dra88b, DELL14, FWZZ21, HA88, HRV08, KP88, Li20, Mar88b, McK04, RAP14, RR16, RSLGR07, RB23, SV18, TSW21, WH16a, WH16b, Wel88, ZJZ^{+}09, vdGM12]. Robustness [DD21, Büh20a, Büh20b, CTB20, Gol16, Kaf03, KS07a, KS07b, LMJ12, RM07, TD07, BW23]. ROC [IRÁ22]. Roguery [Bel93]. Role [AS21, Bel93, Bre87, BOG^{+}19a, BOG^{+}19b, Cor87, Cox90a, Dra87, HJ19, Hil89, HS09, Hol87, IKN09a, Jor89, LL09a, LD18, Mar87a, McC89, Ros87a, Ros87b, SSM23, SH19, ZS09, dP89b, dP89a]. Roles [Dav88, Rei95b]. Root [CP94, Men05, SWY00]. ROS [MvdKD19]. Rosenblatt [BD09]. Ross [HK22]. Roughness [GŠP12]. Roulette [LGA^{+}15]. Roussas [BS10]. Route [AS10]. Route-Length [AS10]. Routine [BH90, Cri90, Fis90, Gow90, Stu90, WS90a, WS90b, You90]. Rubin [Fie06a, LM14b, VK21]. Ruin [DE22]. Rule [WZG24]. Rules [Par23, BW23]. Run [RL92, Yu95]. Rupert [Ano91g]. Russian [LGA^{+}15].

S [CS12c, DT04, EW17, GTWT12, Han11a, Kal11, Rob11, SR12, SX11, WZG24, Zei18, Zha11b, Lub91]. S. [Zei18]. Saddlepoint [Rei88a, LRI23, But00, BN88, Dan88, HW88, Hou88, Kas88, Mit88, Rei88b, Tie88]. Safe [RGVS23]. Safety [Ber12, DuM12a, MT12]. Saga [McC16]. Sake [Gus23]. Salary [Con93, Gas93, Gra93a, Gra93b, Rob93a]. Sales [QOS19]. Salmon [KVP^{+}14]. Samad [MSY16]. Samaniego [RB20]. Sample [BT18, BFW07, CC24, ESS10, GK02, KK15, LZ00, Sha03, ZWS^{+}19, Dav98a, GW02, Leh99]. Sampler [BHJK21, CTSH92, GHSS20]. Samplers [LC08]. Samples [Box87, BD08, DHR^{+}13, EV17, ZWS^{+}19]. Sampling [APSAS17, ABT16, BCPR08, CCFR24, Cra16a, Cra16b, DKSC08a, DKSC08b, DKK10, Duf04, DS92, DG13, EMLB19, FL24, FJ16, Fen16, GHSS20, Han87c, JJ08, Lah03, LQMT13, Let08, Mee11, MH16, Rao11a, Sed11, Slu11, Teh16, TW17, Wol04a, YK16, Zas11, GM98, GL96]. Samuel [Nad02]. SARS [CKLP22]. SARS-CoV-2 [CKLP22]. Savage [Daw86a, Daw86b, Fis86c, Lin86, Pra86, Sam99, Sha86b, Sha86c]. Scalable [BPV20, LM14a]. Scale [BGGM06, DJ15, Pac10, BW23]. Scales [LL04]. Scaling [GGC13, MvdKD19, RR01]. Scan [BHJK21, LC08, PT03]. Scans [ZJZ^{+}09]. Scenarios [Gus05, SKBS20]. Scheme [GHSS20]. Schmetterer [Wit91]. Schmidt [EG22]. Schneiderman [Sim97]. Scholar [BF91]. Scholarly [Fie07]. Schrödinger [HDD24]. Sci [HMRV00]. Sci. [Ano02, Ano03b]. Science [Ano94c, Ash88, Ban88, Bar88b, Ber88, BK17, Bla88, Bre88b, Bri08a, Bri08b, CMN^{+}19, Cha22, Con88, CD21, Dem88b, Dem88c, Dem93, Fie88, Fis88, Gas88, Gew88, Gol88, Goo88a, Goo88b, Gre91, Hol88, JAAH14, Kün08, McC94, OS21, Rob88, SCN^{+}10, Sol03, Ste20, Sup88, TWDM13, WS20, Yan08, YS22, CG02, PSES98, LB22]. Sciences [Böh16, IMS90]. Scientific [And88, Fie07, MO00]. Score [BRM11, LQ03, RW06]. Scoring [Gay06, BW23]. Scott [BF91]. Screening [Gas87a, Gas87b, Gei87b, Gla87, Gol87a, Kay87, KR87, Wit87]. Sea [CP94]. Search [DZ23, FS05, HR11, SKKS14, DWZ97]. Searching [BH90, Cri90, Fis90, Gow90, Stu90, WS90a, WS90b, You90]. Searle [Wel09]. Seber [Bar16, BM16, Nic16]. Second [BC07a, BC07b, Bre87, Cor87, Efr03, Eva07, Gel07a, Hol87, Joh07, LL07, Mar87a, Ros87a, Ros87b]. Secret [Zab23]. Secretary [Fer89a, Fer89b, Fre89, Rob89, Sak89, Sam89]. Section [Ano03a, GWC04, JR18, KM18b, Rub90, SNDS90]. See [JW94]. segmentation [BM98]. Seiden [SC92]. Selecta [Gei86]. Selected [Bar92a]. Selecting [Atk14]. Selection [Boc16, DP07, DG13, Hed92, HBG15, HBM12, IG88a, LPE15, LD24, MGC15, Mor14, MSW13, RR23, SVS11, VMB08, Bay88, DuM88b, Hed88, IG88b, LPT88, Rao88, RR88]. Selective [Fan05a, Fan05b, FMZ21, HBM12, Mee11, MZ05, NHZ16, PY05, Rao11a, Sed11, Slu11, Sør05]. Self [FT96, Mey18b, Oga18, Rei18a, Rei18b, Sch18, WTLW95, Zhu18]. Self-consistency [FT96]. Self-Exciting [Mey18b, Oga18, Rei18a, Rei18b, Sch18, Zhu18]. Self-Similarity [WTLW95]. Seminar [Kle91]. Semiparametric [DTL05, XYY19]. SEMs [RR14a]. Sen [GS08]. Sensing [CK03]. Sensitivity [CP14, CST00, FHS17, GM18, IKY10, RHGF14, VTH14]. Sensor [JCT22]. separating [GW02]. Seq [SJW11]. Sequence [DTZ17, WV94, BM98]. Sequences [GR92a, WRR94]. Sequencing [AWZ20]. Sequential [ALS^{+}01, Jen23, LCL13, QKM20]. Series [BR02, Büh02, CT11, DJ24, Fai89, FK03, GŠP12, Han11b, Ion11, Joe89, Lin19, Pic89, Pol03, Raf89, Raf92, Sin89, Smi89b, Smi89c, Wei89, XT11a, Yao11]. Seroprevalence [DRRS22]. Servers [GKRS05]. Services [Ano87a, Ano89a, Ano92a, Ano93a, Ano94a, Ano95a]. Set [HJJ^{+}21, Wol04a, GL96]. Sethuraman [Hol08]. Sets [Bru18, Whe21, BH96]. Setting [Man02]. Settings [BM24, Fie11a, Han11a, Kal11]. Settle [LM21]. Seven [YS22]. Seymour [CJ07]. Shaarawi [Est03]. Shafer [MZL10]. Shanti [McD98]. Shape [Boo86a, Boo86b, Boo89, Cam86, Cre86, Goo86d, GJ18, GS18, JJ18, Ken86, Ken89b, Ken89c, Ken89a, KM18a, Mar89, MN11, Mey18a, QOS19, Sam86, SS18, Sma89, Sto89, Wat89, AG01]. Shape-Restricted [GS18]. Shayle [Wel09]. Shelemyahu [Muk05]. Sherlock [Kad09]. Shih [YZ06]. Short [DH92, Hal03, RC11]. Short- [DH92]. Should [KEHSY19, Kas10, Sil91]. Shoutir [BM07]. Shrinkage [BJM12, BZ12, CH12, DG12, FW12, GLX12, ML12, Sti90]. Shrunken [THNC03]. Side [FM18]. Sides [Gio23]. Sign [EHX04]. Signal [DS08, MW96]. signal-response [MW96]. Signals [SCtW^{+}09]. Signatures [DS11]. Signed [HP04a]. Signed-Rank [HP04a]. Significance [FHK12, RW09, WV94]. Significant [FHJ08, Hil95]. Significant-Digit [Hil95]. Silver [Cas03b]. Similarity [WTLW95]. Simple [HB08, Pol24, SCN^{+}10]. Simplicity [Tsa92]. Simpson [GM21b]. Simulated [BT93]. Simulating [GM98]. Simulation [FHH23, GR92a, MVA^{+}18, OS22, STG13, Sti91, GW02]. Simulation-Based [MVA^{+}18, GW02]. Simulations [WWZ16]. Single [AWZ20, DK20, LZD^{+}20, Yu95]. Single-Cell [LZD^{+}20]. Single-Molecule [DK20]. Sir [Kat02, Muk24, Rei94]. Sirken [GLW07]. Size [Boo86a, CC24, GGC13, KK15, LZ00, GW02, Boo86b, Cam86, Cre86, Goo86d, Ken86, Sam86]. Slopes [BW07]. Small [BFW07, Box87, BD08, DG12, GR94a, Lah03, Pfe13, Leh99, CK94, GR94b, Hol94b, SC94, Sin94, Sta94, Tho94c]. Small-Area [Lah03]. small-sample [Leh99]. Smoke [WM04]. smoking [GST97]. Smoothers [Sch04]. Smoothing [BDR^{+}22, CJLP10, MMTW01, EM96]. SNPs [CCLC09, PGP09]. Sobel [Muk00]. Social [Böh16, Mol92]. Society [BG99]. Sociological [Bar92b, Clo92a, Clo92b, Hol92, Man92, Mol92]. Soft [AS21]. Software [CC14, Coo16, SDD06, SBCC^{+}24]. Solari [Hel11, Mei11, Wes11]. Solomon [Swi92]. Solutions [EH03]. Solve [Sam89]. Solved [Fer89a, Fer89b, Fre89, Rob89, Sak89, Sam89]. Solving [BHBNKM99, Kas99]. Some [ABB^{+}23, BD04, BDN94, BC22, Box90a, CG92, DeM92, Dem90a, DLS14, Ell92, FWB94, FC92, Fle92a, Fle92b, Goo86a, Gre97b, GJ18, Han87c, Lag92, Lou92, Mar89, Moo90c, MC92, Net90, Pra90, Rob90a, Rob90b, Ste20, Tan23, LRI23, Weg86b, Fre99, DE86, Goo86b, Hey86, Sin86, Sol86]. Source [SDD06]. Sources [Fay94, Gre09, LR17, Men94a, Men94b, Nee88, Sch94a, SV06, Ski94, Zas94]. Southern [BGST04]. Soviet [BK87, Bas87, Ger87, RH87a, RH87b, Sin87, Wei87]. Space [DFN12, FPW16, KDS^{+}15, QOS19, SAC19, Tan23]. Spaces [Boo86a, Boo86b, Cam86, Cre86, Goo86d, Ken86, MSM24, Sam86, vdV14]. Sparse [BPV20, CHP22, HJJ^{+}21, LLW12, RT12, SSS20, YB17, ZZ12]. Sparsity [BJMO12, WZ12]. Spatial [AS10, BRM11, BDR^{+}22, BF16, Cha04, CS12c, CK03, CBD^{+}15, DPR12b, DMRT13, Fri07, GTWT12, GŠP12, GLSS23, JJG17, MSM24, Pac10, SR12, WH16c]. Spatial-Confounding [Pac10]. Spatio [DMRT13, Rei18a, Rei18b, Sch18, Zhu18]. Spatio-Temporal [DMRT13, Rei18a, Rei18b, Sch18, Zhu18]. Spatiotemporal [Oga18]. SPDE [SAY22]. Special [Ano03a, Ano20, CDR23, ES95, GWC04, JS06b, JR18, KM18b, RHC04, RC04, SS18, WZ12, ZMG09, Ano96]. Species [LQMT13]. Specific [GL06, LMJ12]. Specification [DFN12, Leh90a, Sco17]. Specifications [CV08]. Specified [ACS01]. Spectral [ZJ08, vDK04, STW00]. Spectroscopy [Bur92, Kro92, LR92a, LR92b, deL92]. Speed [MO00, WTLW95]. Speedup [FL93]. Spencer [Sam99]. Spent [BF09]. Spherical [Her19, JJG17]. Spherically [BS90, BS12]. Spiegelhalter [Muk24]. Spline [HK02]. Splines [Bre88a, Eub88, HT88, Ram88a, Ram88b, Wah88, EM96]. Split [GHSS20]. Sports [Ald17]. Spurious [Ald95]. Spyglass [Pos23]. squared [PK97]. squared-error [PK97]. Squares [Bel87, HV87, Hil89, Jor89, Mar87b, McC89, Nyb23, Ste87a, Ste87b, Thi87, dP89b, dP89a, Ell98, Hal99]. Stab [Ber94a, Efr94, GCL94, Hin94, MM94, Nav94, Sch94b, You94a, You94b]. Stability [CTB20]. Stabilization [Coh91, Don91, Jag91a, Jag91b, Pol24, Saw91]. Stabilizing [Li20]. Stable [JW94, OS22, Seg12]. Standard [ET86a, ET86b, Har86, vHB21].

Standardization [KC14]. Star [BJM12]. Stat [Lub91]. State [FPW16, Gel14, JLMD16, KDS^{+}15, Tan23, Dav98b]. State-Space [KDS^{+}15]. States [EGG97, CEN^{+}19]. Statist [HMRV00]. Statist. [Ano02, Ano03b]. Statistic [AS10, EHX04, HBG15]. Statistical [ABC^{+}91, Ama94, AY01, And91, Ano89f, Ano94c, BD04, Bai91, BN88, Bar94, BH87a, Bar92b, Bat86, BG99, Ber92a, BM99, Ber03a, BGST04, BG94, Bil86, Bla88, BH02, Boo89, BM98, Bre94c, Bre01, BOG^{+}19a, BOG^{+}19b, BFK86, CMN^{+}19, CKLP22, CC14, CCCH08, Cha91a, Cha91b, CT94a, CT94b, CN91, Clo92a, Clo92b, CWHD22, Con88, CS12c, Cox90a, CD21, Dan88, DS08, DPR12b, DN06, Dem88b, Dem88c, DE86, DGA^{+}20, DJK^{+}20, DWZ97, Dra87, DK20, DS92, Edd86a, Edd86b, ET86a, ET86b, Eud96, FWZZ21, FL24, FHK12, Fie88, Fie08, Fis88, FM18, Fra14, Fri08, FMF13, GTWT12, Gas87a, Gas87b, Gas88, GGM09, Gei87b, Gel11, Gew88, Gil86, Gla87, Gli91, Goe86, Gol87a, Gol88, Goo86a, Goo86b, Goo11, Har86, HR11, Hey86, HJ19, Hil89, Hil95, HW88]. Statistical [Hol88, Hol92, Hou88, Ize01, JS06b, JT90, Jor89, Kas88, Kas11b, Kay87, Ken89b, Ken89c, Ken89a, KR87, KZZW22, LY21, Law16, Leh97, LRdP^{+}04, Lin08, Lin06, LSS^{+}19, MO00, MH93, MP91, Man92, Mar89, Mar88a, MZL10, Mar11, ML14, McC94, McC89, McC11, MMT07, Mit88, Mol92, NS94, NBB^{+}22, NMK08a, NMK08b, NS07, O’S86b, O’S86a, Par04, Peñ06, PW00, PSES98, PB10, Rei88a, Rei88b, RRSM20, Ric86, RTW19, RSK22, Rip94, Rob88, SJW11, Sam91, Sco86, SGKR03, Sen16, SR12, Sin86, Sma89, Sol86, SAL^{+}20, SH19, Ste20, Ste11, Sto89, TWDM13, Tib94, Tie88, Tit86, TOS22, TJL23, Utt86, LRI23, Wah86, Wai03, WLLH21, Wat89, Weg86a, WMHL13, Wit87, Zah91, Zha11a, ZL08, ZJ24, dP89b, dP89a, BBC^{+}02, CG02, Smi02, BH87b, Ber92b, Dav98b]. Statistical [DH92, Fel87, Par92, Por87, Raf92, Smi92a, Bar92a, LB22]. Statistician [Lin22, Muk22]. Statisticians [Che94, PW00, SCN^{+}10, EGG97]. Statistics [AET09a, AET09b, ABB^{+}23, AMV19, Ano03a, Arr88, AS21, Ban88, Ban93a, Ban93b, Bar88b, Bar90b, BC91a, Ber88, Ber90b, Ber92c, Ber90c, Ber04, Bes05, Box90a, Bra88, Bri94, Bya91, CP94, Cha04, CY92a, CO12, Cna93, Con93, Cox90b, Cra86, CK03, Dav98b, DeG86f, Dem88a, Dem90a, Duf93, Dup91, ES95, Eve87a, FWB94, FL24, Fie92, Fre87, Fre90, Fri07, Gas93, Gei87c, Gel09, Gel10, Gil14, Gna94, Goo88a, Goo88b, Gra93a, Gra93b, Gre10, Gup88, Hah93, Hal09, Hod87a, Hod87b, Hog88, Hog93, Hol94a, Hot88c, Hot88b, Hot88a, Hub87b, JJG17, JT04, Kad87b, Kad90b, Kat02, KG05, LS86, Lar11, Leh90b, LLJ09, Lev91, Lin90a, Lin90b, Lin91, LKS04, Lit11a, Lub91, Mad87, Mar07, Men05, MA87, Mol92, Moo88, Moo90c, MM22, Mou90, Mur05, NP93, Net90, Nol92]. Statistics [Olk88, Orc93, PT03, PSW15, PBSO24, Pos23, Pra90, RGVS23, RHC04, Rao92, Rob90a, Rob90b, Rob93a, Rob93b, Roy91a, Roy91b, SY12, Sch11, Sch04, Sim91, Smi87, Smi92b, Sol94, Sol03, SG09, Ste05, Sti94, Sup88, Weg86b, Wei88, Woo93, Wu93, Wyn93, Zel91, Zid88, vS94, Ald01, Ano96, CR96, Dem98, FT96, Fre99, GKT02, Gre97b, PS00, Sti96, Ber92d, CY92b, Cut92, Eve87b, Fle87, Gew92, Gra92, Gre87b, Gre87a, Gri92, Gut87, Hog86b, Kle91, MA87, Smi92b, Tsa92, Tur87, Ver87, Zub87]. Status [FPDE14]. Stein [ABB^{+}23, AGG90a, AGG90b, Bar90a, BS90, BS12, BZ12, Che90, DeG86a, GS12, Men05, PC12, Ste90, Wat90]. Stellar [LRdP^{+}04]. Stephen [BK24, Ell98, HS22, ST13]. Stepwise [HTT20]. Stigler [BK24]. Stochastic [Ald01, BBF^{+}22, BGHM95a, BGHM95b, BDP18, BJP13, Che10, Fri95, GC95, Gey95, JCD^{+}05, KN18, LY21, LW06, MG08, MVA^{+}18, RSG95, STR^{+}17, Sti91, Won95, Yu95]. Stopper [CTSH92]. Stories [Büh14, Kad09, MR14]. Story [Sti07a, Leh97]. Strata [JD21]. Strategies [CC14, Efr14, KS07a, KS07b, KLDR09, LCL13, RL92, RM07, SVS11, TD07, vDM10, MY97]. Strategy [GS06]. Stratification [Rub06a]. Stratified [HB08]. Strength [You90]. Strengthening [Jen19]. Strong [AS21, Bjø14, Daw14, Eva14, Han14b, May14b, Per05, May14a]. Structural [CP94, CPW15, VJ14]. Structure [AB09, ATPP21, BM99, BH90, Cri90, Fis90, Gow90, Ros17, Saw91, Stu90, WS90a, WS90b, You90]. Structured [BJMO12, vDK04]. Structures [Con93, Gas93, Gra93a, Gra93b, KLDR09, Que24, Rob93a]. Struggles [BC07c, BO07, Gel07b, Gel07c, Lit07, Loh07, Pfe07]. Stu [De 20]. Stuart [De 20]. student [Joh97, Leh99]. Studentization [Pol24]. Studies [AB09, BGST04, BH23, CCCH08, CCLC09, CMC^{+}24, DHR^{+}13, Dos08, HB08, KC14, KZI09, LL09a, MH22, NMK08a, NMK08b, PGP09, PC04, Ros88, Ros99, Ros02, Ros17, Ros23a, Rub90, Rub06a, Sav88, SJL^{+}12, SCtW^{+}09, TCC^{+}09, VMC20, WM04, YSR20, Zha11a, ZL08, ZMG09, GL96]. Study [BL20, Bre87, Car11, Cor87, DTL05, DuM12a, FHH23, Hol87, Ize01, Mar87a, MTP^{+}23, QKM20, Ros87a, Ros87b, SFR10, SGKR06, SDD06, BGST04]. Subject [VACC22]. Subjective [Ber86, Fin86, Fis86a, Fis86b, Goo86c, RC11, Sei86, Sha16, Sto86b, Sud86, Sup86]. Subset [HTT20]. Succeed [ZKS19]. Success [FHH23]. Succession [WZG24]. Sue [Twe98]. Sufficientness [BBFT17]. Sujit [Muk97]. Summation [DZ91]. Super [BCvdL20, SKBS20, SAL^{+}20]. Super-Efficient [BCvdL20]. Super-Resolution [SAL^{+}20]. Superefficiency [Vov09]. Superpopulation [GK02]. Support [DGA^{+}20, MM06b]. Supremacy [RSK22]. Sure [GM21b]. Surveillance [PT03]. Survey [Agr92b, Agr92a, BH92, BC07c, BD19, BDPW19, Boo89, BO07, BO17, CEH^{+}17, Duf92, EF92, Gel07b, Gel07c, Han87c, Ken89b, Ken89c, Ken89a, Kre92, Lah03, LRdP^{+}04, LW92, Lit07, Loh07, LR17, LS17, Mar89, Mee11, Meh92, Pfe07, Rao11a, RH87a, Sed11, SW17, Slu11, Sma89, Sto89, Sui92, Wat89, YK16, BK87, Bas87, Ger87, RH87b, Sin87, Wei87]. Survey-Weighted [CEH^{+}17]. Surveys [GK02, HB17, SC17, Sha03]. Survival [Akr04, BCM00, Dos08, LW06, Sin95, Sin16a, SQH18]. Swan [HR11]. Swartz [Ano96]. Swinging [Hod17]. Switching [JHS05]. Sylvia [Muk24]. Symmetric [BS90, BS12]. Symmetrical [LBK21]. Symmetry [Per87]. Synchronization [DKN18]. Syncretism [Gre10]. Synthesis [DPCA14, BW07, BDP18, BL20, Daw04, JT04, POSD13]. Synthesizing [WM04]. Synthetic [DH24]. System [BQW06a, BQW06b, Coo16, FN06, Koe06, Wan06, WGHR06, XR15, MY97, MdL98]. Systems [BGHM95a, BGHM95b, CW93a, DK87, Dem93, Fri95, GC95, Gey95, GS93, Gri92, Hil93, Lin87c, Lin87a, Lin87b, Lin06, Mad93, MAK06, Nor93, Pea93, RSG95, Sha87a, Sha87b, Sha87c, Sob93, Spi87a, Spi87b, Spi87c, SDLC93a, SDLC93b, Wat87, Whi93a, Won95, Yu95, But00, MW96]. Székely [GPW23].

T [Ano96, Hed05]. T. [DeG86e]. Table [Ano16b, Ano23f, Ano23d, Ano23e, Ano24d, Ano24e, Ano24f, How98]. Tables [Agr92b, Agr92a, BH92, Duf92, EF92, Kre92, LW92, Meh92, ROSS18, Sui92]. Tadeusz [AB15]. Taguchi [MW96]. Tailed [KMR^{+}19]. Taiwanese [LRdP^{+}04]. Take [Gio23]. Tale [DGA^{+}20, Lin87a]. Tang [FC11, LP11a]. Tangent [vdV14]. Tavia [Gel97]. Taylor [Twe98]. Teacher [BF91, Har03, Mor03]. Teaching [Hot88c, Hot88a]. Techniques [BO17, CKLR00, Dem93, Tan23]. Technology [Fri06, Gay06, Han06a, Han06b, Hol06, Sti06b, WS20]. Tectonics [CKLR00]. Ted [Ale96]. Telecommunications [Mal02]. Tell [Con93, Gas93, Gra93a, Gra93b, Men05, Rob93a]. Template [VMC20]. Temporal [DMRT13, HT00a, Rei18a, Rei18b, Sch18, Zhu18]. Tercentenary [Bel04]. Term [BGST04]. Test [BT88, EM06, EW17, Gas87a, Gas87b, GGM09, Gei87b, GV88, Gla87, Gol87a, HJM88a, HJM88b, Kay87, KR87, LS23, Not88, Owe88, PQW14, SSVV11, Spu88, Wit87]. Testing [BD87a, Ber03b, CCCH08, CS12b, CWHD22, Coo16, DP07, DSB03, GM21a, GS11a, Hel11, Lan15, Mar04b, Mei11, NNLL18, NMK08a, NMK08b, RWR23, RW09, Vov21, Wes11, ZL08, BBW97, Bay87, BD87b, CB87, Cox87, Eat87, Kad87a, Zel87, Leh97]. Tests [Chu17, DLS14, HP04a, LL04, LGF95, LQ03, LMR20a, Men88, MH16, OR04, WH16c, ZJZ^{+}09, BH96, Ber99, LMR20b, PW99]. Their [AS21, BC20, Oga18, Rei18a, Rei18b, Sch18, Zhu18, But00]. Them [CRSW13]. Theme [DZ91]. Themes [PW00]. Theorem [Doo86, Le 86a, Le 86b, Pol86, Tro86, Gil14, Sei92]. Theoretic [Ber90a, Bro90, Coh90, FWB94, Geo90, Hwa90, MC90a, MC90b, MS90, RGVS23, Ruk90, Str90]. Theoretical [Mur18, SGKR06, Sti08]. Theories [MG20a, GG02, MG20b]. Theory [BFW07, Ber03a, BD19, Boo89, BBK^{+}19, CMN^{+}19, DKN18, DL18b, FK03, GGM09, GJ18, Hol03a, Ize21, Kar92, Ken89b, Ken89c, Ken89a, Lem94, Mar89, MZL10, NNLL18, RCR09a, RCR09b, Rub90, SFR10, Sma89, SNDS90, Sto89, Ver21, Wat89, ZZ12, Fel99, Leh99]. Therapeutic [CG92, DeM92, Ell92, FC92, Fle92a, Fle92b, Lag92, Lou92]. Therapies [AC89, Beg89, Ber89c, Cor89, Har89, KG89, LW89, Roy89, War89a, War89b]. There [Arn07, Gus23]. Thing [Cam91, Gio23, Har91b, Rob91a, Rob91b, Spa91, Spe91, SK91, Tho91]. Think [SH14]. Thinking [PW00, Sun89]. Third [FHJ08]. Thomas [Bel04]. Thompson [Ros23b]. Thought [Kad09]. Thoughts [Efr03]. Three [Bri03, Edw97a, Fie92, Kro92, LBK21]. Three-Way [Kro92, LBK21]. Threshold [BZ21, LW06]. Tiao [PT10]. Til [MFR24b]. Tilting [LRI23]. Time [BR02, Büh02, CT11, DJ24, Fai89, FBPR18, FK03, GŠP12, Han11b, Ion11, JCT22, Joe89, Nic16, Pic89, Pol03, Raf89, Raf92, Sin89, Smi89b, Smi89c, Wei89, XT11a, Yao11]. Time-Reversal [Nic16]. Times [Ans03, LW06, Peñ06]. Timing [BBK06]. Title [Sti13]. Tobacco [WM04]. today [Ald01]. Tom [Wal18]. Tomography [CCL^{+}04]. Tong [CT11, CY14a, Han11b, Ion11, Yao11]. Toolkit [Zas94]. Tools [Lan14, LL09b]. Topic [Ald17]. Topologically [NRB16, NRB17]. Tort [Joh88]. tortoise [PK97]. Tour [OS22]. Toxicities [ZJ24]. Toxicology [Sav88]. Trace [FSB16]. Trace-Contrast [FSB16]. Tracking [CKLP22, Pos23, Sin16a]. traditional [Joh97]. Traffic [BCK^{+}07, CCS04, WTLW95]. traits [DWZ97]. Transfer [Bha87, Bri87b, Dah87, Han87a, Han87b, Ris87, Shi87, Sol87]. Transformation [ARC21]. Transition [BR02]. Translated [BG07]. Transmission [KAX^{+}18]. Transmissions [Mar11]. Transportability [PB14]. Transportation [LRI23]. Treat [PA22]. Treated [SSP^{+}19]. Treatment [ÁEdBCAM17, Ban20, BCvdL20, BCN^{+}23, Chu17, DTL05, Lan15, Li20, Lin87c, QKM20, RHGF14, STLD14, KB98]. Treatments [BT88, GV88, HJM88a, HJM88b, LG17, Not88, Owe88, Spu88]. Tree [Pem95]. Tree-Indexed [Pem95]. Trees [Hol03a, KAX^{+}18, LD24, Ald01]. Trend [Fai89, HT00a, Joe89, Pic89, Raf89, Sin89, Smi89b, Smi89c, Wei89]. Trends [AMV19, FPDE14]. Trial [Ber12, Har89, JBM^{+}06, MT12, RLQ17]. Trials [BC91a, BL10, Ber04, BCN^{+}23, Bya91, Che10, CMC^{+}24, DMJ23, Dup91, Gio23, JT90, LZ00, Lev91, Lin91, MC92, PA22, RLLKV23b, RLLKV23a, RS08b, Roy91a, Roy91b, Sim91, Tha10, TR10, TX23, VBW15, Zel91, BH96, Lin98]. Triangle [Por94, SZ14]. Tribute [Ano91g, Bra07, Cas03a, Geo07, GS12]. Tropical [CP94]. True [Sti13]. Truncated [KJED09, SQH18]. Trust [FHJ08]. Truth [FS05, FBW^{+}16, Pos23]. Tukey [Ans88, Ans03, Cas03a, FM00, FM03, Fer03, Har03, Hoa03, Kaf03, Mal03, Mor03, Wai90, Wai03]. tutorial [HMRV99, HMRV00]. Twin [SJL^{+}12]. Two [Arj16, Ben08, Boo86a, Boo86b, Bre01, Cai08, Cam86, Cre86, DGA^{+}20, Efr08a, Efr08b, Efr14, Goo86d, Gus05, Hol06, Ken86, LK20, Lin87a, Lub91, McC16, Mor08, PS11, RS08a, Sam86, Wal14b, YB17, ZWS^{+}19, ZZS21]. Two-Group [Cai08]. Two-Groups [Ben08, Efr08a, Efr08b, Mor08, RS08a]. Two-Piece [Wal14b]. Two-Sample [ZWS^{+}19]. Type [BBFT17, SGZW10]. Tze [LSY21].

U.K. [Moo90c]. U.S. [Dav98b, DS92]. Ubiquitous [Cra16a, Cra16b, Teh16]. UK [Sil09]. Ulf [Muk06]. Uncertainty [Ber03c, CG93, CG04, DK87, Gew92, GH16, Hod87a, Kay91, Lin87c, Lin87a, Lin87b, Nee88, STG13, Sha87a, Sha87b, Spi87a, Spi87b, Spi87c, Wat87, Win90, Gle98, Fre87, Gei87c, Hod87b, Hub87b, Kad87b, Mad87, Smi87].

Uncluttering [AMV19]. Uncongenial [Fay94, Men94a, Men94b, Sch94a, Ski94, Zas94]. Undercount [Bre94a]. Understanding [AG01, AOW17, FMF13, Tan07a]. undervotes [AP02]. Unified [BBW97, CCFR24, DTZ17, HJJ^{+}21, NRWY12, NNLL18, SQH18]. Uniform [JJN13]. Uniformly [MHdL23, SZ14]. Unifying [CHP22, Hil89, Jor89, McC89, dP89b, dP89a]. Union [AET09a, BH96]. Unit [CP94, Men05]. United [CEN^{+}19, EGG97, RAG14]. Units [SSP^{+}19, ZLS10]. Unity [Aka90, Ald90, Box90b, Dem90b, Efr90, Hac90, Moo90b, Sha90b, Sha90a]. Universal [IKN09a, ZS09]. Universe [VGB14]. universities [Dav98b]. University [Hot88b]. Unknown [GHV12]. Unlabeled [LMW07]. Unlikely [Sin16b]. Unobservables [LN09a, LN09b, Lou09, MKV09]. Unobserved [SRC^{+}16, MY97]. Unreasonable [Owe19]. Unsettling [GM21b, LM21]. Unusual. [SRS91]. Updating [GM21b, Sha21]. USA [PT11]. Use [Hat88, HR16, KT88, KG05, Lew94, LMW07, Mar87a, Sil91]. Used [Kas10]. User [KMR^{+}19]. User-Friendly [KMR^{+}19]. Using [Ber91a, Ber91b, BBK06, Car19, Che91, CPW15, DPL15, Eve91, FPBT22, GR92a, GKLP99, GS06, GL06, GK02, HR16, JD21, JAAH14, Kay91, Lan91, LL07, LL04, LD24, Mar88a, MH16, Mor14, NWEE20, RD06, RW06, STG13, TN13, Yu95, Zas94, ZWS^{+}19, ZT09]. Uspensky [DZ23]. Utility [GV23, GKRS05, Gus23, JL13, MP93]. Utilizing [EH03].

V [DZ23, KK16, Ran07]. V. [Tho02]. Vacuum [Her19]. Valid [MHdL23, NRB16, NRB17, RGVS23]. Validated [LL07]. Validating [BGGM06]. Validation [BC20, OW16]. Validity [PB14, SGKR06]. Valley [SLR15]. Value [Fai89, Joe89, Pic89, Raf89, Sin89, Smi89b, Smi89c, Wei89, HT00a]. Values [Ash88, BC05, BCD05, FPBT22, Gel05, GM05a, GM05b, SSVV11, Tho05, VW23]. Vantage [Vov09]. Varadhan [Zei18]. variability [Dav98a]. Variable [LD24, MGC15, RSLGR07, RR23, SVS11, STR^{+}17, SH14, ZWS^{+}19]. Variables [AWJ^{+}18, BHJK21, CPW15, FHS17, GMS20, Gus05, Imb14b, JW94, JD21, Kit14, LBK21, VBBG11]. Variance [Ber90a, BFWW09, Bro90, Coh90, DTZ17, Geo90, GHV12, Hwa90, JL13, MC90a, MC90b, MS90, Pol24, Ruk90, Str90]. Variances [BB04b, GGM09]. Variants [DLS14, LBK21, ZT09]. Variates [HLO05]. Variation [MRSS15]. Variational [DR19, TN13, WMB19, YNS23]. Variations [DZ91]. Various [LPE15, RR01]. Vasilyevich [She92]. Vecchia [KG21]. Vector [HP04a, MM06b, PQW14, PCT^{+}98]. VEDA [You90]. Venezuela [DS11, HR11, PT11]. Venezuelan [Car11, Mar11, PS11, Jim11]. Verbal [KL08]. Versus [DS11, GvdL19, Pol24, CTB20, DHS^{+}19a, PK97]. Very [RAP14, ZWS21]. Via [KM18a, Ald93, CH89, Dud89, FC11, GT03, GZ89, HPV06, HJ01, LP11a, MT11, PT03, Pol89a, Pol89b, Pyk89, VACC22]. View [LN04, LN09a, LN09b, Lou09, Mar11, MKV09, Wai87, AG01]. viewpoint [How98]. Views [Box90a, Bra88, Dem90a, Leh90a, Moo90c, Net90, Pra90, RLQ17, Rob90a, Rob90b]. Violations [Li20]. Visions [Wai90]. Visual [Tuk90]. Vladimirovich [SS92]. Vol [BN90, Ano02, Ano03b]. Volume [Ano02, Ano03b, Ano05, Ano06, Ano07, Ano08, Ano09, Ano10, Ano11, Ano13]. vote [Smi02]. Voting [CG02, Mar11, GG02, GKT02]. vs [SGKR06].

W [Ans88, Ans03, Ber12, DeG86e, FM00, Fer03, Hoa03, MT12, MM24, Mor03]. W. [Olk91]. Wahba [NMB20]. Waksberg [MM00]. Wald [Bar90b, Ber90b, Ber90c, Cox90b, Fre90, Kad90b, Leh90b, Lin90a, Lin90b, Mou90]. Waldegrave [BF15]. Walk [Rit11, SFR10]. Walker [Kat02]. Wallis [Olk91]. Walter [Hed05]. Want [Con93, Gas93, Gra93a, Gra93b, Rob93a]. Warehousing [Fie06b]. Was [SRS91]. Wassily [FvZ08]. Watson [BF98]. Wave [WWZ16]. Way [EHX04, Kro92, LBK21, Pol24]. Weak [DJ15, EJ11, MZL10, Per05]. Wear [QOS19]. Wearing [Arj16]. Web [SC17]. Weighted [CEH^{+}17, RW09]. Weighting [BC07c, BO07, Gel07b, Gel07c, Lit07, Loh07, Pfe07, RT12]. Weights [HB17, RSLGR07]. Well [Bel87, FPBT22]. Well-Calibrated [FPBT22]. Well-Conditioned [Bel87]. Wellner [BS18]. Wells [Lin87a]. Welsh [Sto14]. Where [Hod17, KEHSY19]. Whetstone [CFL13]. Which [Hol88]. While [Cox91, DP91a, DP91b, KM91, Nor91]. Who [Fer89a, Sam89, Fer89b, Fre89, Rob89, Sak89, Sam89]. Wide [GWVV09, KZI09, LL09a, PGP09, RW09, SCtW^{+}09, TCC^{+}09, VMC20, ZMG09, ZJZ^{+}09]. Wilder [Wai03]. Wildfire [TWDM13]. Wilfrid [Flo93]. Wilk [GB10]. Will [Nic16, ZKS19, JSH07, Sam89]. William [Mar90, CEMR90, Eat07, FST07, Fie07, Han97, Sti07b, Tan07b, Wai90, Zab94]. Wind [Pin13]. Winning [ZKS19]. Wired [Mag93]. Wise [JJN13]. Within [JD21]. Without [FSB16, GKRS05, Shp14, vHB21, GR92b]. Women [Nol92]. Wonderful [Han14a]. Woodroofe [BS16]. Word [DN06]. Word-of-Mouth [DN06]. Work [BK87, Bas87, DE86, Ger87, Goo86a, Goo86b, Hey86, Joh19, RH87a, RH87b, Sin86, Sin87, Sol86, Wei87, Nor99]. Working [Goo07]. Workshop [BBC^{+}02]. World [Arr88, Boo03, Bra88, Daw04, Dem88a, Fie06b, GKRS05, Gup88, Hog88, Hot88c, Hot88b, Hot88a, LL09b, Moo88, Olk88, Zid88]. Wreckage [SKKS14]. Writing [Ano94c]. Wrong [MN11]. Wu [Ber99, CJ16].

Xia [CT11, Han11b, Ion11, Yao11]. Xu [CO12].

Year [BM16, SB16]. Years [DH24, Fie08, GGBS^{+}23, Leh93, SY12, Ste05, BG99]. Yourself [DHS^{+}19a, GvdL19]. Yuan [YZ06]. Yule [Ald01, Ald95]. Yuri [She92].

Z. [Mar90]. Zacks [Muk05]. Zero [LSS^{+}19]. Zero-Inflated [LSS^{+}19]. Zhan [Lit19]. Zhang [Lit19]. Zwet [BF09].

References

Astle:2009:PSC

[AB09] William Astle and David J. Balding. Population structure and cryptic relatedness in genetic association studies. Statistical Science, 24(4):451–471, November 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1271770342.

Atkinson:2015:CPT

[AB15] Anthony C. Atkinson and Barbara Bogacka. A conversation with Professor Tadeusz Caliński. Statistical Science, 30(3): 423–442, August 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1439220721.

Anastasiou:2023:SMM

[ABB^{+}23] Andreas Anastasiou, Alessandro Barp, François-Xavier Briol, Bruno Ebner, Robert E. Gaunt, Fatemeh Ghaderinezhad, Jackson Gorham, Arthur Gretton, Christophe Ley, Qiang Liu, Lester Mackey, Chris J. Oates, Gesine Reinert, and Yvik Swan. Stein’s method meets computational statistics: a review of some recent developments. Statistical Science, 38(1):120–139, February 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/journals/
statistical-science/volume-38/issue-1/Steins-Method-
Meets-Computational-Statistics--A-Review-of-Some/10.
1214/22-STS863.full.

Altman:1991:MNN

[ABC^{+}91] N. Altman, D. Banks, P. Chen, D. Duffy, J. Hardwick, C. Leger, A. Owen, and T. Stukel. Meeting the needs of new statistical researchers. Statistical Science, 6(2):163–174, May 1991. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011821.

Adams:2000:DMF

[ABHK00] Niall M. Adams, Gordon Blunt, David J. Hand, and Mark G. Kelly. Data mining for fun and profit. Statistical Science, 15 (2):111–131, May 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009212753.

Arratia:2016:EFC

[ABT16] Richard Arratia, A. D. Barbour, and Simon Tavaré. Exploiting the Feller coupling for the Ewens sampling formula. Statistical Science, 31(1):27–29, February 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1455115909.

Armitage:1989:ITP

[AC89] Peter Armitage and D. Stephen Coad. [Investigating therapies of potentially great benefit: ECMO]: Comment. Statistical Science, 4(4):322–323, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012389.

Arnold:2001:CSD

[ACS01] Barry C. Arnold, Enrique Castillo, and José María Sarabia. Conditionally specified distributions: An introduction (with comments and a rejoinder by the authors). Statistical Science, 16(3):249–274, August 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1009213728.

Andrieu:2004:CAB

[ADR04] C. Andrieu, A. Doucet, and C. P. Robert. Computational advances for and from Bayesian analysis. Statistical Science, 19 (1):118–127, February 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1089808277.

Alvarez-Esteban:2017:MAT

[ÁEdBCAM17] P. C. Álvarez-Esteban, E. del Barrio, J. A. Cuesta-Albertos, and C. Matrán. Models for the assessment of treatment improvement: The ideal and the feasible. Statistical Science, 32(3):469–485, August 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1504253127.

Adler:2009:CS

[AET09a] Robert Adler, John Ewing, and Peter Taylor. Citation statistics: A report from the International Mathematical Union (IMU) in cooperation with the International Council of Industrial and Applied Mathematics (ICIAM) and the Institute of Mathematical Statistics (IMS). Statistical Science, 24(1):1–14, February 2009. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/

euclid.ss/1255009002. See comments [Hal09, SG09, LLJ09].

Adler:2009:RCS

[AET09b] Robert Adler, John Ewing, and Peter Taylor. Rejoinder: Citation statistics. Statistical Science, 24(1):27–28, February 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1255009007.

Aalen:2001:USH

[AG01] Odd O. Aalen and Håkon K. Gjessing. Understanding the shape of the hazard rate: a process point of view (with comments and a rejoinder by the authors). Statistical Science, 16 (1):1–22, February 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/998929473.

Agresti:2005:CRC

[AG05] Alan Agresti and Anna Gottard. Comment: Randomized confidence intervals and the mid-P approach. Statistical Science, 20(4):367–371, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1137076653. See [GM05a].

Arratia:1990:PACa

[AGG90a] Richard Arratia, Larry Goldstein, and Louis Gordon. Poisson approximation and the Chen-Stein method. Statistical Science, 5(4):403–424, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012015.

Arratia:1990:PACb

[AGG90b] Richard Arratia, Larry Goldstein, and Louis Gordon. [Poisson approximation and the Chen-Stein method]: Rejoinder. Statistical Science, 5(4):432–434, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012020.

Agresti:1992:SEIb

[Agr92a] Alan Agresti. [A survey of exact inference for contingency tables]: Rejoinder. Statistical Science, 7(1):173–177, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011462.

Agresti:1992:SEIa

[Agr92b] Alan Agresti. A survey of exact inference for contingency tables. Statistical Science, 7(1):131–153, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011454.

Abbott:1999:PP

[AH99] D. Abbott and G. P. Harmer. Parrondo’s paradox. Statistical Science, 14(2):206–213, May 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1009212247.

Akaike:1987:PFO

[Aka87] Hirotugu Akaike. [Prediction of future observations in growth curve models]: Comment. Statistical Science, 2(4):464–465, November 1987. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177013124.

Akaike:1990:UDP

[Aka90] Hirotugu Akaike. [The unity and diversity of probability]: Comment. Statistical Science, 5(4):444–446, November 1990. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012022.

Akritas:2004:NSA

[Akr04] Michael G. Akritas. Nonparametric survival analysis. Statistical Science, 19(4):615–623, November 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1113832726.

Aldous:1990:UDP

[Ald90] David Aldous. [The unity and diversity of probability]: Comment. Statistical Science, 5(4):446–447, November 1990. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012023.

Aldous:1993:ACM

[Ald93] David Aldous. Approximate counting via Markov chains. Statistical Science, 8(1):16–19, February 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011078.

Aldrich:1995:CGS

[Ald95] John Aldrich. Correlations genuine and spurious in Pearson and Yule. Statistical Science, 10(4):364–376, November 1995. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177009870.

Aldrich:1997:RFM

[Ald97] John Aldrich. R. A. Fisher and the making of maximum likelihood 1912–1922. Statistical Science, 12(3):162–176, August 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1030037906.

Aldous:2001:SMD

[Ald01] David J. Aldous. Stochastic models and descriptive statistics for phylogenetic trees, from Yule to today. Statistical Science, 16 (1):23–34, February 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/998929474.

Aldrich:2005:FR

[Ald05] John Aldrich. Fisher and regression. Statistical Science, 20(4): 401–417, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1137076660.

Aldous:2013:ACP

[Ald13] David Aldous. Another conversation with Persi Diaconis. Statistical Science, 28(2):269–281, May 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1369147916.

Aldous:2017:ERS

[Ald17] David Aldous. Elo ratings and the sports model: A neglected topic in applied probability? Statistical Science, 32(4):616–629, 11 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Aldous:2018:CJP

[Ald18] David Aldous. A conversation with Jim Pitman. Statistical Science, 33(3):458–467, August 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1534147233.

Alexander:1996:CTH

[Ale96] Kenneth S. Alexander. A conversation with Ted Harris. Statistical Science, 11(2):150–158, May 1996. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1038425658.

Abt:2001:SAI

[ALS^{+}01] Markus Abt, YongBin Lim, Jerome Sacks, Minge Xie, and S. Stanley Young. Sequential approach for identifying lead compounds in large chemical databases. Statistical Science, 16 (2):154–168, May 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009213288.

Atkinson:2019:CPD

[AM19] Peter M. Atkinson and Jorge Mateu. A conversation with Peter Diggle. Statistical Science, 34(3):504–521, August 2019. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/euclid.ss/1570780982.

Amari:1989:GAI

[Ama89] S. I. Amari. [The geometry of asymptotic inference]: Comment. Statistical Science, 4(3):220–222, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012481.

Amari:1994:NNR

[Ama94] S. Amari. [Neural networks: A review from statistical perspective]: Comment. Statistical Science, 9(1):31–32, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010639.

Anderlucci:2019:IBC

[AMV19] Laura Anderlucci, Angela Montanari, and Cinzia Viroli. The importance of being clustered: Uncluttering the trends of statistics from 1970 to 2015. Statistical Science, 34(2):280–300, May 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1563501642.

Anderson:1987:RMA

[And87] T. W. Anderson. [A review of multivariate analysis]: Comment. Statistical Science, 2(4):413–417, November 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013112.

Anderson:1988:SOI

[And88] Paul D. Anderson. Scientific origins of incompatibility in risk assessment. Statistical Science, 3(3):320–327, August 1988. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012833.

Andrews:1991:ASP

[And91] D. F. Andrews. [Avoiding statistical pitfalls]: Comment. Statistical Science, 6(3):253, August 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011687.

Anderson:1996:RFM

[And96] T. W. Anderson. R. A. Fisher and multivariate analysis. Statistical Science, 11(1):20–34, February 1996. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1032209662.

Andersson:2022:ACI

[And22] Per Gösta Andersson. Approximate confidence intervals for a binomial p — once again. Statistical Science, 37(4):598–606, November 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-4/
Approximate-Confidence-Intervals-for-a-Binomial-pOnce-
Again/10.1214/21-STS837.full.

Anonymous:1986:E

[Ano86a] Anonymous. Editorial. Statistical Science, 1(1):1–2, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013803.

Anonymous:1986:Ia

[Ano86b] Anonymous. In this issue. Statistical Science, 1(2):155–156, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013689.

Anonymous:1986:Ib

[Ano86c] Anonymous. In this issue. Statistical Science, 1(3):295–296, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013603.

Anonymous:1986:Ic

[Ano86d] Anonymous. In this issue. Statistical Science, 1(4):417–418, November 1986. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177013507.

Anonymous:1986:PR

[Ano86e] Anonymous. Publications received. Statistical Science, 1(1): 153–154, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013833.

Anonymous:1987:ARS

[Ano87a] Anonymous. Acknowledgment of referees’ services. Statistical Science, 2(1):104, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177013446.

Anonymous:1987:Ia

[Ano87b] Anonymous. In this issue. Statistical Science, 2(1):1–2, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013425.

Anonymous:1987:Ib

[Ano87c] Anonymous. In this issue. Statistical Science, 2(2):105–106, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013332.

Anonymous:1987:Ic

[Ano87d] Anonymous. In this issue. Statistical Science, 2(3):211–212, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013214.

Anonymous:1987:Id

[Ano87e] Anonymous. In this issue. Statistical Science, 2(4):353–354, November 1987. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177013103.

Anonymous:1987:PR

[Ano87f] Anonymous. Publications received. Statistical Science, 2(1): 101–103, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013445.

Anonymous:1988:Ia

[Ano88a] Anonymous. In this issue. Statistical Science, 3(1):1–2, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012992.

Anonymous:1988:Ib

[Ano88b] Anonymous. In this issue. Statistical Science, 3(2):147–148, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012893.

Anonymous:1988:Ic

[Ano88c] Anonymous. In this issue. Statistical Science, 3(3):273, August 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012825.

Anonymous:1988:Id

[Ano88d] Anonymous. In this issue. Statistical Science, 3(4):385, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012753.

Anonymous:1988:PR

[Ano88e] Anonymous. Publications received. Statistical Science, 3(1): 145–146, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013021.

Anonymous:1989:ARS

[Ano89a] Anonymous. Acknowledgment of referees’ services. Statistical Science, 4(2):184, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012604.

Anonymous:1989:DAC

[Ano89b] Anonymous. Discriminant analysis and clustering: Panel on discriminant analysis, classification, and clustering. Statistical Science, 4(1):34–69, February 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012666.

Anonymous:1989:Ia

[Ano89c] Anonymous. In this issue. Statistical Science, 4(1):1, February 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012664.

Anonymous:1989:Ib

[Ano89d] Anonymous. In this issue. Statistical Science, 4(4):297, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012383.

Anonymous:1989:PR

[Ano89e] Anonymous. Publications received. Statistical Science, 4(1): 70–71, February 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012667.

Anonymous:1989:SMA

[Ano89f] Anonymous. Statistical models and analysis in auditing: Panel on nonstandard mixtures of distributions. Statistical Science, 4(1):2–33, February 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012665.

Anonymous:1990:I

[Ano90a] Anonymous. In this issue. Statistical Science, 5(1):1, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012241.

Anonymous:1990:PR

[Ano90b] Anonymous. Publications received. Statistical Science, 5(1): 156–158, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012275.

Anonymous:1991:BMH

[Ano91a] Anonymous. Biography of Morris H. DeGroot. Statistical Science, 6(1):3–14, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1177011925. See correction

[Ano91b].

Anonymous:1991:CNB

[Ano91b] Anonymous. Correction note: Biography of Morris H. DeGroot. Statistical Science, 6(4):448, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011594.  See

correction [Ano91a].

Anonymous:1991:EI

[Ano91c] Anonymous. Editor’s introduction. Statistical Science, 6(4): 339, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011571.

Anonymous:1991:I

[Ano91d] Anonymous. In this issue. Statistical Science, 6(4):319, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011568.

Anonymous:1991:PR

[Ano91e] Anonymous. Publications received. Statistical Science, 6(1): 117–118, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011946.

Anonymous:1991:R

[Ano91f] Anonymous. Rejoinder. Statistical Science, 6(4):360–362, November 1991. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177011576.

Anonymous:1991:RGM

[Ano91g] Anonymous. Rupert G. Miller, Jr., 1933–1986: A tribute. Statistical Science, 6(1):98–99, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011944.

Anonymous:1992:ARS

[Ano92a] Anonymous. Acknowledgment of referees’ services. Statistical Science, 7(2):284–285, May 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011368.

Anonymous:1992:Ia

[Ano92b] Anonymous. In this issue. Statistical Science, 7(1):3–4, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011439.

Anonymous:1992:Ib

[Ano92c] Anonymous. In this issue. Statistical Science, 7(2):181, May 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011352.

Anonymous:1992:PR

[Ano92d] Anonymous. Publications received. Statistical Science, 7(1): 178–179, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011463.

Anonymous:1993:ARS

[Ano93a] Anonymous. Acknowledgment of referees’ services. Statistical Science, 8(1):78, February 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011088.

Anonymous:1993:PR

[Ano93b] Anonymous. Publications received. Statistical Science, 8(1): 79–80, February 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011089.

Anonymous:1994:ARS

[Ano94a] Anonymous. Acknowledgment of referees’ services. Statistical Science, 9(1):146, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177010658.

Anonymous:1994:ACD

[Ano94b] Anonymous. Author’s clarification: [DNA fingerprinting: A review of the controversy]. Statistical Science, 9(3):456, August 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010395.

Anonymous:1994:GWS

[Ano94c] Anonymous. Guidelines on writing for statistical science. Statistical Science, 9(4):591, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177010276.

Anonymous:1994:Ia

[Ano94d] Anonymous. In this issue. Statistical Science, 9(2):149, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010479.

Anonymous:1994:Ib

[Ano94e] Anonymous. In this issue. Statistical Science, 9(3):305–306, August 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010377.

Anonymous:1994:Ic

[Ano94f] Anonymous. In this issue. Statistical Science, 9(4):457, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010258.

Anonymous:1994:PR

[Ano94g] Anonymous. Publications received. Statistical Science, 9(1): 147–148, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010659.

Anonymous:1995:ARS

[Ano95a] Anonymous. Acknowledgment of referees’ services. Statistical Science, 10(1):134, February 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177010135.

Anonymous:1995:Ia

[Ano95b] Anonymous. In this issue. Statistical Science, 10(1):1–2, February 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010122.

Anonymous:1995:Ib

[Ano95c] Anonymous. In this issue. Statistical Science, 10(2):137, May 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010026.

Anonymous:1995:PR

[Ano95d] Anonymous. Publications received. Statistical Science, 10(1): 135–136, February 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010136.

Anonymous:1996:CSA

[Ano96] Anonymous. Comments on: “Methods for approximating integrals in statistics with special emphasis on Bayesian integration problems” by M. Evans and T. Swartz. Statistical Science, 11 (1):54–64, February 1996. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1032209664. See [ES95].

Anonymous:2000:C

[Ano00] Anonymous. Correction. Statistical Science, 15(2):192, May 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1009212758.

Anonymous:2002:VIS

[Ano02] Anonymous. Volume index, Statist. Sci. vol. 17 (2002). Statistical Science, 17(4):468–469, November 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1049993206.

Anonymous:2003:ISS

[Ano03a] Anonymous. Introduction to the special section on statistics and the environment. Statistical Science, 18(4):411, November 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1081443224.

Anonymous:2003:VIS

[Ano03b] Anonymous. Volume index, Statist. Sci. vol. 18 (2003). Statistical Science, 18(4):532–533, November 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1081443234.

Anonymous:2005:IV

[Ano05] Anonymous. Index to volume 20. Statistical Science, 20(4): 431–432, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1137076662.

Anonymous:2006:CV

[Ano06] Anonymous. Contents of volume 21 (2006). Statistical Science, 21(4):586–588, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177334531.

Anonymous:2007:CV

[Ano07] Anonymous. Contents of volume 22 (2007). Statistical Science, 22(4):651–652, November 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1207580177.

Anonymous:2008:CV

[Ano08] Anonymous. Contents of volume 23 (2008). Statistical Science, 23(4):565–566, November 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1242049395.

Anonymous:2009:CV

[Ano09] Anonymous. Contents of volume 24 (2009). Statistical Science, 24(4):574–576, November 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1271770350.

Anonymous:2010:IV

[Ano10] Anonymous. Index of volume 25. Statistical Science, 25(4): 588–589, November 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1300108239.

Anonymous:2011:IV

[Ano11] Anonymous. Index of volume 26 (2011). Statistical Science, 26(4):673–675, November 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1330437944.

Anonymous:2013:CV

[Ano13] Anonymous. Contents of volume 28. Statistical Science, 28(4): 659–660, November 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1386078883.

Anonymous:2016:EB

[Ano16a] Anonymous. Editorial board. Statistical Science, 31(4):??, November 2016. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1484816570.

Anonymous:2016:TC

[Ano16b] Anonymous. Table of contents. Statistical Science, 31(4):??, November 2016. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1484816571.

Anonymous:2020:ISI

[Ano20] Anonymous. Introduction to the special issue. Statistical Science, 35(1):1, February 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1583226021.

Anonymous:2023:EBa

[Ano23a] Anonymous. Editorial board. Statistical Science, 38(2): ??, May 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-2/
Editorial-Board/sts382c2.full.

Anonymous:2023:EBb

[Ano23b] Anonymous. Editorial board. Statistical Science, 38(3):??, August 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-3/
Editorial-Board/sts383c2.full.

Anonymous:2023:EBc

[Ano23c] Anonymous. Editorial board. Statistical Science, 38(4):??, November 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-4/
Editorial-Board/sts384c2.full.

Anonymous:2023:TCa

[Ano23d] Anonymous. Table of contents. Statistical Science, 38(2): ??, May 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-2/
Table-of-Contents/sts382p1.full.

Anonymous:2023:TCb

[Ano23e] Anonymous. Table of contents. Statistical Science, 38(3):??, August 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-3/
Table-of-Contents/sts383p1.full.

Anonymous:2023:TC

[Ano23f] Anonymous. Table of contents. Statistical Science, 38(4):??, November 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-4/
Table-of-Contents/sts384p1.full.

Anonymous:2024:EBa

[Ano24a] Anonymous. Editorial board. Statistical Science, 39(1):??, February 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-1/
Editorial-Board/sts391c2.full.

Anonymous:2024:EBb

[Ano24b] Anonymous. Editorial board. Statistical Science, 39(2): ??, May 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-2/
Editorial-Board/sts392c2.full.

Anonymous:2024:EBc

[Ano24c] Anonymous. Editorial board. Statistical Science, 39(3):??, August 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-3/
Editorial-Board/sts393c2.full.

Anonymous:2024:TCa

[Ano24d] Anonymous. Table of contents. Statistical Science, 39(1):??, February 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-1/
Table-of-Contents/sts391p1.full.

Anonymous:2024:TCb

[Ano24e] Anonymous. Table of contents. Statistical Science, 39(2): ??, May 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-2/
Table-of-Contents/sts392p1.full.

Anonymous:2024:TCc

[Ano24f] Anonymous. Table of contents. Statistical Science, 39(3):??, August 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-3/
Table-of-Contents/sts393p1.full.

Anscombe:1988:FMJ

[Ans88] Francis J. Anscombe. Frederick Mosteller and John W. Tukey: A conversation. Statistical Science, 3(1):136–144, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013020;   http://www.jstor.org/stable/2245933.

Anscombe:2003:QCC

[Ans03] F. R. Anscombe. Quiet contributor: The civic career and times of John W. Tukey. Statistical Science, 18(3):287–310, August 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1076102417.

Aronow:2017:UDA

[AOW17] Peter M. Aronow and Molly R. Offer-Westort. Understanding Ding’s apparent paradox. Statistical Science, 32(3):346– 348, August 2017. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1504253117.

Agresti:2002:MUO

[AP02] Alan Agresti and Brett Presnell. Misvotes, undervotes and overvotes: The 2000 presidential election in Florida. Statistical Science, 17(4):436–440, November 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1049993202.

Agapiou:2017:ISI

[APSAS17] S. Agapiou, O. Papaspiliopoulos, D. Sanz-Alonso, and A. M. Stuart. Importance sampling: Intrinsic dimension and computational cost. Statistical Science, 32(3):405–431, August 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1504253124.

Atkinson:2021:BCT

[ARC21] Anthony C. Atkinson, Marco Riani, and Aldo Corbellini. The Box–Cox transformation: Review and extensions. Statistical Science, 36(2):239–255, May 2021. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-2/The-BoxCox-Transformation-
Review-and-Extensions/10.1214/20-STS778.full.

Arjas:2016:HAW

[Arj16] Elja Arjas. How about wearing two hats, first Popper’s and then de Finetti’s? Statistical Science, 31(4):545–548, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1484816582.

Armitage:1990:BBC

[Arm90] Peter Armitage. [Biostatistics and Bayes]: Comment. Statistical Science, 5(3):284–286, August 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012093.

Arnold:2007:MHT

[Arn07] Barry C. Arnold. Majorization: Here, there and everywhere. Statistical Science, 22(3):407–413, August 2007. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1199285040.

Arrow:1988:GOC

[Arr88] Kenneth J. Arrow. [Golden oldies: Classic articles from the world of statistics and probability]: Comment. Statistical Science, 3(1):90–91, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013005. See [Hot88c].

Aldous:1993:IIP

[AS93] David Aldous and J. Michael Steele. Introduction to the interface of probability and algorithms. Statistical Science, 8 (1):3–9, February 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011076.

Aldous:2010:CSN

[AS10] David J. Aldous and Julian Shun. Connected spatial networks over random points and a route-length statistic. Statistical Science, 25(3):275–288, August 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1294167960.

Augustin:2021:CFS

[AS21] Thomas Augustin and Georg Schollmeyer. Comment: On focusing, soft and strong revision of Choquet capacities and their role in statistics. Statistical Science, 36(2):205–209, May 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-36/issue-2/
Comment--On-Focusing-Soft-and-Strong-Revision-of-Choquet/
10.1214/21-STS765D.full.

Ashford:1988:SVR

[Ash88] Nicholas A. Ashford. Science and values in the regulatory process. Statistical Science, 3(3):377–383, August 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012842.

Antonelli:2016:MBG

[ATH16] Joseph Antonelli, Lorenzo Trippa, and Sebastien Haneuse. Mitigating bias in generalized linear mixed models: The case for Bayesian nonparametrics. Statistical Science, 31(1):80–95, February 2016. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1455115915.

Atkinson:1986:IOH

[Atk86] A. C. Atkinson. [Influential observations, high leverage points, and outliers in linear regression]: Comment: Aspects of diagnostic regression analysis. Statistical Science, 1(3):397–402, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013624.

Atkinson:2014:SBC

[Atk14] Anthony C. Atkinson. Selecting a biased-coin design. Statistical Science, 29(1):144–163, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1399645742.

Athreya:2021:EIL

[ATPP21] Avanti Athreya, Minh Tang, Youngser Park, and Carey E. Priebe. On estimation and inference in latent structure random graphs. Statistical Science, 36(1):68–88, February 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/journals/
statistical-science/volume-36/issue-1/On-Estimation-
and-Inference-in-Latent-Structure-Random-Graphs/10.
1214/20-STS787.full.

Audigier:2018:MIM

[AWJ^{+}18] Vincent Audigier, Ian R. White, Shahab Jolani, Thomas P. A. Debray, Matteo Quartagno, James Carpenter, Stef van Buuren, and Matthieu Resche-Rigon. Multiple imputation for multilevel data with continuous and binary variables. Statistical Science, 33(2):160–183, May 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1525313140.

Agarwal:2020:DDP

[AWZ20] Divyansh Agarwal, Jingshu Wang, and Nancy R. Zhang. Data denoising and post-denoising corrections in single cell RNA sequencing. Statistical Science, 35(1):112–128, February 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1583226032.

Amemiya:2001:NFA

[AY01] Yasuo Amemiya and Ilker Yalcin. Nonlinear factor analysis as a statistical method. Statistical Science, 16(3):275–294, August 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1009213729.

Bailar:1988:F

[Bai88] John C. Bailar. Foreword. Statistical Science, 3(3):274, August 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012826.

Bailey:1991:ASP

[Bai91] R. A. Bailey. [Avoiding statistical pitfalls]: Comment. Statistical Science, 6(3):254–255, August 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011688.

Bailey:2017:IRF

[Bai17] R. A. Bailey. Inference from randomized (factorial) experiments. Statistical Science, 32(3):352–355, August 2017. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/euclid.ss/1504253119.

Banks:1988:IBS

[Ban88] David L. Banks. [The interface between statistics and philosophy of science]: Comment. Statistical Science, 3(4):404– 406, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012758.

Banks:1993:ISCa

[Ban93a] David Banks. Is industrial statistics out of control? Statistical Science, 8(4):356–377, November 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177010773.

Banks:1993:ISCb

[Ban93b] David Banks. [Is industrial statistics out of control?]: Rejoinder. Statistical Science, 8(4):402–409, November 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010784.

Banks:1996:CJG

[Ban96] David L. Banks. A conversation with I. J. Good. Statistical Science, 11(1):1–19, February 1996. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1032209661.

Bansak:2020:GAP

[Ban20] Kirk Bansak. A generalized approach to power analysis for local average treatment effects. Statistical Science, 35(2):254– 271, May 2020. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1591171230.

Baram:1988:IHM

[Bar88a] Michael S. Baram. Insurability of hazardous materials activities. Statistical Science, 3(3):339–345, August 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012835.

Barnard:1988:IBS

[Bar88b] George A. Barnard. [The interface between statistics and philosophy of science]: Comment. Statistical Science, 3(4):401– 403, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012756.

Barnard:1989:RFH

[Bar89] G. A. Barnard. [R. A. Fisher on the history of inverse probability]: Comment. Statistical Science, 4(3):258–260, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012490.

Barbour:1990:PAC

[Bar90a] A. D. Barbour. [Poisson approximation and the Chen-Stein method]: Comment. Statistical Science, 5(4):425–427, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012017.

Barnard:1990:WML

[Bar90b] George A. Barnard. [The 1988 Wald memorial lectures: The present position in Bayesian statistics]: Comment. Statistical Science, 5(1):65–71, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012254.

Barnard:1992:RSI

[Bar92a] George A. Barnard. Review of Statistical Inference and Analysis: Selected Correspondence of R. A. Fisher. Statistical Science, 7(1):5–12, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011440.

Bartholomew:1992:ISM

[Bar92b] David J. Bartholomew. [The impact of sociological methodology on statistical methodology]: Comment. Statistical Science, 7(2):196–198, May 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011355.

Barron:1994:NNR

[Bar94] Andrew R. Barron. [Neural networks: A review from statistical perspective]: Comment. Statistical Science, 9(1):33–35, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010640.

Barry:2005:CJH

[Bar05] Daniel Barry. A conversation with John Hartigan. Statistical Science, 20(4):418–430, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1137076661.

Barker:2016:CGF

[Bar16] Richard Barker. A conversation with G. A. F. Seber. Statistical Science, 31(2):151–160, May 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1464105034.

Basu:1987:SSW

[Bas87] Asit P. Basu. [Survey of Soviet work in reliability]: Comment. Statistical Science, 2(4):502–503, November 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013133.

Bates:1986:CSR

[Bat86] Douglas M. Bates. [Computers in statistical research]: Comment. Statistical Science, 1(4):448–449, November 1986. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013514.

See [Edd86a].

Bather:1996:CHC

[Bat96] John Bather. A conversation with Herman Chernoff. Statistical Science, 11(4):335–350, November 1996. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1032280306.

Bayarri:1987:TPH

[Bay87] M. J. Bayarri. [Testing precise hypotheses]: Comment. Statistical Science, 2(3):342–344, August 1987. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013242.

Bayarri:1988:SMF

[Bay88] M. J. Bayarri. [Selection models and the file drawer problem]: Comment. Statistical Science, 3(1):128–131, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013016.

Bayarri:1991:RMA

[BB91] M. J. Bayarri and James Berger. [Replication and meta-analysis in parapsychology]: Comment. Statistical Science, 6(4):379– 382, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011578.

Bayarri:2004:IBF

[BB04a] M. J. Bayarri and J. O. Berger. The interplay of Bayesian and frequentist analysis. Statistical Science, 19(1):58–80, February 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1089808273.

Boos:2004:CVO

[BB04b] Dennis D. Boos and Cavell Brownie. Comparing variances and other measures of dispersion. Statistical Science, 19(4): 571–578, November 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1113832721.

Bru:2018:DG

[BB18] Marie-France Bru and Bernard Bru. Dice games. Statistical Science, 33(2):285–297, May 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1525313147.

Buja:2019:MAC

[BBB^{+}19] Andreas Buja, Lawrence Brown, Richard Berk, Edward George, Emil Pitkin, Mikhail Traskin, Kai Zhang, and Linda Zhao. Models as approximations I: Consequences illustrated with linear regression. Statistical Science, 34(4):523–544, November 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/

1578474016. See Part II [BBK^{+}19], discussion [vdG19, Law19, FJL^{+}19, Tjø19, Lit19, DKK19, WSC19, GK19], and rejoinder [BKB^{+}19].

Berk:2002:WSA

[BBC^{+}02] Richard A. Berk, Peter Bickel, Katherine Campbell, Robert Fovell, Sallie Keller-McNulty, Elizabeth Kelly, Rodman Linn, Byungkyu Park, Alan Perelson, Nagui Rouphail, Jerome Sacks, and Frederic Schoenberg. Workshop on statistical approaches for the evaluation of complex computer models. Statistical Science, 17(2):173–192, May 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1030550860.

Baker:2022:ASC

[BBF^{+}22] Evan Baker, Pierre Barbillon, Arindam Fadikar, Robert B. Gramacy, Radu Herbei, David Higdon, Jiangeng Huang, Leah R. Johnson, Pulong Ma, Anirban Mondal, Bianica Pires, Jerome Sacks, and Vadim Sokolov. Analyzing stochastic computer models: a review with opportunities. Statistical Science, 37(1):64–89, February 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-37/issue-1/Analyzing-Stochastic-Computer-
Models-A-Review-with-Opportunities/10.1214/21-STS822.
full.

Bacallado:2017:SPG

[BBFT17] S. Bacallado, M. Battiston, S. Favaro, and L. Trippa. Sufficientness postulates for Gibbs-type priors and hierarchical generalizations. Statistical Science, 32(4):487–500, 11 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Bajari:2023:EDM

[BBI^{+}23] Patrick Bajari, Brian Burdick, Guido W. Imbens, Lorenzo Masoero, James McQueen, Thomas S. Richardson, and Ido M. Rosen. Experimental design in marketplaces. Statistical Science, 38(3):458–476, August 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-3/Experimental-Design-in-Marketplaces/
10.1214/23-STS883.full.

Borle:2006:TBP

[BBK06] Sharad Borle, Peter Boatwright, and Joseph B. Kadane. The timing of bid placement and extent of multiple bidding: An empirical investigation using eBay online auctions. Statistical Science, 21(2):194–205, May 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1154979821.

Buja:2019:MAI

[BBK^{+}19] Andreas Buja, Lawrence Brown, Arun Kumar Kuchibhotla, Richard Berk, Edward George, and Linda Zhao. Models as approximations II: a model-free theory of parametric regression. Statistical Science, 34(4):545–565, November 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1578474017. See

Part I [BBB^{+}19], discussion [vdG19, Law19, FJL^{+}19, Tjø19, Lit19, DKK19, WSC19, GK19], and rejoinder [BKB^{+}19].

Bar:2010:LAE

[BBSW10] Haim Bar, James Booth, Elizabeth Schifano, and Martin T. Wells. Laplace approximated EM microarray analysis: An empirical Bayes approach for comparative microarray experiments. Statistical Science, 25(3):388–407, August 2010. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1294167966.

Berger:1997:UFB

[BBW97] J. O. Berger, B. Boukai, and Y. Wang. Unified frequentist and Bayesian testing of a precise hypothesis. Statistical Science, 12 (3):133–160, August 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1030037904.

Bartlett:1991:ESR

[BC91a] Robert H. Bartlett and Richard G. Cornell. [Ethics and statistics in randomized clinical trials]: Comment. Statistical Science, 6(1):63–65, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011935.

Baxter:1991:C

[BC91b] Ron Baxter and Murray Cameron. Comment. Statistical Science, 6(4):339–343, November 1991. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011572.

Berger:2005:CFR

[BC05] Roger L. Berger and George Casella. Comment: Fuzzy and randomized confidence intervals and P -values. Statistical Science, 20(4):372–374, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1137076654. See [GM05a].

Bayarri:2007:BCS

[BC07a] M. J. Bayarri and M. E. Castellanos. Bayesian checking of the second levels of hierarchical models. Statistical Science, 22 (3):322–343, August 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid. org/euclid.ss/1199285031. See comments [Eva07, Gel07a, Joh07, BC07b, LL07].

Bayarri:2007:RBC

[BC07b] M. J. Bayarri and M. E. Castellanos. Rejoinder: Bayesian checking of the second levels of hierarchical models. Statistical Science, 22(3):363–367, August 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1199285036. See [BC07a].

Bell:2007:CSS

[BC07c] Robert M. Bell and Michael L. Cohen. Comment: Struggles with survey weighting and regression modeling. Statistical Science, 22(2):165–167, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1190905512. See [Gel07c].

Brentnall:2020:RMB

[BC20] Adam R. Brentnall and Jack Cuzick. Risk models for breast cancer and their validation. Statistical Science, 35(1):14–30, February 2020. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1583226026.

Battey:2022:SPI

[BC22] H. S. Battey and D. R. Cox. Some perspectives on inference in high dimensions. Statistical Science, 37(1):110–122, February 2022. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
journals/statistical-science/volume-37/issue-1/Some-
Perspectives-on-Inference-in-High-Dimensions/10.1214/
21-STS824.full.

Brown:2001:IEB

[BCD01] Lawrence D. Brown, T. Tony Cai, and Anirban DasGupta. Interval estimation for a binomial proportion. Statistical Science, 16(2):101–133, May 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009213286.

Brown:2005:CFR

[BCD05] Lawrence D. Brown, T. Tony Cai, and Anirban DasGupta. Comment: Fuzzy and randomized confidence intervals and P - values. Statistical Science, 20(4):375–379, November 2005. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1137076655.

See [GM05a].

Bickel:2007:MT

[BCK^{+}07] Peter J. Bickel, Chao Chen, Jaimyoung Kwon, John Rice, Erik van Zwet, and Pravin Varaiya. Measuring traffic. Statistical Science, 22(4):581–597, November 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1207580173.

Brooks:2000:BAS

[BCM00] S. P. Brooks, E. A. Catchpole, and B. J. T. Morgan. Bayesian animal survival estimation. Statistical Science, 15(4):357–376, November 2000. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1009213003.

Brantner:2023:MIT

[BCN^{+}23] Carly Lupton Brantner, Ting-Hsuan Chang, Trang Quynh Nguyen, Hwanhee Hong, Leon Di Stefano, and Elizabeth A. Stuart. Methods for integrating trials and non-experimental data to examine treatment effect heterogeneity. Statistical Science, 38(4):640–654, November 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-4/Methods-for-Integrating-Trials-

and-Non-experimental-Data-to-Examine/10.1214/23-STS890.
full.

Berti:2008:CGS

[BCPR08] Patrizia Berti, Guido Consonni, Luca Pratelli, and Pietro Rigo. Comment: Gibbs sampling, exponential families and orthogonal polynomials. Statistical Science, 23(2):179–182, May 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1219339108. See [DKSC08a].

Benkeser:2020:NSE

[BCvdL20] David Benkeser, Weixin Cai, and Mark J. van der Laan. A nonparametric super-efficient estimator of the average treatment effect. Statistical Science, 35(3):484–495, August 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-35/issue-3/
A-Nonparametric-Super-Efficient-Estimator-of-the-Average-
Treatment-Effect/10.1214/19-STS735.full.

Becker:1987:DGDa

[BCW87a] Richard A. Becker, William S. Cleveland, and Allan R. Wilks. Dynamic graphics for data analysis. Statistical Science, 2(4): 355–383, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013104.

Becker:1987:DGDb

[BCW87b] Richard A. Becker, William S. Cleveland, and Allan R. Wilks. [Dynamic graphics for data analysis]: Rejoinder. Statistical Science, 2(4):392–395, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177013110.

Berger:1987:TPHa

[BD87a] James O. Berger and Mohan Delampady. Testing precise hypotheses. Statistical Science, 2(3):317–335, August 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013238.

Berger:1987:TPHb

[BD87b] James O. Berger and Mohan Delampady. [Testing precise hypotheses]: Rejoinder. Statistical Science, 2(3):348–352, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013245.

Billera:2001:GIM

[BD01] Louis J. Billera and Persi Diaconis. A geometric interpretation of the Metropolis–Hastings algorithm. Statistical Science, 16(4):335–339, November 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1015346318.

Babu:2004:SSC

[BD04] G. Jogesh Babu and S. George Djorgovski. Some statistical and computational challenges, and opportunities in astronomy. Statistical Science, 19(2):322–332, May 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1105714166.

Brazzale:2008:API

[BD08] Alessandra R. Brazzale and Anthony C. Davison. Accurate parametric inference for small samples. Statistical Science, 23(4):465–484, November 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1242049390.

Brillinger:2009:CMR

[BD09] David R. Brillinger and Richard A. Davis. A conversation with Murray Rosenblatt. Statistical Science, 24(1):116–140, February 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1255009014.

Berger:2019:LBC

[BD19] James O. Berger and Anirban DasGupta. Larry Brown’s contributions to parametric inference, decision theory and foundations: a survey. Statistical Science, 34(4):621–634, November 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1578474028.

Balding:1994:DFR

[BDN94] David J. Balding, Peter Donnelly, and Richard A. Nichols. [DNA fingerprinting: A review of the controversy]: Comment: Some causes for concern about DNA profiles. Statistical Science, 9(2):248–251, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177010489.

Birrell:2018:ESS

[BDP18] Paul J. Birrell, Daniela De Angelis, and Anne M. Presanis. Evidence synthesis for stochastic epidemic models. Statistical Science, 33(1):34–43, 02 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Bhadra:2019:LMH

[BDPW19] Anindya Bhadra, Jyotishka Datta, Nicholas G. Polson, and Brandon Willard. Lasso meets horseshoe: a survey. Statistical Science, 34(3):405–427, August 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1570780977.

Baddeley:2022:DSS

[BDR^{+}22] Adrian Baddeley, Tilman M. Davies, Suman Rakshit, Gopalan Nair, and Greg McSwiggan. Diffusion smoothing for spatial point patterns. Statistical Science, 37(1):123–142, February 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-1/
Diffusion-Smoothing-for-Spatial-Point-Patterns/10.1214/
21-STS825.full.

Becker:2007:MMA

[Bec07] Betsy Jane Becker. Multivariate meta-analysis: Contributions of Ingram Olkin. Statistical Science, 22(3):401–406, August 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1199285039.

Becker:2009:CLG

[Bec09] Mark P. Becker. A conversation with Leo Goodman. Statistical Science, 24(3):361–385, August 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1270041261.

Begg:1989:ITP

[Beg89] Colin B. Begg. [Investigating therapies of potentially great benefit: ECMO]: Comment. Statistical Science, 4(4):320– 322, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012388.

Behseta:2019:CRK

[Beh19] Sam Behseta. A conversation with Robert E. Kass. Statistical Science, 34(2):334–348, May 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1563501645.

Belsley:1987:CLS

[Bel87] David A. Belsley. [Collinearity and least squares regression]: Comment: Well-conditioned collinearity indices. Statistical Science, 2(1):86–91, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013441.

Bellhouse:1991:GL

[Bel91] D. R. Bellhouse. The genoese lottery. Statistical Science, 6 (2):141–148, May 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011819.

Bellhouse:1993:RRH

[Bel93] David Bellhouse. The role of roguery in the history of probability. Statistical Science, 8(4):410–420, November 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010785.

Bellhouse:2004:RTB

[Bel04] D. R. Bellhouse. The Reverend Thomas Bayes, FRS: A biography to celebrate the tercentenary of his birth. Statistical Science, 19(1):3–43, February 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1089808271.

Benjamini:2008:CME

[Ben08] Yoav Benjamini. Comment: Microarrays, empirical Bayes and the two-groups model. Statistical Science, 23(1):23–28, February 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1215441277. See [Efr08a].

Berger:1986:ASP

[Ber86] James O. Berger. [The axioms of subjective probability]: Comment. Statistical Science, 1(3):351–352, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013614. See [Fis86a].

Berger:1988:IBS

[Ber88] James O. Berger. [The interface between statistics and philosophy of science]: Comment. Statistical Science, 3(4):403– 404, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012757.

Bernardo:1989:HGY

[Ber89a] Jose M. Bernardo. [How to get your first research grant]: Comment. Statistical Science, 4(2):138–139, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012594.

Bernardo:1989:GAI

[Ber89b] Jose M. Bernardo. [The geometry of asymptotic inference]: Comment: On multivariate Jeffreys’ priors. Statistical Science, 4(3):227–229, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012483.

Berry:1989:ITP

[Ber89c] Donald A. Berry. [Investigating therapies of potentially great benefit: ECMO]: Comment: Ethics and ECMO. Statistical Science, 4(4):306–310, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012385.

Berger:1990:DDT

[Ber90a] James O. Berger. [Developments in decision-theoretic variance estimation]: Comment. Statistical Science, 5(1):102–103, February 1990. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177012264.

Berger:1990:WML

[Ber90b] James O. Berger. [The 1988 Wald memorial lectures: The present position in Bayesian statistics]: Comment. Statistical Science, 5(1):71–75, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012255.

Bernardo:1990:WML

[Ber90c] Jose M. Bernardo. [The 1988 Wald memorial lectures: The present position in Bayesian statistics]: Comment. Statistical Science, 5(1):75–76, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012256.

Berry:1991:IUDa

[Ber91a] Donald A. Berry. Inferences using DNA profiling in forensic identification and paternity cases. Statistical Science, 6 (2):175–189, May 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011822.

Berry:1991:IUDb

[Ber91b] Donald A. Berry. [Inferences using DNA profiling in forensic identification and paternity cases]: Rejoinder. Statistical Science, 6(2):202–205, May 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011827.

Beran:1992:SMDa

[Ber92a] Jan Beran. Statistical methods for data with long-range dependence. Statistical Science, 7(4):404–416, November 1992. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011122.

Beran:1992:SMDb

[Ber92b] Jan Beran. [Statistical methods for data with long-range dependence]: Rejoinder. Statistical Science, 7(4):425–427, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011127.

Berliner:1992:SPCa

[Ber92c] L. Mark Berliner. Statistics, probability and chaos. Statistical Science, 7(1):69–90, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011444.

Berliner:1992:SPCb

[Ber92d] L. Mark Berliner. [Statistics, probability and chaos]: Rejoinder (part 2). Statistical Science, 7(1):118–122, February 1992. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011452.

Beran:1994:BMT

[Ber94a] Rudolf Beran. [Bootstrap: More than a stab in the dark?]: Comment. Statistical Science, 9(3):395–396, August 1994. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010384.

Berry:1994:DFR

[Ber94b] Donald A. Berry. [DNA fingerprinting: A review of the controversy]: Comment. Statistical Science, 9(2):252–255, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010490.

Berger:1999:CPW

[Ber99] Roger Berger. Comment on Perlman and Wu, “The Emperor’s new tests” (with rejoinder by authors). Statistical Science, 14(4):370–381, November 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1009212518. See [PW99].

Beran:2003:IBS

[Ber03a] Rudolf Beran. The impact of the bootstrap on statistical algorithms and theory. Statistical Science, 18(2):175–184, May 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1063994972.

Berger:2003:CFJ

[Ber03b] James O. Berger. Could Fisher, Jeffreys and Neyman have agreed on testing? Statistical Science, 18(1):1–32, February 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1056397485.

Berliner:2003:UCC

[Ber03c] L. Mark Berliner. Uncertainty and climate change. Statistical Science, 18(4):430–435, November 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1081443227.

Berry:2004:BSE

[Ber04] Donald A. Berry. Bayesian statistics and the efficiency and ethics of clinical trials. Statistical Science, 19(1):175–187, February 2004. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1089808281.

Bernardo:2009:C

[Ber09] José M. Bernardo. Comment. Statistical Science, 24(2):173– 175, May 2009. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1263478374.

Bernardo:2011:DOP

[Ber11] José M. Bernardo. Discussion of “Objective Priors: An Introduction for Frequentists” by M. Ghosh. Statistical Science, 26 (2):203–205, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1312204007. See [Gho11a].

Berry:2012:DMB

[Ber12] Don Berry. Discussion of “Multivariate Bayesian Logistic Regression for Analysis of Clinical Trial Safety Issues” by W. DuMouchel. Statistical Science, 27(3):344–345, August 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1346849941. See [DuM12a].

Best:2005:LCC

[Bes05] Joel Best. Lies, calculations and constructions: Beyond how to lie with statistics. Statistical Science, 20(3):210–214, August 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1124891286.

Billard:1991:ESS

[BF91] L. Billard and Marianne A. Ferber. Elizabeth Scott: Scholar, teacher, administrator. Statistical Science, 6(2):206–216, May 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011828.

Beran:1998:CGW

[BF98] R. J. Beran and N. I. Fisher. A conversation with Geoff Watson. Statistical Science, 13(1):75–93, February 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1028905975.

Beran:2009:ESB

[BF09] R. J. Beran and N. I. Fisher. An evening spent with Bill van Zwet. Statistical Science, 24(1):87–115, February 2009. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1255009013.

Bellhouse:2015:HOP

[BF15] David R. Bellhouse and Nicolas Fillion. Le Her and other problems in probability discussed by Bernoulli, Montmort and Waldegrave. Statistical Science, 30(1):26–39, February 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1425492438.

Borchers:2016:SCR

[BF16] David Borchers and Rachel Fewster. Spatial capture–recapture models. Statistical Science, 31(2):219–232, May 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1464105039.

Buja:1986:CSR

[BFK86] Andreas Buja, E. B. Fowlkes, and J. R. Kettenring. [Computers in statistical research]: Comment. Statistical Science, 1(4): 440–442, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013510. See [Edd86a].

Bedard:2007:HAB

[BFW07] M. Bédard, D. A. S. Fraser, and A. Wong. Higher accuracy for Bayesian and frequentist inference: Large sample theory for small sample likelihood. Statistical Science, 22(3):301–321, August 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1199285030.

Booth:2009:MVC

[BFWW09] James G. Booth, Walter T. Federer, Martin T. Wells, and Russell D. Wolfinger. A multivariate variance components model for analysis of covariance in designed experiments. Statistical Science, 24(2):223–237, May 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1263478383.

Bienenstock:1994:NNR

[BG94] Elie Bienenstock and Stuart Geman. [Neural networks: A review from statistical perspective]: Comment. Statistical Science, 9(1):36–38, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177010641.

Bellhouse:1999:HSS

[BG99] David R. Bellhouse and Christian Genest. A history of the statistical society of Canada: the formative years (with discussion). Statistical Science, 14(1):80–125, February 1999. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1009211807.

Bellhouse:2005:PHC

[BG05] David R. Bellhouse and Christian Genest. A public health controversy in 19th century Canada. Statistical Science, 20 (2):178–192, May 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1121347639.

Bellhouse:2007:MBA

[BG07] David R. Bellhouse and Christian Genest. Maty’s biography of Abraham De moivre, translated, annotated and augmented. Statistical Science, 22(1):109–136, February 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1185975640.

Bellhouse:2024:AGC

[BG24] David R. Bellhouse and Christian Genest. Antoine Gombaud, Chevalier de Méré. Statistical Science, 39(3):493–507, August 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-3/
Antoine-Gombaud-Chevalier-de-M%c3%a9r%c3%a9/10.1214/
24-STS926.full.

Bapna:2006:MDC

[BGGM06] Ravi Bapna, Paulo Goes, Ram Gopal, and James R. Marsden. Moving from data-constrained to data-enabled research: Experiences and challenges in collecting, validating and analyzing large-scale e-commerce data. Statistical Science, 21(2):116–130, May 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1154979815.

Besag:1995:BCSa

[BGHM95a] Julian Besag, Peter Green, David Higdon, and Kerrie Mengersen. Bayesian computation and stochastic systems. Statistical Science, 10(1):3–41, February 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177010123.

Besag:1995:BCSb

[BGHM95b] Julian Besag, Peter Green, David Higdon, and Kerrie Mengersen. [Bayesian computation and stochastic systems]: Rejoinder. Statistical Science, 10(1):58–66, February 1995. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010130.

Benedetti:1993:RAH

[BGLC93] Jacqueline Benedetti, Stephanie Green, Mei-Ling Lee, and John Crowley. Report of the ad hoc committee on design of an experiment on double-blind refereeing. Statistical Science, 8(3): 318–320, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010905.

Berhane:2004:SIS

[BGST04] Kiros Berhane, W. James Gauderman, Daniel O. Stram, and Duncan C. Thomas. Statistical issues in studies of the longterm effects of air Pollution: The Southern California Children’s Health Study. Statistical Science, 19(3):414–449, August 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1110999309.

Barry:1987:SAHa

[BH87a] Daniel Barry and J. A. Hartigan. Statistical analysis of hominoid molecular evolution. Statistical Science, 2(2):191–207, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013353.

Barry:1987:SAHb

[BH87b] Daniel Barry and J. A. Hartigan. [Statistical analysis of hominoid molecular evolution]: Rejoinder. Statistical Science, 2 (2):209–210, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013356.

Buja:1990:OOM

[BH90] A. Buja and C. Hurley. [OMEGA (online multivariate exploratory graphical analysis): Routine searching for structure]: Comment. Statistical Science, 5(2):208–211, May 1990. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012167.

Bedrick:1992:SEI

[BH92] Edward J. Bedrick and Joe R. Hill. [A survey of exact inference for contingency tables]: Comment. Statistical Science, 7(1): 153–157, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011455.

Berger:1996:BTI

[BH96] Roger L. Berger and Jason C. Hsu. Bioequivalence trials, intersection-union tests and equivalence confidence sets. Statistical Science, 11(4):283–319, November 1996. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1032280304.

Brown:1999:CLM

[BH99] Byron Wm. Brown, Jr. and Myles Hollander. A conversation with Lincoln E. Moses. Statistical Science, 14(3):338–354, August 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1009212412.

Bolton:2002:SFD

[BH02] Richard J. Bolton and David J. Hand. Statistical fraud detection: A review. Statistical Science, 17(3):235–255, August 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1042727940.

Buhlmann:2007:BAR

[BH07a] Peter Bühlmann and Torsten Hothorn. Boosting algorithms: Regularization, prediction and model fitting. Statistical Science, 22(4):477–505, November 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1207580163. See comments

[BMW07, Has07, BH07b].

Buhlmann:2007:RBA

[BH07b] Peter Bühlmann and Torsten Hothorn. Rejoinder: Boosting algorithms: Regularization, prediction and model fitting. Statistical Science, 22(4):516–522, November 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1207580166. See [BH07a].

Bogomolov:2023:RAM

[BH23] Marina Bogomolov and Ruth Heller. Replicability across multiple studies. Statistical Science, 38(4):602–620, November 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/journals/
statistical-science/volume-38/issue-4/Replicability-
Across-Multiple-Studies/10.1214/23-STS892.full.

Bhansali:1987:RTF

[Bha87] R. J. Bhansali. [Rational transfer function approximation]: Comment. Statistical Science, 2(2):151–152, May 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013344.

Bhamidi:2022:CDJ

[Bha22] Shankar Bhamidi. A conversation with David J. Aldous. Statistical Science, 37(4):607–624, November 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-37/issue-4/A-Conversation-with-David-
J-Aldous/10.1214/22-STS849.full.

Bar-Hillel:1999:SBC

[BHBNKM99] Maya Bar-Hillel, Dror Bar-Natan, Gil Kalai, and Brendan McKay. Solving the Bible code puzzle. Statistical Science, 14 (2):150–173, May 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009212243.

Backlund:2021:HSG

[BHJK21] Grant Backlund, James P. Hobert, Yeun Ji Jung, and Kshitij Khare. A hybrid scan Gibbs sampler for Bayesian models with latent variables. Statistical Science, 36(3):379–399, August 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/journals/
statistical-science/volume-36/issue-3/A-Hybrid-Scan-
Gibbs-Sampler-for-Bayesian-Models-with-Latent/10.1214/
20-STS788.full.

Bevilacqua:2015:FMC

[BHP15] Moreno Bevilacqua, Amanda S. Hering, and Emilio Porcu. On the flexibility of multivariate covariance models: Comment on the paper by Genton and Kleiber. Statistical Science, 30 (2):167–169, May 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1433341473. See [GK15a].

Bickel:1988:RBR

[Bic88] Peter J. Bickel. [Rank-based robust analysis of linear models. I. Exposition and review]: Comment. Statistical Science, 3 (2):263–264, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012919.

Billard:1986:CSR

[Bil86] Lynne Billard. [Computers in statistical research]: Comment. Statistical Science, 1(4):446–448, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013513.  See

[Edd86a].

Billard:1993:RAH

[Bil93] L. Billard. [Report of the ad hoc committee on design of an experiment on double-blind refereeing]: Comment. Statistical Science, 8(3):320–322, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177010906.

Bingham:1996:CDK

[Bin96] N. H. Bingham. A conversation with David Kendall. Statistical Science, 11(3):159–188, August 1996. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1032280213.

Bartlett:2006:C

[BJM06] Peter L. Bartlett, Michael I. Jordan, and Jon D. McAuliffe. Comment. Statistical Science, 21(3):341–346, August 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1166642437.

Berger:2012:BNS

[BJM12] James Berger, William H. Jefferys, and Peter Müller. Bayesian nonparametric shrinkage applied to Cepheid star oscillations. Statistical Science, 27(1):3–10, February 2012. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1331729978.

Bach:2012:SST

[BJMO12] Francis Bach, Rodolphe Jenatton, Julien Mairal, and Guillaume Obozinski. Structured sparsity through convex optimization. Statistical Science, 27(4):450–468, November 2012. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1356098550.

Bjornstad:1990:PLRa

[Bjø90a] Jan F. Bjørnstad. Predictive likelihood: A review. Statistical Science, 5(2):242–254, May 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012175.

Bjornstad:1990:PLRb

[Bjø90b] Jan F. Bjørnstad. [Predictive likelihood: A review]: Rejoinder. Statistical Science, 5(2):262–265, May 1990. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012178.

Bjornstad:2014:DBA

[Bjø14] Jan F. Bjørnstad. Discussion of “On the Birnbaum Argument for the Strong Likelihood Principle”. Statistical Science, 29 (2):259–260, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1408368579. See [May14a].

Broderick:2013:CFM

[BJP13] Tamara Broderick, Michael I. Jordan, and Jim Pitman. Cluster and feature modeling from combinatorial stochastic processes. Statistical Science, 28(3):289–312, August 2013. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1377696938.

Barlow:1987:SSW

[BK87] Richard E. Barlow and Zohel S. Khalil. [Survey of Soviet work in reliability]: Comment. Statistical Science, 2(4):495– 497, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013129.

Buja:2008:CPH

[BK08] Andreas Buja and Hans R. Künsch. A conversation with Peter Huber. Statistical Science, 23(1):120–135, February 2008. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1215441288.

Bertsimas:2017:LRA

[BK17] Dimitris Bertsimas and Angela King. Logistic regression: From art to science. Statistical Science, 32(3):367–384, August 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1504253122.

Behseta:2024:CSM

[BK24] Sam Behseta and Robert E. Kass. A conversation with Stephen M. Stigler. Statistical Science, 39(1):192–208, February 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-1/
A-Conversation-with-Stephen-M-Stigler/10.1214/22-STS878.
full.

Buja:2019:MAR

[BKB^{+}19] Andreas Buja, Arun Kumar Kuchibhotla, Richard Berk, Edward George, Eric Tchetgen Tchetgen, and Linda Zhao. Models as approximations — rejoinder. Statistical Science, 34 (4):606–620, November 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1578474027. See [BBB^{+}19,

BBK^{+}19].

Bickel:1990:CII

[BL90] Peter J. Bickel and Lucien Le Cam. A conversation with Ildar Ibragimov. Statistical Science, 5(3):347–355, August 1990. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012103.

Bartroff:2010:ADP

[BL10] Jay Bartroff and Tze Leung Lai. Approximate dynamic programming and its applications to the design of Phase I cancer trials. Statistical Science, 25(2):245–257, May 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1290175845.

Bowen:2020:CSD

[BL20] Claire McKay Bowen and Fang Liu. Comparative study of differentially private data synthesis methods. Statistical Science, 35(2):280–307, May 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1591171232.

Blattenberger:1988:EDS

[Bla88] Gail Blattenberger. [Employment discrimination and statistical science]: Comment. Statistical Science, 3(2):183–185, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012900.

Bura:2021:CDC

[BLL^{+}21] Efstathia Bura, Bing Li, Lexin Li, Christopher Nachtsheim, Daniel Pena, Claude Setodji, and Robert E. Weiss. A conversation with Dennis Cook. Statistical Science, 36(2):328– 337, May 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-36/issue-2/
A-Conversation-with-Dennis-Cook/10.1214/20-STS801.full.

Barrios:2013:MNR

[BLNBP13] Ernesto Barrios, Antonio Lijoi, Luis E. Nieto-Barajas, and Igor Prünster. Modeling with normalized random measure mixture models. Statistical Science, 28(3):313–334, August 2013. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1377696939.

Block:2001:CRB

[Blo01a] Henry W. Block. A conversation with Richard Barlow. Statistical Science, 16(4):368–388, November 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1015346321.

Block:2001:C

[Blo01b] Henry W. Block. Correction. Statistical Science, 16(4):389, November 2001. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1015346322.

Berger:1999:ILM

[BLW99] James O. Berger, Brunero Liseo, and Robert L. Wolpert. Integrated likelihood methods for eliminating nuisance parameters. Statistical Science, 14(1):1–28, February 1999. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1009211804.

Braun:1998:SMD

[BM98] Jerome V. Braun and Hans-Georg Müller. Statistical methods for DNA sequence segmentation. Statistical Science, 13 (2):142–162, May 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1028905933.

Beran:1999:AMS

[BM99] Jan Beran and Guerino Mazzola. Analyzing musical structure and performance — a statistical approach. Statistical Science, 14(1):47–79, February 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009211806.

Banerjee:2007:CSK

[BM07] Tathagata Banerjee and Rahul Mukerjee. A conversation with Shoutir Kishore Chatterjee. Statistical Science, 22(2):279–290, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1190905529.

Buckland:2016:YAP

[BM16] Stephen T. Buckland and Byron J. T. Morgan. 50-year anniversary of papers by Cormack, Jolly and Seber. Statistical Science, 31(2):141, May 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1464105032.

Brazzale:2024:LAN

[BM24] Alessandra R. Brazzale and Valentina Mameli. Likelihood asymptotics in nonregular settings: a review with emphasis on the likelihood ratio. Statistical Science, 39(2):322– 345, May 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-2/
Likelihood-Asymptotics-in-Nonregular-Settings--A-Review-
with-Emphasis/10.1214/23-STS910.full.

Buja:2007:CBA

[BMW07] Andreas Buja, David Mease, and Abraham J. Wyner. Comment: Boosting algorithms: Regularization, prediction and model fitting. Statistical Science, 22(4):506–512, November 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1207580164. See [BH07a].

Barndorff-Nielsen:1988:SMS

[BN88] O. E. Barndorff-Nielsen. [Saddlepoint methods and statistical inference]: Comment. Statistical Science, 3(2):228–229, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012907.

Barndorff-Nielsen:1989:GAI

[BN89] O. E. Barndorff-Nielsen. [The geometry of asymptotic inference]: Comment. Statistical Science, 4(3):222–227, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177012482. See correction [BN90].

Barndorff-Nielsen:1990:CNC

[BN90] O. E. Barndorff-Nielsen. Correction note: Comment by O. E. Barndorff-Nielsen on “The Geometry of Asymptotic Inference,” by R. E. Kass, 1989, Vol. 4, 222–227. Statistical Science, 5(3):370, August 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012105. See [BN89].

Balakrishnan:2004:CHD

[BN04] Narayanaswamy Balakrishnan and Haikady N. Nagaraja. A conversation with H. A. David. Statistical Science, 19(4):720– 734, November 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1113832736.

Buckland:2007:EPD

[BNF^{+}07] Stephen T. Buckland, Ken B. Newman, Carmen Fernández, Len Thomas, and John Harwood. Embedding population dynamics models in inference. Statistical Science, 22(1):44–58, February 2007. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1185975636.

Blum:2013:CRD

[BNPS13] M. G. B. Blum, M. A. Nunes, D. Prangle, and S. A. Sisson. A comparative review of dimension reduction methods in approximate Bayesian computation. Statistical Science, 28(2):189–208, May 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1369147911.

Breidt:2007:CSS

[BO07] F. Jay Breidt and Jean D. Opsomer. Comment: Struggles with survey weighting and regression modeling. Statistical Science, 22(2):168–170, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1190905513. See [Gel07c].

Breidt:2017:MAS

[BO17] F. Jay Breidt and Jean D. Opsomer. Model-assisted survey estimation with modern prediction techniques. Statistical Science, 32(2):190–205, May 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1494489811.

Bochkina:2016:SKN

[Boc16] Natalia A. Bochkina. Selection of KL neighbourhood in robust Bayesian inference. Statistical Science, 31(4):499–502, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1484816575.

Briol:2019:PIR

[BOG^{+}19a] François-Xavier Briol, Chris J. Oates, Mark Girolami, Michael A. Osborne, and Dino Sejdinovic. Probabilistic integration: A role in statistical computation? Statistical Science, 34(1):1–22, February 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1555056025. See comment

[HJ19] and rejoinder [BOG^{+}19b].

Briol:2019:RPI

[BOG^{+}19b] François-Xavier Briol, Chris J. Oates, Mark Girolami, Michael A. Osborne, and Dino Sejdinovic. Rejoinder: Probabilistic integration: A role in statistical computation? Statistical Science, 34(1):38–42, February 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1555056029.

Bohning:2016:RPR

[Böh16] Dankmar Böhning. Ratio plot and ratio regression with applications to social and medical sciences. Statistical Science, 31 (2):205–218, May 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1464105038.

Bookstein:1986:SSSa

[Boo86a] Fred L. Bookstein. Size and shape spaces for Landmark data in two dimensions. Statistical Science, 1(2):181–222, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013696.

Bookstein:1986:SSSb

[Boo86b] Fred L. Bookstein. [Size and shape spaces for Landmark data in two dimensions]: Rejoinder. Statistical Science, 1(2):238–242, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013702.

Bookstein:1989:SST

[Boo89] Fred L. Bookstein. [A survey of the statistical theory of shape]: Comment. Statistical Science, 4(2):99–105, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012583.

Boos:2003:IBW

[Boo03] Dennis D. Boos. Introduction to the bootstrap world. Statistical Science, 18(2):168–174, May 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1063994971.

Box:1987:GGF

[Box87] Joan Fisher Box. Guinness, Gosset, Fisher, and small samples. Statistical Science, 2(1):45–52, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013437.

Box:1990:ABE

[Box90a] George Box. [Applications in business and economic statistics: Some personal views]: Comment. Statistical Science, 5(4): 390–391, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012009.

Box:1990:UDP

[Box90b] George Box. [The unity and diversity of probability]: Comment. Statistical Science, 5(4):448–449, November 1990. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012024.

Bertsimas:2020:SRS

[BPV20] Dimitris Bertsimas, Jean Pauphilet, and Bart Van Parys. Sparse regression: Scalable algorithms and empirical performance. Statistical Science, 35(4):555–578, November 2020. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-35/issue-4/Sparse-Regression-Scalable-
Algorithms-and-Empirical-Performance/10.1214/19-STS701.

full. See comments [CTB20].

Bedford:2006:EER

[BQW06a] Tim Bedford, John Quigley, and Lesley Walls. Expert elicitation for reliable system design. Statistical Science, 21(4): 428–450, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177334515. See comments [FN06, Koe06,

Wan06, BQW06b].

Bedford:2006:REE

[BQW06b] Tim Bedford, John Quigley, and Lesley Walls. Rejoinder: Expert elicitation for reliable system design. Statistical Science, 21(4):460–462, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177334519. See [BQW06a].

Belin:1994:CWRa

[BR94a] Thomas R. Belin and John E. Rolph. Can we reach consensus on census adjustment? Statistical Science, 9(4):486– 508, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010261.

Belin:1994:CWRb

[BR94b] Thomas R. Belin and John E. Rolph. [Can we reach consensus on census adjustment?]: Rejoinder. Statistical Science, 9(4): 520–521, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010266.

Berchtold:2002:MTD

[BR02] André Berchtold and Adrian Raftery. The mixture transition distribution model for high-order Markov chains and non-Gaussian time series. Statistical Science, 17(3):328–356, August 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1042727943.

Brant:1986:IOH

[Bra86] Rollin Brant. [Influential observations, high leverage points, and outliers in linear regression]: Comment. Statistical Science, 1(3):405–407, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013626.

Bradley:1988:GOC

[Bra88] Ralph A. Bradley. [Golden oldies: Classic articles from the world of statistics and probability]: Comment: Harold Hotelling’s views on statistics. Statistical Science, 3(1):98– 103, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013008. See [Hot88c].

Bradburn:2007:TBK

[Bra07] Norman M. Bradburn. A tribute to bill Kruskal. Statistical Science, 22(2):262–263, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1190905522.

Breslow:1987:RSC

[Bre87] Norman Breslow. [The role of a second control group in an observational study]: Comment. Statistical Science, 2(3):311–312, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013236.

Breiman:1988:MRS

[Bre88a] Leo Breiman. [Monotone regression splines in action]: Comment. Statistical Science, 3(4):442–445, November 1988. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012762.

Breslow:1988:MMQ

[Bre88b] Norman Breslow. [From mouse-to-man: The quantitative assessment of cancer risks]: Comment: Risk assessment: Science or policy? Statistical Science, 3(1):28–33, February 1988. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012994.

Breslow:1990:BB

[Bre90a] Norman Breslow. Biostatistics and Bayes. Statistical Science, 5 (3):269–284, August 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012092.

Breslow:1990:BBR

[Bre90b] Norman Breslow. [Biostatistics and Bayes]: Rejoinder. Statistical Science, 5(3):295–298, August 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012098.

Breiman:1994:CAU

[Bre94a] Leo Breiman. The 1991 census adjustment: Undercount or bad data? Statistical Science, 9(4):458–475, November 1994. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010259.

Breiman:1994:CWR

[Bre94b] Leo Breiman. [Can we reach consensus on census adjustment?]: Rejoinder. Statistical Science, 9(4):521–527, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010267.

Breiman:1994:NNR

[Bre94c] Leo Breiman. [Neural networks: A review from statistical perspective]: Comment. Statistical Science, 9(1):38–42, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010642.

Breiman:2001:SMT

[Bre01] Leo Breiman. Statistical modeling: The two cultures (with comments and a rejoinder by the author). Statistical Science, 16 (3):199–231, August 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009213726.

Bret:2018:MII

[Bre18] Carles Bretó. Modeling and inference for infectious disease dynamics: A likelihood-based approach. Statistical Science, 33 (1):57–69, 02 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Brillinger:1986:GAM

[Bri86] David R. Brillinger. [Generalized additive models]: Comment. Statistical Science, 1(3):310–312, August 1986. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013605.

Brillinger:1987:PFO

[Bri87a] David R. Brillinger. [Prediction of future observations in growth curve models]: Comment. Statistical Science, 2(4): 448–450, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013120.

Brillinger:1987:RTF

[Bri87b] David R. Brillinger. [Rational transfer function approximation]: Comment. Statistical Science, 2(2):152–154, May 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013345.

Brillinger:1994:RSP

[Bri94] David R. Brillinger. [Report on statistics and physical oceanography]: Comment. Statistical Science, 9(2):201–202, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010482.

Brillinger:2003:TEP

[Bri03] David R. Brillinger. Three environmental probabilistic risk problems. Statistical Science, 18(4):412–421, November 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1081443225.

Brillinger:2008:NLD

[Bri08a] David R. Brillinger. The 2005 Neyman Lecture: Dynamic indeterminism in science. Statistical Science, 23(1):48–64, February 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.

ss/1215441282. See comment [Kün08, Yan08, Bri08b].

Brillinger:2008:RNL

[Bri08b] David R. Brillinger. Rejoinder: The 2005 Neyman lecture: Dynamic indeterminism in science. Statistical Science, 23(1): 76–77, February 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1215441285. See [Bri08a].

Baddeley:2011:SPS

[BRM11] Adrian Baddeley, Ege Rubak, and Jesper Møller. Score, pseudoscore and residual diagnostics for spatial point process models. Statistical Science, 26(4):613–646, November 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1330437942.

Brown:1990:DDT

[Bro90] Lawrence D. Brown. [Developments in decision-theoretic variance estimation]: Comment. Statistical Science, 5(1):103–106, February 1990. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177012265.

Brookmeyer:1993:BHI

[Bro93] Ron Brookmeyer. [Backcalculation of HIV infection rates]: Comment. Statistical Science, 8(2):102–104, May 1993. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010995.

Brunel:2018:MEC

[Bru18] Victor-Emmanuel Brunel. Methods for estimation of convex sets. Statistical Science, 33(4):615–632, November 2018. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/euclid.ss/1543482061.

Brandwein:1990:SES

[BS90] Ann Cohen Brandwein and William E. Strawderman. Stein estimation: The spherically symmetric case. Statistical Science, 5(3):356–369, August 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012104.

Block:1997:B

[BS97] Henry W. Block and Thomas H. Savits. Burn-in. Statistical Science, 12(1):1–19, February 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1029963258.

Banks:2006:DME

[BS06a] David L. Banks and Yasmin H. Said. Data mining in electronic commerce. Statistical Science, 21(2):234–246, May 2006. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1154979824.

Bousquet:2006:C

[BS06b] Olivier Bousquet and Bernhard Schölkopf. Comment. Statistical Science, 21(3):337–340, August 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1166642436.

Bhattacharya:2010:CGG

[BS10] Debasis Bhattacharya and Francisco J. Samaniego. A conversation with George G. Roussas. Statistical Science, 25(4): 566–587, November 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1300108238.

Brandwein:2012:SES

[BS12] Ann Cohen Brandwein and William E. Strawderman. Stein estimation for spherically symmetric distributions: Recent developments. Statistical Science, 27(1):11–23, February 2012. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1331729979.

Bray:2013:APP

[BS13] Andrew Bray and Frederic Paik Schoenberg. Assessment of point process models for earthquake forecasting. Statistical Science, 28(4):510–520, November 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1386078876.

Banerjee:2016:CMW

[BS16] Moulinath Banerjee and Bodhisattva Sen. A conversation with Michael Woodroofe. Statistical Science, 31(3):433–441, August 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1475001237.

Banerjee:2018:CJW

[BS18] Moulinath Banerjee and Richard J. Samworth. A conversation with Jon Wellner. Statistical Science, 33(4):633–651, November 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1543482062.

Bravington:2016:CKM

[BSA16] Mark V. Bravington, Hans J. Skaug, and Eric C. Anderson. Close-kin mark–recapture. Statistical Science, 31(2):259–274, May 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1464105042.

Bacchetti:1993:BHIa

[BSJ93a] Peter Bacchetti, Mark R. Segal, and Nicholas P. Jewell. Backcalculation of HIV infection rates. Statistical Science, 8(2):82– 101, May 1993. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177010994.

Bacchetti:1993:BHIb

[BSJ93b] Peter Bacchetti, Mark R. Segal, and Nicholas P. Jewell. [Backcalculation of HIV infection rates]: Rejoinder. Statistical Science, 8(2):114–119, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011001.

Bechhofer:1988:ODC

[BT88] Robert E. Bechhofer and Ajit C. Tamhane. [Optimal designs for comparing test treatments with controls]: Comment. Statistical Science, 3(4):477–480, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012768.

Bertsimas:1993:SA

[BT93] Dimitris Bertsimas and John Tsitsiklis. Simulated annealing. Statistical Science, 8(1):10–15, February 1993. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011077.

Bai:2018:HIR

[BT18] Shuyang Bai and Murad S. Taqqu. How the instability of ranks under long memory affects large-sample inference. Statistical Science, 33(1):96–116, 02 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Buckland:2016:CRM

[Buc16] Stephen T. Buckland. A conversation with Richard M. Cormack. Statistical Science, 31(2):142–150, May 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1464105033.

Buhlmann:2002:BTS

[Büh02] Peter Bühlmann. Bootstraps for time series. Statistical Science, 17(2):52–72, May 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1023798998.

Buhlmann:2014:DBB

[Büh14] Peter Bühlmann. Discussion of big Bayes stories and Bayes-Bag. Statistical Science, 29(1):91–94, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1399645732.

Buhlmann:2020:ICR

[Büh20a] Peter Bühlmann. Invariance, causality and robustness. Statistical Science, 35(3):404–426, August 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-35/issue-3/Invariance-Causality-and-
Robustness/10.1214/19-STS721.full. See rejoinder [Büh20b].

Buhlmann:2020:RIC

[Büh20b] Peter Bühlmann. Rejoinder: Invariance, causality and robustness. Statistical Science, 35(3):434–436, August 2020. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-35/issue-3/Rejoinder-Invariance-Causality-
and-Robustness/10.1214/20-STS797.full. See [Büh20a].

Burridge:1989:IGC

[Bur89] James Burridge. [Inference from grouped continuous data: A review]: Comment. Statistical Science, 4(2):179–181, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012602.

Burdick:1992:MMA

[Bur92] Donald S. Burdick. [Multilinear models: Applications in spectroscopy]: Comment. Statistical Science, 7(3):314–315, August 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011228.

Butler:1990:PLR

[But90] Ronald W. Butler. [Predictive likelihood: A review]: Comment. Statistical Science, 5(2):255–259, May 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012176.

Butler:2000:RFS

[But00] Ronald W. Butler. Reliabilities for feedback systems and their saddlepoint approximation. Statistical Science, 15(3):279–298, August 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1009212818.

Berry:2023:CRA

[BV23] Scott M. Berry and Kert Viele. Comment: Response adaptive randomization in practice. Statistical Science, 38(2):229– 232, May 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-2/
Comment-Response-Adaptive-Randomization-in-Practice/
10.1214/23-STS865F.full.

Becker:2007:SRS

[BW07] Betsy Jane Becker and Meng-Jia Wu. The synthesis of regression slopes in meta-analysis. Statistical Science, 22(3):414–429, August 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1199285041.

Bolin:2023:LSI

[BW23] David Bolin and Jonas Wallin. Local scale invariance and robustness of proper scoring rules. Statistical Science, 38(1):140– 159, February 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-1/
Local-scale-invariance-and-robustness-of-proper-scoring-
rules/10.1214/22-STS864.full.

Byar:1991:ESR

[Bya91] David P. Byar. [Ethics and statistics in randomized clinical trials]: Comment. Statistical Science, 6(1):65–68, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011936.

Brown:2012:GES

[BZ12] Lawrence D. Brown and Linda H. Zhao. A geometrical explanation of Stein shrinkage. Statistical Science, 27(1):24–30, February 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1331729980.

Bucher:2021:HRB

[BZ21] Axel Bücher and Chen Zhou. A horse race between the block maxima method and the peak-over-threshold approach. Statistical Science, 36(3):360–378, August 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-3/A-Horse-Race-between-the-
Block-Maxima-Method-and-the/10.1214/20-STS795.full.

Cai:2008:CME

[Cai08] T. Tony Cai. Comment: Microarrays, empirical Bayes and the two-group model. Statistical Science, 23(1):29–33, February 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1215441278. See [GM05a].

Cai:2012:MAI

[Cai12] T. Tony Cai. Minimax and adaptive inference in nonparametric function estimation. Statistical Science, 27(1):31–50, February 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1331729981.

Cai:2019:GMN

[Cai19] T. Tony Cai. Gaussianization machines for non-Gaussian function estimation models. Statistical Science, 34(4):635–656, November 2019. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1578474029.

Campbell:1986:SSS

[Cam86] Gregory Campbell. [Size and shape spaces for Landmark data in two dimensions]: Comment. Statistical Science, 1(2):227–228, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013699.

Campbell:1991:BGT

[Cam91] Katherine Campbell. [That BLUP is a good thing: The estimation of random effects]: Comment. Statistical Science, 6 (1):32–34, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011927.

Carroll:1993:RAH

[Car93] R. J. Carroll. [Report of the ad hoc committee on design of an experiment on double-blind refereeing]: Comment. Statistical Science, 8(3):323, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177010907.

Carriquiry:2011:EFV

[Car11] Alicia L. Carriquiry. Election forensics and the 2004 Venezuelan Presidential Recall Referendum as a case study. Statistical Science, 26(4):471–478, November 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1330437930.

Carroll:2014:EDD

[Car14a] Raymond J. Carroll. Estimating the distribution of dietary consumption patterns. Statistical Science, 29(1):2–8, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.

ss/1399645722. See discussion [FS14, Car14b].

Carroll:2014:RDE

[Car14b] Raymond J. Carroll. Reply to the discussion of “Estimating the Distribution of Dietary Consumption Patterns”. Statistical Science, 29(1):103, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1399645737. See [Car14a].

Carnegie:2019:CCM

[Car19] Nicole Bohme Carnegie. Comment: Contributions of model features to BART causal inference performance using ACIC 2016 competition data. Statistical Science, 34(1):90–93, February 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1555056036.

Casella:2003:FTJ

[Cas03a] George Casella. Foreword [to tribute to John Tukey]. Statistical Science, 18(3):283–284, August 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1076102415.

Casella:2003:ISA

[Cas03b] George Casella. Introduction to the Silver anniversary of the bootstrap. Statistical Science, 18(2):133–134, May 2003. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1063994967.

Cattelan:2012:MPC

[Cat12] Manuela Cattelan. Models for paired comparison data: A review with emphasis on dependent data. Statistical Science, 27 (3):412–433, August 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1346849947.

Casella:1987:TPH

[CB87] George Casella and Roger L. Berger. [Testing precise hypotheses]: Comment. Statistical Science, 2(3):344–347, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013243.

Cressie:2015:CMS

[CBD^{+}15] Noel Cressie, Sandy Burden, Walter Davis, Pavel N. Krivitsky, Payam Mokhtarian, Thomas Suesse, and Andrew Zammit-Mangion. Capturing multivariate spatial dependence: Model, estimate and then predict. Statistical Science, 30(2):170–175, May 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1433341474.

Carey:2014:FPC

[CC14] Vincent Carey and Dianne Cook. Four papers on contemporary software design strategies for statistical methodologists. Statistical Science, 29(2):165–166, May 2014. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1408368568.

Castelletti:2024:BSS

[CC24] Federico Castelletti and Guido Consonni. Bayesian sample size determination for causal discovery. Statistical Science, 39(2): 305–321, May 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-2/
Bayesian-Sample-Size-Determination-for-Causal-Discovery/
10.1214/23-STS905.full.

Chang:2008:CQF

[CCCH08] I-Shou Chang, Chung-Hsing Chen, Li-Chu Chien, and Chao A. Hsiung. Comment: Quantifying the fraction of missing information for hypothesis testing in statistical and genetic studies. Statistical Science, 23(3):318–320, August 2008. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1233153059.  See

[NMK08a].

Casarin:2024:LEU

[CCFR24] Roberto Casarin, Radu V. Craiu, Lorenzo Frattarolo, and Christian P. Robert. Living on the edge: an unified approach to antithetic sampling. Statistical Science, 39(1):115–136, February 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/journals/
statistical-science/volume-39/issue-1/Living-on-the-
Edge--An-Unified-Approach-to-Antithetic/10.1214/23-
STS888.full.

Castro:2004:NTR

[CCL^{+}04] Rui Castro, Mark Coates, Gang Liang, Robert Nowak, and Bin Yu. Network tomography: Recent developments. Statistical Science, 19(3):499–517, August 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1110999312.

Chatterjee:2009:ACC

[CCLC09] Nilanjan Chatterjee, Yi-Hau Chen, Sheng Luo, and Raymond J. Carroll. Analysis of case-control association studies: SNPs, imputation and haplotypes. Statistical Science, 24(4):489– 502, November 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1271770344.

Cao:2004:BEB

[CCS04] Jin Cao, William S. Cleveland, and Don X. Sun. Bandwidth estimation for best-effort Internet traffic. Statistical Science, 19 (3):518–543, August 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1110999313.

Crane:2021:SFM

[CD21] Harry Crane and Walter Dempsey. A statistical framework for modern network science. Statistical Science, 36(1):51–67, February 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-36/issue-1/
A-Statistical-Framework-for-Modern-Network-Science/
10.1214/19-STS759.full.

Carriquiry:2023:ESI

[CDR23] Alicia L. Carriquiry, Michael J. Daniels, and Nancy Reid. Editorial: Special issue on reproducibility and replicability. Statistical Science, 38(4):525–526, November 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-4/Editorial-Special-Issue-on-
Reproducibility-and-Replicability/10.1214/23-STS909.
full.

Casella:1995:IBE

[CDW95] George Casella, Thomas J. DiCiccio, and Martin T. Wells. [Inference based on estimating functions in the presence of nuisance parameters]: Comment: Alternative aspects of conditional inference. Statistical Science, 10(2):179–185, May 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010032.

Chen:2017:AIS

[CEH^{+}17] Qixuan Chen, Michael R. Elliott, David Haziza, Ye Yang, Malay Ghosh, Roderick J. A. Little, Joseph Sedransk, and Mary Thompson. Approaches to improving survey-weighted estimates. Statistical Science, 32(2):227–248, May 2017. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1494489813.

Celmins:1998:MG

[Cel98] Aivars Celmiņš. The method of Gauss in 1799. Statistical Science, 13(2):123–135, May 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1028905931.

Costigan-Eaves:1990:WP

[CEMR90] Patricia Costigan-Eaves and Michael Macdonald-Ross. William Playfair (1759–1823). Statistical Science, 5(3):318–326, August 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012100.

Cruze:2019:POC

[CEN^{+}19] Nathan B. Cruze, Andreea L. Erciulescu, Balgobin Nandram, Wendy J. Barboza, and Linda J. Young. Producing official county-level agricultural estimates in the United States: Needs and challenges. Statistical Science, 34(2):301–316, May 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1563501643.

Cook:2008:PFC

[CF08] R. Dennis Cook and Liliana Forzani. Principal fitted components for dimension reduction in regression. Statistical Science, 23(4):485–501, November 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1242049391. See [Coo07a].

Consonni:2013:WAB

[CFL13] Guido Consonni, Jonathan J. Forster, and Luca La Rocca. The whetstone and the alum block: Balanced objective Bayesian comparison of nested models for discrete data. Statistical Science, 28(3):398–423, August 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1377696943.

Crowley:1992:ETI

[CG92] John Crowley and Stephanie Green. [Evaluating therapeutic interventions: Some issues and experiences]: Comment. Statistical Science, 7(4):441–443, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011129.

Carlin:1993:BHI

[CG93] John B. Carlin and Andrew Gelman. [Backcalculation of HIV infection rates]: Comment: Assessing uncertainty in backprojection. Statistical Science, 8(2):104–106, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010996.

Casella:2002:VES

[CG02] George Casella and Jeff Gill. Voting, elections, and statistical science. Statistical Science, 17(4):381–382, November 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1049993198.

Clyde:2004:MU

[CG04] Merlise Clyde and Edward I. George. Model uncertainty. Statistical Science, 19(1):81–94, February 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1089808274.

Cox:1993:RAH

[CGP^{+}93] David Cox, Leon Gleser, Michael Perlman, Nancy Reid, and Kathryn Roeder. Report of the ad hoc committee on doubleblind refereeing. Statistical Science, 8(3):310–317, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010904.

Chatterjee:1986:IOHa

[CH86a] Samprit Chatterjee and Ali S. Hadi. Influential observations, high leverage points, and outliers in linear regression. Statistical Science, 1(3):379–393, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013622.

Chatterjee:1986:IOHb

[CH86b] Samprit Chatterjee and Ali S. Hadi. [Influential observations, high leverage points, and outliers in linear regression]: Rejoinder. Statistical Science, 1(3):415–416, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013630.

Csorgo:1989:AEP

[CH89] Miklós Csörgő and Lájos Horváth. [Asymptotics via empirical processes]: Comment. Statistical Science, 4(4):360–365, November 1989. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177012398.

Casella:2012:SCP

[CH12] George Casella and J. T. Gene Hwang. Shrinkage confidence procedures. Statistical Science, 27(1):51–60, February 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1331729982.

Carlin:2015:CAG

[CH15] Bradley P. Carlin and Amy H. Herring. A conversation with Alan Gelfand. Statistical Science, 30(3):413–422, August 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1439220720.

Chambers:1986:CED

[Cha86] John M. Chambers. [Computing environments for data analysis]: Comment. Statistical Science, 1(2):271–272, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013705.

Chatfield:1991:ASPa

[Cha91a] Christopher Chatfield. Avoiding statistical pitfalls. Statistical Science, 6(3):240–252, August 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011686.

Chatfield:1991:ASPb

[Cha91b] Christopher Chatfield. [Avoiding statistical pitfalls]: Rejoinder. Statistical Science, 6(3):267–268, August 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011693.

Chang:2004:SS

[Cha04] Ted Chang. Spatial statistics. Statistical Science, 19(4):624– 635, November 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1113832727.

Chambers:2014:OOP

[Cha14] John M. Chambers. Object-oriented programming, functional programming and R. Statistical Science, 29(2):167–180, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1408368569.

Chacon:2015:PBN

[Cha15] José E. Chacón. A population background for nonparametric density-based clustering. Statistical Science, 30(4):518– 532, November 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1449670856.

Chambers:2022:DSG

[Cha22] John M. Chambers. Data, science, and global disasters. Statistical Science, 37(2):284–288, May 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-37/issue-2/Data-Science-and-Global-Disasters/
10.1214/22-STS858.full.

Chen:1990:PAC

[Che90] Louis H. Y. Chen. [Poisson approximation and the Chen-Stein method]: Comment. Statistical Science, 5(4):429–432, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012019.

Chernoff:1991:IUD

[Che91] Herman Chernoff. [Inferences using DNA profiling in forensic identification and paternity cases]: Comment. Statistical Science, 6(2):192–196, May 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011824.

Chelton:1994:POB

[Che94] Dudley B. Chelton. Physical oceanography: A brief overview for statisticians. Statistical Science, 9(2):150–166, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010480.

Chernoff:1999:GEI

[Che99] Herman Chernoff. Gustav Elfving’s impact on experimental design. Statistical Science, 14(2):201–205, May 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1009212246.

Cheung:2010:SAM

[Che10] Ying Kuen Cheung. Stochastic approximation and modern model-based designs for dose-finding clinical trials. Statistical Science, 25(2):191–201, May 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1290175841.

Chen:2017:CMU

[Che17] Jiahua Chen. Consistency of the MLE under mixture models. Statistical Science, 32(1):47–63, February 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1491465627.

Cao:2022:UFH

[CHP22] Shanshan Cao, Xiaoming Huo, and Jong-Shi Pang. A unifying framework of high-dimensional sparse estimation with difference-of-convex (DC) regularizations. Statistical Science, 37(3):411–424, August 2022. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-37/issue-3/A-Unifying-Framework-of-High-
Dimensional-Sparse-Estimation-with-Difference/10.1214/
21-STS832.full.

Christensen:2007:CFL

[Chr07] Ronald Christensen. Comment: Fisher lecture: Dimension reduction in regression. Statistical Science, 22(1):27–31, February 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1185975632. See [Coo07a].

Chung:2017:RBT

[Chu17] EunYi Chung. Randomization-based tests for “No Treatment Effects”. Statistical Science, 32(3):349–351, August 2017. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/euclid.ss/1504253118.

Chen:2000:PPD

[CI00] Ming-Hui Chen and Joseph G. Ibrahim. Power prior distributions for regression models. Statistical Science, 15(1): 46–60, February 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009212673.

Christensen:2007:CSG

[CJ07] Ronald Christensen and Wesley Johnson. A conversation with Seymour Geisser. Statistical Science, 22(4):621–636, November 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1207580175.

Chipman:2016:CJW

[CJ16] Hugh A. Chipman and V. Roshan Joseph. A conversation with Jeff Wu. Statistical Science, 31(4):624–636, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1484816590.

Carvalho:2010:PLS

[CJLP10] Carlos M. Carvalho, Michael S. Johannes, Hedibert F. Lopes, and Nicholas G. Polson. Particle learning and smoothing. Statistical Science, 25(1):88–106, February 2010. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1280841735.

Coffman:1993:PAP

[CJLS93] E. G. Coffman, D. S. Johnson, G. S. Lueker, and P. W. Shor. Probabilistic analysis of packing and related partitioning problems. Statistical Science, 8(1):40–47, February 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011082.

Chambers:2006:MNA

[CJLV06a] John M. Chambers, David A. James, Diane Lambert, and Scott Vander Wiel. Monitoring networked applications with incremental quantile estimation. Statistical Science, 21(4):463– 475, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177334520. See [DLM06, LMN06, Yu06,

CJLV06b].

Chambers:2006:RMN

[CJLV06b] John M. Chambers, David A. James, Diane Lambert, and Scott Vander Wiel. Rejoinder: Monitoring networked applications with incremental quantile estimation. Statistical Science, 21(4): 485–486, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177334524. See [CJLV06a].

Cressie:1994:SAE

[CK94] Noel Cressie and Mark S. Kaiser. [Small area estimation: An appraisal]: Comment. Statistical Science, 9(1):76–80, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010648.

Cressie:2003:SSP

[CK03] Noel Cressie and John Kornak. Spatial statistics in the presence of location error with an application to remote sensing of the environment. Statistical Science, 18(4):436–456, November 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1081443228.

Cappello:2022:SCT

[CKLP22] Lorenzo Cappello, Jaehee Kim, Sifan Liu, and Julia A. Palacios. Statistical challenges in tracking the evolution of SARS-CoV-2. Statistical Science, 37(2):162–182, May 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-2/
Statistical-Challenges-in-Tracking-the-Evolution-of-
SARS-CoV-2/10.1214/22-STS853.full.

Chang:2000:RTP

[CKLR00] Ted Chang, Daijin Ko, Jiandong Lu, and Jean-Yves Royer. Regression techniques in plate tectonics. Statistical Science, 15(4): 342–356, November 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009213002.

Clark:1990:QPE

[Cla90] Herbert H. Clark. [Quantifying probabilistic expressions]: Comment. Statistical Science, 5(1):12–16, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012243.

Cliff:1990:QPE

[Cli90] Norman Cliff. [Quantifying probabilistic expressions]: Comment. Statistical Science, 5(1):16–18, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012244.

Clogg:1992:ISMa

[Clo92a] Clifford C. Clogg. The impact of sociological methodology on statistical methodology. Statistical Science, 7(2):183–196, May 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011354.

Clogg:1992:ISMb

[Clo92b] Clifford C. Clogg. [The impact of sociological methodology on statistical methodology]: Rejoinder. Statistical Science, 7 (2):205–207, May 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011359.

Chen:2016:AMC

[CLSW16] Hao Chen, Jason L. Loeppky, Jerome Sacks, and William J. Welch. Analysis methods for computer experiments: How to assess and what counts? Statistical Science, 31(1):40–60, February 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1455115913.

Chu:1991:CKRa

[CM91a] C.-K. Chu and J. S. Marron. Choosing a kernel regression estimator. Statistical Science, 6(4):404–419, November 1991. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011586.

Chu:1991:CKRb

[CM91b] C.-K. Chu and J. S. Marron. [Choosing a kernel regression estimator]: Rejoinder. Statistical Science, 6(4):433–436, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011592.

Colnet:2024:CIM

[CMC^{+}24] Bénédicte Colnet, Imke Mayer, Guanhua Chen, Awa Dieng, Ruohong Li, Gaël Varoquaux, Jean-Philippe Vert, Julie Josse, and Shu Yang. Causal inference methods for combining randomized trials and observational studies: a review. Statistical Science, 39(1):165–191, February 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-39/issue-1/Causal-Inference-Methods-
for-Combining-Randomized-Trials-and-Observational-Studies/
10.1214/23-STS889.full.

Cai:2019:STP

[CMN^{+}19] Junhui Cai, Avishai Mandelbaum, Chaitra H. Nagaraja, Haipeng Shen, and Linda Zhao. Statistical theory powering data science. Statistical Science, 34(4):669–691, November 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1578474031.

Clayton:1991:ASP

[CN91] Murray K. Clayton and Erik V. Nordheim. [Avoiding statistical pitfalls]: Comment. Statistical Science, 6(3):255–257, August 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011689.

Cnaan:1993:ISC

[Cna93] Avital Cnaan. [Is industrial statistics out of control?]: Comment. Statistical Science, 8(4):378–379, November 1993. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010774.

Chen:2012:PLH

[CO12] Dayue Chen and Ingram Olkin. Pao-Lu Hsu (Xu, Bao-lu): The grandparent of probability and statistics in China. Statistical Science, 27(3):434–445, August 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1346849948.

Cohen:1990:DDT

[Coh90] Arthur Cohen. [Developments in decision-theoretic variance estimation]: Comment. Statistical Science, 5(1):106–107, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012266.

Cohen:1991:GSP

[Coh91] Joel E. Cohen. [The growth and stabilization of populations]: Comment: Partially observed Markov chains and genetic demography. Statistical Science, 6(3):275–277, August 1991. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011695.

Conway:1988:EDS

[Con88] Delores A. Conway. [Employment discrimination and statistical science]: Comment. Statistical Science, 3(2):171–175, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012898.

Conway:1993:CST

[Con93] Delores A. Conway. [Can statistics tell us what we do not want to hear? the case of complex salary structures]: Comment. Statistical Science, 8(2):158–165, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011007.

Cook:1986:IOH

[Coo86] R. Dennis Cook. [Influential observations, high leverage points, and outliers in linear regression]: Comment. Statistical Science, 1(3):393–397, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013623.

Cook:2007:FLD

[Coo07a] R. Dennis Cook. Fisher lecture: Dimension reduction in regression. Statistical Science, 22(1):1–26, February 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1185975631.  See

comments [Chr07, Li07, LN07, Coo07b, CF08].

Cook:2007:RFL

[Coo07b] R. Dennis Cook. Rejoinder: Fisher lecture: Dimension reduction in regression. Statistical Science, 22(1):40–43, February 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1185975635. See [Coo07a].

Coolen:2016:SSR

[Coo16] Frank P. A. Coolen. On software and system reliability growth and testing. Statistical Science, 31(4):541–544, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1484816581.

Cornell:1987:RSC

[Cor87] Richard G. Cornell. [The role of a second control group in an observational study]: Comment. Statistical Science, 2(3): 310–311, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013235.

Cornell:1989:ITP

[Cor89] Richard G. Cornell. [Investigating therapies of potentially great benefit: ECMO]: Comment. Statistical Science, 4(4): 326–327, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012391.

Cox:1987:TPH

[Cox87] D. R. Cox. [Testing precise hypotheses]: Comment. Statistical Science, 2(3):335–336, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013239.

Cox:1990:RMS

[Cox90a] D. R. Cox. Role of models in statistical analysis. Statistical Science, 5(2):169–174, May 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012165.

Cox:1990:WML

[Cox90b] David R. Cox. [The 1988 Wald memorial lectures: The present position in Bayesian statistics]: Comment. Statistical Science, 5(1):76–78, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012257.

Cox:1991:EAM

[Cox91] Lawrence H. Cox. [Enhancing access to microdata while protecting confidentiality: Prospects for the future]: Comment. Statistical Science, 6(3):232–234, August 1991. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011682.

Chan:1994:RSP

[CP94] Ngai Hang Chan and Wilfredo Palma. [Report on statistics and physical oceanography]: Comment: Unit root and structural changes in tropical sea levels. Statistical Science, 9(2):203–207, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010483.

Cameron:2014:RPM

[CP14] Ewan Cameron and Anthony Pettitt. Recursive pathways to marginal likelihood estimation with prior-sensitivity analysis. Statistical Science, 29(3):397–419, August 2014. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1411437520.

Clarke:2015:ESM

[CPW15] Paul S. Clarke, Tom M. Palmer, and Frank Windmeijer. Estimating structural mean models with multiple instrumental variables using the generalised method of moments. Statistical Science, 30(1):96–117, February 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1425492442.

Cifarelli:1996:FCP

[CR96] Donato Michele Cifarelli and Eugenio Regazzini. De Finetti’s contribution to probability and statistics. Statistical Science, 11(4):253–282, November 1996. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1032280303.

Chopin:2017:LPI

[CR17] Nicolas Chopin and James Ridgway. Leave Pima Indians alone: Binary regression as a benchmark for Bayesian computation. Statistical Science, 32(1):64–87, February 2017. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1491465628.

Craig:1986:EDS

[Cra86] Cecil C. Craig. Early days in statistics at Michigan. Statistical Science, 1(2):292–293, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013710.

Crane:2016:RUE

[Cra16a] Harry Crane. Rejoinder: The ubiquitous Ewens sampling formula. Statistical Science, 31(1):37–39, February 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1455115912. See [Cra16b, Teh16].

Crane:2016:UES

[Cra16b] Harry Crane. The ubiquitous Ewens sampling formula. Statistical Science, 31(1):1–19, February 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1455115906. See [Teh16] and

rejoinder [Cra16a].

Cressie:1986:SSS

[Cre86] Noel Cressie. [Size and shape spaces for Landmark data in two dimensions]: Comment. Statistical Science, 1(2):226, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013698.

Critchley:1990:OOM

[Cri90] Frank Critchley. [OMEGA (online multivariate exploratory graphical analysis): Routine searching for structure]: Comment. Statistical Science, 5(2):211–213, May 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012168.

Cotter:2013:MMF

[CRSW13] S. L. Cotter, G. O. Roberts, A. M. Stuart, and D. White. MCMC methods for functions: Modifying old algorithms to make them faster. Statistical Science, 28(3):424–446, August 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1377696944.

Choirat:2012:EDP

[CS12a] Christine Choirat and Raffaello Seri. Estimation in discrete parameter models. Statistical Science, 27(2):278–293, May 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1340110873.

Cohen:2012:IPM

[CS12b] Arthur Cohen and Harold Sackrowitz. The interval property in multiple testing of pairwise differences. Statistical Science, 27(2):294–307, May 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1340110874.

Cooley:2012:DSM

[CS12c] D. Cooley and S. R. Sain. Discussion of “Statistical Modeling of Spatial Extremes” by A. C. Davison, S. A. Padoan and M. Ribatet. Statistical Science, 27(2):187–188, May 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1340110865. See [DPR12b].

Campolongo:2000:SAI

[CST00] F. Campolongo, A. Saltelli, and S. Tarantola. Sensitivity anaysis as an ingredient of modeling. Statistical Science, 15(4): 377–395, November 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009213004.

Cheng:1994:NNRa

[CT94a] Bing Cheng and D. M. Titterington. Neural networks: A review from a statistical perspective. Statistical Science, 9(1): 2–30, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010638.

Cheng:1994:NNRb

[CT94b] Bing Cheng and D. M. Titterington. [Neural networks: A review from statistical perspective]: Rejoinder. Statistical Science, 9(1):49–54, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177010646.

Chen:1998:CCC

[CT98] T. Timothy Chen and John Jen Tai. A conversation with C. C. Li. Statistical Science, 13(4):378–387, November 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1028905832.

Chan:2011:DFM

[CT11] Kung-Sik Chan and Ruey S. Tsay. Discussion of “Feature Matching in Time Series Modeling” by Y. Xia and H. Tong.

Statistical Science, 26(1):53–56, February 2011. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1307626563.

Chen:2020:LRS

[CTB20] Yuansi Chen, Armeen Taeb, and Peter Bühlmann. A look at robustness and stability of ℓ_{1}-versus ℓ_{0}-regularization: Discussion of papers by Bertsimas et al. and Hastie et al. Statistical Science, 35(4):614–622, November 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-35/issue-4/A-Look-at-Robustness-and-
Stability-of-ell_1-versus-ell_0/10.1214/20-STS809.full.

See [BPV20, HTT20].

Cui:1992:PMC

[CTSH92] Lu Cui, Martin A. Tanner, Debajyoti Sinha, and W. J. Hall. [Practical Markov chain Monte Carlo]: Comment: Monitoring convergence of the Gibbs sampler: Further experience with the Gibbs stopper. Statistical Science, 7(4):483–486, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011138.

Cutler:1992:SPC

[Cut92] Colleen D. Cutler. [Statistics, probability and chaos]: Comment. Statistical Science, 7(1):91–94, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011445.

Chaloner:1995:BED

[CV95] Kathryn Chaloner and Isabella Verdinelli. Bayesian experimental design: A review. Statistical Science, 10(3):273–304, August 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177009939.

Consonni:2008:CPS

[CV08] Guido Consonni and Piero Veronese. Compatibility of prior specifications across linear models. Statistical Science, 23(3): 332–353, August 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1233153062.

Cox:1993:BAE

[CW93a] D. R. Cox and Nanny Wermuth. [Bayesian analysis in expert systems]: Rejoinder. Statistical Science, 8(3):276–277, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010897.

Cox:1993:LDR

[CW93b] D. R. Cox and Nanny Wermuth. Linear dependencies represented by chain graphs. Statistical Science, 8(3):204–218, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010887.

Comess:2022:SMP

[CWHD22] Saskia Comess, Hannah Wang, Susan Holmes, and Claire Donnat. Statistical modeling for practical pooled testing during the COVID-19 pandemic. Statistical Science, 37(2):229– 250, May 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-2/
Statistical-Modeling-for-Practical-Pooled-Testing-During-
the-COVID-19/10.1214/22-STS857.full.

Chatterjee:1992:CFS

[CY92a] Sangit Chatterjee and Mustafa R. Yilmaz. Chaos, fractals and statistics. Statistical Science, 7(1):49–68, February 1992. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011443.

Chatterjee:1992:SPC

[CY92b] Sangit Chatterjee and Mustafa R. Yilmaz. [Statistics, probability and chaos]: Rejoinder (part 1). Statistical Science, 7(1): 114–117, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011451.

Chan:2014:CHT

[CY14a] Kung-Sik Chan and Qiwei Yao. A conversation with Howell Tong. Statistical Science, 29(3):425–438, August 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1411437522.

Chan:2014:OMR

[CY14b] Kwun Chuen Gary Chan and Sheung Chi Phillip Yam. Oracle, multiple robust and multipurpose calibration in a missing response problem. Statistical Science, 29(3):380–396, August 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1411437519.

Cho:2020:CDK

[CZL^{+}20] Hunyong Cho, Joshua P. Zitovsky, Xinyi Li, Minxin Lu, Kushal Shah, John Sperger, Matthew C. B. Tsilimigras, and Michael R. Kosorok. Comment: Diagnostics and kernel-based extensions for linear mixed effects models with endogenous covariates. Statistical Science, 35(3):396–399, August 2020. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-35/issue-3/Comment--Diagnostics-and-
Kernel-based-Extensions-for-Linear-Mixed/10.1214/20-
STS782.full.

Dahlhaus:1987:RTF

[Dah87] R. Dahlhaus. [Rational transfer function approximation]: Comment. Statistical Science, 2(2):154–156, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013346.

DallAglio:1989:HGY

[Dal89] Giorgio Dall’Aglio. [How to get your first research grant]: Comment. Statistical Science, 4(2):136–138, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012593.

Daniels:1988:SMS

[Dan88] H. E. Daniels. [Saddlepoint methods and statistical inference]: Comment. Statistical Science, 3(2):229, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012908.

Darnell:1988:HH

[Dar88] Adrian C. Darnell. Harold Hotelling 1895–1973. Statistical Science, 3(1):57–62, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177013000.

DasGupta:2005:CLB

[Das05] Anirban DasGupta. A conversation with Larry Brown. Statistical Science, 20(2):193–203, May 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1121347640.

Davis:1988:CPR

[Dav88] Devra Lee Davis. Changing policy roles of environmental epidemiology. Statistical Science, 3(3):281–285, August 1988. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012828.

David:1998:ESM

[Dav98a] H. A. David. Early sample measures of variability. Statistical Science, 13(4):368–377, November 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1028905831.

David:1998:SUU

[Dav98b] Herbert A. David. Statistics in U.S. universities in 1933 and the establishment of the Statistical Laboratory at Iowa State. Statistical Science, 13(1):66–74, February 1998. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1028905974.

Dawes:1986:SRC

[Daw86a] Robyn M. Dawes. [Savage revisited]: Comment. Statistical Science, 1(4):495–497, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013522.

Dawid:1986:SRC

[Daw86b] A. P. Dawid. [Savage revisited]: Comment. Statistical Science, 1(4):488–492, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177013520.

Dawson:1991:RMA

[Daw91] Ree Dawson. [Replication and meta-analysis in parapsychology]: Comment. Statistical Science, 6(4):382–385, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011579.

Dawid:2004:PCE

[Daw04] A. P. Dawid. Probability, causality and the empirical world: A Bayes–de Finetti–Popper–Borel synthesis. Statistical Science, 19(1):44–57, February 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1089808272.

Dawid:2014:DBA

[Daw14] A. P. Dawid. Discussion of “On the Birnbaum Argument for the Strong Likelihood Principle”. Statistical Science, 29 (2):240–241, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1408368574. See [May14a].

Dear:1992:AAP

[DB92] Keith B. G. Dear and Colin B. Begg. An approach for assessing publication bias prior to performing a meta-analysis. Statistical Science, 7(2):237–245, May 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011363.

Brabanter:2021:RRK

[DD21] Kris De Brabanter and Jos De Brabanter. Robustness by reweighting for kernel estimators: an overview. Statistical Science, 36(4):578–594, November 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-4/Robustness-by-Reweighting-
for-Kernel-Estimators-An-Overview/10.1214/20-STS816.
full.

DeVeaux:2020:CJS

[De 20] Richard D. De Veaux. A conversation with J. Stuart (Stu) Hunter. Statistical Science, 35(4):663–671, November 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-35/issue-4/
A-Conversation-with-J-Stuart-Stu-Hunter/10.1214/19-
STS766.full.

Diaconis:1986:SSA

[DE86] Persi Diaconis and Eduardo Engel. [Some statistical applications of Poisson’s work]: Comment. Statistical Science, 1 (2):171–174, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013691.

DiCiccio:1996:BCI

[DE96] Thomas J. DiCiccio and Bradley Efron. Bootstrap confidence intervals. Statistical Science, 11(3):189–228, August 1996. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1032280214.

Diaconis:2022:GRI

[DE22] Persi Diaconis and Stewart N. Ethier. Gambler’s Ruin and the ICM. Statistical Science, 37(3):289–305, August 2022. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-37/issue-3/Gamblers-Ruin-and-the-ICM/
10.1214/21-STS826.full.

DeGroot:1986:CCS

[DeG86a] Morris H. DeGroot. A conversation with Charles Stein. Statistical Science, 1(4):454–462, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013517.

DeGroot:1986:CDB

[DeG86b] Morris H. DeGroot. A conversation with David Blackwell. Statistical Science, 1(1):40–53, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013814.

DeGroot:1986:CEL

[DeG86c] Morris H. DeGroot. A conversation with Erich L. Lehmann. Statistical Science, 1(2):243–258, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013703.

DeGroot:1986:CPD

[DeG86d] Morris H. DeGroot. A conversation with Persi Diaconis. Statistical Science, 1(3):319–334, August 1986. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013610.

DeGroot:1986:CWA

[DeG86e] Morris H. DeGroot. A conversation with T. W. Anderson. Statistical Science, 1(1):97–105, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013823.

DeGroot:1986:FIA

[DeG86f] Morris H. DeGroot. The first issue of the annals of mathematical statistics. Statistical Science, 1(1):149–152, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013832.

DeGroot:1987:CCR

[DeG87a] Morris H. DeGroot. A conversation with C. R. Rao. Statistical Science, 2(1):53–67, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013438.

DeGroot:1987:CGB

[DeG87b] Morris H. DeGroot. A conversation with George Box. Statistical Science, 2(3):239–258, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013223.

DeGroot:1988:CGB

[DeG88] Morris H. DeGroot. A conversation with George A. Barnard. Statistical Science, 3(2):196–212, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012905.

deLeeuw:1992:MMA

[deL92] Jan deLeeuw. [Multilinear models: Applications in spectroscopy]: Comment. Statistical Science, 7(3):310–311, August 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011226.

Dudik:2014:DRP

[DELL14] Miroslav Dudík, Dumitru Erhan, John Langford, and Lihong Li. Doubly robust policy evaluation and optimization. Statistical Science, 29(4):485–511, November 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1421330544.

Dempster:1986:RMA

[Dem86] A. P. Dempster. [Regression models for adjusting the 1980 Census]: Comment. Statistical Science, 1(1):21–23, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013807. See [FN86a].

Deming:1988:GOC

[Dem88a] W. Edwards Deming. [Golden oldies: Classic articles from the world of statistics and probability]: Comment: Recollections about Harold Hotelling. Statistical Science, 3(1):103– 104, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013009.

Dempster:1988:EDSa

[Dem88b] Arthur P. Dempster. Employment discrimination and statistical science. Statistical Science, 3(2):149–161, May 1988. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012894.

Dempster:1988:EDSb

[Dem88c] Arthur P. Dempster. [Employment discrimination and statistical science]: Rejoinder. Statistical Science, 3(2):191–195, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012904.

Deming:1990:ABE

[Dem90a] W. Edwards Deming. [Applications in business and economic statistics: Some personal views]: Comment. Statistical Science, 5(4):391–392, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012010.

Dempster:1990:UDP

[Dem90b] A. P. Dempster. [The unity and diversity of probability]: Comment. Statistical Science, 5(4):449, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012025.

DeMets:1992:ETI

[DeM92] David L. DeMets. [Evaluating therapeutic interventions: Some issues and experiences]: Comment. Statistical Science, 7(4): 443–444, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011130.

Dempster:1993:BAE

[Dem93] A. P. Dempster. [Bayesian analysis in expert systems]: Comment: Assessing the science behind graphical modelling techniques. Statistical Science, 8(3):247–250, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010889.

Dempster:1998:LSM

[Dem98] A. P. Dempster. Logicist statistics. I. Models and modeling. Statistical Science, 13(3):248–276, August 1998. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1028905887.

Dellaportas:2012:JSM

[DFN12] Petros Dellaportas, Jonathan J. Forster, and Ioannis Ntzoufras. Joint specification of model space and parameter space prior distributions. Statistical Science, 27(2):232–246, May 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1340110871.

Dawid:1995:IBE

[DG95] A. P. Dawid and C. Goutis. [Inference based on estimating functions in the presence of nuisance parameters]: Comment. Statistical Science, 10(2):185–186, May 1995. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010033.

Datta:2012:SAS

[DG12] G. Datta and M. Ghosh. Small area shrinkage estimation. Statistical Science, 27(1):95–114, February 2012. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1331729985.

Dutta:2013:BMS

[DG13] Ritabrata Dutta and Jayanta K. Ghosh. Bayes model selection with path sampling: Factor models and other examples. Statistical Science, 28(1):95–115, February 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1359468410.

Diggle:2020:TTP

[DGA^{+}20] Peter J. Diggle, Emanuele Giorgi, Julienne Atsame, Sylvie Ntsame Ella, Kisito Ogoussan, and Katherine Gass. A tale of two parasites: Statistical modelling to support disease control programmes in Africa. Statistical Science, 35(1):42–50, February 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1583226028.

Dempster:1992:SMD

[DH92] Arthur P. Dempster and Jing-Shiang Hwang. [Statistical methods for data with long-range dependence]: Comment: Short- range consequences of long-range dependence. Statistical Science, 7(4):416–420, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011123.

DeVeaux:2005:HLB

[DH05] Richard D. De Veaux and David J. Hand. How to lie with bad data. Statistical Science, 20(3):231–238, August 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1124891289.

Drechsler:2024:YSD

[DH24] Jörg Drechsler and Anna-Carolina Haensch. 30 years of synthetic data. Statistical Science, 39(2):221–242, May 2024. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-39/issue-2/30-Years-of-Synthetic-Data/
10.1214/24-STS927.full.

Drovandi:2017:PED

[DHM^{+}17] Christopher C. Drovandi, Christopher C. Holmes, James M. McGree, Kerrie Mengersen, Sylvia Richardson, and Elizabeth G. Ryan. Principles of experimental design for big data analysis. Statistical Science, 32(3):385–404, August 2017. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/euclid.ss/1504253123.

Deng:2013:HAL

[DHR^{+}13] Yiting Deng, D. Sunshine Hillygus, Jerome P. Reiter, Yajuan Si, and Siyu Zheng. Handling attrition in longitudinal studies: The case for refreshment samples. Statistical Science, 28(2):238–256, May 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1369147914.

Dorie:2019:AVD

[DHS^{+}19a] Vincent Dorie, Jennifer Hill, Uri Shalit, Marc Scott, and Dan Cervone. Automated versus do-it-yourself methods for causal inference: Lessons learned from a data analysis competition. Statistical Science, 34(1):43–68, February 2019. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1555056030. See

comment [GvdL19].

Dorie:2019:RRD

[DHS^{+}19b] Vincent Dorie, Jennifer Hill, Uri Shalit, Marc Scott, and Dan Cervone. Rejoinder: Response to discussions and a look ahead. Statistical Science, 34(1):94–99, February 2019. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1555056037.

Davison:2003:RDB

[DHY03] A. C. Davison, D. V. Hinkley, and G. A. Young. Recent developments in bootstrap methodology. Statistical Science, 18 (2):141–157, May 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1063994969.

Diaconis:1991:RMA

[Dia91] Persi Diaconis. [Replication and meta-analysis in parapsychology]: Comment. Statistical Science, 6(4):386, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011580.

Ding:2017:PRB

[Din17a] Peng Ding. A paradox from randomization-based causal inference. Statistical Science, 32(3):331–345, August 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1504253116.

Ding:2017:RPR

[Din17b] Peng Ding. Rejoinder: A paradox from randomization-based causal inference. Statistical Science, 32(3):362–366, August 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/

1504253121. See [LRR17].

Donoho:2015:HCL

[DJ15] David Donoho and Jiashun Jin. Higher criticism for largescale inference, especially for rare and weak effects. Statistical Science, 30(1):1–25, February 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1425492437.

Deb:2024:NQR

[DJ24] Soudeep Deb and Kaushik Jana. Nonparametric quantile regression for time series with replicated observations and its application to climate data. Statistical Science, 39(3):428–448, August 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-3/
Nonparametric-Quantile-Regression-for-Time-Series-with-
Replicated-Observations-and/10.1214/23-STS918.full.

DuMouchel:2001:CID

[DJK^{+}01] William DuMouchel, Wen-Hua Ju, Alan F. Karr, Matthias Schonlau, Martin Theusan, and Yehuda Vardi. Computer intrusion: Detecting masquerades. Statistical Science, 16(1): 58–74, February 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/998929476.

Dinh:2020:SIE

[DJK^{+}20] Khanh N. Dinh, Roman Jaksik, Marek Kimmel, Amaury Lambert, and Simon Tavaré. Statistical inference for the evolutionary history of cancer genomes. Statistical Science, 35(1):129– 144, February 2020. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1583226033.

Dempster:1987:PES

[DK87] A. P. Dempster and Augustine Kong. [Probabilistic expert systems in medicine: Practical issues in handling uncertainty]: Comment. Statistical Science, 2(1):32–36, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013430. See [Spi87a].

Du:2020:SMS

[DK20] Chao Du and S. C. Kou. Statistical methodology in singlemolecule experiments. Statistical Science, 35(1):75–91, February 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1583226030.

Didelez:2010:GMI

[DKK10] Vanessa Didelez, Svend Kreiner, and Niels Keiding. Graphical models for inference under outcome-dependent sampling. Statistical Science, 25(3):368–387, August 2010. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1294167965.

Davison:2019:CMA

[DKK19] Anthony C. Davison, Erwan Koch, and Jonathan Koh. Comment: Models are approximations! Statistical Science, 34 (4):584–590, November 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1578474023. See [BBB^{+}19,

BBK^{+}19].

Dahlhaus:2018:RBT

[DKN18] Rainer Dahlhaus, István Z. Kiss, and Jan C. Neddermeyer. On the relationship between the theory of cointegration and the theory of phase synchronization. Statistical Science, 33(3):334– 357, August 2018. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1534147227.

Diaconis:2008:GSE

[DKSC08a] Persi Diaconis, Kshitij Khare, and Laurent Saloff-Coste. Gibbs sampling, exponential families and orthogonal polynomials. Statistical Science, 23(2):151–178, May 2008. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1219339107.  See

comments [JJ08, BCPR08].

Diaconis:2008:RGS

[DKSC08b] Persi Diaconis, Kshitij Khare, and Laurent Saloff-Coste. Rejoinder: Gibbs sampling, exponential families and orthogonal polynomials. Statistical Science, 23(2):196–200, May 2008. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1219339112.

Ding:2018:CIM

[DL18a] Peng Ding and Fan Li. Causal inference: A missing data perspective. Statistical Science, 33(2):214–237, May 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1525313143.

Durot:2018:LTM

[DL18b] Cécile Durot and Hendrik P. Lopuhaä. Limit theory in monotone function estimation. Statistical Science, 33(4):547–567, November 2018. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1543482058.

Denby:2006:CMN

[DLM06] Lorraine Denby, James M. Landwehr, and Jean Meloche. Comment: Monitoring networked applications with incremental quantile estimation. Statistical Science, 21(4):476–478, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.

ss/1177334521. See [CJLV06a].

Derkach:2014:PAT

[DLS14] Andriy Derkach, Jerry F. Lawless, and Lei Sun. Pooled association tests for rare genetic variants: A review and some new results. Statistical Science, 29(2):302–321, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1408368583.

Drum:1993:RMD

[DM93] Melinda Drum and Peter McCullagh. [Regression models for discrete longitudinal responses]: Comment. Statistical Science, 8(3):300–301, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010900.

Duan:2023:CRA

[DMJ23] Yunshan Duan, Peter Müller, and Yuan Ji. Comment: Response-adaptive randomization in clinical trials: From myths to practical considerations. Statistical Science, 38(2):212– 215, May 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-2/
Comment--Response-Adaptive-Randomization-in-Clinical-
Trials--From/10.1214/23-STS865B.full.

Diggle:2013:SST

[DMRT13] Peter J. Diggle, Paula Moraga, Barry Rowlingson, and Benjamin M. Taylor. Spatial and spatio-temporal log-Gaussian Cox processes: Extending the geostatistical paradigm. Statistical Science, 28(4):542–563, November 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1386078878.

Didelez:2010:AIM

[DMS10] Vanessa Didelez, Sha Meng, and Nuala A. Sheehan. Assumptions of IV methods for observational epidemiology. Statistical Science, 25(1):22–40, February 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1280841731.

Dellarocas:2006:SMP

[DN06] Chrysanthos Dellarocas and Ritu Narayan. A statistical measure of a Population’s propensity to engage in post-purchase online word-of-mouth. Statistical Science, 21(2):277–285, May 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1154979827.

Donnelly:1991:GSP

[Don91] Peter Donnelly. [The growth and stabilization of populations]: Comment. Statistical Science, 6(3):277–279, August 1991. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011696.

Doob:1986:CLT

[Doo86] J. L. Doob. [The central limit theorem around 1935]: Comment. Statistical Science, 1(1):93–94, February 1986. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013820.  See

[Le 86a].

Doss:2008:CQI

[Dos08] Hani Doss. Comment: Quantifying information loss in survival studies. Statistical Science, 23(3):313–317, August 2008. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1233153058.

delPino:1989:URIb

[dP89a] Guido del Pino. [The unifying role of iterative generalized least squares in statistical algorithms]: Rejoinder. Statistical Science, 4(4):407–408, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012412.

delPino:1989:URIa

[dP89b] Guido del Pino. The unifying role of iterative generalized least squares in statistical algorithms. Statistical Science, 4(4): 394–403, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012408.

Duncan:1991:EAMa

[DP91a] George T. Duncan and Robert W. Pearson. Enhancing access to microdata while protecting confidentiality: Prospects for the future. Statistical Science, 6(3):219–232, August 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011681.

Duncan:1991:EAMb

[DP91b] George T. Duncan and Robert W. Pearson. [Enhancing access to microdata while protecting confidentiality: Prospects for the future]: Rejoinder. Statistical Science, 6(3):237–239, August 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011685.

DeGruttola:1993:BHI

[DP93] Victor De Gruttola and Marcello Pagano. [Backcalculation of HIV infection rates]: Comment. Statistical Science, 8(2):109, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010998.

Drton:2007:MTE

[DP07] Mathias Drton and Michael D. Perlman. Multiple testing and error control in Gaussian graphical model selection. Statistical Science, 22(3):430–449, August 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1199285042.

Angelis:2014:EHB

[DPCA14] Daniela De Angelis, Anne M. Presanis, Stefano Conti, and A. E. Ades. Estimation of HIV burden through Bayesian evidence synthesis. Statistical Science, 29(1):9–17, February 2014. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1399645723.

Drovandi:2015:BII

[DPL15] Christopher C. Drovandi, Anthony N. Pettitt, and Anthony Lee. Bayesian indirect inference using a parametric auxiliary model. Statistical Science, 30(1):72–95, February 2015. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1425492441.

Dissanayake:2018:FDG

[DPP18] G. S. Dissanayake, M. S. Peiris, and T. Proietti. Fractionally differenced Gegenbauer processes with long memory: A review. Statistical Science, 33(3):413–426, August 2018. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1534147230.

Davison:2012:R

[DPR12a] A. C. Davison, S. A. Padoan, and M. Ribatet. Rejoinder. Statistical Science, 27(2):199–201, May 2012. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1340110869.

Davison:2012:SMS

[DPR12b] A. C. Davison, S. A. Padoan, and M. Ribatet. Statistical modeling of spatial extremes. Statistical Science, 27(2):161–186, May 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.

ss/1340110864. See discussion [CS12c, GTWT12, SR12].

Durante:2019:CCM

[DR19] Daniele Durante and Tommaso Rigon. Conditionally conjugate mean-field variational Bayes for logistic models. Statistical Science, 34(3):472–485, August 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1570780980.

Draper:1987:PFO

[Dra87] David Draper. [Prediction of future observations in growth curve models]: Comment: On exchangeability judgments in predictive modeling and the role of data in statistical research. Statistical Science, 2(4):454–461, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013122.

Draper:1988:RBRa

[Dra88a] David Draper. Rank-based robust analysis of linear models. I. Exposition and review. Statistical Science, 3(2):239–257, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012915.

Draper:1988:RBRb

[Dra88b] David Draper. [Rank-based robust analysis of linear models. I. Exposition and review]: Rejoinder. Statistical Science, 3 (2):266–271, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012921.

DiCiccio:2022:CIS

[DRRS22] Thomas J. DiCiccio, David M. Ritzwoller, Joseph P. Romano, and Azeem M. Shaikh. Confidence intervals for seroprevalence. Statistical Science, 37(3):306–321, August 2022. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-37/issue-3/Confidence-Intervals-for-
Seroprevalence/10.1214/21-STS844.full.

Duncan:1992:UGC

[DS92] Joseph W. Duncan and William C. Shelton. U.S. Government contributions to probability sampling and statistical analysis. Statistical Science, 7(3):320–338, August 1992. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011230.

Diamond:1994:CWR

[DS94] Ian Diamond and Chris Skinner. [Can we reach consensus on census adjustment?]: Comment. Statistical Science, 9(4): 508–510, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010262.

Davison:2008:BCS

[DS08] A. C. Davison and N. Sartori. The Banff challenge: Statistical detection of a noisy signal. Statistical Science, 23(3):354–364, August 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1233153063.

Delfino:2011:AVR

[DS11] Gustavo Delfino and Guillermo Salas. Analysis of the 2004 Venezuela Referendum: The official results versus the petition signatures. Statistical Science, 26(4):479–501, November 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1330437931.

Dudoit:2003:MHT

[DSB03] Sandrine Dudoit, Juliet Popper Shaffer, and Jennifer C. Boldrick. Multiple hypothesis testing in microarray experiments. Statistical Science, 18(1):71–103, February 2003. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1056397487.

Dekking:2013:E

[DSW13] Michael Dekking, Michel Stein, and Jon Wellner. Editorial. Statistical Science, 28(4):465, November 2013. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1386078873.

DiCiccio:2004:CDF

[DT04] Thomas J. DiCiccio and Mary E. Thompson. A conversation with Donald A. S. Fraser. Statistical Science, 19(2):370–386, May 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1105714168.

Davidian:2005:SET

[DTL05] Marie Davidian, Anastasios A. Tsiatis, and Selene Leon. Semiparametric estimation of treatment effect in a pretest–posttest study with missing data. Statistical Science, 20(3):261–301, August 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1124891293.

Dai:2017:CDS

[DTZ17] Wenlin Dai, Tiejun Tong, and Lixing Zhu. On the choice of difference sequence in a unified framework for variance estimation in nonparametric regression. Statistical Science, 32(3):455– 468, August 2017. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1504253126.

Dudley:1989:AEP

[Dud89] R. M. Dudley. [Asymptotics via empirical processes]: Comment. Statistical Science, 4(4):354, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012395.

Duffy:1992:SEI

[Duf92] Diana E. Duffy. [A survey of exact inference for contingency tables]: Comment. Statistical Science, 7(1):157–160, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011456.

Duffy:1993:ISC

[Duf93] Diane E. Duffy. [Is industrial statistics out of control?]: Comment. Statistical Science, 8(4):380–384, November 1993. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010775.

Duffield:2004:SPI

[Duf04] Nick Duffield. Sampling for passive Internet measurement: A review. Statistical Science, 19(3):472–498, August 2004. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1110999311.

DuMouchel:1988:MMQ

[DuM88a] William DuMouchel. [From mouse-to-man: The quantitative assessment of cancer risks]: Comment. Statistical Science, 3 (1):43–44, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012998.

DuMouchel:1988:SMF

[DuM88b] William DuMouchel. [Selection models and the file drawer problem]: Comment. Statistical Science, 3(1):132–133, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013018.

DuMouchel:2012:MBL

[DuM12a] William DuMouchel. Multivariate Bayesian logistic regression for analysis of clinical study safety issues. Statistical Science, 27 (3):319–339, August 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid. org/euclid.ss/1346849939. See discussion [MT12, Ber12].

DuMouchel:2012:R

[DuM12b] William DuMouchel. Rejoinder. Statistical Science, 27(3): 348–349, August 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1346849943.

Dunson:2017:TAP

[Dun17] David B. Dunson. Toward automated prior choice. Statistical Science, 32(1):41–43, February 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1491465625.

Dupont:1991:ESR

[Dup91] William D. Dupont. [Ethics and statistics in randomized clinical trials]: Comment. Statistical Science, 6(1):69–71, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011937.

Delaigle:2016:CPH

[DW16] Aurore Delaigle and Matt P. Wand. A conversation with Peter Hall. Statistical Science, 31(2):275–304, May 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1464105043.

Doerge:1997:SIS

[DWZ97] R. W. Doerge, B. S. Weir, and Z-B. Zeng. Statistical issues in the search for genes affecting quantitative traits in experimental populations. Statistical Science, 12(3):195–219, August 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1030037909.

Diaconis:1991:CFS

[DZ91] Persi Diaconis and Sandy Zabell. Closed form summation for classical distributions: Variations on a theme of De moivre. Statistical Science, 6(3):284–302, August 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011699.

Diaconis:2023:PSJ

[DZ23] Persi Diaconis and Sandy Zabell. In praise (and search) of J. V. Uspensky. Statistical Science, 38(1):160–183, February 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/journals/
statistical-science/volume-38/issue-1/In-Praise-and-
Search-of-J-V-Uspensky/10.1214/22-STS866.full.

Easterling:1989:DAC

[Eas89] Robert G. Easterling. [Design and analysis of computer experiments]: Comment. Statistical Science, 4(4):425–427, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012415.

Eaton:1987:TPH

[Eat87] Morris L. Eaton. [Testing precise hypotheses]: Comment. Statistical Science, 2(3):337–338, August 1987. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013240.

Eaton:2001:AI

[Eat01] Morris L. Eaton. About this issue. Statistical Science, 16 (4):311, November 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1015346316.

Eaton:2007:WHK

[Eat07] Morris L. Eaton. William H. Kruskal and the development of coordinate-free methods. Statistical Science, 22(2):264–265, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1190905523.

Eddy:1986:CSRa

[Edd86a] William F. Eddy. Computers in statistical research. Statistical Science, 1(4):419–437, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1177013508. See comments

[Utt86, BFK86, Sco86, Goe86, Bil86, Bat86, Weg86a, Edd86b].

Eddy:1986:CSRb

[Edd86b] William F. Eddy. [Computers in statistical research]: Rejoinder. Statistical Science, 1(4):451–453, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013516.  See

[Edd86a].

Eddy:1987:DGD

[Edd87] William F. Eddy. [Dynamic graphics for data analysis]: Comment. Statistical Science, 2(4):386–387, November 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013107.

Edwards:1997:TEP

[Edw97a] A. W. F. Edwards. Three early papers on efficient parametric estimation. Statistical Science, 12(1):35–47, February 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1029963260.

Edwards:1997:WDF

[Edw97b] A. W. F. Edwards. What did Fisher mean by “inverse probability” in 1912–1922? Statistical Science, 12(3):177–184, August 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1030037907.

Epstein:1992:SEI

[EF92] Leonardo D. Epstein and Stephen E. Fienberg. [A survey of exact inference for contingency tables]: Comment. Statistical Science, 7(1):160–162, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011457.

Ericksen:1994:CWR

[EFK94] Eugene P. Ericksen, Stephen E. Fienberg, and Joseph B. Kadane. [Can we reach consensus on census adjustment?]: Comment. Statistical Science, 9(4):511–515, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010263.

Efron:1990:UDP

[Efr90] B. Efron. [The unity and diversity of probability]: Comment. Statistical Science, 5(4):450, November 1990. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012026.

Efron:1994:BMT

[Efr94] B. Efron. [Bootstrap: More than a stab in the dark?]: Comment. Statistical Science, 9(3):396–398, August 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010385.

Efron:1998:RFC

[Efr98] Bradley Efron. R. A. Fisher in the 21st century (invited paper presented at the 1996 R. A. Fisher Lecture). Statistical Science, 13(2):95–122, May 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1028905930.

Efron:2003:STB

[Efr03] Bradley Efron. Second thoughts on the bootstrap. Statistical Science, 18(2):135–140, May 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1063994968.

Efron:2008:MEB

[Efr08a] Bradley Efron. Microarrays, empirical Bayes and the twogroups model. Statistical Science, 23(1):1–22, February 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1215441276. See comments [Ben08, Cai08, Mor08, RS08a, Efr08b].

Efron:2008:RME

[Efr08b] Bradley Efron. Rejoinder: Microarrays, empirical Bayes and the two-groups model. Statistical Science, 23(1):45–47, February 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1215441281. See [GM05a].

Efron:2010:FIE

[Efr10a] Bradley Efron. The future of indirect evidence. Statistical Science, 25(2):145–157, May 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1290175835.

Efron:2010:RFI

[Efr10b] Bradley Efron. Rejoinder: The future of indirect evidence. Statistical Science, 25(2):170–171, May 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1290175839.

Efron:2014:TMS

[Efr14] Bradley Efron. Two modeling strategies for empirical Bayes estimation. Statistical Science, 29(2):285–301, May 2014. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1408368582.

Efron:2019:BOB

[Efr19a] Bradley Efron. Bayes, oracle Bayes and empirical Bayes. Statistical Science, 34(2):177–201, May 2019. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1563501631. See comments

[Lou19, Lai19, vdV19, Jia19, GR19, WMB19].

Efron:2019:RBO

[Efr19b] Bradley Efron. Rejoinder: Bayes, oracle Bayes, and empirical Bayes. Statistical Science, 34(2):234–235, May 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1563501639.

Edelmann:2022:RPG

[EG22] Dominic Edelmann and Jelle Goeman. A regression perspective on generalized distance covariance and the Hilbert–Schmidt independence criterion. Statistical Science, 37(4):562–579, November 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-4/
A-Regression-Perspective-on-Generalized-Distance-Covariance-
and-the-HilbertSchmidt/10.1214/21-STS841.full.

Ellenberg:1997:CNS

[EGG97] Jonas H. Ellenberg, Mitchell H. Gail, and Nancy L. Geller. Conversations with NIH statisticians: interviews with the pioneers of biostatistics at the United States National Institute of Health. Statistical Science, 12(2):77–81, May 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1029963422.

Ernst:2003:UQF

[EH03] Michael D. Ernst and Alan D. Hutson. Utilizing a quantile function approach to obtain exact bootstrap solutions. Statistical Science, 18(2):231–240, May 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1063994978.

Elmore:2004:SSO

[EHX04] R. T. Elmore, T. P. Hettmansperger, and F. Xuan. The sign statistic, one-way layouts and mixture models. Statistical Science, 19(4):579–587, November 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1113832722.

Evans:2011:WII

[EJ11] Michael Evans and Gun Ho Jang. Weak informativity and the information in one prior relative to another. Statistical Science, 26(3):423–439, August 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1320066929.

Ellenberg:1992:ETI

[Ell92] Susan S. Ellenberg. [Evaluating therapeutic interventions: Some issues and experiences]: Comment. Statistical Science, 7(4):445–446, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011131.

Ellenberg:1997:CMK

[Ell97a] Jonas H. Ellenberg. A conversation with Morton Kramer. Statistical Science, 12(2):103–107, May 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1029963426.

Ellenberg:1997:CJC

[Ell97b] Susan S. Ellenberg. A conversation with John C. Bailar III. Statistical Science, 12(2):119–124, May 1997. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1029963429.

Ellis:1998:ILS

[Ell98] Steven P. Ellis. Instability of least squares, least absolute deviation and least median of squares linear regression, with a comment by Stephen Portnoy and Ivan Mizera and a rejoinder by the author. Statistical Science, 13(4):337–350, November 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1028905829.

Eilers:1996:FSB

[EM96] Paul H. C. Eilers and Brian D. Marx. Flexible smoothing with B-splines and penalties. Statistical Science, 11(2):89–121, May 1996. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1038425655.

Escobar:2006:RAT

[EM06] Luis A. Escobar and William Q. Meeker. A review of accelerated test models. Statistical Science, 21(4):552–577, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177334529.

Elvira:2019:GMI

[EMLB19] Víctor Elvira, Luca Martino, David Luengo, and Mónica F. Bugallo. Generalized multiple importance sampling. Statistical Science, 34(1):129–155, February 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1555056039.

Ericksen:1986:RMA

[Eri86] Eugene P. Ericksen. [Regression models for adjusting the 1980 Census]: Comment. Statistical Science, 1(1):18–21, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013806. See [FN86a].

Ernst:2004:PMB

[Ern04] Michael D. Ernst. Permutation methods: A basis for exact inference. Statistical Science, 19(4):676–685, November 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1113832732.

Evans:1995:MAI

[ES95] Michael Evans and Tim Swartz. Methods for approximating integrals in statistics with special emphasis on Bayesian integration problems. Statistical Science, 10(3):254–272, August 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177009938. See comments [Ano96].

Ebrahimi:2010:SIA

[ESS10] Nader Ebrahimi, Ehsan S. Soofi, and Refik Soyer. On the sample information about parameter and prediction. Statistical Science, 25(3):348–367, August 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1294167964.

Esterby:2003:CAH

[Est03] Sylvia R. Esterby. A conversation with Abdel H. El-Shaarawi. Statistical Science, 18(4):481–488, November 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1081443231.

Efron:1986:BMSa

[ET86a] B. Efron and R. Tibshirani. Bootstrap methods for standard errors, confidence intervals, and other measures of statistical accuracy. Statistical Science, 1(1):54–75, February 1986. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013815.

Efron:1986:BMSb

[ET86b] B. Efron and R. Tibshirani. [Bootstrap methods for standard errors, confidence intervals, and other measures of statistical accuracy]: Rejoinder. Statistical Science, 1(1):77, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013817.

Eubank:1988:MRS

[Eub88] Randy Eubank. [Monotone regression splines in action]: Comment. Statistical Science, 3(4):446–450, November 1988. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012763.

Eudey:1996:SCD

[Eud96] T. Lynn Eudey. Statistical considerations in DNA flow cytometry. Statistical Science, 11(4):320–334, November 1996. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1032280305.

Elliott:2017:INS

[EV17] Michael R. Elliott and Richard Valliant. Inference for nonprobability samples. Statistical Science, 32(2):249–264, May 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1494489814.

Evans:2007:CBC

[Eva07] M. Evans. Comment: Bayesian checking of the second levels of hierarchical models. Statistical Science, 22(3):344–348, August 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1199285032. See [BC07a].

Evans:2012:AMP

[Eva12] Stephen Evans. An answer to multiple problems with analysis of data on harms? Statistical Science, 27(3):346–347, August 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1346849942.

Evans:2014:DBA

[Eva14] Michael Evans. Discussion of “On the Birnbaum Argument for the Strong Likelihood Principle”. Statistical Science, 29 (2):242–246, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1408368575. See [May14a].

Everitt:1987:SP

[Eve87a] B. S. Everitt. Statistics in psychiatry. Statistical Science, 2 (2):107–116, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid. org/euclid.ss/1177013333. See comments [Gut87, Gre87b, Fle87, Zub87, MA87, Gre87a, Tur87, Ver87, Eve87b].

Everitt:1987:SPR

[Eve87b] B. S. Everitt. [Statistics in psychiatry]: Rejoinder. Statistical Science, 2(2):134, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177013342.

Evett:1991:IUD

[Eve91] Ian Evett. [Inferences using DNA profiling in forensic identification and paternity cases]: Comment. Statistical Science, 6 (2):200–202, May 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011826.

Etz:2017:JBH

[EW17] Alexander Etz and Eric-Jan Wagenmakers. J. B. S. Haldane’s contribution to the Bayes factor hypothesis test. Statistical Science, 32(2):313–329, May 2017. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1494489818.

Fairley:1989:EVA

[Fai89] David Fairley. [Extreme value analysis of environmental time series: An application to trend detection in ground-level ozone]: Comment. Statistical Science, 4(4):381–383, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012402.

Fan:2005:RSO

[Fan05a] Jianqing Fan. Rejoinder: A selective overview of nonparametric methods in financial econometrics. Statistical Science, 20(4): 351–357, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1137076651. See [Fan05b].

Fan:2005:SON

[Fan05b] Jianqing Fan. A selective overview of nonparametric methods in financial econometrics. Statistical Science, 20(4):317– 337, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1137076647. See comments [PY05, Sør05,

MZ05, Fan05a].

Fay:1994:MII

[Fay94] Robert E. Fay. [Multiple-imputation inferences with uncongenial sources of input]: Comment. Statistical Science, 9(4): 558–560, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010270.

Fearnhead:2018:PDM

[FBPR18] Paul Fearnhead, Joris Bierkens, Murray Pollock, and Gareth O. Roberts. Piecewise deterministic Markov processes for continuous-time Monte Carlo. Statistical Science, 33(3):386– 412, August 2018. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1534147229.

Fraser:2016:BRQ

[FBW^{+}16] D. A. S. Fraser, M. Bédard, A. Wong, Wei Lin, and A. M. Fraser. Bayes, reproducibility and the quest for truth. Statistical Science, 31(4):578–590, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1484816587.

Farewell:1992:ETI

[FC92] Vern T. Farewell and Richard J. Cook. [Evaluating therapeutic interventions: Some issues and experiences]: Comment. Statistical Science, 7(4):446–448, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011132.

Fuentes:2011:DER

[FC11] Claudio Fuentes and George Casella. Discussion of “Estimating Random Effects via Adjustment for Density Maximization” by C. Morris and R. Tang. Statistical Science, 26(2):288–290, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1312204021. See [MT11].

Feigenbaum:1993:PAD

[Fei93] Joan Feigenbaum. Probabilistic algorithms for defeating adversaries. Statistical Science, 8(1):26–30, February 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011080.

Fellegi:1986:RMA

[Fel86] I. P. Fellegi. [Regression models for adjusting the 1980 Census]: Comment. Statistical Science, 1(1):30–33, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013810. See [FN86a].

Felsenstein:1987:SAH

[Fel87] Joseph Felsenstein. [Statistical analysis of hominoid molecular evolution]: Comment. Statistical Science, 2(2):208–209, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013355.

Fellman:1999:GEC

[Fel99] J. Fellman. Gustav Elfving’s contribution to the emergence of the optimal experimental design theory. Statistical Science, 14 (2):197–200, May 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009212245.

Feng:2016:DPE

[Fen16] Shui Feng. Diffusion processes and the Ewens sampling formula. Statistical Science, 31(1):20–22, February 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1455115907.

Ferguson:1989:WSSa

[Fer89a] Thomas S. Ferguson. Who solved the Secretary Problem? Statistical Science, 4(3):282–289, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012493.

Ferguson:1989:WSSb

[Fer89b] Thomas S. Ferguson. [Who solved the Secretary Problem?]: Rejoinder. Statistical Science, 4(3):294–296, August 1989. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012498.

Fernholz:2003:RJW

[Fer03] Luisa Turrin Fernholz. Remembering John W. Tukey. Statistical Science, 18(3):336–340, August 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1076102421.

Freuli:2023:RSU

[FHH23] Francesca Freuli, Leonhard Held, and Rachel Heyard. Replication success under questionable research practices-a simulation study. Statistical Science, 38(4):621–639, November 2023. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-4/Replication-Success-Under-
Questionable-Research-Practicesa-Simulation-Study/10.
1214/23-STS904.full.

Flegal:2008:MCM

[FHJ08] James M. Flegal, Murali Haran, and Galin L. Jones. Markov chain Monte Carlo: Can we trust the third significant figure? Statistical Science, 23(2):250–260, May 2008. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1219339116.

Feuerverger:2012:SSN

[FHK12] Andrey Feuerverger, Yu He, and Shashi Khatri. Statistical significance of the Netflix challenge. Statistical Science, 27 (2):202–231, May 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1340110870.

Ferkingstad:2015:MCN

[FHS15] Egil Ferkingstad, Lars Holden, and Geir Kjetil Sandve. Monte Carlo null models for genomic data. Statistical Science, 30 (1):59–71, February 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1425492440.

Flynn:2017:SLN

[FHS17] Cheryl J. Flynn, Clifford M. Hurvich, and Jeffrey S. Simonoff. On the sensitivity of the lasso to the number of predictor variables. Statistical Science, 32(1):88–105, February 2017. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1491465629.

Fienberg:1988:EDS

[Fie88] Stephen E. Fienberg. [Employment discrimination and statistical science]: Comment. Statistical Science, 3(2):190–191, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012903.

Fienberg:1992:BHS

[Fie92] Stephen E. Fienberg. A brief history of statistics in three and one-half chapters: A review essay. Statistical Science, 7 (2):208–225, May 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011360.

Fienberg:1994:CJN

[Fie94] Stephen E. Fienberg. A conversation with Janet L. Norwood. Statistical Science, 9(4):574–590, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010275.

Fienberg:1997:IRF

[Fie97] Stephen E. Fienberg. Introduction to R. A. Fisher on inverse probability and likelihood. Statistical Science, 12(3):161, August 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1030037905.

Fienberg:2006:CCC

[Fie06a] Stephen E. Fienberg. Comment: Complex causal questions require careful model formulation: Discussion of Rubin on experiments with “censoring” due to death. Statistical Science, 21 (3):317–318, August 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1166642433. See [Rub06a].

Fienberg:2006:PCC

[Fie06b] Stephen E. Fienberg. Privacy and confidentiality in an ecommerce world: Data mining, data warehousing, matching and disclosure limitation. Statistical Science, 21(2):143–154, May 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1154979817.

Fienberg:2007:WKM

[Fie07] Stephen E. Fienberg. William Kruskal: My scholarly and scientific model. Statistical Science, 22(2):266–268, May 2007. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1190905524.

Fienberg:2008:ESY

[Fie08] Stephen E. Fienberg. The early statistical years: 1947–1967. a conversation with Howard Raiffa. Statistical Science, 23(1): 136–149, February 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1215441289.

Fienberg:2011:BMM

[Fie11a] Stephen E. Fienberg. Bayesian models and methods in public policy and government settings. Statistical Science, 26 (2):212–226, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid. org/euclid.ss/1312204010. See discussion [Han11a, Kal11].

Fienberg:2011:R

[Fie11b] Stephen E. Fienberg. Rejoinder. Statistical Science, 26(2): 238–239, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1312204014.

Fine:1986:ASP

[Fin86] Terrence L. Fine. [The axioms of subjective probability]: Comment. Statistical Science, 1(3):352–354, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013615. See [Fis86a].

Fishburn:1986:ASPa

[Fis86a] Peter C. Fishburn. The axioms of subjective probability. Statistical Science, 1(3):335–345, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1177013611. See comments

[Sup86, Ber86, Fin86, Sei86, Sto86b, Sud86, Fis86b].

Fishburn:1986:ASPb

[Fis86b] Peter C. Fishburn. The axioms of subjective probability: Rejoinder. Statistical Science, 1(3):358, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013619.  See

[Fis86a].

Fishburn:1986:SRC

[Fis86c] Peter C. Fishburn. [Savage revisited]: Comment. Statistical Science, 1(4):492–495, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013521.

Fisher:1988:EDS

[Fis88] Franklin M. Fisher. [Employment discrimination and statistical science]: Comment. Statistical Science, 3(2):161–165, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012895.

Fisher:1990:OOM

[Fis90] N. I. Fisher. [OMEGA (online multivariate exploratory graphical analysis): Routine searching for structure]: Comment. Statistical Science, 5(2):213–215, May 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012169.

Fisher:2015:CJF

[Fis15] N. I. Fisher. A conversation with Jerry Friedman. Statistical Science, 30(2):268–295, May 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1433341483.

Favaro:2016:RES

[FJ16] Stefano Favaro and Lancelot F. James. Relatives of the Ewens sampling formula in Bayesian nonparametrics. Statistical Science, 31(1):30–33, February 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1455115910.

Fusaro:1989:ABQ

[FJH^{+}89] R. E. Fusaro, N. P. Jewell, W. W. Hauck, D. C. Heilbron, J. D. Kalbfleisch, and J. M. Neuhaus. An annotated bibliography of quantitative methodology relating to the AIDS epidemic. Statistical Science, 4(3):264–281, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012492.

Freeman:2019:CMA

[FJL^{+}19] Nikki L. B. Freeman, Xiaotong Jiang, Owen E. Leete, Daniel J. Luckett, Teeranan Pokaprakarn, and Michael R. Kosorok. Comment: Models as approximations. Statistical Science, 34(4):572–574, November 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1578474020. See [BBB^{+}19,

BBK^{+}19].

Frankel:1996:CLK

[FK96] Martin Frankel and Benjamin King. A conversation with Leslie Kish. Statistical Science, 11(1):65–87, February 1996. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1032209665.

Fokianos:2003:RTC

[FK03] Konstantinos Fokianos and Benjamin Kedem. Regression theory for categorical time series. Statistical Science, 18(3):357– 376, August 2003. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1076102425.

Feigenbaum:1993:PAS

[FL93] Joan Feigenbaum and Jeffrey C. Lagarias. Probabilistic algorithms for speedup. Statistical Science, 8(1):20–25, February 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011079.

Faulkner:2024:SAS

[FL24] Michael F. Faulkner and Samuel Livingstone. Sampling algorithms in statistical physics: a guide for statistics and machine learning. Statistical Science, 39(1):137–164, February 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-1/
Sampling-Algorithms-in-Statistical-Physics--A-Guide-
for-Statistics/10.1214/23-STS893.full.

Fleiss:1987:SPC

[Fle87] Joseph L. Fleiss. [Statistics in psychiatry]: Comment. Statistical Science, 2(2):120–121, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013336. See [Eve87a].

Fleming:1992:ETIa

[Fle92a] Thomas R. Fleming. Evaluating therapeutic interventions: Some issues and experiences. Statistical Science, 7(4):428– 441, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011128.

Fleming:1992:ETIb

[Fle92b] Thomas R. Fleming. [Evaluating therapeutic interventions: Some issues and experiences]: Rejoinder. Statistical Science, 7(4):452–456, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011135.

Flournoy:1989:HGY

[Flo89] N. Flournoy. [How to get your first research grant]: Comment. Statistical Science, 4(2):139–141, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012595.

Flournoy:1993:CWJ

[Flo93] Nancy Flournoy. A conversation with Wilfrid J. Dixon. Statistical Science, 8(4):458–477, November 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177010788.

Fitzmaurice:1993:RMDa

[FLR93a] Garrett M. Fitzmaurice, Nan M. Laird, and Andrea G. Rotnitzky. Regression models for discrete longitudinal responses. Statistical Science, 8(3):284–299, August 1993. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010899.

Fitzmaurice:1993:RMDb

[FLR93b] Garrett M. Fitzmaurice, Nan M. Laird, and Andrea G. Rotnitzky. [Regression models for discrete longitudinal responses]: Rejoinder. Statistical Science, 8(3):306–309, August 1993. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010903.

Fluhler:1990:BBC

[Flu90] H. Fluhler. [Biostatistics and Bayes]: Comment. Statistical Science, 5(3):286–287, August 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012094.

Fan:1993:LRA

[FM93] J. Fan and J. S. Marron. [Local regression: Automatic kernel carpentry]: Comment. Statistical Science, 8(2):129–134, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011003.

Fernholz:2000:CJW

[FM00] Luisa T. Fernholz and Stephan Morgenthaler. A conversation with John W. Tukey and Elizabeth Tukey. Statistical Science, 15(1):79–94, February 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1009212675;   http://
www.jstor.org/stable/2676678.

Fernholz:2003:CJT

[FM03] Luisa Turrin Fernholz and Stephan Morgenthaler. A conversation with John Tukey. Statistical Science, 18(3):346–356, August 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1076102424.

Fithian:2018:FLR

[FM18] William Fithian and Rahul Mazumder. Flexible low-rank statistical modeling with missing data and side information. Statistical Science, 33(2):238–260, May 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1525313144.

Friendly:2013:EIU

[FMF13] Michael Friendly, Georges Monette, and John Fox. Elliptical insights: Understanding statistical methods through elliptical geometry. Statistical Science, 28(1):1–39, February 2013. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1359468407.

Fan:2021:SOD

[FMZ21] Jianqing Fan, Cong Ma, and Yiqiao Zhong. A selective overview of deep learning. Statistical Science, 36(2):264– 290, May 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-36/issue-2/
A-Selective-Overview-of-Deep-Learning/10.1214/20-STS783.
full.

Freedman:1986:RMAa

[FN86a] D. A. Freedman and W. C. Navidi. Regression models for adjusting the 1980 Census. Statistical Science, 1(1):3–11, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.

ss/1177013804. See comments [Kad86, Eri86, Dem86, Wol86, Mad86, Fel86, Mos86, Nat86, FN86b].

Freedman:1986:RMAb

[FN86b] D. A. Freedman and W. C. Navidi. [Regression models for adjusting the 1980 Census]: Rejoinder. Statistical Science, 1 (1):35–39, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013813. See [FN86a].

Fenton:2006:CEE

[FN06] Norman Fenton and Martin Neil. Comment: Expert elicitation for reliable system design. Statistical Science, 21(4): 451–453, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177334516. See [BQW06a].

Folks:1995:COK

[Fol95] J. Leroy Folks. A conversation with Oscar Kempthorne. Statistical Science, 10(4):321–336, November 1995. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177009867.

Findley:1995:CHA

[FP95] David F. Findley and Emanuel Parzen. A conversation with Hirotugu Akaike. Statistical Science, 10(1):104–117, February 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010133.

Fay:2022:IPV

[FPBT22] Michael P. Fay, Michael A. Proschan, Erica H. Brittain, and Ram Tiwari. Interpreting p-values and confidence intervals using well-calibrated null preference priors. Statistical Science, 37(4):455–472, November 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-37/issue-4/Interpreting-p-Values-and-
Confidence-Intervals-Using-Well-Calibrated-Null/10.
1214/21-STS833.full.

Finucane:2014:BEP

[FPDE14] Mariel M. Finucane, Christopher J. Paciorek, Goodarz Danaei, and Majid Ezzati. Bayesian estimation of population-level trends in measures of health status. Statistical Science, 29 (1):18–25, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1399645724.

Fasiolo:2016:CIM

[FPW16] Matteo Fasiolo, Natalya Pya, and Simon N. Wood. A comparison of inferential methods for highly nonlinear state space models in ecology and epidemiology. Statistical Science, 31(1): 96–118, February 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1455115916.

Fraser:2004:ACI

[Fra04] D. A. S. Fraser. Ancillaries and conditional inference. Statistical Science, 19(2):333–369, May 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1105714167.

Fraser:2011:BPJ

[Fra11a] D. A. S. Fraser. Is Bayes posterior just quick and dirty confidence? Statistical Science, 26(3):299–316, August 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1320066918. See discussion [Rob11, SX11, Zha11b].

Fraser:2011:R

[Fra11b] D. A. S. Fraser. Rejoinder. Statistical Science, 26(3):329–331, August 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1320066923.

Fraser:2014:DAC

[Fra14] D. A. S. Fraser. Discussion: On arguments concerning statistical principles. Statistical Science, 29(2):252–253, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1408368577.

French:1986:CPD

[Fre86] Simon French. [Combining probability distributions: A critique and an annotated bibliography]: Comment. Statistical Science, 1(1):138, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013827. See [GZ86a].

Freedman:1987:UPA

[Fre87] David Freedman. [Uncertainty, policy analysis and statistics]: Comment. Statistical Science, 2(3):276–277, August 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013225.

Freeman:1989:WSS

[Fre89] Peter R. Freeman. [Who solved the Secretary Problem?]: Comment. Statistical Science, 4(3):294, August 1989. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012497.

French:1990:WML

[Fre90] Simon French. [The 1988 Wald memorial lectures: The present position in Bayesian statistics]: Comment. Statistical Science, 5(1):78–80, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012258.

Freedman:1999:ACS

[Fre99] David Freedman. From association to causation: some remarks on the history of statistics. Statistical Science, 14(3):243–258, August 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1009212409.

Freedman:2008:RDJ

[Fre08] David A. Freedman. Randomization does not justify logistic regression. Statistical Science, 23(2):237–249, May 2008. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1219339115.

Frigessi:1995:BCS

[Fri95] Arnoldo Frigessi. [Bayesian computation and stochastic systems]: Comment. Statistical Science, 10(1):41–43, February 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010124.

Friedman:2006:CCT

[Fri06] Jerome H. Friedman. Comment: Classifier technology and the illusion of progress. Statistical Science, 21(1):15–18, February 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.

ss/1149600840. See comments [Han06a].

Friendly:2007:MGM

[Fri07] Michael Friendly. A.-M. Guerry’s moral statistics of France: Challenges for multivariable spatial analysis. Statistical Science, 22(3):368–399, August 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1199285037.

Friendly:2008:GAS

[Fri08] Michael Friendly. The Golden age of statistical graphics. Statistical Science, 23(4):502–535, November 2008. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1242049392.

Fienberg:2005:SML

[FS05] Stephen E. Fienberg and Paul C. Stern. In search of the magic lasso: The truth about the polygraph. Statistical Science, 20 (3):249–260, August 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1124891292.

Fienberg:2014:DED

[FS14] Stephen E. Fienberg and Rebecca C. Steorts. Discussion of “Estimating the Distribution of Dietary Consumption Patterns”. Statistical Science, 29(1):95–96, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1399645733. See [Car14a].

Fewster:2016:TCM

[FSB16] R. M. Fewster, B. C. Stevenson, and D. L. Borchers. Tracecontrast models for capture–recapture without capture histories. Statistical Science, 31(2):245–258, May 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1464105041.

Fienberg:2007:WKL

[FST07] Stephen E. Fienberg, Stephen M. Stigler, and Judith M. Tanur. The William Kruskal legacy: 1919–2005. Statistical Science, 22 (2):255–261, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1190905521.

Flury:1996:SCF

[FT96] Bernard Flury and Thaddeus Tarpey. Self-consistency: a fundamental concept in statistics. Statistical Science, 11(3):229–243, August 1996. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1032280215.

Favaro:2013:MNR

[FT13] Stefano Favaro and Yee Whye Teh. MCMC for normalized random measure mixture models. Statistical Science, 28(3): 335–359, August 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1377696940.

Fisher:2008:RWH

[FvZ08] Nicholas I. Fisher and Willem R. van Zwet. Remembering Wassily Hoeffding. Statistical Science, 23(4):536–547, November 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1242049393.

Freedman:1994:CWR

[FW94a] D. Freedman and K. Wachter. [Can we reach consensus on census adjustment?]: Rejoinder. Statistical Science, 9(4):527– 537, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010268.

Freedman:1994:HCA

[FW94b] D. Freedman and K. Wachter. Heterogeneity and census adjustment for the intercensal base. Statistical Science, 9(4): 476–485, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010260.

Fourdrinier:2012:ILE

[FW12] Dominique Fourdrinier and Martin T. Wells. On improved loss estimation for shrinkage estimators. Statistical Science, 27(1): 61–81, February 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1331729983.

Fang:1994:SAN

[FWB94] Kai-Tai Fang, Yuan Wang, and Peter M. Bentler. Some applications of number-theoretic methods in statistics. Statistical Science, 9(3):416–428, August 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177010392.

Fan:2021:RHD

[FWZZ21] Jianqing Fan, Kaizheng Wang, Yiqiao Zhong, and Ziwei Zhu. Robust high-dimensional factor models with applications to statistical machine learning. Statistical Science, 36(2):303– 327, May 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-36/issue-2/
Robust-High-Dimensional-Factor-Models-with-Applications-
to-Statistical-Machine/10.1214/20-STS785.full.

Fan:2010:EAR

[FYL10] Xiaodan Fan, Yuan Yuan, and Jun S. Liu. The EM algorithm and the rise of computational biology. Statistical Science, 25(4): 476–491, November 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1300108232.

Freedman:1988:MMQa

[FZ88a] D. A. Freedman and H. Zeisel. From mouse-to-man: The quantitative assessment of cancer risks. Statistical Science, 3 (1):3–28, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012993.

Freedman:1988:MMQb

[FZ88b] D. A. Freedman and H. Zeisel. [From mouse-to-man: The quantitative assessment of cancer risks]: Rejoinder. Statistical Science, 3(1):45–56, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012999.

Gail:1997:CNM

[Gai97] Mitchell H. Gail. A conversation with Nathan Mantel. Statistical Science, 12(2):88–97, May 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1029963424.

Galton:1989:KC

[Gal89] Francis Galton. Kinship and correlation. Statistical Science, 4(2):81–86, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012581.

Gastwirth:1987:SPMa

[Gas87a] Joseph L. Gastwirth. The statistical precision of medical screening procedures: Application to polygraph and AIDS antibodies test data. Statistical Science, 2(3):213–222, August 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013215.

Gastwirth:1987:SPMb

[Gas87b] Joseph L. Gastwirth. [The statistical precision of medical screening procedures: Application to polygraph and AIDS antibodies test data]: Rejoinder. Statistical Science, 2(3):234–238, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013222.

Gastwirth:1988:EDS

[Gas88] Joseph L. Gastwirth. [Employment discrimination and statistical science]: Comment. Statistical Science, 3(2):175–183, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012899.

Gastwirth:1993:CST

[Gas93] Joseph L. Gastwirth. [Can statistics tell us what we do not want to hear? the case of complex salary structures]: Comment. Statistical Science, 8(2):165–171, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011008.

Gayler:2006:CCT

[Gay06] Ross W. Gayler. Comment: Classifier technology and the illusion of progress — credit scoring. Statistical Science, 21(1): 19–23, February 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1149600841. See comments [Han06a].

Genest:2010:CMB

[GB10] Christian Genest and Gordon Brackstone. A conversation with Martin Bradbury Wilk. Statistical Science, 25(2):258–273, May 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1290175846.

Gelfand:1995:BCS

[GC95] Alan E. Gelfand and Bradley P. Carlin. [Bayesian computation and stochastic systems]: Comment. Statistical Science, 10 (1):43–46, February 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010125.

Gramsbsch:1994:BMT

[GCL94] Patricia M. Gramsbsch, Mary Kathryn Cowles, and Thomas A. Louis. [Bootstrap: More than a stab in the dark?]: Comment. Statistical Science, 9(3):398–400, August 1994. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010386.

Geisser:1986:OSB

[Gei86] Seymour Geisser. Opera selecta boxi. Statistical Science, 1(1): 106–113, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013824.

Geisser:1987:PFO

[Gei87a] Seymour Geisser. [Prediction of future observations in growth curve models]: Comment. Statistical Science, 2(4):465–467, November 1987. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177013125.

Geisser:1987:SPM

[Gei87b] Seymour Geisser. [The statistical precision of medical screening procedures: Application to polygraph and AIDS antibodies test data]: Comment. Statistical Science, 2(3):231–232, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013220.

Geisser:1987:UPA

[Gei87c] Seymour Geisser. [Uncertainty, policy analysis and statistics]: Comment. Statistical Science, 2(3):277–279, August 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013226.

Gelfand:1992:PMC

[Gel92] Alan E. Gelfand. [Practical Markov chain Monte Carlo]: Comment. Statistical Science, 7(4):486–487, November 1992. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011139.

Geller:1997:CTG

[Gel97] Nancy L. Geller. A conversation with Tavia Gordon. Statistical Science, 12(2):113–118, May 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1029963428.

Gelman:2005:CFB

[Gel05] Andrew Gelman. Comment: Fuzzy and Bayesian p-values and u-values. Statistical Science, 20(4):380–381, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1137076656.

Gelman:2007:CBC

[Gel07a] Andrew Gelman. Comment: Bayesian checking of the second levels of hierarchical models. Statistical Science, 22(3):349–352, August 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1199285033. See [BC07a].

Gelman:2007:RSS

[Gel07b] Andrew Gelman. Rejoinder: Struggles with survey weighting and regression modeling. Statistical Science, 22(2):184–188, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1190905517. See [Gel07c].

Gelman:2007:SSW

[Gel07c] Andrew Gelman. Struggles with survey weighting and regression modeling. Statistical Science, 22(2):153–164, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1190905511. See comments [BC07c, BO07, Lit07, Loh07, Pfe07, Gel07b].

Gelman:2009:BJP

[Gel09] Andrew Gelman. Bayes, Jeffreys, prior distributions and the philosophy of statistics. Statistical Science, 24(2):176–178, May 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1263478375.

Gelman:2010:BST

[Gel10] Andrew Gelman. Bayesian statistics Then and now. Statistical Science, 25(2):162–165, May 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1290175837.

Gelman:2011:BSP

[Gel11] Andrew Gelman. Bayesian statistical pragmatism. Statistical Science, 26(1):10–11, February 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1307626555.

Gelman:2014:HBA

[Gel14] Andrew Gelman. How Bayesian analysis cracked the red-state, blue-state problem. Statistical Science, 29(1):26–35, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1399645725.

Genest:1993:RAH

[Gen93] Christian Genest. [Report of the ad hoc committee on design of an experiment on double-blind refereeing]: Comment. Statistical Science, 8(3):323–327, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177010908.

George:1990:DDT

[Geo90] Edward I. George. [Developments in decision-theoretic variance estimation]: Comment. Statistical Science, 5(1):107–109, February 1990. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177012267.

George:2007:TIO

[Geo07] Edward I. George. A tribute to Ingram Olkin. Statistical Science, 22(3):400, August 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1199285038.

Gertsbakh:1987:SSW

[Ger87] Llya Gertsbakh. [Survey of Soviet work in reliability]: Comment. Statistical Science, 2(4):501–502, November 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013132.

Geweke:1988:EDS

[Gew88] John Geweke. [Employment discrimination and statistical science]: Comment: Statistical science and economic science. Statistical Science, 3(2):188–189, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012902.

Geweke:1992:SPC

[Gew92] John Geweke. [Statistics, probability and chaos]: Comment: Inference and prediction in the presence of uncertainty and determinism. Statistical Science, 7(1):94–101, February 1992. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011446.

Geyer:1992:PMCa

[Gey92a] Charles J. Geyer. Practical Markov chain Monte Carlo. Statistical Science, 7(4):473–483, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011137.

Geyer:1992:PMCb

[Gey92b] Charles J. Geyer. [Practical Markov chain Monte Carlo]: Rejoinder. Statistical Science, 7(4):502–503, November 1992. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011147.

Geyer:1995:BCS

[Gey95] Charles J. Geyer. [Bayesian computation and stochastic systems]: Comment. Statistical Science, 10(1):46–48, February 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010126.

Gainous:2002:WDV

[GG02] Jason Gainous and Jeff Gill. Why does voting get so complicated? a review of theories for analyzing democratic participation. Statistical Science, 17(4):383–404, November 2002. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1049993199.

Gustafson:2009:IEM

[GG09] Paul Gustafson and Sander Greenland. Interval estimation for messy observational data. Statistical Science, 24(3):328–342, August 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1270041259.

Greenacre:2023:ACD

[GGBS^{+}23] Michael Greenacre, Eric Grunsky, John Bacon-Shone, Ionas Erb, and Thomas Quinn. Aitchison’s compositional data analysis 40 years on: a reappraisal. Statistical Science, 38(3):386– 410, August 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-3/
Aitchisons-Compositional-Data-Analysis-40-Years-on-
A-Reappraisal/10.1214/22-STS880.full.

Gelfand:2013:SIP

[GGC13] Alan E. Gelfand, Souparno Ghosh, and James S. Clark. Scaling integral projection models for analyzing size demography. Statistical Science, 28(4):641–658, November 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1386078882.

Gastwirth:2009:ILT

[GGM09] Joseph L. Gastwirth, Yulia R. Gel, and Weiwen Miao. The impact of Levene’s test of equality of variances on statistical theory and practice. Statistical Science, 24(3):343–360, August 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1270041260.

Grund:1991:CKR

[GH91] Birgit Grund and Wolfgang Hardle. [Choosing a kernel regression estimator]: Comment. Statistical Science, 6(4):421– 425, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011588.

Glad:2016:MUF

[GH16] Ingrid Glad and Nils Lid Hjort. Model uncertainty first, not afterwards. Statistical Science, 31(4):490–494, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1484816573.

Ghosh:2011:OPI

[Gho11a] Malay Ghosh. Objective priors: An introduction for frequentists. Statistical Science, 26(2):187–202, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1312204006.  See

discussion [Sch11, Swe11].

Ghosh:2011:R

[Gho11b] Malay Ghosh. Rejoinder. Statistical Science, 26(2):210–211, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1312204009.

Geirsson:2020:LSS

[GHSS20] Óli Páll Geirsson, Birgir Hrafnkelsson, Daniel Simpson, and Helgi Sigurdarson. LGM split sampler: An efficient MCMC sampling scheme for latent Gaussian models. Statistical Science, 35(2):218–233, May 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1591171228.

Giraud:2012:HDR

[GHV12] Christophe Giraud, Sylvie Huet, and Nicolas Verzelen. Highdimensional regression with unknown variance. Statistical Science, 27(4):500–518, November 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1356098553.

Gilbert:1986:SPI

[Gil86] Freeman Gilbert. [A statistical perspective on ill-posed inverse problems]: Comment. Statistical Science, 1(4):523, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013529. See [O’S86b].

Gill:2014:SCB

[Gil14] Richard D. Gill. Statistics, causality and Bell’s Theorem. Statistical Science, 29(4):512–528, November 2014. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1421330545.

Ginevan:1988:RIA

[Gin88] Michael E. Ginevan. Radon as an indoor air pollutant. Statistical Science, 3(3):371–373, August 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012840.

Giovagnoli:2023:CRA

[Gio23] Alessandra Giovagnoli. Comment: Is response-adaptive randomization a “Good Thing” or not in clinical trials? Why we cannot take sides. Statistical Science, 38(2):224–228, May 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-2/

Comment--Is-Response-Adaptive-Randomization-a-Good-
Thing-or/10.1214/23-STS865E.full.

Girolami:2014:CMG

[Gir14] Mark A. Girolami. Contribution by M. A. Girolami. Statistical Science, 29(1):97, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1399645734.

Groeneboom:2018:SDT

[GJ18] Piet Groeneboom and Geurt Jongbloed. Some developments in the theory of shape constrained inference. Statistical Science, 33(4):473–492, November 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1543482054.

Gasser:1991:CKR

[GJSE91] Theo Gasser, Christine Jennen-Steinmetz, and Joachim Engel. [Choosing a kernel regression estimator]: Comment. Statistical Science, 6(4):419–421, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011587.

Gnanadesikan:1987:RMA

[GK87] R. Gnanadesikan and J. R. Kettenring. [A review of multivariate analysis]: Comment. Statistical Science, 2(4):428– 430, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013116.

Graubardand:2002:ISP

[GK02] Barry I. Graubardand and Edward L. Korn. Inference for superpopulation parameters using sample surveys. Statistical Science, 17(2):73–96, May 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1023798999.

Glasserman:2006:CCH

[GK06] Paul Glasserman and Steven Kou. A conversation with Chris Heyde. Statistical Science, 21(2):286–298, May 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1154979828.

Genton:2015:CCF

[GK15a] Marc G. Genton and William Kleiber. Cross-covariance functions for multivariate geostatistics. Statistical Science, 30(2): 147–163, May 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1433341471. See comment [BHP15].

Genton:2015:R

[GK15b] Marc G. Genton and William Kleiber. Rejoinder. Statistical Science, 30(2):181–183, May 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1433341476.

Ghanem:2019:DMA

[GK19] Dalia Ghanem and Todd A. Kuffner. Discussion: Models as approximations. Statistical Science, 34(4):604–605, November 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/

1578474026. See [BBB^{+}19, BBK^{+}19].

Gelman:1999:ALD

[GKLP99] Andrew Gelman, David H. Krantz, Chiayu Lin, and Phillip N. Price. Analysis of local decisions using hierarchical modeling, applied to home Radon measurement and remediation. Statistical Science, 14(3):305–337, August 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1009212411.

Gomatam:2005:DDD

[GKRS05] S. Gomatam, A. F. Karr, J. P. Reiter, and A. P. Sanil. Data dissemination and disclosure limitation in a world without microdata: A risk–utility framework for remote access analysis servers. Statistical Science, 20(2):163–177, May 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1121347638.

Gelman:2002:MSV

[GKT02] Andrew Gelman, Jonathan N. Katz, and Francis Tuerlinckx. The mathematics and statistics of voting power. Statistical Science, 17(4):420–435, November 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1049993201.

Goldstein:1996:RSS

[GL96] Larry Goldstein and Bryan Langholz. Risk set sampling in epidemiologic cohort studies. Statistical Science, 11(1):35–53, February 1996. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1032209663.

Goldfarb:2006:HSR

[GL06] Avi Goldfarb and Qiang Lu. Household-specific regressions using clickstream data. Statistical Science, 21(2):247–255, May 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1154979825.

Gladen:1987:SPM

[Gla87] Beth C. Gladen. [The statistical precision of medical screening procedures: Application to polygraph and AIDS antibodies test data]: Comment. Statistical Science, 2(3):233, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013221.

Gleser:1998:AUM

[Gle98] Leon Jay Gleser. Assessing uncertainty in measurement. Statistical Science, 13(3):277–290, August 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1028905888.

Glick:1991:ASP

[Gli91] Ned Glick. [Avoiding statistical pitfalls]: Comment. Statistical Science, 6(3):258–262, August 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011690.

Guhaniyogi:2023:DBI

[GLSS23] Rajarshi Guhaniyogi, Cheng Li, Terrance Savitsky, and Sanvesh Srivastava. Distributed Bayesian inference in massive spatial data. Statistical Science, 38(2):262–284, May 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-2/
Distributed-Bayesian-Inference-in-Massive-Spatial-Data/
10.1214/22-STS868.full.

Graubard:2007:CMS

[GLW07] Barry I. Graubard, Paul S. Levy, and Gordon B. Willis. A conversation with Monroe Sirken. Statistical Science, 22(4): 637–650, November 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1207580176.

George:2012:MSE

[GLX12] Edward I. George, Feng Liang, and Xinyi Xu. From minimax shrinkage estimation to minimax shrinkage prediction. Statistical Science, 27(1):82–94, February 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1331729984.

Gelman:1998:SNC

[GM98] Andrew Gelman and Xiao-Li Meng. Simulating normalizing constants: from importance sampling to bridge sampling to path sampling. Statistical Science, 13(2):163–185, May 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1028905934.

Geyer:2005:FRC

[GM05a] Charles J. Geyer and Glen D. Meeden. Fuzzy and randomized confidence intervals and P -values. Statistical Science, 20(4): 358–366, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1137076652. See comments [AG05, BC05,

BCD05, Tho05, GM05b].

Geyer:2005:RFR

[GM05b] Charles J. Geyer and Glen D. Meeden. Rejoinder: Fuzzy and randomized confidence intervals and P -values. Statistical Science, 20(4):384–387, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1137076658. See [GM05a].

Gustafson:2018:WSP

[GM18] Paul Gustafson and Lawrence C. McCandless. When is a sensitivity parameter exactly that? Statistical Science, 33(1):86–95, 02 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Gao:2021:MRN

[GM21a] Chao Gao and Zongming Ma. Minimax rates in network analysis: Graphon estimation, community detection and hypothesis testing. Statistical Science, 36(1):16–33, February 2021. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-1/Minimax-Rates-in-Network-
Analysis--Graphon-Estimation-Community-Detection/10.
1214/19-STS736.full.

Gong:2021:JJM

[GM21b] Ruobin Gong and Xiao-Li Meng. Judicious judgment meets unsettling updating: Dilation, sure loss and Simpson’s Paradox. Statistical Science, 36(2):169–190, May 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-2/Judicious-Judgment-Meets-
Unsettling-Updating--Dilation-Sure-Loss-and/10.1214/
19-STS765.full.

Genovese:2004:NIC

[GMN^{+}04] Christopher R. Genovese, Christopher J. Miller, Robert C. Nichol, Mihir Arjunwadkar, and Larry Wasserman. Nonparametric inference for the cosmic microwave background. Statistical Science, 19(2):308–321, May 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1105714165.

Gerchinovitz:2020:FIR

[GMS20] Sébastien Gerchinovitz, Pierre Ménard, and Gilles Stoltz. Fano’s inequality for random variables. Statistical Science, 35(2):178–201, May 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1591171226.

Gijbels:2017:GDD

[GN17] Irène Gijbels and Stanislav Nagy. On a general definition of depth for functional data. Statistical Science, 32(4):630–639, 11 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Gnanadesikan:1994:RSP

[Gna94] Anand Gnanadesikan. [Report on statistics and physical oceanography]: Comment. Statistical Science, 9(2):208–212, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010484.

Gallego:2024:PCA

[GNR^{+}24] Víctor Gallego, Roi Naveiro, Alberto Redondo, David Ríos Insua, and Fabrizio Ruggeri. Protecting classifiers from attacks. Statistical Science, 39(3):449–468, August 2024. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-39/issue-3/Protecting-Classifiers-from-
Attacks/10.1214/24-STS922.full.

Godambe:1995:IBE

[God95] V. P. Godambe. [Inference based on estimating functions in the presence of nuisance parameters]: Comment. Statistical Science, 10(2):173–174, May 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177010029.

Goel:1986:CSR

[Goe86] Prem K. Goel. [Computers in statistical research]: Comment. Statistical Science, 1(4):444–445, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013512.  See

[Edd86a].

Goldberg:1987:SPM

[Gol87a] Judith D. Goldberg. [The statistical precision of medical screening procedures: Application to polygraph and AIDS antibodies test data]: Comment. Statistical Science, 2(3):230–231, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013219.

Goldstein:1987:RMA

[Gol87b] Matthew Goldstein. [A review of multivariate analysis]: Comment. Statistical Science, 2(4):418–420, November 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013113.

Goldberger:1988:EDS

[Gol88] Arthur S. Goldberger. [Employment discrimination and statistical science]: Comment. Statistical Science, 3(2):165–166, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012896.

Goldstein:2016:IRA

[Gol16] Michael Goldstein. Issues in robustness analysis. Statistical Science, 31(4):503–505, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1484816576.

Good:1986:SSAa

[Goo86a] I. J. Good. Some statistical applications of Poisson’s work. Statistical Science, 1(2):157–170, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1177013690. See comments

[Sin86, Goo86b].

Good:1986:SSAb

[Goo86b] I. J. Good. [Some statistical applications of Poisson’s work]: Rejoinder. Statistical Science, 1(2):179–180, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013695. See [Goo86a].

Good:1986:ASP

[Goo86c] I. J. Good. [The axioms of subjective probability]: Comment. Statistical Science, 1(3):346–347, August 1986. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013612.

Goodall:1986:SSS

[Goo86d] Colin Goodall. [Size and shape spaces for Landmark data in two dimensions]: Comment. Statistical Science, 1(2):234–238, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013701.

Good:1988:IBSa

[Goo88a] I. J. Good. The interface between statistics and philosophy of science. Statistical Science, 3(4):386–397, November 1988. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012754.

Good:1988:IBSb

[Goo88b] I. J. Good. [The interface between statistics and philosophy of science]: Rejoinder. Statistical Science, 3(4):406–412, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012759.

Goodman:2007:WBK

[Goo07] Leo A. Goodman. Working with bill Kruskal: From 1950 onward. Statistical Science, 22(2):269–272, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1190905525.

Goodman:2011:DSI

[Goo11] Steven N. Goodman. Discussion of “Statistical Inference: The Big Picture” by R. E. Kass. Statistical Science, 26(1): 12–14, February 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1307626556. See [Kas11b].

Gower:1990:OOM

[Gow90] J. C. Gower. [OMEGA (online multivariate exploratory graphical analysis): Routine searching for structure]: Comment. Statistical Science, 5(2):216–217, May 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012170.

Greenland:1999:CCC

[GPR99] Sander Greenland, Judea Pearl, and James M. Robins. Confounding and collapsibility in causal inference. Statistical Science, 14(1):29–46, February 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1009211805.

Gel:2023:CGJ

[GPW23] Yulia R. Gel, Edsel A. Peña, and Huixia Judy Wang. Conversations with Gábor J. Székely. Statistical Science, 38(2):355– 367, May 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-2/
Conversations-with-G%c3%a1bor-J-Sz%c3%a9kely/10.1214/
22-STS873.full.

Gelman:1992:IIS

[GR92a] Andrew Gelman and Donald B. Rubin. Inference from iterative simulation using multiple sequences. Statistical Science, 7(4): 457–472, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011136.

Gelman:1992:PMC

[GR92b] Andrew Gelman and Donald B. Rubin. [Practical Markov chain Monte Carlo]: Rejoinder: Replication without contrition. Statistical Science, 7(4):503–511, November 1992. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011148.

Gail:1993:BHI

[GR93] Mitchell H. Gail and Philip S. Rosenberg. [Backcalculation of HIV infection rates]: Comment. Statistical Science, 8(2): 107–109, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010997.

Ghosh:1994:SAEa

[GR94a] M. Ghosh and J. N. K. Rao. Small area estimation: An appraisal. Statistical Science, 9(1):55–76, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010647.

Ghosh:1994:SAEb

[GR94b] M. Ghosh and J. N. K. Rao. [Small area estimation: An appraisal]: Rejoinder. Statistical Science, 9(1):90–93, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010654.

Gilliland:2010:CJH

[GR10] Dennis Gilliland and R. V. Ramamoorthi. A conversation with James Hannan. Statistical Science, 25(1):126–144, February 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1280841737.

Greenshtein:2019:CEB

[GR19] Eitan Greenshtein and Ya’acov Ritov. Comment: Empirical Bayes, compound decisions and exchangeability. Statistical Science, 34(2):224–228, May 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1563501637. See [Efr19a].

Granger:1992:SPC

[Gra92] Clive W. J. Granger. [Statistics, probability and chaos]: Comment. Statistical Science, 7(1):102–104, February 1992. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011447.

Gray:1993:CSTa

[Gra93a] Mary W. Gray. Can statistics tell us what we do not want to hear? the case of complex salary structures. Statistical Science, 8(2):144–158, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011006.

Gray:1993:CSTb

[Gra93b] Mary W. Gray. [Can statistics tell us what we do not want to hear? the case of complex salary structures]: Rejoinder. Statistical Science, 8(2):177–179, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011010.

Greenhouse:1987:SPCb

[Gre87a] Joel B. Greenhouse. [Statistics in psychiatry]: Comment. Statistical Science, 2(2):127–129, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013339. See [Eve87a].

Greenhouse:1987:SPCa

[Gre87b] Samuel W. Greenhouse. [Statistics in psychiatry]: Comment. Statistical Science, 2(2):118–120, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013335. See [Eve87a].

Greenhouse:1991:RMA

[Gre91] Joel B. Greenhouse. [Replication and meta-analysis in parapsychology]: Comment: Parapsychology — on the margins of science? Statistical Science, 6(4):386–389, November 1991. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011581.

Green:1997:CFE

[Gre97a] Sylvan B. Green. A conversation with Fred Ederer. Statistical Science, 12(2):125–131, May 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1029963430.

Greenhouse:1997:SRB

[Gre97b] Samuel W. Greenhouse. Some reflections on the beginnings and development of statistics in “Your Father’s NIH”. Statistical Science, 12(2):82–87, May 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1029963423.

Greenland:2009:RPP

[Gre09] Sander Greenland. Relaxation penalties and priors for plausible modeling of nonidentified bias sources. Statistical Science, 24 (2):195–210, May 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1263478381.

Greenland:2010:CNS

[Gre10] Sander Greenland. Comment: The need for syncretism in applied statistics. Statistical Science, 25(2):158–161, May 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1290175836.

Griffeath:1992:SPC

[Gri92] David Griffeath. [Statistics, probability and chaos]: Comment: Randomness in complex systems. Statistical Science, 7(1): 104–108, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011448.

Grunwald:2016:CMD

[Grü16] Peter Grünwald. Contextuality of misspecification and datadependent losses. Statistical Science, 31(4):495–498, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1484816574.

Glymour:1993:BAE

[GS93] Clark Glymour and Peter Spirtes. [Bayesian analysis in expert systems]: Comment: Conditional independence and causal inference. Statistical Science, 8(3):250–257, August 1993. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010890.

Ghose:2006:EPS

[GS06] Anindya Ghose and Arun Sundararajan. Evaluating pricing strategy using e-commerce data: Evidence and estimation challenges. Statistical Science, 21(2):131–142, May 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1154979816.

Ghosh:2008:CPK

[GS08] Malay Ghosh and Michael J. Schell. A conversation with Pranab Kumar Sen. Statistical Science, 23(4):548–564, November 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1242049394.

Goeman:2011:MTE

[GS11a] Jelle J. Goeman and Aldo Solari. Multiple testing for exploratory research. Statistical Science, 26(4):584–597, November 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.

ss/1330437937. See discussion [Hel11, Mei11, Wes11].

Goeman:2011:R

[GS11b] Jelle J. Goeman and Aldo Solari. Rejoinder. Statistical Science, 26(4):608–612, November 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1330437941.

George:2012:TCS

[GS12] Edward I. George and William E. Strawderman. A tribute to Charles Stein. Statistical Science, 27(1):1–2, February 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1331729977.

Goeman:2013:CN

[GS13] Jelle J. Goeman and Aldo Solari. Correction note. Statistical Science, 28(3):464, August 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1377696946.

Guntuboyina:2018:NSR

[GS18] Adityanand Guntuboyina and Bodhisattva Sen. Nonparametric shape-restricted regression. Statistical Science, 33(4):568–594, November 2018. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1543482059.

Gneiting:2012:EFD

[GŠP12] Tilmann Gneiting, Hana Ševčíková, and Donald B. Percival. Estimators of fractal dimension: Assessing the roughness of time series and spatial data. Statistical Science, 27(2):247–277, May 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1340110872.

Givens:1997:PBM

[GST97] Geof H. Givens, D. D. Smith, and R. L. Tweedie. Publication bias in meta-analysis: a Bayesian data-augmentation approach to account for issues exemplified in the passive smoking debate. Statistical Science, 12(4):221–250, November 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1030037958.

Gibson:2018:CAE

[GST18] Gavin J. Gibson, George Streftaris, and David Thong. Comparison and assessment of epidemic models. Statistical Science, 33 (1):19–33, 02 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Griffiths:1994:AIP

[GT94] R. C. Griffiths and Simon Tavaré. Ancestral inference in population genetics. Statistical Science, 9(3):307–319, August 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010378.

George:2003:DDG

[GT03] A. W. George and E. A. Thompson. Discovering disease genes: Multipoint linkage analysis via a new Markov chain Monte Carlo approach. Statistical Science, 18(4):515–531, November 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1081443233.

Gabda:2012:DSM

[GTWT12] Darmesah Gabda, Ross Towe, Jennifer Wadsworth, and Jonathan Tawn. Discussion of “Statistical Modeling of Spatial Extremes” by A. C. Davison, S. A. Padoan and M. Ribatet. Statistical Science, 27(2):189–192, May 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1340110866. See [DPR12b].

Gupta:1988:GOC

[Gup88] Shanti S. Gupta. [Golden oldies: Classic articles from the world of statistics and probability]: Comment. Statistical Science, 3 (1):104–106, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013010.

Gustafson:2005:MEM

[Gus05] Paul Gustafson. On model expansion, model contraction, identifiability and prior information: Two illustrative scenarios involving mismeasured variables. Statistical Science, 20 (2):111–140, May 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1121347636.

Gustafson:2023:PRS

[Gus23] Paul Gustafson. Parameter restrictions for the sake of identification: Is there utility in asserting that perhaps a restriction holds? Statistical Science, 38(3):477–489, August 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-3/
Parameter-Restrictions-for-the-Sake-of-Identification-
-Is-There/10.1214/23-STS885.full.

Guthrie:1987:SPC

[Gut87] Donald Guthrie. [Statistics in psychiatry]: Comment. Statistical Science, 2(2):116–117, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013334. See [Eve87a].

Giovagnoli:1988:ODC

[GV88] A. Giovagnoli and I. Verdinelli. [Optimal designs for comparing test treatments with controls]: Comment. Statistical Science, 3(4):482–484, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012770.

Giovagnoli:2023:BAR

[GV23] Alessandra Giovagnoli and Isabella Verdinelli. Bayesian adaptive randomization with compound utility functions. Statistical Science, 38(1):52–67, February 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-1/Bayesian-Adaptive-Randomization-
with-Compound-Utility-Functions/10.1214/21-STS848.full.

Gruber:2019:CAV

[GvdL19] Susan Gruber and Mark J. van der Laan. Comment on “Automated Versus Do-It-Yourself Methods for Causal Inference: Lessons Learned from a Data Analysis Competition”. Statistical Science, 34(1):82–85, February 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1555056034. See [DHS^{+}19a].

Gelfand:2002:SBA

[GW02] Alan E. Gelfand and Fei Wang. A simulation-based approach to Bayesian sample size determination for performance under a given model and for separating models. Statistical Science, 17(2):193–208, May 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1030550861.

Genovese:2004:ISS

[GWC04] Chris Genovese, Larry Wasserman, and George Casella. Introduction to the special section on astrostatistics. Statistical Science, 19(2):264, May 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1105714161.

Goddard:2009:EEM

[GWVV09] Michael E. Goddard, Naomi R. Wray, Klara Verbyla, and Peter M. Visscher. Estimating effects and making predictions from genome-wide marker data. Statistical Science, 24(4):517– 529, November 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1271770346.

Genest:1986:CPDa

[GZ86a] Christian Genest and James V. Zidek. Combining probability distributions: A critique and an annotated bibliography. Statistical Science, 1(1):114–135, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1177013825. See comments

[Sha86a, Fre86, Win86, Mor86, Hog86a, GZ86b].

Genest:1986:CPDb

[GZ86b] Christian Genest and James V. Zidek. [Combining probability distributions: A critique and an annotated bibliography]: Rejoinder. Statistical Science, 1(1):147–148, February 1986. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013831.

Gine:1989:AEP

[GZ89] Evarist Gine and Joel Zinn. [Asymptotics via empirical processes]: Comment. Statistical Science, 4(4):355–356, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012396.

Guo:2020:MIC

[GZZ20] Ruosi Guo, Chunming Zhang, and Zhengjun Zhang. Maximum independent component analysis with application to EEG data. Statistical Science, 35(1):145–157, February 2020. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/euclid.ss/1583226034.

Hettmansperger:1988:RBR

[HA88] T. P. Hettmansperger and James C. Aubuchon. [Rank-based robust analysis of linear models. I. Exposition and review]: Comment. Statistical Science, 3(2):262–263, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012918.

Hacking:1990:UDP

[Hac90] Ian Hacking. [The unity and diversity of probability]: Comment: In praise of the diversity of probabilities. Statistical Science, 5(4):450–454, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012027.

Hahn:1993:ISC

[Hah93] Gerald J. Hahn. [Is industrial statistics out of control?]: Comment. Statistical Science, 8(4):384–387, November 1993. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010776.

Hahn:1995:CDM

[Hah95] Gerald J. Hahn. A conversation with Donald Marquardt. Statistical Science, 10(4):377–393, November 1995. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177009871.

Hald:1999:HML

[Hal99] Anders Hald. On the history of maximum likelihood in relation to inverse probability and least squares. Statistical Science, 14 (2):214–222, May 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009212248.

Hall:2003:SPB

[Hal03] Peter Hall. A short prehistory of the bootstrap. Statistical Science, 18(2):158–167, May 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1063994970.

Hall:2009:CCS

[Hal09] Peter Gavin Hall. Comment: Citation statistics. Statistical Science, 24(1):25–26, February 2009. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1255009006. See [Hal09].

Hannan:1987:RTFa

[Han87a] E. J. Hannan. Rational transfer function approximation. Statistical Science, 2(2):135–151, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013343.

Hannan:1987:RTFb

[Han87b] E. J. Hannan. [Rational transfer function approximation]: Rejoinder. Statistical Science, 2(2):160–161, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013350.

Hansen:1987:SHR

[Han87c] Morris H. Hansen. Some history and reminiscences on survey sampling. Statistical Science, 2(2):180–190, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013352.

Hankey:1997:CWM

[Han97] Benjamin Hankey. A conversation with William M. Haenszel. Statistical Science, 12(2):108–112, May 1997. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1029963427.

Hand:2006:CTI

[Han06a] David J. Hand. Classifier technology and the illusion of progress. Statistical Science, 21(1):1–14, February 2006. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1149600839.

See comments [Fri06, Gay06, Hol06, Sti06b, Han06b].

Hand:2006:RCT

[Han06b] David J. Hand. Rejoinder: Classifier technology and the illusion of progress. Statistical Science, 21(1):30–34, February 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1149600844. See comments [Han06a].

Hand:2011:DBM

[Han11a] David J. Hand. Discussion of “Bayesian Models and Methods in Public Policy and Government Settings” by S. E. Fienberg. Statistical Science, 26(2):227–230, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1312204011. See [Fie11a].

Hansen:2011:DFM

[Han11b] Bruce E. Hansen. Discussion of “Feature Matching in Time Series Modeling” by Y. Xia and H. Tong. Statistical Science, 26(1):47–48, February 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1307626561. See [XT11a].

Hand:2014:WEL

[Han14a] David J. Hand. Wonderful examples, but Let’s not close our eyes. Statistical Science, 29(1):98–100, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1399645735.

Hannig:2014:DBA

[Han14b] Jan Hannig. Discussion of “On the Birnbaum Argument for the Strong Likelihood Principle”. Statistical Science, 29(2): 254–258, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1408368578. See [May14a].

Hartigan:1986:BMS

[Har86] J. A. Hartigan. [Bootstrap methods for standard errors, confidence intervals, and other measures of statistical accuracy]: Comment. Statistical Science, 1(1):75–77, February 1986. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013816.

Hardwick:1989:ITP

[Har89] Janis Hardwick. [Investigating therapies of potentially great benefit: ECMO]: Comment: Recent progress in clinical trial designs that adapt for ethical purposes. Statistical Science, 4(4): 327–336, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012392.

Hart:1991:CKR

[Har91a] Jeffrey D. Hart. [Choosing a kernel regression estimator]: Comment. Statistical Science, 6(4):425–427, November 1991. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011589.

Harville:1991:BGT

[Har91b] David A. Harville. [That BLUP is a good thing: The estimation of random effects]: Comment. Statistical Science, 6 (1):35–39, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011928.

Hartigan:2003:MJT

[Har03] J. A. Hartigan. A memory of John Tukey as a teacher. Statistical Science, 18(3):341, August 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1076102422.

Haseman:1988:MMQ

[Has88] J. K. Haseman. [From mouse-to-man: The quantitative assessment of cancer risks]: Comment. Statistical Science, 3(1): 33–39, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012995.

Hastie:2007:CBA

[Has07] Trevor Hastie. Comment: Boosting algorithms: Regularization, prediction and model fitting. Statistical Science, 22(4): 513–515, November 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1207580165. See [BH07a].

Hattis:1988:UBM

[Hat88] Dale Hattis. The use of biological markers in risk assessment. Statistical Science, 3(3):358–366, August 1988. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012838.

Hansen:2008:CBS

[HB08] Ben B. Hansen and Jake Bowers. Covariate balance in simple, stratified and clustered comparative studies. Statistical Science, 23(2):219–236, May 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1219339114.

Haziza:2017:CWS

[HB17] David Haziza and Jean-François Beaumont. Construction of weights in surveys: A review. Statistical Science, 32(2):206–226, May 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1494489812.

Held:2015:ABM

[HBG15] Leonhard Held, Daniel Sabanés Bové, and Isaac Gravestock. Approximate Bayesian model selection with the deviance statistic. Statistical Science, 30(2):242–257, May 2015. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1433341481.

Huang:2012:SRG

[HBM12] Jian Huang, Patrick Breheny, and Shuangge Ma. A selective review of group selection in high-dimensional models. Statistical Science, 27(4):481–499, November 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1356098552.

Heng:2024:DSB

[HDD24] Jeremy Heng, Valentin De Bortoli, and Arnaud Doucet. Diffusion Schrödinger bridges for Bayesian computation. Statistical Science, 39(1):90–99, February 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-39/issue-1/Diffusion-Schr%c3%b6dinger-
Bridges-for-Bayesian-Computation/10.1214/23-STS908.
full.

Hedges:1988:SMF

[Hed88] Larry V. Hedges. [Selection models and the file drawer problem]: Comment. Statistical Science, 3(1):118–120, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013013.

Hedges:1992:MPS

[Hed92] Larry V. Hedges. Modeling publication selection effects in metaanalysis. Statistical Science, 7(2):246–255, May 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011364.

Hedayat:2005:CWF

[Hed05] A. S. Hedayat. A conversation with Walter T. Federer. Statistical Science, 20(3):302–315, August 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1124891294.

Heitjan:1989:IGCa

[Hei89a] Daniel F. Heitjan. Inference from grouped continuous data: A review. Statistical Science, 4(2):164–179, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012601.

Heitjan:1989:IGCb

[Hei89b] Daniel F. Heitjan. [Inference from grouped continuous data: A review]: Rejoinder. Statistical Science, 4(2):182–183, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012603.

Heller:2011:DMT

[Hel11] Ruth Heller. Discussion of “Multiple Testing for Exploratory Research” by J. J. Goeman and A. Solari. Statistical Science, 26(4):598–600, November 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1330437938. See [GS11a].

Hernan:2019:CSC

[Her19] Miguel A. Hernán. Comment: Spherical cows in a vacuum: Data analysis competitions for causal inference. Statistical Science, 34(1):69–71, February 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1555056031.

Heyde:1986:SSA

[Hey86] C. C. Heyde. [Some statistical applications of Poisson’s work]: Comment. Statistical Science, 1(2):176–177, May 1986. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013693.

Heyde:1995:CJG

[Hey95] Chris Heyde. A conversation with Joe Gani. Statistical Science, 10(2):214–230, May 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010039.

Hill:1989:URI

[Hil89] Joe R. Hill. [The unifying role of iterative generalized least squares in statistical algorithms]: Comment. Statistical Science, 4(4):406, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012411.

Hill:1993:BAE

[Hil93] Joe R. Hill. [Bayesian analysis in expert systems]: Comment. Statistical Science, 8(3):258–261, August 1993. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010891.

Hill:1995:SDS

[Hil95] Theodore P. Hill. A statistical derivation of the significantdigit law. Statistical Science, 10(4):354–363, November 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (elec-

tronic). URL 2-P&origin=MSN;   http://links.jstor.org/
sici?sici=0883-4237(199511)10:4<354:ASDOTS>2.0.CO;
http://projecteuclid.org/euclid.ss/1177009869.

Hinkley:1994:BMT

[Hin94] David Hinkley. [Bootstrap: More than a stab in the dark?]: Comment. Statistical Science, 9(3):400–403, August 1994. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010387.

Hinkelmann:2001:ROK

[Hin01] Klaus Hinkelmann. Remembering Oscar Kempthorne (1919– 2000). Statistical Science, 16(2):169–183, May 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1009213289.

Hobert:2001:HEI

[HJ01] James P. Hobert and Galin L. Jones. Honest exploration of intractable probability distributions via Markov chain Monte Carlo. Statistical Science, 16(4):312–334, November 2001. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1015346317.

Hickernell:2019:CPI

[HJ19] Fred J. Hickernell and R. Jagadeeswaran. Comment on “Probabilistic Integration: A Role in Statistical Computation?”. Statistical Science, 34(1):23–28, February 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1555056026. See [BOG^{+}19a]

and rejoinder [BOG^{+}19b].

Huang:2021:UPD

[HJJ^{+}21] Jian Huang, Yuling Jiao, Bangti Jin, Jin Liu, Xiliang Lu, and Can Yang. A unified primal dual active set algorithm for nonconvex sparse recovery. Statistical Science, 36(2):215– 238, May 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-36/issue-2/
A-Unified-Primal-Dual-Active-Set-Algorithm-for-Nonconvex-
Sparse/10.1214/19-STS758.full.

Hedayat:1988:ODCa

[HJM88a] A. S. Hedayat, Mike Jacroux, and Dibyen Majumdar. Optimal designs for comparing test treatments with controls. Statistical Science, 3(4):462–476, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012767.

Hedayat:1988:ODCb

[HJM88b] A. S. Hedayat, Mike Jacroux, and Dibyen Majumdar. [Optimal designs for comparing test treatments with controls]: Rejoinder. Statistical Science, 3(4):487–491, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012773.

Hoaglin:1986:IOH

[HK86] David C. Hoaglin and Peter J. Kempthorne. [Influential observations, high leverage points, and outliers in linear regression]:

Comment. Statistical Science, 1(3):408–412, August 1986. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013627.

Hansen:2002:SAE

[HK02] Mark H. Hansen and Charles Kooperberg. Spline adaptation in extended linear models (with comments and a rejoinder by the authors). Statistical Science, 17(2):2–51, May 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1023798997.

Hsu:2022:CRP

[HK22] Li Hsu and Charles Kooperberg. A conversation with Ross Prentice. Statistical Science, 37(1):143–158, February 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-1/
A-Conversation-with-Ross-Prentice/10.1214/21-STS829.
full.

Habermann:2017:CRG

[HKL17] Hermann Habermann, Courtney Kennedy, and Partha Lahiri. A conversation with Robert Groves. Statistical Science, 32(1): 128–137, February 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1491465631.

Hu:2022:MEM

[HKL22] Zhirui Hu, Zheng Tracy Ke, and Jun S. Liu. Measurement error models: From nonparametric methods to deep neural networks. Statistical Science, 37(4):473–493, November 2022. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-37/issue-4/Measurement-Error-Models-
-From-Nonparametric-Methods-to-Deep-Neural/10.1214/
21-STS834.full.

Hastie:1993:LRAa

[HL93a] Trevor Hastie and Clive Loader. Local regression: Automatic kernel carpentry. Statistical Science, 8(2):120–129, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011002.

Hastie:1993:LRAb

[HL93b] Trevor Hastie and Clive Loader. [Local regression: Automatic kernel carpentry]: Rejoinder. Statistical Science, 8(2):139–143, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011005.

Hickernell:2005:CVQ

[HLO05] Fred J. Hickernell, Christiane Lemieux, and Art B. Owen. Control variates for quasi-Monte Carlo. Statistical Science, 20 (1):1–31, February 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1118065040.

Hollander:1995:CFP

[HM95] Myles Hollander and Albert W. Marshall. A conversation with Frank Proschan. Statistical Science, 10(1):118–133, February 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010134.

Hansen:2016:AAM

[HM16] Lars Peter Hansen and Massimo Marinacci. Ambiguity aversion and model misspecification: An economic perspective. Statistical Science, 31(4):511–515, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1484816578.

Hoeting:1999:BMA

[HMRV99] Jennifer A. Hoeting, David Madigan, Adrian E. Raftery, and Chris T. Volinsky. Bayesian model averaging: a tutorial (with comments by M. Clyde, David Draper and E. I. George, and a rejoinder by the authors). Statistical Science, 14(4):382– 417, November 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009212519. See correction [HMRV00].

Hoeting:2000:CBM

[HMRV00] Jennifer A. Hoeting, David Madigan, Adrian E. Raftery, and Chris T. Volinsky. Correction to: “Bayesian model averaging: a tutorial” [Statist. Sci. 14 (1999), no. 4, 382–417; MR 2001a:62033]. Statistical Science, 15(3):193–195, August 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1009212814. See [HMRV99].

Hoaglin:2003:JWT

[Hoa03] David C. Hoaglin. John W. Tukey and data analysis. Statistical Science, 18(3):311–318, August 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1076102418.

Hodges:1987:UPAa

[Hod87a] James S. Hodges. Uncertainty, policy analysis and statistics. Statistical Science, 2(3):259–275, August 1987. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013224.

Hodges:1987:UPAb

[Hod87b] James S. Hodges. [Uncertainty, policy analysis and statistics]: Rejoinder. Statistical Science, 2(3):288–291, August 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013231.

Hodges:2017:SFL

[Hod17] James S. Hodges. Swinging for the fence in a league where everyone bunts. Statistical Science, 32(1):33–35, February 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1491465623.

Hoff:2021:AME

[Hof21] Peter Hoff. Additive and multiplicative effects network models. Statistical Science, 36(1):34–50, February 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/journals/
statistical-science/volume-36/issue-1/Additive-and-
Multiplicative-Effects-Network-Models/10.1214/19-STS757.
full.

Hogarth:1986:CPD

[Hog86a] Robin M. Hogarth. [Combining probability distributions: A critique and an annotated bibliography]: Comment. Statistical Science, 1(1):145–147, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013830. See [GZ86a].

Hogg:1986:OIM

[Hog86b] Robert V. Hogg. On the origins of the Institute of Mathematical Statistics. Statistical Science, 1(2):285–291, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013709.

Hogg:1988:GOC

[Hog88] Robert V. Hogg. [Golden oldies: Classic articles from the world of statistics and probability]: Comment. Statistical Science, 3 (1):95–97, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013007. See [Hot88c].

Hogg:1993:ISC

[Hog93] Robert V. Hogg. [Is industrial statistics out of control?]: Comment. Statistical Science, 8(4):387–391, November 1993. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010777.

Holland:1987:RSC

[Hol87] Paul W. Holland. [The role of a second control group in an observational study]: Comment. Statistical Science, 2(3):306–308, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013233.

Holland:1988:EDS

[Hol88] Paul W. Holland. [Employment discrimination and statistical science]: Comment: Causal mechanism or causal effect: Which is best for statistical science? Statistical Science, 3(2):186–188, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012901.

Holland:1992:ISM

[Hol92] Paul W. Holland. [The impact of sociological methodology on statistical methodology]: Comment: It’s the interplay That’s important. Statistical Science, 7(2):198–201, May 1992. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011356.

Holloway:1994:RSP

[Hol94a] Greg Holloway. [Report on statistics and physical oceanography]: Comment. Statistical Science, 9(2):212–213, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010485.

Holt:1994:SAE

[Hol94b] D. Holt. [Small area estimation: An appraisal]: Comment. Statistical Science, 9(1):80–82, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177010649.

Hollander:2001:CRB

[Hol01] Myles Hollander. A conversation with Ralph A. Bradley. Statistical Science, 16(1):75–100, February 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/998929483.

Holmes:2003:BPT

[Hol03a] Susan Holmes. Bootstrapping phylogenetic trees: Theory and methods. Statistical Science, 18(2):241–255, May 2003. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1063994979.

Holmes:2003:BEC

[Hol03b] Susan Holmes. Bradley Efron: A conversation with good friends. Statistical Science, 18(2):268–281, May 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1063994981.

Holte:2006:ETP

[Hol06] Robert C. Holte. Elaboration on two points raised in “Classifier Technology and the Illusion of Progress”. Statistical Science, 21 (1):24–26, February 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1149600842. See comments [Han06a].

Hollander:2008:CJS

[Hol08] Myles Hollander. A conversation with Jayaram Sethuraman. Statistical Science, 23(2):272–285, May 2008. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1219339118.

Horowitz:2003:BE

[Hor03] Joel L. Horowitz. The bootstrap in econometrics. Statistical Science, 18(2):211–218, May 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1063994976.

Hotelling:1988:GOCc

[Hot88a] Harold Hotelling. [Golden oldies: Classic articles from the world of statistics and probability]: Comment: Academic politics and the teaching of statistics. Statistical Science, 3(1):92–95, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013006. See [Hot88c].

Hotelling:1988:GOCb

[Hot88b] Harold Hotelling. Golden oldies: Classic articles from the world of statistics and probability: The place of statistics in the University. Statistical Science, 3(1):72–83, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013002. See [Hot88c].

Hotelling:1988:GOCa

[Hot88c] Harold Hotelling. Golden oldies: Classic articles from the world of statistics and probability: The teaching of statistics. Statistical Science, 3(1):63–71, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1177013001. See comments

[Hot88b, Moo88, Zid88, Arr88, Hot88a, Hog88, Bra88, Dem88a, Gup88, Olk88].

Hougaard:1988:SMS

[Hou88] Philip Hougaard. [Saddlepoint methods and statistical inference]: Comment. Statistical Science, 3(2):230–231, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012909.

Howard:1998:TDB

[How98] J. V. Howard. The 2 × 2 table: a discussion from a Bayesian viewpoint. Statistical Science, 13(4):351–367, November 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1028905830.

Hallin:2004:MSR

[HP04a] Marc Hallin and Davy Paindaveine. Multivariate signed-rank tests in vector autoregressive order identification. Statistical Science, 19(4):697–711, November 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1113832734.

Hollander:2004:NMR

[HP04b] Myles Hollander and Edsel A. Peña. Nonparametric methods in reliability. Statistical Science, 19(4):644–651, November 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1113832729.

Hill:2006:NBM

[HPV06] Shawndra Hill, Foster Provost, and Chris Volinsky. Networkbased marketing: Identifying likely adopters via consumer networks. Statistical Science, 21(2):256–276, May 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1154979826.

Hausmann:2011:SBS

[HR11] Ricardo Hausmann and Roberto Rigobon. In search of the black swan: Analysis of the statistical evidence of electoral fraud in Venezuela. Statistical Science, 26(4):543–563, November 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1330437935.

Hay:2016:EPD

[HR16] Gordon Hay and Clive Richardson. Estimating the prevalence of drug use using mark–recapture methods. Statistical Science, 31(2):191–204, May 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1464105037.

Hobert:2011:ICP

[HRR11] James P. Hobert, Vivekananda Roy, and Christian P. Robert. Improving the convergence properties of the data augmentation algorithm with an application to Bayesian mixture modeling.

Statistical Science, 26(3):332–351, August 2011. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1320066924.

Hubert:2008:HBR

[HRV08] Mia Hubert, Peter J. Rousseeuw, and Stefan Van Aelst. Highbreakdown robust multivariate methods. Statistical Science, 23 (1):92–119, February 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1215441287.

Hastie:1998:MMH

[HS98] Trevor Hastie and Patrice Y. Simard. Metrics and models for handwritten character recognition. Statistical Science, 13(1): 54–65, February 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1028905973.

Hill:2009:CER

[HS09] Jennifer Hill and Marc Scott. Comment: The essential role of pair matching. Statistical Science, 24(1):54–58, February 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1255009009. See [IKN09a].

He:2022:CSP

[HS22] Xuming He and Xiaofeng Shao. A conversation with Stephen Portnoy. Statistical Science, 37(3):443–454, August 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-3/
A-Conversation-with-Stephen-Portnoy/10.1214/21-STS845.
full.

Hastie:1986:GAMa

[HT86a] Trevor Hastie and Robert Tibshirani. Generalized additive models. Statistical Science, 1(3):297–310, August 1986. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013604.

Hastie:1986:GAMb

[HT86b] Trevor Hastie and Robert Tibshirani. [Generalized additive models]: Rejoinder. Statistical Science, 1(3):314–318, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013609.

Hastie:1988:MRS

[HT88] Trevor Hastie and Robert Tibshirani. [Monotone regression splines in action]: Comment. Statistical Science, 3(4):450– 456, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012764.

Hall:2000:NAT

[HT00a] Peter Hall and Nader Tajvidi. Nonparametric analysis of temporal trend when fitting parametric models to Extreme-Value data. Statistical Science, 15(2):153–167, May 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1009212755.

Hastie:2000:BBC

[HT00b] Trevor Hastie and Robert Tibshirani. Bayesian backfitting (with comments and a rejoinder by the authors). Statistical Science, 15(3):196–223, August 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1009212815.

Hastie:2020:BSF

[HTT20] Trevor Hastie, Robert Tibshirani, and Ryan Tibshirani. Best subset, forward stepwise or lasso? analysis and recommendations based on extensive comparisons. Statistical Science, 35(4):579–592, November 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-35/issue-4/Best-Subset-Forward-Stepwise-
or-Lasso-Analysis-and-Recommendations-Based/10.1214/

19-STS733.full. See comments [CTB20].

Huber:1986:CED

[Hub86] Peter J. Huber. [Computing environments for data analysis]: Comment. Statistical Science, 1(2):273, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013706.

Huber:1987:DGD

[Hub87a] Peter J. Huber. [Dynamic graphics for data analysis]: Comment. Statistical Science, 2(4):385–386, November 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013106.

Huber:1987:UPA

[Hub87b] Peter J. Huber. [Uncertainty, policy analysis and statistics]: Comment. Statistical Science, 2(3):279–281, August 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013227.

Hutton:2016:WDP

[Hut16] Jane Hutton. What does “propensity” add? Statistical Science, 31(4):549–551, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1484816583.

Hadi:1987:CLS

[HV87] Ali S. Hadi and Paul F. Velleman. [Collinearity and least squares regression]: Comment: Diagnosing near collinearities in least squares regression. Statistical Science, 2(1):93–98, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013443.

Hinkley:1988:SMS

[HW88] D. V. Hinkley and S. Wang. [Saddlepoint methods and statistical inference]: Comment. Statistical Science, 3(2):232–233, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012910.

Hwang:1990:DDT

[Hwa90] Jiunn T. Hwang. [Developments in decision-theoretic variance estimation]: Comment: How much can the improvements be realized? Statistical Science, 5(1):110–111, February 1990. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012268.

Hyman:1991:RMA

[Hym91] Ray Hyman. [Replication and meta-analysis in parapsychology]: Comment. Statistical Science, 6(4):389–392, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011582.

Heagerty:2000:MMM

[HZ00] Patrick J. Heagerty and Scott L. Zeger. Marginalized multilevel models and likelihood inference (with comments and a rejoinder by the authors). Statistical Science, 15(1):1–26, February 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1009212671.

Hastie:2006:C

[HZ06] Trevor Hastie and Ji Zhu. Comment. Statistical Science, 21 (3):352–357, August 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1166642439.

Iyengar:1988:SMFa

[IG88a] Satish Iyengar and Joel B. Greenhouse. Selection models and the file drawer problem. Statistical Science, 3(1):109– 117, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013012.

Iyengar:1988:SMFb

[IG88b] Satish Iyengar and Joel B. Greenhouse. [Selection models and the file drawer problem]: Rejoinder. Statistical Science, 3(1): 133–135, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013019.

Imai:2023:PFH

[IJ23] Kosuke Imai and Zhichao Jiang. Principal fairness for human and algorithmic decision-making. Statistical Science, 38(2):317– 328, May 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-2/
Principal-Fairness-for-Human-and-Algorithmic-Decision-
Making/10.1214/22-STS872.full.

Imai:2009:ERP

[IKN09a] Kosuke Imai, Gary King, and Clayton Nall. The essential role of pair matching in cluster-randomized experiments, with application to the Mexican universal health insurance evaluation. Statistical Science, 24(1):29–53, February 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1255009008. See comments

[HS09, ZS09].

Imai:2009:RMP

[IKN09b] Kosuke Imai, Gary King, and Clayton Nall. Rejoinder: Matched pairs and the future of cluster-randomized experiments. Statistical Science, 24(1):65–72, February 2009. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1255009011.

Imai:2010:IIS

[IKY10] Kosuke Imai, Luke Keele, and Teppei Yamamoto. Identification, inference and sensitivity analysis for causal mediation effects. Statistical Science, 25(1):51–71, February 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1280841733.

Imbens:2014:R

[Imb14a] Guido Imbens. Rejoinder. Statistical Science, 29(3):375–379, August 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1411437518.

Imbens:2014:IVE

[Imb14b] Guido W. Imbens. Instrumental variables: An econometrician’s perspective. Statistical Science, 29(3):323–358, August 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1411437513.

IMS:1990:CDR

[IMS90] IMS Panel on Cross-Disciplinary Research in the Statistical Sciences. Cross-disciplinary research in the statistical sciences. Statistical Science, 5(1):121–146, February 1990. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012273.

Ionides:2011:DFM

[Ion11] Edward L. Ionides. Discussion of “Feature Matching in Time Series Modeling” by Y. Xia and H. Tong. Statistical Science, 26 (1):49–52, February 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1307626562.

Ivanova:2023:CQC

[IR23] Anastasia Ivanova and William F. Rosenberger. Comment: A quarter century of methodological research in responseadaptive randomization. Statistical Science, 38(2):209–211, May 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-2/
Comment--A-Quarter-Century-of-Methodological-Research-
in-Response/10.1214/23-STS865A.full.

Inacio:2022:CAR

[IRÁ22] Vanda Inácio and María Xosé Rodríguez-Álvarez. The covariate-adjusted ROC curve: The concept and its importance, review of inferential methods, and a new Bayesian estimator. Statistical Science, 37(4):541–561, November 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/journals/
statistical-science/volume-37/issue-4/The-Covariate-
Adjusted-ROC-Curve--The-Concept-and-Its/10.1214/21-
STS839.full.

Izenman:1987:PFO

[Ize87] Alan Julian Izenman. [Prediction of future observations in growth curve models]: Comment. Statistical Science, 2(4): 461–463, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013123.

Izenman:2001:SLA

[Ize01] Alan Julian Izenman. Statistical and legal aspects of the forensic study of illicit drugs. Statistical Science, 16(1):35–57, February 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/998929475.

Izenman:2021:RMT

[Ize21] Alan Julian Izenman. Random matrix theory and its applications. Statistical Science, 36(3):421–442, August 2021. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-3/Random-Matrix-Theory-and-
Its-Applications/10.1214/20-STS799.full.

Johnson:2014:SMU

[JAAH14] Sandra Johnson, Eva Abal, Kathleen Ahern, and Grant Hamilton. From science to management: Using Bayesian networks to learn about Lyngbya. Statistical Science, 29(1):36–41, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1399645726.

Jagers:1991:GSPa

[Jag91a] Peter Jagers. The growth and stabilization of populations. Statistical Science, 6(3):269–274, August 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011694.

Jagers:1991:GSPb

[Jag91b] Peter Jagers. [The growth and stabilization of populations]: Rejoinder. Statistical Science, 6(3):282–283, August 1991. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011698.

Jansen:2006:AID

[JBM^{+}06] Ivy Jansen, Caroline Beunckens, Geert Molenberghs, Geert Verbeke, and Craig Mallinckrodt. Analyzing incomplete discrete longitudinal clinical trial data. Statistical Science, 21 (1):52–69, February 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1149600846.

Jones:2005:ESC

[JCD^{+}05] Beatrix Jones, Carlos Carvalho, Adrian Dobra, Chris Hans, Chris Carter, and Mike West. Experiments in stochastic computation for high-dimensional graphical models. Statistical Science, 20(4):388–400, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1137076659.

Jahja:2022:RTE

[JCT22] Maria Jahja, Andrew Chin, and Ryan J. Tibshirani. Real-time estimation of COVID-19 infections: Deconvolution and sensor fusion. Statistical Science, 37(2):207–228, May 2022. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-37/issue-2/Real-Time-Estimation-of-COVID-
19-Infections--Deconvolution-and/10.1214/22-STS856.
full.

Jiang:2021:ICE

[JD21] Zhichao Jiang and Peng Ding. Identification of causal effects within principal strata using auxiliary variables. Statistical Science, 36(4):493–508, November 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-4/Identification-of-Causal-
Effects-Within-Principal-Strata-Using-Auxiliary-Variables/
10.1214/20-STS810.full.

Jennison:1990:BBC

[Jen90] C. Jennison. [Biostatistics and Bayes]: Comment. Statistical Science, 5(3):288–291, August 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012095.

Jensen:2019:CSE

[Jen19] David Jensen. Comment: Strengthening empirical evaluation of causal inference methods. Statistical Science, 34(1):77–81, February 2019. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1555056033.

Jennison:2023:CGS

[Jen23] Christopher Jennison. Comment: Group sequential designs with response-adaptive randomisation. Statistical Science, 38(2):219–223, May 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/journals/statistical-science/volume-
38/issue-2/Comment-Group-Sequential-Designs-with-Response-
Adaptive-Randomisation/10.1214/23-STS865D.full.

Jasra:2005:MCM

[JHS05] A. Jasra, C. C. Holmes, and D. A. Stephens. Markov chain Monte Carlo methods and the label switching problem in Bayesian mixture modeling. Statistical Science, 20(1):50–67, February 2005. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1118065042.

Jiang:2019:CEB

[Jia19] Wenhua Jiang. Comment: Empirical Bayes interval estimation. Statistical Science, 34(2):219–223, May 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1563501636. See

[Efr19a].

Jimenez:2011:FAV

[Jim11] Raúl Jiménez. Forensic analysis of the Venezuelan Recall Referendum. Statistical Science, 26(4):564–583, November 2011. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1330437936.

Jones:2008:CGS

[JJ08] Galin L. Jones and Alicia A. Johnson. Comment: Gibbs sampling, exponential families, and orthogonal polynomials. Statistical Science, 23(2):183–186, May 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1219339109. See [DKSC08a].

Johnson:2018:SCE

[JJ18] Andrew L. Johnson and Daniel R. Jiang. Shape constraints in economics and operations research. Statistical Science, 33(4):527–546, November 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1543482057.

Jeong:2017:SPM

[JJG17] Jaehong Jeong, Mikyoung Jun, and Marc G. Genton. Spherical process models for global spatial statistics. Statistical Science, 32(4):501–513, 11 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Johnson:2013:CWM

[JJN13] Alicia A. Johnson, Galin L. Jones, and Ronald C. Neath. Component-wise Markov chain Monte Carlo: Uniform and geometric ergodicity under mixing and composition. Statistical Science, 28(3):360–375, August 2013. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1377696941.

Johnstone:2013:MVE

[JL13] David Johnstone and Dennis Lindley. Mean–variance and expected utility: The Borch Paradox. Statistical Science, 28 (2):223–237, May 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1369147913.

Johnson:2016:MSH

[JLMD16] Devin S. Johnson, Jeff L. Laake, Sharon R. Melin, and Robert L. DeLong. Multivariate state hidden Markov models for mark–recapture data. Statistical Science, 31(2):233–244, May 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1464105040.

Jensen:2004:CDG

[JLZL04] Shane T. Jensen, X. Shirley Liu, Qing Zhou, and Jun S. Liu. Computational discovery of gene regulatory binding motifs: A Bayesian perspective. Statistical Science, 19(1):188–204, February 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1089808282.

Joe:1989:EVA

[Joe89] Harry Joe. [Extreme value analysis of environmental time series: An application to trend detection in ground-level ozone]: Comment. Statistical Science, 4(4):384–385, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012403.

Johnson:1988:BMT

[Joh88] Ralph H. Johnson. Biological markers in tort litigation. Statistical Science, 3(3):367–370, August 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012839.

Johnson:1997:ATG

[Joh97] Valen E. Johnson. An alternative to traditional GPA for evaluating student performance. Statistical Science, 12(4):251– 278, November 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1030037959.

Johnson:2007:CBC

[Joh07] Valen E. Johnson. Comment: Bayesian checking of the second levels of hierarchical models. Statistical Science, 22(3):353–358, August 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1199285034. See [BC07a].

Johnstone:2019:LBW

[Joh19] Iain M. Johnstone. Larry Brown’s work on admissibility. Statistical Science, 34(4):657–668, November 2019. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1578474030.

Jones:1991:CKR

[Jon91] M. C. Jones. [Choosing a kernel regression estimator]: Comment. Statistical Science, 6(4):427–430, November 1991. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011590.

Jongbloed:2019:CPG

[Jon19] Geurt Jongbloed. A conversation with Piet Groeneboom. Statistical Science, 34(1):156–168, February 2019. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1555056040.

Jorgensen:1989:URI

[Jor89] Bent Jorgensen. [The unifying role of iterative generalized least squares in statistical algorithms]: Comment. Statistical Science, 4(4):403–404, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012409.

Jordan:2004:GM

[Jor04] Michael I. Jordan. Graphical models. Statistical Science, 19(1): 140–155, February 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1089808279.

Josse:2018:ISS

[JR18] Julie Josse and Jerome P. Reiter. Introduction to the special section on missing data. Statistical Science, 33(2):139–141, May 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1525313138.

Jank:2006:FDA

[JS06a] Wolfgang Jank and Galit Shmueli. Functional data analysis in electronic commerce research. Statistical Science, 21 (2):155–166, May 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1154979818.

Jank:2006:SIS

[JS06b] Wolfgang Jank and Galit Shmueli. A special issue on statistical challenges and opportunities in electronic commerce research. Statistical Science, 21(2):113–115, May 2006. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1154979814.

Joffe:2007:DEI

[JSH07] Marshall M. Joffe, Dylan Small, and Chi-Yuan Hsu. Defining and estimating intervention effects for groups that will develop an auxiliary outcome. Statistical Science, 22(1):74–97, February 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1185975638.

Jennison:1990:SAI

[JT90] Christopher Jennison and Bruce W. Turnbull. Statistical approaches to interim monitoring of medical trials: A review and commentary. Statistical Science, 5(3):299–317, August 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012099.

Jiang:2004:IMI

[JT04] Wenxin Jiang and Bruce Turnbull. The indirect method: Inference based on intermediate statistics — a synthesis and examples. Statistical Science, 19(2):239–263, May 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1105714160.

Janicki:1994:COS

[JW94] Aleksander Janicki and Aleksander Weron. Can one see αstable variables and processes? Statistical Science, 9(1):109– 126, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010656.

Johnson:1989:DAC

[JY89] Mark E. Johnson and Donald Ylvisaker. [Design and analysis of computer experiments]: Comment. Statistical Science, 4 (4):428, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012416.

Kadane:1986:RMA

[Kad86] Joseph B. Kadane. [Regression models for adjusting the 1980 Census]: Comment. Statistical Science, 1(1):12–17, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013805. See [FN86a].

Kadane:1987:TPH

[Kad87a] Joseph B. Kadane. [Testing precise hypotheses]: Comment. Statistical Science, 2(3):347–348, August 1987. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013244.

Kadane:1987:UPA

[Kad87b] Joseph B. Kadane. [Uncertainty, policy analysis and statistics]: Comment. Statistical Science, 2(3):281–282, August 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013228.

Kadane:1990:QPE

[Kad90a] Joseph B. Kadane. [Quantifying probabilistic expressions]: Comment: Codifying chance. Statistical Science, 5(1):18–20, February 1990. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177012245.

Kadane:1990:WML

[Kad90b] Joseph B. Kadane. [The 1988 Wald memorial lectures: The present position in Bayesian statistics]: Comment. Statistical Science, 5(1):80–82, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012259.

Kadane:2009:BTE

[Kad09] Joseph B. Kadane. Bayesian thought in early modern detective stories: Monsieur Lecoq, C. Auguste Dupin and Sherlock Holmes. Statistical Science, 24(2):238–243, May 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1263478384.

Kafadar:2003:JTR

[Kaf03] Karen Kafadar. John Tukey and robustness. Statistical Science, 18(3):319–331, August 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1076102419.

Kalton:2011:DBM

[Kal11] Graham Kalton. Discussion of “Bayesian Models and Methods in Public Policy and Government Settings” by S. E. Fienberg. Statistical Science, 26(2):231–234, May 2011. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1312204012.  See

[Fie11a].

Karlin:1992:RFE

[Kar92] Samuel Karlin. R. A. Fisher and evolutionary theory. Statistical Science, 7(1):13–33, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011441.

Kass:1988:SMS

[Kas88] Robert E. Kass. [Saddlepoint methods and statistical inference]: Comment. Statistical Science, 3(2):234–236, May 1988. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012912.

Kass:1989:GAIa

[Kas89a] Robert E. Kass. The geometry of asymptotic inference. Statistical Science, 4(3):188–219, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012480.

Kass:1989:GAIb

[Kas89b] Robert E. Kass. [The geometry of asymptotic inference]: Rejoinder. Statistical Science, 4(3):233–234, August 1989. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012486.

Kass:1992:E

[Kas92a] Robert E. Kass. Editorial. Statistical Science, 7(1):1–2, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011438.

Kass:1992:Ia

[Kas92b] Robert E. Kass. In this issue. Statistical Science, 7(3):287–288, August 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011224.

Kass:1992:Ib

[Kas92c] Robert E. Kass. In this issue. Statistical Science, 7(4):403, November 1992. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177011121.

Kass:1993:Ia

[Kas93a] Robert E. Kass. In this issue. Statistical Science, 8(2):81, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010993.

Kass:1993:Ib

[Kas93b] Robert E. Kass. In this issue. Statistical Science, 8(3):203, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010886.

Kass:1993:Ic

[Kas93c] Robert E. Kass. In this issue. Statistical Science, 8(4):355, November 1993. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177010772.

Kass:1994:I

[Kas94] Robert E. Kass. In this issue. Statistical Science, 9(1):1, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010637.

Kass:1999:ISB

[Kas99] Robert E. Kass. Introduction to “Solving the Bible Code Puzzle” by Brendan McKay, Dror Bar-Natan, Maya Bar-Hillel and Gil Kalai. Statistical Science, 14(2):149, May 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1009212242.

Kass:2009:CIJ

[Kas09] Robert Kass. Comment: The importance of Jeffreys’s legacy. Statistical Science, 24(2):179–182, May 2009. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1263478376.

Kass:2010:CHS

[Kas10] Robert E. Kass. Comment: How should indirect evidence be used? Statistical Science, 25(2):166–169, May 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1290175838.

Kass:2011:R

[Kas11a] Robert E. Kass. Rejoinder. Statistical Science, 26(1):19–20, February 2011. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1307626559.

Kass:2011:SIB

[Kas11b] Robert E. Kass. Statistical inference: The big picture. Statistical Science, 26(1):1–9, February 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1307626554. See discussion

[Goo11, McC11, Ste11].

Katz:2002:SGW

[Kat02] Richard W. Katz. Sir Gilbert Walker and a connection between El Niño and statistics. Statistical Science, 17(2):97–112, May 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1023799000.

Kendall:2018:ETG

[KAX^{+}18] Michelle Kendall, Diepreye Ayabina, Yuanwei Xu, James Stimson, and Caroline Colijn. Estimating transmission from genetic and epidemiological data: A metric to compare transmission trees. Statistical Science, 33(1):70–85, 02 2018. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Kaye:1987:SPM

[Kay87] D. H. Kaye. [The statistical precision of medical screening procedures: Application to polygraph and AIDS antibodies test data]: Comment: The polygraph and the PVP. Statistical Science, 2(3):223–226, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013216.

Kaye:1991:IUD

[Kay91] D. H. Kaye. [Inferences using DNA profiling in forensic identification and paternity cases]: Comment: Uncertainty in DNA profile evidence. Statistical Science, 6(2):196–200, May 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011825.

Korn:1998:CPE

[KB98] Edward L. Korn and Sheldon Baumrind. Clinician preferences and the estimation of causal treatment differences. Statistical Science, 13(3):209–235, August 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1028905885.

Kahneman:2020:CLC

[KBH20] Daniel Kahneman and Maya Bar-Hillel. Comment: Laplace and cognitive illusions. Statistical Science, 35(2):171–172, May 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/

1591171223. See [MG20a].

Keiding:2014:SCC

[KC14] Niels Keiding and David Clayton. Standardization and control for confounding in observational studies: A historical perspective. Statistical Science, 29(4):529–558, November 2014. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1421330546.

Kantas:2015:PMP

[KDS^{+}15] Nikolas Kantas, Arnaud Doucet, Sumeetpal S. Singh, Jan Maciejowski, and Nicolas Chopin. On particle methods for parameter estimation in state-space models. Statistical Science, 30 (3):328–351, August 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1439220716.

Karavani:2019:CCI

[KEHSY19] Ehud Karavani, Tal El-Hay, Yishai Shimoni, and Chen Yanover. Comment: Causal inference competitions: Where should we aim? Statistical Science, 34(1):86–89, February 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1555056035.

Kendall:1986:SSS

[Ken86] David G. Kendall. [Size and shape spaces for Landmark data in two dimensions]: Comment. Statistical Science, 1(2):222–226, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013697.

Kendall:1989:SSTc

[Ken89a] David G. Kendall. [A survey of the statistical theory of shape]: Rejoinder. Statistical Science, 4(2):116–120, May 1989. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012589.

Kendall:1989:SSTa

[Ken89b] David G. Kendall. A survey of the statistical theory of shape. Statistical Science, 4(2):87–99, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012582.

Kendall:1989:SSTb

[Ken89c] Wilfrid S. Kendall. [A survey of the statistical theory of shape]: Comment. Statistical Science, 4(2):111–113, May 1989. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012586.

Kendall:1991:KRH

[Ken91] David G. Kendall. Kolmogorov as i remember him. Statistical Science, 6(3):303–312, August 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011700.

Kettenring:2001:CRG

[Ket01] Jon R. Kettenring. A conversation with Ramanathan Gnanadesikan. Statistical Science, 16(3):295–309, August 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1009213730.

Kass:1989:ITP

[KG89] Robert E. Kass and Joel B. Greenhouse. [Investigating therapies of potentially great benefit: ECMO]: Comment: A Bayesian perspective. Statistical Science, 4(4):310–317, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012386.

Kramer:2005:HCS

[KG05] Walter Krämer and Gerd Gigerenzer. How to confuse with statistics or: The use and misuse of conditional probabilities. Statistical Science, 20(3):223–230, August 2005. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1124891288.

Katzfuss:2021:GFV

[KG21] Matthias Katzfuss and Joseph Guinness. A general framework for Vecchia approximations of Gaussian processes. Statistical Science, 36(1):124–141, February 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-1/A-General-Framework-for-Vecchia-
Approximations-of-Gaussian-Processes/10.1214/19-STS755.
full.

Kitagawa:2014:IVB

[Kit14] Toru Kitagawa. Instrumental variables before and LATEr. Statistical Science, 29(3):359–362, August 2014. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1411437514.

Kurland:2009:LDF

[KJED09] Brenda F. Kurland, Laura L. Johnson, Brian L. Egleston, and Paula H. Diehr. Longitudinal data with follow-up truncated by death: Match the analysis method to research aims. Statistical Science, 24(2):211–222, May 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1263478382.

Krivitsky:2015:QES

[KK15] Pavel N. Krivitsky and Eric D. Kolaczyk. On the question of effective sample size in network modeling: An asymptotic inquiry. Statistical Science, 30(2):184–198, May 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1433341477.

Koul:2016:CEV

[KK16] Hira L. Koul and Roger Koenker. A conversation with Estate V. Khmaladze. Statistical Science, 31(3):453–464, August 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1475001239.

King:2008:VAM

[KL08] Gary King and Ying Lu. Verbal autopsy methods with multiple causes of death. Statistical Science, 23(1):78–91, February 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1215441286.

Kooperberg:2009:SAS

[KLDR09] Charles Kooperberg, Michael LeBlanc, James Y. Dai, and Indika Rajapakse. Structures and assumptions: Strategies to harness gene–gene and gene–environment interactions in GWAS. Statistical Science, 24(4):472–488, November 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1271770343.

Klein:1991:SS

[Kle91] Lawrence Klein. The Statistics Seminar, MIT, 1942–1943. Statistical Science, 6(4):320–330, November 1991. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011569.

Keller-McNulty:1991:EAM

[KM91] Sallie Keller-McNulty. [Enhancing access to microdata while protecting confidentiality: Prospects for the future]: Comment. Statistical Science, 6(3):234–235, August 1991. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011683.

Kenward:1998:LBF

[KM98] M. G. Kenward and G. Molenberghs. Likelihood based frequentist inference when data are missing at random. Statistical Science, 13(3):236–247, August 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1028905886.

Koenker:2018:SCD

[KM18a] Roger Koenker and Ivan Mizera. Shape constrained density estimation via penalized Rényi divergence. Statistical Science, 33(4):510–526, November 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1543482056.

Kypraios:2018:ISS

[KM18b] Theodore Kypraios and Vladimir N. Minin. Introduction to the special section on inference for infectious disease dynamics. Statistical Science, 33(1):1–3, 02 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Ke:2019:UFC

[KMR^{+}19] Yuan Ke, Stanislav Minsker, Zhao Ren, Qiang Sun, and Wen-Xin Zhou. User-friendly covariance estimation for heavytailed distributions. Statistical Science, 34(3):454–471, August 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1570780979.

Khuri:2006:DIG

[KMSG06] André I. Khuri, Bhramar Mukherjee, Bikas K. Sinha, and Malay Ghosh. Design issues for generalized linear models: A review. Statistical Science, 21(3):376–399, August 2006. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1166642442.

Kempthorne:1991:RHD

[KMSZ91] Peter Kempthorne, Nitis Mukhopadhyay, Pranab K. Sen, and Shelemyahu Zacks. Research — how to do it: A panel discussion. Statistical Science, 6(2):149–162, May 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011820.

Keller-McNulty:2006:R

[KMWAC06] Sallie Keller-McNulty, Alyson Wilson, and Christine Anderson-Cook. Reliability. Statistical Science, 21(4):427, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177334514.

Kypraios:2018:BNS

[KN18] Theodore Kypraios and Philip D. O. Neill. Bayesian nonparametrics for stochastic epidemic models. Statistical Science, 33 (1):44–56, 02 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Koltchinskii:2019:CDD

[KNR19] Vladimir Koltchinskii, Richard Nickl, and Philippe Rigollet. A conversation with Dick Dudley. Statistical Science, 34(1):169– 175, February 2019. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1555056041.

Koehler:2006:CEE

[Koe06] Andrew Koehler. Comment: Expert elicitation for reliable system design. Statistical Science, 21(4):454–455, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177334517. See [BQW06a].

Korn:2006:CCI

[Kor06] Edward L. Korn. Comment: Causal inference in the medical area. Statistical Science, 21(3):310–312, August 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1166642431. See comments [Han06a].

Koenker:1988:RBR

[KP88] Roger Koenker and Stephen Portnoy. [Rank-based robust analysis of linear models. I. Exposition and review]: Comment. Statistical Science, 3(2):259–261, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012917.

Kircher:1987:SPM

[KR87] John C. Kircher and David C. Raskin. [The statistical precision of medical screening procedures: Application to polygraph and AIDS antibodies test data]: Comment: Base rates and the statistical precision. Statistical Science, 2(3):226–228, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013217.

Kreiner:1992:SEI

[Kre92] Svend Kreiner. [A survey of exact inference for contingency tables]: Comment: Exact inference in multidimensional tables. Statistical Science, 7(1):163–165, February 1992. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011458.

Kroonenberg:1992:MMA

[Kro92] Pieter M. Kroonenberg. [Multilinear models: Applications in spectroscopy]: Comment: PARAFAC in three-way land. Statistical Science, 7(3):312–314, August 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011227.

Kruskal:1990:QPE

[Kru90] William Kruskal. [Quantifying probabilistic expressions]: Comment. Statistical Science, 5(1):20–21, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012246.

Karon:1993:BHI

[KS93] John M. Karon and Glen A. Satten. [Backcalculation of HIV infection rates]: Comment. Statistical Science, 8(2):109–112, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010999.

Kang:2007:DDR

[KS07a] Joseph D. Y. Kang and Joseph L. Schafer. Demystifying double robustness: A comparison of alternative strategies for estimating a population mean from incomplete data. Statistical Science, 22(4):523–539, November 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1207580167. See comments

[RM07, TD07].

Kang:2007:RDD

[KS07b] Joseph D. Y. Kang and Joseph L. Schafer. Rejoinder: Demystifying double robustness: A comparison of alternative strategies for estimating a population mean from incomplete data. Statistical Science, 22(4):574–580, November 2007. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1207580172.

Kaldor:1988:MMQ

[KT88] J. Kaldor and L. Tomatis. [From mouse-to-man: The quantitative assessment of cancer risks]: Comment: The use of animal experiments in cancer risk assessment. Statistical Science, 3 (1):41–43, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012997.

Kunsch:2008:CNL

[Kün08] Hans R. Künsch. Comment: The 2005 Neyman lecture: Dynamic indeterminism in science. Statistical Science, 23(1): 65–68, February 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1215441283. See [Bri08a].

Kvam:2006:CHM

[Kva06] Paul Kvam. A conversation with Harry Martz. Statistical Science, 21(4):578–585, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177334530.

Kuikka:2014:EBI

[KVP^{+}14] Sakari Kuikka, Jarno Vanhatalo, Henni Pulkkinen, Samu Mäntyniemi, and Jukka Corander. Experiences in Bayesian inference in Baltic salmon management. Statistical Science, 29 (1):42–49, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1399645727.

Kraft:2009:RGW

[KZI09] Peter Kraft, Eleftheria Zeggini, and John P. A. Ioannidis. Replication in genome-wide association studies. Statistical Science, 24(4):561–573, November 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1271770349.

Ko:2022:HPS

[KZZW22] Seyoon Ko, Hua Zhou, Jin J. Zhou, and Joong-Ho Won. High-performance statistical computing in the computing environments of the 2020s. Statistical Science, 37(4):494–518, November 2022. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
journals/statistical-science/volume-37/issue-4/High-
Performance-Statistical-Computing-in-the-Computing-
Environments-of-the/10.1214/21-STS835.full.

Lagakos:1992:ETI

[Lag92] Stephen Lagakos. [Evaluating therapeutic interventions: Some issues and experiences]: Comment. Statistical Science, 7(4): 449–450, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011133.

Lagarias:1993:PN

[Lag93] Jeffrey C. Lagarias. Pseudorandom numbers. Statistical Science, 8(1):31–39, February 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011081;   http://
www.jstor.org/stable/pdfplus/2246038.pdf.

Lahiri:2003:IBS

[Lah03] P. Lahiri. On the impact of bootstrap in survey sampling and small-area estimation. Statistical Science, 18(2):199–210, May 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1063994975.

Laird:1989:CFD

[Lai89] Nan M. Laird. A conversation with F. N. David. Statistical Science, 4(3):235–246, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012487.

Laird:2010:EAG

[Lai10] Nan M. Laird. The EM algorithm in genetics, genomics and public health. Statistical Science, 25(4):450–457, November 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1300108230.

Laird:2019:CBO

[Lai19] Nan Laird. Comment: Bayes, oracle Bayes, and empirical Bayes. Statistical Science, 34(2):206–208, May 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1563501633. See

[Efr19a].

Lange:1991:IUD

[Lan91] Kenneth Lange. [Inferences using DNA profiling in forensic identification and paternity cases]: Comment. Statistical Science, 6(2):190–192, May 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011823.

Lang:2014:ERA

[Lan14] Duncan Temple Lang. Enhancing R with advanced compilation tools and methods. Statistical Science, 29(2):181–200, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1408368570.

Lang:2015:CLT

[Lan15] Joseph B. Lang. A closer look at testing the “No-Treatment-Effect” hypothesis in a comparative experiment. Statistical Science, 30(3):352–371, August 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1439220717.

Laplace:1986:MPC

[Lap86] Pierre Simon Laplace. Memoir on the probability of the causes of events. Statistical Science, 1(3):364–378, August 1986. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013621.

Larsen:2011:DCB

[Lar11] Michael D. Larsen. Discussion of “Calibrated Bayes, for Statistics in General, and Missing Data in Particular” by R. J. A. Little. Statistical Science, 26(2):175–178, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1312204003. See [Lar11].

Lawrance:2016:CCC

[Law16] Anthony J. Lawrance. Chaos communication: A case of statistical engineering. Statistical Science, 31(4):558–577, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1484816586.

Lawless:2019:CMA

[Law19] Jerald F. Lawless. Comment on models as approximations, parts I and II, by Buja et al. Statistical Science, 34(4):569–571, November 2019. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1578474019. See [BBB^{+}19, BBK^{+}19].

Lavine:2022:CCL

[LB22] Michael Lavine and Jan F. Bjørnstad. Comments on confidence as likelihood by Pawitan and Lee in Statistical Science, November 2021. Statistical Science, 37(4):625–627, November 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-4/
Comments-on-Confidence-as-Likelihood-by-Pawitan-and-
Lee-in/10.1214/22-STS862.full. See [PL21, PL22].

Lombardo:2021:SNS

[LBK21] Rosaria Lombardo, Eric J. Beh, and Pieter M. Kroonenberg. Symmetrical and non-symmetrical variants of threeway correspondence analysis for ordered variables. Statistical Science, 36(4):542–561, November 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-4/Symmetrical-and-Non-symmetrical-
Variants-of-Three-Way-Correspondence-Analysis/10.1214/
20-STS814.full.

Levine:2008:CRS

[LC08] Richard A. Levine and George Casella. Comment: On random scan Gibbs samplers. Statistical Science, 23(2):192–195, May 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1219339111.

Lin:2024:CRJ

[LC24] Xihong Lin and Nilanjan Chatterjee. A conversation with Raymond J. Carroll. Statistical Science, 39(3):508–517, August 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-3/
A-Conversation-with-Raymond-J-Carroll/10.1214/24-STS931.
full.

Lin:2013:LSS

[LCL13] Ming Lin, Rong Chen, and Jun S. Liu. Lookahead strategies for sequential Monte Carlo. Statistical Science, 28(1):69–94, February 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1359468409.

Li:2021:CRM

[LCX21] Xiaodong Li, Yudong Chen, and Jiaming Xu. Convex relaxation methods for community detection. Statistical Science, 36(1):2–15, February 2021. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-1/Convex-Relaxation-Methods-
for-Community-Detection/10.1214/19-STS715.full.

Linero:2018:BAM

[LD18] Antonio R. Linero and Michael J. Daniels. Bayesian approaches for missing not at random outcome data: The role of identifying restrictions. Statistical Science, 33(2):198–213, May 2018. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/euclid.ss/1525313142.

Luo:2024:VSU

[LD24] Chuji Luo and Michael J. Daniels. Variable selection using Bayesian additive regression trees. Statistical Science, 39(2): 286–304, May 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-2/
Variable-Selection-Using-Bayesian-Additive-Regression-
Trees/10.1214/23-STS900.full.

LeCam:1986:CLTa

[Le 86a] L. Le Cam. The central limit theorem around 1935. Statistical Science, 1(1):78–91, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1177013818. See comments

[Tro86, Doo86, Pol86, Le 86b].

LeCam:1986:CLTb

[Le 86b] L. Le Cam. [The central limit theorem around 1935]: Rejoinder. Statistical Science, 1(1):95–96, February 1986. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013822.  See

[Le 86a].

Lehmann:1990:MSV

[Leh90a] E. L. Lehmann. Model specification: The views of Fisher and Neyman, and later developments. Statistical Science, 5 (2):160–168, May 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012164.

Lehmann:1990:WML

[Leh90b] E. L. Lehmann. [The 1988 Wald memorial lectures: The present position in Bayesian statistics]: Comment. Statistical Science, 5(1):82–83, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012260.

Lehmann:1993:MEC

[Leh93] E. L. Lehmann. Mentors and early collaborators: Reminiscences from the years 1940–1956 with an epilogue. Statistical Science, 8(3):331–341, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177010910.

Lehmann:1997:TSH

[Leh97] E. L. Lehmann. Testing Statistical Hypotheses: the story of a book. Statistical Science, 12(1):48–52, February 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1029963261.

Lehmann:1999:SSS

[Leh99] E. L. Lehmann. “Student” and small-sample theory. Statistical Science, 14(4):418–426, November 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1009212520.

Lele:2003:IBE

[Lel03] Subhash R. Lele. Impact of bootstrap on the estimating functions. Statistical Science, 18(2):185–190, May 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1063994973.

Lempert:1994:DFR

[Lem94] Richard Lempert. [DNA fingerprinting: A review of the controversy]: Comment: Theory and practice in DNA fingerprinting. Statistical Science, 9(2):255–258, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177010491.

Letac:2008:CLP

[Let08] Gérard Letac. Comment: Lancaster probabilities and Gibbs sampling. Statistical Science, 23(2):187–191, May 2008. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1219339110.

Levine:1991:ESR

[Lev91] Robert J. Levine. [Ethics and statistics in randomized clinical trials]: Comment. Statistical Science, 6(1):71–74, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011938.

Lewontin:1994:DFR

[Lew94] R. C. Lewontin. [DNA fingerprinting: A review of the controversy]: Comment: The use of DNA profiles in forensic contexts. Statistical Science, 9(2):259–262, May 1994. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177010492.

Lopez:2017:ECE

[LG17] Michael J. Lopez and Roee Gutman. Estimation of causal effects with multiple treatments: A review and new ideas. Statistical Science, 32(3):432–454, August 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1504253125.

Lyne:2015:RRE

[LGA^{+}15] Anne-Marie Lyne, Mark Girolami, Yves Atchadé, Heiko Strathmann, and Daniel Simpson. On Russian roulette estimates for Bayesian inference with doubly-intractable likelihoods. Statistical Science, 30(4):443–467, November 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1449670853.

Lin:1995:CTI

[LGF95] Tony Lin, Lois Swirsky Gold, and David Freedman. Carcinogenicity tests and interspecies concordance. Statistical Science, 10(4):337–353, November 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177009868.

Li:1999:CCL

[Li99] Zhaohai Li. A conversation with Chin Long Chiang. Statistical Science, 14(4):457–470, November 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1009212522.

Li:2007:CFLa

[Li07] Bing Li. Comment: Fisher lecture: Dimension reduction in regression. Statistical Science, 22(1):32–35, February 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1185975633. See [Coo07a].

Li:2020:CSD

[Li20] Fan Li. Comment: Stabilizing the doubly-robust estimators of the average treatment effect under positivity violations. Statistical Science, 35(3):503–510, August 2020. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-35/issue-3/Comment--Stabilizing-the-
Doubly-Robust-Estimators-of-the-Average/10.1214/20-
STS774.full.

Lilienfeld:1988:CRM

[Lil88] David E. Lilienfeld. Changing research methods in environmental epidemiology. Statistical Science, 3(3):275–280, August 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012827.

Lindley:1986:SRC

[Lin86] D. V. Lindley. [Savage revisited]: Comment. Statistical Science, 1(4):486–488, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177013519.

Lindley:1987:PESa

[Lin87a] Dennis V. Lindley. [Probabilistic expert systems in medicine: Practical issues in handling uncertainty]: Comment: A tale of two Wells. Statistical Science, 2(1):38–40, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013432. See [Spi87a].

Lindley:1987:PESb

[Lin87b] Dennis V. Lindley. [Probabilistic expert systems in medicine: Practical issues in handling uncertainty]: Rejoinder. Statistical Science, 2(1):42–43, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013435. See [Spi87a].

Lindley:1987:PAT

[Lin87c] Dennis V. Lindley. The probability approach to the treatment of uncertainty in artificial intelligence and expert systems. Statistical Science, 2(1):17–24, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013427.

Lindley:1990:WMLa

[Lin90a] Dennis V. Lindley. The 1988 Wald memorial lectures: The present position in Bayesian statistics. Statistical Science, 5 (1):44–65, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012253.

Lindley:1990:WMLb

[Lin90b] Dennis V. Lindley. [The 1988 Wald memorial lectures: The present position in Bayesian statistics]: Rejoinder. Statistical Science, 5(1):85–89, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012262.

Lindley:1991:ESR

[Lin91] Foster Lindley. [Ethics and statistics in randomized clinical trials]: Comment: Personal and impersonal care. Statistical Science, 6(1):74–77, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011939.

Lindley:1998:DAB

[Lin98] Dennis V. Lindley. Decision analysis and bioequivalence trials. Statistical Science, 13(2):136–141, May 1998. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1028905932.

Lindqvist:2006:SMA

[Lin06] Bo Henry Lindqvist. On the statistical modeling and analysis of repairable systems. Statistical Science, 21(4):532–551, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177334528.

Lindquist:2008:SAF

[Lin08] Martin A. Lindquist. The statistical analysis of fMRI data. Statistical Science, 23(4):439–464, November 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1242049389.

Lindley:2009:C

[Lin09] Dennis Lindley. Comment. Statistical Science, 24(2):183–184, May 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1263478377.

Lindgren:2019:GIR

[Lin19] Georg Lindgren. Gaussian integrals and Rice series in crossing distributions-to compute the distribution of maxima and other features of Gaussian processes. Statistical Science, 34(1):100– 128, February 2019. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1555056038.

Linn:2020:MTR

[Lin20] Kristin A. Linn. Moving toward rigorous evaluation of mobile health interventions. Statistical Science, 35(3):394–395, August 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/journals/
statistical-science/volume-35/issue-3/Moving-Toward-
Rigorous-Evaluation-of-Mobile-Health-Interventions/
10.1214/20-STS781.full.

Lin:2022:LLC

[Lin22] Xihong Lin. Lessons learned from the COVID-19 pandemic: a statistician’s reflection. Statistical Science, 37(2):278–283, May 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-2/
Lessons-Learned-from-the-COVID-19-Pandemic--A-Statisticians/
10.1214/22-STS860.full.

Little:2007:CSS

[Lit07] Roderick J. Little. Comment: Struggles with survey weighting and regression modeling. Statistical Science, 22(2):171–174, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1190905514. See [Gel07c].

Little:2011:CBS

[Lit11a] Roderick Little. Calibrated Bayes, for statistics in general, and missing data in particular. Statistical Science, 26(2):162–174, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.

ss/1312204002. See discussion [Lar11, Sch11].

Little:2011:R

[Lit11b] Roderick Little. Rejoinder. Statistical Science, 26(2):185–186, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1312204005.

Little:2019:CMA

[Lit19] Roderick J. Little. Comment: “Models as Approximations I: Consequences Illustrated with Linear Regression” by A. Buja, R. Berk, L. Brown, E. George, E. Pitkin, L. Zhan and K. Zhang. Statistical Science, 34(4):580–583, November 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/

1578474022. See [BBB^{+}19, BBK^{+}19].

Lei:2020:PTR

[LK20] Jing Lei and Joseph B. Kadane. On the probability that two random integers are coprime. Statistical Science, 35(2):272– 279, May 2020. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1591171231.

Lindsay:2004:RFS

[LKS04] Bruce G. Lindsay, Jon Kettenring, and David O. Siegmund. A report on the future of statistics. Statistical Science, 19(3): 387–413, August 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1110999308.

Laird:1987:PFO

[LL87] Nan Laird and Nick Lange. [Prediction of future observations in growth curve models]: Comment. Statistical Science, 2(4): 451–454, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013121.

Lyberg:1994:CWR

[LL94] Lars Lyberg and Sixten Lundstrom. [Can we reach consensus on census adjustment?]: Comment. Statistical Science, 9(4): 515–517, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010264.

Lindsay:1995:IBE

[LL95] Bruce G. Lindsay and Bing Li. [Inference based on estimating functions in the presence of nuisance parameters]: Comment.

Statistical Science, 10(2):175–177, May 1995. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010030.

Li:2004:NNT

[LL04] Jun Li and Regina Y. Liu. New nonparametric tests of multivariate locations and scales using data depth. Statistical Science, 19(4):686–696, November 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1113832733.

Larsen:2007:CBC

[LL07] Michael D. Larsen and Lu Lu. Comment: Bayesian checking of the second level of hierarchical models: Cross-validated posterior predictive checks using discrepancy measures. Statistical Science, 22(3):359–362, August 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1199285035. See comments

[BC07a].

Laird:2009:RFB

[LL09a] Nan M. Laird and Christoph Lange. The role of familybased designs in genome-wide association studies. Statistical Science, 24(4):388–397, November 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1271770338.

Lindsay:2009:MAT

[LL09b] Bruce Lindsay and Jiawei Liu. Model assessment tools for a model false world. Statistical Science, 24(3):303–318, August 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1270041257.

Lehmann:2009:CCS

[LLJ09] Sune Lehmann, Benny E. Lautrup, and Andrew D. Jackson. Comment: Citation statistics. Statistical Science, 24 (1):17–20, February 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1255009004. See [Hal09].

Landon:2011:PPP

[LLS11] Joshua Landon, Frank X. Lee, and Nozer D. Singpurwalla. A problem in particle physics and its Bayesian analysis. Statistical Science, 26(3):352–368, August 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1320066925.

Lewandowski:2010:PEE

[LLV10] Andrew Lewandowski, Chuanhai Liu, and Scott Vander Wiel. Parameter expansion and efficient inference. Statistical Science, 25(4):533–544, November 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1300108236.

Lafferty:2012:SNG

[LLW12] John Lafferty, Han Liu, and Larry Wasserman. Sparse nonparametric graphical models. Statistical Science, 27(4):519– 537, November 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1356098554.

Lawrence:2014:SGR

[LM14a] Michael Lawrence and Martin Morgan. Scalable genomics with R and Bioconductor. Statistical Science, 29(2):214–226, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1408368572.

Li:2014:CDB

[LM14b] Fan Li and Fabrizia Mealli. A conversation with Donald B. Rubin. Statistical Science, 29(3):439–457, August 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1411437523.

Liu:2021:CSU

[LM21] Chuanhai Liu and Ryan Martin. Comment: Settle the unsettling: An inferential models perspective. Statistical Science, 36(2):196–200, May 2021. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-2/Comment-Settle-the-Unsettling-
An-Inferential-Models-Perspective/10.1214/21-STS765B.
full.

Lee:2012:RCS

[LMJ12] Yoonkyung Lee, Steven N. MacEachern, and Yoonsuh Jung. Regularization of case-specific parameters for robustness and efficiency. Statistical Science, 27(3):350–372, August 2012. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1346849944.

Lawrence:2006:CMN

[LMN06] Earl Lawrence, George Michailidis, and Vijayan N. Nair. Comment: Monitoring networked applications with incremental quantile estimation. Statistical Science, 21(4):479–482, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.

ss/1177334522. See [CJLV06a].

Liu:2020:NAF

[LMR20a] Lin Liu, Rajarshi Mukherjee, and James M. Robins. On nearly assumption-free tests of nominal confidence interval coverage for causal parameters estimated by machine learning. Statistical Science, 35(3):518–539, August 2020. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-35/issue-3/On-Nearly-Assumption-Free-
Tests-of-Nominal-Confidence-Interval-Coverage/10.1214/

20-STS786.full. See rejoinder [LMR20b].

Liu:2020:RNA

[LMR20b] Lin Liu, Rajarshi Mukherjee, and James M. Robins. Rejoinder: On nearly assumption-free tests of nominal confidence interval coverage for causal parameters estimated by machine learning. Statistical Science, 35(3):545–554, August 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/journals/
statistical-science/volume-35/issue-3/Rejoinder--On-
nearly-assumption-free-tests-of-nominal-confidence/
10.1214/20-STS804.full. See [LMR20a].

Liang:2007:UUD

[LMW07] Feng Liang, Sayan Mukherjee, and Mike West. The use of unlabeled data in predictive modeling. Statistical Science, 22 (2):189–205, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1190905518.

Lee:2004:CMM

[LN04] Youngjo Lee and John A. Nelder. Conditional and marginal models: Another view. Statistical Science, 19(2):219–238, May 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1105714159.

Li:2007:CFLb

[LN07] Lexin Li and Christopher J. Nachtsheim. Comment: Fisher lecture: Dimension reduction in regression. Statistical Science, 22(1):36–39, February 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1185975634. See [Coo07a].

Lee:2009:LIM

[LN09a] Youngjo Lee and John A. Nelder. Likelihood inference for models with unobservables: Another view. Statistical Science, 24 (3):255–269, August 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1270041252.

Lee:2009:RLI

[LN09b] Youngjo Lee and John A. Nelder. Rejoinder: Likelihood inference for models with unobservables another view. Statistical Science, 24(3):294–302, August 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1270041256.

Lohr:2007:CSS

[Loh07] Sharon L. Lohr. Comment: Struggles with survey weighting and regression modeling. Statistical Science, 22(2):175–178, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1190905515. See [Gel07c].

Louis:1992:ETI

[Lou92] Thomas A. Louis. [Evaluating therapeutic interventions: Some issues and experiences]: Comment. Statistical Science, 7(4): 450–452, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011134.

Louis:2009:DLI

[Lou09] Thomas A. Louis. Discussion of likelihood inference for models with unobservables: Another view. Statistical Science, 24(3): 270–272, August 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1270041253.

Louis:2019:CBO

[Lou19] Thomas A. Louis. Comment: Bayes, oracle Bayes, and empirical Bayes. Statistical Science, 34(2):202–205, May 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/

1563501632. See [Efr19a].

Lahiri:2011:DER

[LP11a] P. Lahiri and Santanu Pramanik. Discussion of “Estimating Random Effects via Adjustment for Density Maximization” by C. Morris and R. Tang. Statistical Science, 26(2):291–295, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1312204022. See [MT11].

Lijoi:2011:CER

[LP11b] Antonio Lijoi and Igor Prünster. A conversation with Eugenio Regazzini. Statistical Science, 26(4):647–672, November 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1330437943.

Leeb:2015:VCI

[LPE15] Hannes Leeb, Benedikt M. Pötscher, and Karl Ewald. On various confidence intervals post-model-selection. Statistical Science, 30(2):216–227, May 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1433341479.

Laird:1988:SMF

[LPT88] Nan Laird, G. P. Patil, and C. Taillie. [Selection models and the file drawer problem]: Comment. Statistical Science, 3(1): 126–128, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013015.

Lindsay:2003:IFQ

[LQ03] Bruce G. Lindsay and Annie Qu. Inference functions and quadratic score tests. Statistical Science, 18(3):394–410, August 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1076102427.

Lee:2013:DPP

[LQMT13] Jaeyong Lee, Fernando A. Quintana, Peter Müller, and Lorenzo Trippa. Defining predictive probability functions for species sampling models. Statistical Science, 28(2):209–222, May 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1369147912.

Leurgans:1992:MMA

[LR92a] Sue Leurgans and Robert T. Ross. Multilinear models: Applications in spectroscopy. Statistical Science, 7(3):289–310, August 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011225.

Leurgens:1992:MMA

[LR92b] Sue Leurgens and Robert T. Ross. [Multilinear models: Applications in spectroscopy]: Rejoinder. Statistical Science, 7 (3):315–319, August 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011229.

Lohr:2017:CSD

[LR17] Sharon L. Lohr and Trivellore E. Raghunathan. Combining survey data with other data sources. Statistical Science, 32 (2):293–312, May 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1494489817.

Liang:2004:SMD

[LRdP^{+}04] Chyng-Lan Liang, John A. Rice, Imke de Pater, Charles Alcock, Tim Axelrod, Andrew Wang, and Stuart Marshall. Statistical methods for detecting stellar occultations by Kuiper belt objects: The Taiwanese–American occultation survey. Statistical Science, 19(2):265–274, May 2004. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1105714162.

Vecchia:2023:SCB

[LRI23] Davide La Vecchia, Elvezio Ronchetti, and Andrej Ilievski. On some connections between Esscher’s tilting, saddlepoint approximations, and optimal transportation: a statistical perspective. Statistical Science, 38(1):30–51, February 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-1/
On-Some-Connections-Between-Esschers-Tilting-Saddlepoint-
Approximations-and-Optimal/10.1214/21-STS847.full.

Loh:2017:APE

[LRR17] Wen Wei Loh, Thomas S. Richardson, and James M. Robins. An apparent paradox explained. Statistical Science, 32(3):356– 361, August 2017. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/

euclid.ss/1504253120. See rejoinder [Din17b].

Lai:1986:CHR

[LS86] Tze Leung Lai and David Siegmund. The contributions of Herbert Robbins to mathematical statistics. Statistical Science, 1 (2):276–284, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013708.

Lindley:2002:ECC

[LS02] Dennis V. Lindley and Nozer D. Singpurwalla. On exchangeable, causal and cascading failures. Statistical Science, 17 (2):209–219, May 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1030550862.

Lauritzen:2003:GMG

[LS03] Steffen L. Lauritzen and Nuala A. Sheehan. Graphical models for genetic analyses. Statistical Science, 18(4):489–514, November 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1081443232.

Lahiri:2011:I

[LS11] P. Lahiri and Eric Slud. Introduction. Statistical Science, 26(2):161, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1312204001.

Lumley:2017:FRM

[LS17] Thomas Lumley and Alastair Scott. Fitting regression models to survey data. Statistical Science, 32(2):265–278, May 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1494489815.

Li:2023:RBT

[LS23] Xinran Li and Dylan S. Small. Randomization-based test for censored outcomes: a new look at the logrank test. Statistical Science, 38(1):92–107, February 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-1/Randomization-Based-Test-
for-Censored-Outcomes--A-New-Look/10.1214/22-STS851.
full.

Lad:2015:ECD

[LSA15] Frank Lad, Giuseppe Sanfilippo, and Gianna Agrò. Extropy: Complementary dual of entropy. Statistical Science, 30(1): 40–58, February 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1425492439.

Liu:2019:SAZ

[LSS^{+}19] Lei Liu, Ya-Chen Tina Shih, Robert L. Strawderman, Daowen Zhang, Bankole A. Johnson, and Haitao Chai. Statistical analysis of zero-inflated nonnegative continuous data: A review. Statistical Science, 34(2):253–279, May 2019. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1563501641.

Lu:2021:CTL

[LSY21] Ying Lu, Dylan S. Small, and Zhiliang Ying. A conversation with Tze Leung Lai. Statistical Science, 36(1):158–167, February 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-36/issue-1/
A-Conversation-with-Tze-Leung-Lai/10.1214/20-STS775.
full.

Leonard:1990:PLR

[LTH90] Tom Leonard, Kam-Wah Tsui, and John S. J. Hsu. [Predictive likelihood: A review]: Comment. Statistical Science, 5 (2):259–262, May 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012177.

Lerch:2017:FDE

[LTRG17] Sebastian Lerch, Thordis L. Thorarinsdottir, Francesco Ravazzolo, and Tilmann Gneiting. Forecaster’s dilemma: Extreme events and forecast evaluation. Statistical Science, 32(1):106– 127, February 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1491465630.

Lubinsky:1991:CTF

[Lub91] David J. Lubinsky. Comment: Two functional programming environments for statistics — Lisp-Stat and S. Statistical Science, 6(4):352–360, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011575.

Lin:1989:ITP

[LW89] D. Y. Lin and L. J. Wei. [Investigating therapies of potentially great benefit: ECMO]: Comment. Statistical Science, 4(4): 324–325, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012390.

Lin:1992:SEI

[LW92] D. Y. Lin and L. J. Wei. [A survey of exact inference for contingency tables]: Comment. Statistical Science, 7(1):166– 167, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011459.

Lee:2006:TRS

[LW06] Mei-Ling Ting Lee and G. A. Whitmore. Threshold regression for survival analysis: Modeling event times by a stochastic process reaching a boundary. Statistical Science, 21(4): 501–513, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177334526.

Lai:2021:SAS

[LY21] Tze Leung Lai and Hongsong Yuan. Stochastic approximation: From statistical origin to big-data, multidisciplinary applications. Statistical Science, 36(2):291–302, May 2021. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-2/Stochastic-Approximation--
From-Statistical-Origin-to-Big-Data-Multidisciplinary/
10.1214/20-STS784.full.

Li:2022:BAR

[LYL22] Xinran Li, Dingdong Yi, and Jun S. Liu. Bayesian analysis of rank data with covariates and heterogeneous rankers. Statistical Science, 37(1):1–23, February 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-37/issue-1/Bayesian-Analysis-of-Rank-
Data-with-Covariates-and-Heterogeneous-Rankers/10.1214/
20-STS818.full.

Liang:1995:IBEa

[LZ95a] Kung-Yee Liang and Scott L. Zeger. Inference based on estimating functions in the presence of nuisance parameters. Statistical Science, 10(2):158–173, May 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177010028.

Liang:1995:IBEb

[LZ95b] Kung-Yee Liang and Scott L. Zeger. [Inference based on estimating functions in the presence of nuisance parameters]: Rejoinder. Statistical Science, 10(2):196–199, May 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010037.

Lee:2000:CTS

[LZ00] Sandra J. Lee and Marvin Zelen. Clinical trials and sample size considerations: Another perspective. Statistical Science, 15(2):95–110, May 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009212752.

Lin:2020:MBA

[LZD^{+}20] Zhixiang Lin, Mahdi Zamanighomi, Timothy Daley, Shining Ma, and Wing Hung Wong. Model-based approach to the joint analysis of single-cell data on chromatin accessibility and gene expression. Statistical Science, 35(1):2–13, February 2020. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/euclid.ss/1583226022.

Mezzich:1987:SPC

[MA87] Juan E. Mezzich and Chul Woo Ahn. [Statistics in psychiatry]: Comment: Psychiatric statistics and clinical information. Statistical Science, 2(2):125–127, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013338. See [Eve87a].

MacNeill:1993:CDJ

[Mac93] Ian MacNeill. A conversation with David J. Finney. Statistical Science, 8(2):187–201, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011012.

Madansky:1986:RMA

[Mad86] Albert Madansky. [Regression models for adjusting the 1980 Census]: Comment. Statistical Science, 1(1):28–30, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013809. See [FN86a].

Madansky:1987:UPA

[Mad87] Albert Madansky. [Uncertainty, policy analysis and statistics]: Comment. Statistical Science, 2(3):282–286, August 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013229.

Madras:1992:PMC

[Mad92] Neal Madras. [Practical Markov chain Monte Carlo]: Comment. Statistical Science, 7(4):488–489, November 1992. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011140.

Madigan:1993:BAE

[Mad93] David Madigan. [Bayesian analysis in expert systems]: Comment: What’s next? Statistical Science, 8(3):261–263, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010892.

Maggs:1993:RWM

[Mag93] Bruce M. Maggs. Randomly wired multistage networks. Statistical Science, 8(1):70–75, February 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011086.

Mithas:2006:DCS

[MAK06] Sunil Mithas, Daniel Almirall, and M. S. Krishnan. Do CRM systems cause one-to-one marketing effectiveness? Statistical Science, 21(2):223–233, May 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1154979823.

Mallows:2002:PIT

[Mal02] Colin Mallows. Parity: Implementing the telecommunications act of 1996. Statistical Science, 17(3):256–285, August 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1042727941.

Mallows:2003:JTB

[Mal03] Colin Mallows. John Tukey at Bell labs. Statistical Science, 18 (3):332–335, August 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1076102420.

Manski:1992:ISM

[Man92] Charles F. Manski. [The impact of sociological methodology on statistical methodology]: Comment. Statistical Science, 7 (2):201–203, May 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011357.

Mandelkern:2002:SCI

[Man02] Mark Mandelkern. Setting confidence intervals for bounded parameters. Statistical Science, 17(2):149–172, May 2002. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1030550859.

Margolin:1987:RSC

[Mar87a] Barry H. Margolin. [The role of a second control group in an observational study]: Comment: The use of multiple control groups in designed experiments. Statistical Science, 2(3): 308–310, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013234.

Marquardt:1987:CLS

[Mar87b] Donald W. Marquardt. [Collinearity and least squares regression]: Comment. Statistical Science, 2(1):84–85, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013440.

Margolin:1988:SAU

[Mar88a] Barry H. Margolin. Statistical aspects of using biologic markers. Statistical Science, 3(3):351–357, August 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012837.

Martin:1988:RBR

[Mar88b] R. Douglas Martin. [Rank-based robust analysis of linear models. I. Exposition and review]: Comment. Statistical Science, 3(2):264–266, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012920.

Mardia:1989:SST

[Mar89] Kanti V. Mardia. [A survey of the statistical theory of shape]: Comment: Some contributions to shape analysis. Statistical Science, 4(2):108–111, May 1989. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012585.

Marshall:1990:CZW

[Mar90] Albert W. Marshall. A conversation with Z. William Birnbaum. Statistical Science, 5(2):227–241, May 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012174.

Marden:2004:PQP

[Mar04a] John I. Marden. Positions and QQ plots. Statistical Science, 19(4):606–614, November 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1113832725.

Marinucci:2004:TNG

[Mar04b] Domenico Marinucci. Testing for non-Gaussianity on cosmic microwave background radiation: A review. Statistical Science, 19(2):294–307, May 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1105714164.

Martin:2007:BKC

[Mar07] Margaret E. Martin. Bill Kruskal and the committee on national statistics. Statistical Science, 22(2):273–274, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1190905526.

Martin:2011:VPR

[Mar11] Isbelia Martín. 2004 Venezuelan Presidential Recall Referendum (2004 PRR): A statistical analysis from the point of view of electronic voting data transmissions. Statistical Science, 26(4): 528–542, November 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1330437934.

Marschner:2021:GFA

[Mar21] Ian C. Marschner. A general framework for the analysis of adaptive experiments. Statistical Science, 36(3):465–492, August 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-36/issue-3/
A-General-Framework-for-the-Analysis-of-Adaptive-Experiments/
10.1214/20-STS803.full.

Mayo:2014:BAS

[May14a] Deborah G. Mayo. On the Birnbaum argument for the Strong Likelihood Principle. Statistical Science, 29(2):227–239, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1408368573. See discussion [Daw14, Eva14, Han14b, Bjø14] and rejoinder [May14b].

Mayo:2014:RBA

[May14b] Deborah G. Mayo. Rejoinder: “On the Birnbaum Argument for the Strong Likelihood Principle”. Statistical Science, 29 (2):261–266, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1408368580. See [May14a].

Mazliak:2015:GEN

[Maz15] Laurent Mazliak. The ghosts of the École Normale. Statistical Science, 30(3):391–412, August 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1439220719.

Maatta:1990:DDTa

[MC90a] Jon M. Maatta and George Casella. Developments in decisiontheoretic variance estimation. Statistical Science, 5(1):90–101, February 1990. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177012263.

Maatta:1990:DDTb

[MC90b] Jon M. Maatta and George Casella. [Developments in decisiontheoretic variance estimation]: Rejoinder. Statistical Science, 5 (1):118–120, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012272.

Mosteller:1992:SPP

[MC92] Frederick Mosteller and Thomas C. Chalmers. Some progress and problems in meta-analysis of clinical trials. Statistical Science, 7(2):227–236, May 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011362.

McCullagh:1986:GAM

[McC86] Peter McCullagh. [Generalized additive models]: Comment. Statistical Science, 1(3):314, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013608.

McCullagh:1989:URI

[McC89] Peter McCullagh. [The unifying role of iterative generalized least squares in statistical algorithms]: Comment. Statistical Science, 4(4):404–405, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012410.

McClelland:1994:NNR

[McC94] James L. McClelland. [Neural networks: A review from statistical perspective]: Comment: Neural networks and cognitive science: Motivations and applications. Statistical Science, 9 (1):42–45, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010643.

McCullagh:1995:IBE

[McC95] Peter McCullagh. [Inference based on estimating functions in the presence of nuisance parameters]: Comment. Statistical Science, 10(2):177–179, May 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177010031.

McCulloch:2011:DSI

[McC11] Robert McCulloch. Discussion of “Statistical Inference: The Big Picture” by R. E. Kass. Statistical Science, 26(1):15–16, February 2011. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1307626557. See [Kas11b].

McCullagh:2016:TEC

[McC16] Peter McCullagh. Two early contributions to the Ewens saga. Statistical Science, 31(1):23–26, February 2016. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1455115908.

McDonald:1998:CSG

[McD98] Gary C. McDonald. A conversation with Shanti Gupta. Statistical Science, 13(3):291–305, August 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1028905889.

McKean:2004:RAL

[McK04] Joseph W. McKean. Robust analysis of linear models. Statistical Science, 19(4):562–570, November 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1113832720.

McKeague:2021:NPM

[McK21] Ian W. McKeague. Noncommutative probability and multiplicative cascades. Statistical Science, 36(2):256–263, May 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-36/issue-2/
Noncommutative-Probability-and-Multiplicative-Cascades/
10.1214/20-STS780.full.

Moolgavkar:1988:MMQ

[MD88] Suresh H. Moolgavkar and Anup Dewanji. [From mouse-to-man: The quantitative assessment of cancer risks]: Comment. Statistical Science, 3(1):39–41, February 1988. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012996.

Michailidis:1998:GSD

[MdL98] George Michailidis and Jan de Leeuw. The Gifi system of descriptive multivariate analysis. Statistical Science, 13(4):307– 336, November 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1028905828.

Michailidis:2000:AA

[MdL00] George Michailidis and Jan de Leeuw. Acknowledgment and apology. Statistical Science, 15(2):191, May 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1009212757.

Meeden:2011:DIF

[Mee11] Glen Meeden. Discussion of “Impact of Frequentist and Bayesian Methods on Survey Sampling Practice: A Selective Appraisal” by J. N. K. Rao. Statistical Science, 26(2):257–259, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1312204016. See [Rao11a].

Mehta:1992:SEI

[Meh92] Cyrus R. Mehta. [A survey of exact inference for contingency tables]: Comment: An interdisciplinary approach to exact inference for contingency tables. Statistical Science, 7(1): 167–170, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011460.

Meinshausen:2011:DMT

[Mei11] Nicolai Meinshausen. Discussion of “Multiple Testing for Exploratory Research” by J. J. Goeman and A. Solari. Statistical Science, 26(4):601–603, November 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1330437939. See [GS11a].

Mendelsohn:1988:TBM

[Men88] Mortimer L. Mendelsohn. Tests for biologic markers of genotoxic exposure and effect. Statistical Science, 3(3):346–350, August 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012836.

Meng:1994:MIIa

[Men94a] Xiao-Li Meng. Multiple-imputation inferences with uncongenial sources of input. Statistical Science, 9(4):538–558, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010269.

Meng:1994:MIIb

[Men94b] Xiao-Li Meng. [Multiple-imputation inferences with uncongenial sources of input]: Rejoinder. Statistical Science, 9(4): 566–573, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010274.

Meng:2005:URS

[Men05] Xiao-Li Meng. From unit root to Stein’s estimator to Fisher’s k statistics: If you have a moment, I can tell you more. Statistical Science, 20(2):141–162, May 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1121347637.

Meng:2009:DHL

[Men09] Xiao-Li Meng. Decoding the h-likelihood. Statistical Science, 24 (3):280–293, August 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1270041255.

Meyer:2018:FEI

[Mey18a] Mary C. Meyer. A framework for estimation and inference in generalized additive models with shape and order restrictions. Statistical Science, 33(4):595–614, November 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1543482060.

Meyer:2018:SEP

[Mey18b] Sebastian Meyer. Self-exciting point processes: Infections and implementations. Statistical Science, 33(3):327–329, August 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1534147225.

Martin:2024:ABC

[MFR24a] Gael M. Martin, David T. Frazier, and Christian P. Robert. Approximating Bayes in the 21st century. Statistical Science, 39(1):20–45, February 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/journals/statistical-science/volume-
39/issue-1/Approximating-Bayes-in-the-21st-Century/
10.1214/22-STS875.full.

Martin:2024:CBT

[MFR24b] Gael M. Martin, David T. Frazier, and Christian P. Robert. Computing Bayes: From then ‘til now. Statistical Science, 39(1):3–19, February 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/journals/statistical-science/volume-

39/issue-1/Computing-Bayes-From-Then-Til-Now/10.1214/
22-STS876.full.

Martin:2008:SAN

[MG08] Ryan Martin and Jayanta K. Ghosh. Stochastic approximation and Newton’s estimate of a mixing distribution. Statistical Science, 23(3):365–382, August 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1233153064.

Miller:2020:LTC

[MG20a] Joshua B. Miller and Andrew Gelman. Laplace’s theories of cognitive illusions, heuristics and biases. Statistical Science, 35(2): 159–170, May 2020. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/

euclid.ss/1591171222. See comments [KBH20, Sha20] and rejoinder [MG20b].

Miller:2020:RLT

[MG20b] Joshua B. Miller and Andrew Gelman. Rejoinder: Laplace’s theories of cognitive illusions, heuristics and biases. Statistical Science, 35(2):175–177, May 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:

//projecteuclid.org/euclid.ss/1591171225. See [MG20a,

KBH20, Sha20].

Moreno:2015:PMC

[MGC15] Elías Moreno, Javier Girón, and George Casella. Posterior model consistency in variable selection as the model dimension grows. Statistical Science, 30(2):228–241, May 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1433341480.

Malley:1993:QSI

[MH93] James D. Malley and John Hornstein. Quantum statistical inference. Statistical Science, 8(4):433–457, November 1993. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010787.

McElroy:2012:CDF

[MH12] Tucker S. McElroy and Scott H. Holan. A conversation with David Findley. Statistical Science, 27(4):594–606, November 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1356098558.

Mondal:2016:RTP

[MH16] Debashis Mondal and Nina Hinrichs. Rank tests from partially ordered data using importance and MCMC sampling methods. Statistical Science, 31(3):325–347, August 2016. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1475001232.

Micheloud:2022:PCR

[MH22] Charlotte Micheloud and Leonhard Held. Power calculations for replication studies. Statistical Science, 37(3):369–379, August 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-3/
Power-Calculations-for-Replication-Studies/10.1214/
21-STS828.full.

Moosavi:2023:CBU

[MHdL23] Niloofar Moosavi, Jenny Häggström, and Xavier de Luna. The costs and benefits of uniformly valid causal inference with high-dimensional nuisance parameters. Statistical Science, 38(1):1–12, February 2023. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-1/The-Costs-and-Benefits-of-
Uniformly-Valid-Causal-Inference-with/10.1214/21-STS843.
full.

Mira:2001:OIP

[Mir01] Antonietta Mira. Ordering and improving the performance of Monte Carlo Markov chains. Statistical Science, 16(4):340– 350, November 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1015346319.

Mitchell:1988:SMS

[Mit88] Ann F. S. Mitchell. [Saddlepoint methods and statistical inference]: Comment. Statistical Science, 3(2):237, May 1988. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012913.

Mittal:1989:HGY

[Mit89] Yashaswini Mittal. [How to get your first research grant]: Comment. Statistical Science, 4(2):141–143, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012596.

Molenberghs:2009:DLI

[MKV09] Geert Molenberghs, Michael G. Kenward, and Geert Verbeke. Discussion of likelihood inference for models with unobservables: Another view. Statistical Science, 24(3):273–279, August 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1270041254.

Morris:2012:SEM

[ML12] Carl N. Morris and Martin Lysy. Shrinkage estimation in multilevel normal models. Statistical Science, 27(1):115–134, February 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1331729986.

Martin:2014:DFS

[ML14] Ryan Martin and Chuanhai Liu. Discussion: Foundations of statistical inference, revisited. Statistical Science, 29(2):247– 251, May 2014. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1408368576.

Meredith:1994:BMT

[MM94] Michael P. Meredith and Jorge G. Morel. [Bootstrap: More than a stab in the dark?]: Comment. Statistical Science, 9 (3):404–406, August 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010388.

Marker:2000:CJW

[MM00] David Marker and David Morganstein. A conversation with Joseph Waksberg. Statistical Science, 15(3):299–312, August 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1009212819.

Moguerza:2006:R

[MM06a] Javier M. Moguerza and Alberto Muñoz. Rejoinder. Statistical Science, 21(3):358–362, August 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1166642440.

Moguerza:2006:SVM

[MM06b] Javier M. Moguerza and Alberto Muñoz. Support vector machines with applications. Statistical Science, 21(3):322–336, August 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1166642435.

Mosler:2022:CAN

[MM22] Karl Mosler and Pavlo Mozharovskyi. Choosing among notions of multivariate depth statistics. Statistical Science, 37(3):348– 368, August 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-3/
Choosing-Among-Notions-of-Multivariate-Depth-Statistics/
10.1214/21-STS827.full.

Mealli:2024:CGW

[MM24] Fabrizia Mealli and Julie Holland Mortimer. A conversation with Guido W. Imbens. Statistical Science, 39(2):357– 373, May 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-2/
A-Conversation-with-Guido-W-Imbens/10.1214/23-STS906.
full.

Mengersen:2007:CAS

[MMT07] K. Mengersen, S. A. Moynihan, and R. L. Tweedie. Causality and association: The statistical and legal approaches. Statistical Science, 22(2):227–254, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1190905520.

Mammen:2001:GPF

[MMTW01] E. Mammen, J. S. Marron, B. A. Turlach, and M. P. Wand. A general projection framework for constrained smoothing. Statistical Science, 16(3):232–248, August 2001. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1009213727.

McCulloch:2011:MSR

[MN11] Charles E. McCulloch and John M. Neuhaus. Misspecifying the shape of a random effects distribution: Why getting it wrong May not matter. Statistical Science, 26(3):388–402, August 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1320066927.

MacKay:2000:SMS

[MO00] R. J. MacKay and R. W. Oldford. Scientific method, statistical method and the speed of light. Statistical Science, 15(3): 254–278, August 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009212817.

Molenaar:1992:ISM

[Mol92] Ivo W. Molenaar. [The impact of sociological methodology on statistical methodology]: Comment: The fence between statistics and social research. Statistical Science, 7(2):203–205, May 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011358.

Monmonier:2005:LM

[Mon05] Mark Monmonier. Lying with maps. Statistical Science, 20 (3):215–222, August 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1124891287.

Moore:1988:GOC

[Moo88] David S. Moore. [Golden oldies: Classic articles from the world of statistics and probability]: Comment. Statistical Science, 3 (1):84–87, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013003. See [Hot88c].

Mood:1990:MR

[Moo90a] Alexander M. Mood. Miscellaneous reminiscences. Statistical Science, 5(1):35–43, February 1990. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012252.

Moore:1990:UDP

[Moo90b] David S. Moore. [The unity and diversity of probability]: Comment. Statistical Science, 5(4):454–456, November 1990. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012028.

Moore:1990:ABE

[Moo90c] Peter G. Moore. [Applications in business and economic statistics: Some personal views]: Comment: A U.K. perspective on applications in business and economic statistics. Statistical Science, 5(4):392–395, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012011.

Morris:1986:CPD

[Mor86] Peter A. Morris. [Combining probability distributions: A critique and an annotated bibliography]: Comment. Statistical Science, 1(1):141–144, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013829. See [GZ86a].

Morgan:1988:QRA

[Mor88] M. Granger Morgan. Quantitative risk assessment: Low frequency electromagnetic fields as an example. Statistical Science, 3(3):314–319, August 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012832.

Morris:1989:I

[Mor89a] C. N. Morris. In this issue. Statistical Science, 4(3):187, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012479.

Morris:1989:DAC

[Mor89b] Max D. Morris. [Design and analysis of computer experiments]: Comment. Statistical Science, 4(4):423–425, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012414.

Morris:1990:Ia

[Mor90a] C. N. Morris. In this issue. Statistical Science, 5(2):159, May 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012163.

Morris:1990:Ib

[Mor90b] C. N. Morris. In this issue. Statistical Science, 5(3):267–268, August 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012091.

Morris:1990:Ic

[Mor90c] C. N. Morris. In this issue. Statistical Science, 5(4):371, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012007.

Morris:1991:Ia

[Mor91a] C. N. Morris. In this issue. Statistical Science, 6(1):1, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011923.

Morris:1991:Ib

[Mor91b] C. N. Morris. In this issue. Statistical Science, 6(2):119–120, May 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011817.

Morris:1991:Ic

[Mor91c] C. N. Morris. In this issue. Statistical Science, 6(3):217–218, August 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011680.

Morris:1991:D

[Mor91d] Carl Morris. Dedication. Statistical Science, 6(1):2, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011924.

Morris:1991:RMA

[Mor91e] Robert L. Morris. [Replication and meta-analysis in parapsychology]: Comment. Statistical Science, 6(4):393–395, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011583.

Morris:1992:PRO

[Mor92] Carl N. Morris. Parting remarks from the outgoing Editor. Statistical Science, 7(2):182, May 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011353.

Morgenthaler:2003:JWT

[Mor03] Stephan Morgenthaler. John W. Tukey as teacher. Statistical Science, 18(3):342–345, August 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1076102423.

Morton:2005:E

[Mor05] Sally C. Morton. Ephedra. Statistical Science, 20(3):242–248, August 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1124891291.

Morris:2008:CME

[Mor08] Carl N. Morris. Comment: Microarrays, empirical Bayes and the two-groups model. Statistical Science, 23(1):34–40, February 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1215441279. See [GM05a].

Morris:2011:R

[Mor11] Carl Morris. Rejoinder. Statistical Science, 26(2):296–298, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1312204023.

Mortlock:2014:FMD

[Mor14] Daniel Mortlock. Finding the most distant quasars using Bayesian selection methods. Statistical Science, 29(1):50–57, February 2014. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1399645728.

Moses:1986:RMA

[Mos86] Lincoln E. Moses. [Regression models for adjusting the 1980 Census]: Comment. Statistical Science, 1(1):33–34, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013811. See [FN86a].

Mosteller:1991:RMA

[Mos91] Frederick Mosteller. [Replication and meta-analysis in parapsychology]: Comment. Statistical Science, 6(4):395–396, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011584.

Mouchart:1990:WML

[Mou90] Michel Mouchart. [The 1988 Wald memorial lectures: The present position in Bayesian statistics]: Comment. Statistical Science, 5(1):84–85, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012261.

Mallows:1991:ASP

[MP91] C. L. Mallows and D. Pregibon. [Avoiding statistical pitfalls]: Comment. Statistical Science, 6(3):263–264, August 1991. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011691.

Muliere:1993:UM

[MP93] Pietro Muliere and Giovanni Parmigiani. Utility and means in the 1930s. Statistical Science, 8(4):421–432, November 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010786.

Muller:2004:NBD

[MQ04] Peter Müller and Fernando A. Quintana. Nonparametric Bayesian data analysis. Statistical Science, 19(1):95–110, February 2004. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1089808275.

Mengersen:2014:BBS

[MR14] Kerrie L. Mengersen and Christian P. Robert. Big Bayes stories — foreword. Statistical Science, 29(1):1, February 2014. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1399645721.

Marron:2015:FDA

[MRSS15] J. S. Marron, James O. Ramsay, Laura M. Sangalli, and Anuj Srivastava. Functional data analysis of amplitude and phase variation. Statistical Science, 30(4):468–484, November 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1449670854.

MacGibbon:1990:DDT

[MS90] K. Brenda MacGibbon and Glenn E. Shorrock. [Developments in decision-theoretic variance estimation]: Comment. Statistical Science, 5(1):112–113, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012269.

Muller:2024:AMS

[MSM24] Raoul Müller, Dominic Schuhmacher, and Jorge Mateu. ANOVA for metric spaces, with applications to spatial data. Statistical Science, 39(2):262–285, May 2024. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-39/issue-2/ANOVA-for-Metric-Spaces-with-
Applications-to-Spatial-Data/10.1214/23-STS898.full.

Muller:2013:MSL

[MSW13] Samuel Müller, J. L. Scealy, and A. H. Welsh. Model selection in linear mixed models. Statistical Science, 28(2):135–167, May 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1369147909.

Martin:2016:CSH

[MSY16] Ryan Martin, John Stufken, and Min Yang. A conversation with Samad Hedayat. Statistical Science, 31(4):637–647, November 2016. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1484816591.

Morris:2011:ERE

[MT11] Carl Morris and Ruoxi Tang. Estimating random effects via adjustment for density maximization. Statistical Science, 26 (2):271–287, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid. org/euclid.ss/1312204020. See discussion [FC11, LP11a].

McEvoy:2012:DMB

[MT12] Bradley W. McEvoy and Ram C. Tiwari. Discussion of “Multivariate Bayesian Logistic Regression for Analysis of Clinical Trial Safety Issues” by W. DuMouchel. Statistical Science, 27 (3):340–343, August 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1346849940. See [DuM12a].

Miettinen:2015:FMI

[MTNO15] Jari Miettinen, Sara Taskinen, Klaus Nordhausen, and Hannu Oja. Fourth moments and independent component analysis. Statistical Science, 30(3):372–390, August 2015. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1439220718.

Masoero:2023:CSR

[MTP^{+}23] Lorenzo Masoero, Emma Thomas, Giovanni Parmigiani, Svitlana Tyekucheva, and Lorenzo Trippa. Cross-study replicability in cluster analysis. Statistical Science, 38(2):303– 316, May 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-2/
Cross-Study-Replicability-in-Cluster-Analysis/10.1214/
22-STS871.full.

Mukhopadhyay:1997:CSK

[Muk97] Nitis Mukhopadhyay. A conversation with Sujit Kumar Mitra. Statistical Science, 12(1):61–75, February 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1029963263.

Mukhopadhyay:2000:CMS

[Muk00] Nitis Mukhopadhyay. A conversation with Milton Sobel. Statistical Science, 15(2):168–190, May 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1009212756.

Mukhopadhyay:2002:CKM

[Muk02] Nitis Mukhopadhyay. A conversation with Kanti Mardia. Statistical Science, 17(2):113–148, May 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1023799001.

Mukhopadhyay:2005:CSZ

[Muk05] Nitis Mukhopadhyay. A conversation with Shelemyahu Zacks. Statistical Science, 20(1):89–110, February 2005. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1118065044.

Mukhopadhyay:2006:CUG

[Muk06] Nitis Mukhopadhyay. A conversation with Ulf Grenander. Statistical Science, 21(3):404–426, August 2006. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1166642444.

Mukhopadhyay:2017:CLB

[Muk17] Nitis Mukhopadhyay. A conversation with Lynne Billard. Statistical Science, 32(1):138–164, February 2017. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1491465632.

Mukherjee:2022:BPH

[Muk22] Bhramar Mukherjee. Being a public health statistician during a global pandemic. Statistical Science, 37(2):270–277, May 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-2/
Being-a-Public-Health-Statistician-During-a-Global-
Pandemic/10.1214/22-STS859.full.

Mukherjee:2024:CSD

[Muk24]  Bhramar Mukherjee.  In conversation with Sir David

Spiegelhalter and Professor Sylvia Richardson. Statistical Science, 39(1):209–220, February 2024. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-39/issue-1/In-Conversation-with-Sir-
David-Spiegelhalter-and-Professor-Sylvia-Richardson/
10.1214/23-STS897.full.

Muller:1993:LRA

[Mul93] Hans-Georg Muller. [Local regression: Automatic kernel carpentry]: Comment. Statistical Science, 8(2):134–139, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011004.

Murray:2005:HAO

[Mur05] Charles Murray. How to accuse the other guy of lying with statistics. Statistical Science, 20(3):239–241, August 2005. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1124891290.

Murray:2018:MIR

[Mur18] Jared S. Murray. Multiple imputation: A review of practical and theoretical findings. Statistical Science, 33(2):142– 159, May 2018. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1525313139.

McKinley:2018:ABC

[MVA^{+}18] Trevelyan J. McKinley, Ian Vernon, Ioannis Andrianakis, Nicky McCreesh, Jeremy E. Oakley, Rebecca N. Nsubuga, Michael Goldstein, and Richard G. White. Approximate Bayesian computation and simulation-based inference for complex stochastic epidemic models. Statistical Science, 33(1):4–18, 02 2018. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Meulman:2019:RRI

[MvdKD19] Jacqueline J. Meulman, Anita J. van der Kooij, and Kevin L. W. Duisters. ROS regression: Integrating regularization with optimal scaling regression. Statistical Science, 34(3):361– 390, August 2019. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1570780975.

Molenberghs:2010:FGL

[MVDV10] Geert Molenberghs, Geert Verbeke, Clarice G. B. Demétrio, and Afrânio M. C. Vieira. A family of generalized linear models for repeated measures with normal and conjugate random effects. Statistical Science, 25(3):325–347, August 2010. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1294167963.

Miller:1996:PDS

[MW96] Arden Miller and C. F. J. Wu. Parameter design for signalresponse systems: a different look at Taguchi’s dynamic parameter design. Statistical Science, 11(2):122–136, May 1996. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1038425656.

Mosteller:1990:QPEa

[MY90a] Frederick Mosteller and Cleo Youtz. Quantifying probabilistic expressions. Statistical Science, 5(1):2–12, February 1990. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012242.

Mosteller:1990:QPEb

[MY90b] Frederick Mosteller and Cleo Youtz. [Quantifying probabilistic expressions]: Rejoinder. Statistical Science, 5(1):32–34, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012251.

Manton:1997:EUP

[MY97] Kenneth G. Manton and Anatoli I. Yashin. Effects of unobserved and partially observed covariate processes on system failure: a review of models and estimation strategies. Statistical Science, 12(1):20–34, February 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1029963259.

Mykland:2005:CSO

[MZ05] Per A. Mykland and Lan Zhang. Comment: A selective overview of nonparametric methods in financial econometrics. Statistical Science, 20(4):347–350, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1137076650.  See

[Fan05b].

Martin:2010:DST

[MZL10] Ryan Martin, Jianchun Zhang, and Chuanhai Liu. Dempster– Shafer theory and statistical inference with weak beliefs. Statistical Science, 25(1):72–87, February 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1280841734.

Navidi:1994:AGD

[NA94] W. Navidi and N. Arnheim. Analysis of genetic data from the polymerase chain reaction. Statistical Science, 9(3):320–333, August 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010379.

Nadarajah:2002:CSK

[Nad02] Saralees Nadarajah. A conversation with Samuel Kotz. Statistical Science, 17(2):220–233, May 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1030550863.

Nathan:1986:RMA

[Nat86] Gad Nathan. [Regression models for adjusting the 1980 Census]: Comment. Statistical Science, 1(1):34–35, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013812. See [FN86a].

Naus:2000:CJH

[Nau00] Joseph I. Naus. A conversation with Johannes H. B. Kemperman. Statistical Science, 15(4):396–408, November 2000. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1009213005.

Naus:2016:CAC

[Nau16] Joseph Naus. A conversation with Arthur Cohen. Statistical Science, 31(3):442–452, August 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1475001238.

Navidi:1994:BMT

[Nav94] William Navidi. [Bootstrap: More than a stab in the dark?]: Comment. Statistical Science, 9(3):407–408, August 1994. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010389.

Nicholson:2022:ISM

[NBB^{+}22] George Nicholson, Marta Blangiardo, Mark Briers, Peter J. Diggle, Tor Erlend Fjelde, Hong Ge, Robert J. B. Goudie, Radka Jersakova, Ruairidh E. King, Brieuc C. L.

Lehmann, Ann-Marie Mallon, Tullia Padellini, Yee Whye Teh, Chris Holmes, and Sylvia Richardson. Interoperability of statistical models in pandemic preparedness: Principles and reality. Statistical Science, 37(2):183–206, May 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-2/
Interoperability-of-Statistical-Models-in-Pandemic-
Preparedness--Principles-and/10.1214/22-STS854.full.

Needleman:1988:SPI

[Nee88] Jack Needleman. Sources and policy implications of uncertainty in risk assessment. Statistical Science, 3(3):328–338, August 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012834.

Nelder:1986:GAM

[Nel86] J. A. Nelder. [Generalized additive models]: Comment. Statistical Science, 1(3):312, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177013606.

Neter:1990:ABE

[Net90] John Neter. [Applications in business and economic statistics: Some personal views]: Comment. Statistical Science, 5(4): 395–397, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012012.

Newton:2002:CEP

[New02a] H. Joseph Newton. A conversation with Emanuel Parzen. Statistical Science, 17(3):357–378, August 2002. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1042727944.  See

correction [New02b].

Newton:2002:CSC

[New02b] H. Joseph Newton. Correction to “A conversation with Emanuel Parzen”. Statistical Science, 17(4):467, November 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1049993205. See [New02a].

Niu:2016:MCP

[NHZ16] Yue S. Niu, Ning Hao, and Heping Zhang. Multiple changepoint detection: A selective overview. Statistical Science, 31(4): 611–623, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1484816589.

Nichols:2016:FON

[Nic16] James D. Nichols. And the first one now will later be last: Timereversal in Cormack–Jolly–Seber models. Statistical Science, 31(2):175–190, May 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1464105036.

Nychka:2020:CGW

[NMB20] Douglas Nychka, Ping Ma, and Douglas Bates. A conversation with Grace Wahba. Statistical Science, 35(2):308–320, May 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1591171233.

Nicolae:2008:QFM

[NMK08a] Dan L. Nicolae, Xiao-Li Meng, and Augustine Kong. Quantifying the fraction of missing information for hypothesis testing in statistical and genetic studies. Statistical Science, 23 (3):287–312, August 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid. org/euclid.ss/1233153057. See comments [ZL08, CCCH08, NMK08b].

Nicolae:2008:RQF

[NMK08b] Dan L. Nicolae, Xiao-Li Meng, and Augustine Kong. Rejoinder: Quantifying the fraction of missing information for hypothesis testing in statistical and genetic studies. Statistical Science, 23 (3):325–331, August 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1233153061. See [NMK08a].

Neykov:2018:UTC

[NNLL18] Matey Neykov, Yang Ning, Jun S. Liu, and Han Liu. A unified theory of confidence regions and testing for high-dimensional estimating equations. Statistical Science, 33(3):427–443, August 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1534147231.

Nolan:1992:WSA

[Nol92] Deborah Nolan. Women in statistics in academe: Mentors matter. Statistical Science, 7(2):267–272, May 1992. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011366.

Norwood:1991:EAM

[Nor91] Janet L. Norwood. [Enhancing access to microdata while protecting confidentiality: Prospects for the future]: Comment. Statistical Science, 6(3):236–237, August 1991. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011684.

Normand:1993:BAE

[Nor93] Sharon-Lise Normand. [Bayesian analysis in expert systems]: Comment. Statistical Science, 8(3):263–265, August 1993. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010893.

Nordstrom:1999:LWG

[Nor99] Kenneth Nordström. The life and work of Gustav Elfving. Statistical Science, 14(2):174–196, May 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1009212244.

Notz:1988:ODC

[Not88] William I. Notz. [Optimal designs for comparing test treatments with controls]: Comment. Statistical Science, 3(4):480– 482, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012769.

Nair:1993:ISC

[NP93] Vijayan N. Nair and Daryl Pregibon. [Is industrial statistics out of control?]: Comment. Statistical Science, 8(4):391– 394, November 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010778.

Nieto-Reyes:2016:TVD

[NRB16] Alicia Nieto-Reyes and Heather Battey. A topologically valid definition of depth for functional data. Statistical Science, 31 (1):61–79, February 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1455115914. See correction [NRB17].

Nieto-Reyes:2017:CTV

[NRB17] Alicia Nieto-Reyes and Heather Battey. Correction to A Topologically Valid Definition of Depth for Functional Data. Statistical Science, 32(4):640, 11 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). See [NRB16].

Nasri:2021:CDD

[NRSV21] Bouchra R. Nasri, Bruno N. Rémillard, Barbara Szyszkowicz, and Jean Vaillancourt. A conversation with Don Dawson. Statistical Science, 36(4):612–622, November 2021. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-4/A-Conversation-with-Don-Dawson/
10.1214/21-STS821.full.

Negahban:2012:UFH

[NRWY12] Sahand N. Negahban, Pradeep Ravikumar, Martin J. Wainwright, and Bin Yu. A unified framework for high-dimensional analysis of M-estimators with decomposable regularizers. Statistical Science, 27(4):538–557, November 2012. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1356098555.

Nelson:1994:SIC

[NS94] David O. Nelson and Terence P. Speed. Statistical issues in constructing high resolution physical maps. Statistical Science, 9(3):334–354, August 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010380.

Normand:2007:SCA

[NS07] Sharon-Lise T. Normand and David M. Shahian. Statistical and clinical aspects of hospital outcomes profiling. Statistical Science, 22(2):206–226, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1190905519.

Nott:2020:CPD

[NWEE20] David J. Nott, Xueou Wang, Michael Evans, and Berthold-Georg Englert. Checking for prior-data conflict using prior-to-posterior divergences. Statistical Science, 35(2):234–253, May 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1591171229.

Nyblom:2023:NLM

[Nyb23] Jukka Nyblom. Note on Legendre’s method of least squares. Statistical Science, 38(3):510–513, August 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-3/Note-on-Legendres-Method-
of-Least-Squares/10.1214/23-STS887.full.

OQuigley:2010:CRR

[OC10] John O’Quigley and Mark Conaway. Continual reassessment and related dose-finding designs. Statistical Science, 25(2): 202–216, May 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1290175842.

Ogata:2013:PEP

[Oga13] Yosihiko Ogata. A prospect of earthquake prediction research. Statistical Science, 28(4):521–541, November 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1386078877.

Ogata:2018:CRS

[Oga18] Yosihiko Ogata. Comment on “A Review of Self-Exciting Spatiotemporal Point Process and Their Applications” by Alex Reinhart. Statistical Science, 33(3):319–322, August 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/

1534147222. See [Rei18b, Rei18a].

OHagan:1989:DAC

[O’H89] Anthony O’Hagan. [Design and analysis of computer experiments]: Comment. Statistical Science, 4(4):430–432, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012418.

Owen:1989:DAC

[OKS89] A. Owen, J. Koehler, and S. Sharifzadeh. [Design and analysis of computer experiments]: Comment. Statistical Science, 4(4): 429–430, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012417.

Olkin:1987:CAH

[Olk87a] Ingram Olkin. A conversation with Albert H. Bowker. Statistical Science, 2(4):472–483, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013127.

Olkin:1987:CMH

[Olk87b] Ingram Olkin. A conversation with Morris Hansen. Statistical Science, 2(2):162–179, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1177013351. See correction

[Olk89b].

Olkin:1988:GOC

[Olk88] Ingram Olkin. [Golden oldies: Classic articles from the world of statistics and probability]: Comment. Statistical Science, 3(1): 107–108, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013011.

Olkin:1989:CMB

[Olk89a] Ingram Olkin. A conversation with Maurice Bartlett. Statistical Science, 4(2):151–163, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012600.

Olkin:1989:CNC

[Olk89b] Ingram Olkin. Correction note: A conversation with Morris Hansen. Statistical Science, 4(2):185, May 1989. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012605.  See

[Olk87b].

Olkin:1991:CWA

[Olk91] Ingram Olkin. A conversation with W. Allen Wallis. Statistical Science, 6(2):121–140, May 1991. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011818.

Olkin:1992:CCE

[Olk92a] Ingram Olkin. A conversation with Churchill Eisenhart. Statistical Science, 7(4):512–530, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011149.

Olkin:1992:EI

[Olk92b] Ingram Olkin. Editor’s introduction. Statistical Science, 7(2): 226, May 1992. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177011361.

Olshen:2001:CLB

[Ols01] Richard Olshen. A conversation with Leo Breiman. Statistical Science, 16(2):184–198, May 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1009213290.

Oja:2004:MNT

[OR04] Hannu Oja and Ronald H. Randles. Multivariate nonparametric tests. Statistical Science, 19(4):598–605, November 2004. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1113832724.

Orchard:1993:ISC

[Orc93] T. J. Orchard. [Is industrial statistics out of control?]: Comment. Statistical Science, 8(4):394–395, November 1993. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010779.

OSullivan:1986:SPIb

[O’S86a] Finbarr O’Sullivan. [A statistical perspective on ill-posed inverse problems]: Rejoinder. Statistical Science, 1(4):523–527, November 1986. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177013530. See [O’S86b].

OSullivan:1986:SPIa

[O’S86b] Finbarr O’Sullivan. A statistical perspective on ill-posed inverse problems. Statistical Science, 1(4):502–518, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.

ss/1177013525. See comments [Tit86, Wah86, Ric86, Gil86, O’S86a].

Ommen:2021:PFS

[OS21] Danica M. Ommen and Christopher P. Saunders. A problem in forensic science highlighting the differences between the Bayes factor and likelihood ratio. Statistical Science, 36(3):344– 359, August 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-36/issue-3/
A-Problem-in-Forensic-Science-Highlighting-the-Differences-
between-the/10.1214/20-STS805.full.

Oesting:2022:CTT

[OS22]  Marco Oesting and Kirstin Strokorb.  A comparative

tour through the simulation algorithms for max-stable processes. Statistical Science, 37(1):42–63, February 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/journals/
statistical-science/volume-37/issue-1/A-Comparative-
Tour-through-the-Simulation-Algorithms-for-Max-Stable/
10.1214/20-STS820.full.

Ogburn:2014:CDI

[OV14] Elizabeth L. Ogburn and Tyler J. VanderWeele. Causal diagrams for interference. Statistical Science, 29(4):559–578, November 2014. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1421330547.

Owen:2016:BCV

[OW16] Art B. Owen and Jingshu Wang. Bi-cross-validation for factor analysis. Statistical Science, 31(1):119–139, February 2016. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1455115917.

Owen:1988:ODC

[Owe88] R. J. Owen. [Optimal designs for comparing test treatments with controls]: Comment. Statistical Science, 3(4):486–487, November 1988. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177012772.

Owen:2019:CUE

[Owe19] Art B. Owen. Comment: Unreasonable effectiveness of Monte Carlo. Statistical Science, 34(1):29–33, February 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1555056027.

Opsomer:2001:NRC

[OWY01] Jean Opsomer, Yuedong Wang, and Yuhong Yang. Nonparametric regressin with correlated errors. Statistical Science, 16 (2):134–153, May 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009213287.

Prentice:2022:ITC

[PA22] Ross L. Prentice and Aaron K. Aragaki. Intention-to-treat comparisons in randomized trials. Statistical Science, 37(3):380– 393, August 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-3/
Intention-to-Treat-Comparisons-in-Randomized-Trials/
10.1214/21-STS830.full.

Paciorek:2010:ISS

[Pac10] Christopher J. Paciorek. The importance of scale for spatialconfounding bias and precision of spatial regression estimators. Statistical Science, 25(1):107–125, February 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1280841736.

PSO:1994:RSP

[Pan94] Panel on Statistics and Oceanography. Report on statistics and physical oceanography. Statistical Science, 9(2):167–201, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010481.

Panaretos:2011:CDR

[Pan11] Victor M. Panaretos. A conversation with David R. Brillinger. Statistical Science, 26(3):440–469, August 2011. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1320066930.

Parzen:1992:SMD

[Par92] Emanuel Parzen. [Statistical methods for data with long-range dependence]: Comment. Statistical Science, 7(4):420, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011124.

Parzen:2004:QPS

[Par04] Emanuel Parzen. Quantile probability and statistical data modeling. Statistical Science, 19(4):652–662, November 2004. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1113832730.

Parmigiani:2023:DRP

[Par23] Giovanni Parmigiani. Defining replicability of prediction rules. Statistical Science, 38(4):543–556, November 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-4/Defining-Replicability-of-
Prediction-Rules/10.1214/23-STS891.full.

Pires:2010:SME

[PB10] Ana M. Pires and João A. Branco. A statistical model to explain the Mendel–Fisher controversy. Statistical Science, 25(4): 545–565, November 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1300108237.

Pearl:2014:EVD

[PB14] Judea Pearl and Elias Bareinboim. External validity: From do-calculus to transportability across populations. Statistical Science, 29(4):579–595, November 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1421330548.

Porcu:2024:MMJ

[PBSO24] Emilio Porcu, Moreno Bevilacqua, Robert Schaback, and Chris J. Oates. The Matérn model: a journey through statistics, numerical analysis and machine learning. Statistical Science, 39(3):469–492, August 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-39/issue-3/The-Mat%c3%a9rn-Model--A-
Journey-Through-Statistics-Numerical-Analysis/10.1214/
24-STS923.full.

Putt:2004:NAA

[PC04] Mary E. Putt and Vernon M. Chinchilli. Nonparametric approaches to the analysis of crossover studies. Statistical Science, 19(4):712–719, November 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1113832735.

Perlman:2012:RSE

[PC12] Michael D. Perlman and Sanjay Chaudhuri. Reversing the Stein effect. Statistical Science, 27(1):135–143, February 2012. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1331729987.

Perlmutter:1998:MIC

[PCT^{+}98] Sharon M. Perlmutter, Pamela C. Cosman, Chien-Wen Tseng, Richard A. Olshen, Robert M. Gray, King C. P. Li, and Colleen J. Bergin. Medical image compression and vector quantization. Statistical Science, 13(1):30–53, February 1998. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1028905972.

Pearl:1993:BAE

[Pea93] Judea Pearl. [Bayesian analysis in expert systems]: Comment: Graphical models, causality and intervention. Statistical Science, 8(3):266–269, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177010894.

Pemantle:1995:TIP

[Pem95] Robin Pemantle. Tree-indexed processes. Statistical Science, 10(2):200–213, May 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010038.

Pena:2006:DMS

[Peñ06] Edsel A. Peña. Dynamic modeling and statistical analysis of event times. Statistical Science, 21(4):487–500, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177334525.

Perlman:1987:RMA

[Per87] Michael D. Perlman. [A review of multivariate analysis]: Comment: Group symmetry covariance models. Statistical Science, 2(4):421–425, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177013114.

Perline:2005:SWF

[Per05] Richard Perline. Strong, weak and false inverse power laws. Statistical Science, 20(1):68–88, February 2005. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1118065043.

Pfeffermann:2007:CSS

[Pfe07] Danny Pfeffermann. Comment: Struggles with survey weighting and regression modeling. Statistical Science, 22(2):179–183, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1190905516. See [Gel07c].

Pfeffermann:2013:NID

[Pfe13] Danny Pfeffermann. New important developments in small area estimation. Statistical Science, 28(1):40–68, February 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1359468408.

Pfeiffer:2009:CDG

[PGP09] Ruth M. Pfeiffer, Mitchell H. Gail, and David Pee. On combining data from genome-wide association studies to discover disease-associated SNPs. Statistical Science, 24(4):547–560, November 2009. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1271770348.

Papamarkou:2022:CMC

[PHYW22] Theodore Papamarkou, Jacob Hinkle, M. Todd Young, and David Womble. Challenges in Markov chain Monte Carlo for Bayesian neural networks. Statistical Science, 37(3):425–442, August 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-3/
Challenges-in-Markov-Chain-Monte-Carlo-for-Bayesian-
Neural-Networks/10.1214/21-STS840.full.

Pickands:1989:EVA

[Pic89] James Pickands III. [Extreme value analysis of environmental time series: An application to trend detection in groundlevel ozone]: Comment. Statistical Science, 4(4):388, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012406.

Pinson:2013:WEF

[Pin13] Pierre Pinson. Wind energy: Forecasting challenges for its operational management. Statistical Science, 28(4):564–585, November 2013. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1386078879.

Portnoy:1997:GHL

[PK97] Stephen Portnoy and Roger Koenker. The Gaussian hare and the Laplacian tortoise: computability of squared-error versus absolute-error estimators. Statistical Science, 12(4):279– 300, November 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1030037960.

Pawitan:2021:CL

[PL21] Yudi Pawitan and Youngjo Lee. Confidence as likelihood. Statistical Science, 36(4):509–517, November 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-4/Confidence-as-Likelihood/

10.1214/20-STS811.full. See comments [LB22] and rejoinder [PL22].

Pawitan:2022:RCL

[PL22] Yudi Pawitan and Youngjo Lee. Rejoinder: Confidence as likelihood. Statistical Science, 37(4):628–629, November 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-4/

Rejoinder-Confidence-as-Likelihood/10.1214/22-STS869.

full. See [PL21, LB22].

Plackett:1989:RFH

[Pla89] Robin L. Plackett. [R. A. Fisher on the history of inverse probability]: Comment. Statistical Science, 4(3):256–258, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012489.

Prentice:1993:RMD

[PM93] Ross L. Prentice and Lloyd A. Mancl. [Regression models for discrete longitudinal responses]: Comment. Statistical Science, 8(3):302–304, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010901.

Pollard:1986:CLT

[Pol86] David Pollard. [The central limit theorem around 1935]: Comment. Statistical Science, 1(1):94–95, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013821.  See

[Le 86a].

Pollard:1989:AEPa

[Pol89a] David Pollard. Asymptotics via empirical processes. Statistical Science, 4(4):341–354, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012394.

Pollard:1989:AEPb

[Pol89b] David Pollard. [Asymptotics via empirical processes]: Rejoinder. Statistical Science, 4(4):365–366, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012399.

Polson:1992:PMC

[Pol92] Nicholas G. Polson. [Practical Markov chain Monte Carlo]: Comment. Statistical Science, 7(4):490–491, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011141.

Politis:2003:IBM

[Pol03] Dimitris N. Politis. The impact of bootstrap methods on time series analysis. Statistical Science, 18(2):219–230, May 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1063994977.

Politis:2024:SVV

[Pol24] Dimitris N. Politis. Studentization versus variance stabilization: a simple way out of an old dilemma. Statistical Science, 39(3):409–427, August 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-39/issue-3/Studentization-Versus-Variance-
Stabilization--A-Simple-Way-Out-of/10.1214/23-STS917.
full.

Portnoy:1987:SAH

[Por87] Stephen Portnoy. [Statistical analysis of hominoid molecular evolution]: Comment. Statistical Science, 2(2):207–208, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013354.

Portnoy:1994:LCP

[Por94] Stephen Portnoy. A Lewis Carroll pillow problem: Probability of an obtuse triangle. Statistical Science, 9(2):279–284, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010497.

Possolo:2023:TTT

[Pos23] Antonio Possolo. Tracking truth through measurement and the spyglass of statistics. Statistical Science, 38(4):655–671, November 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-4/
Tracking-Truth-Through-Measurement-and-the-Spyglass-
of-Statistics/10.1214/23-STS899.full.

Presanis:2013:CDD

[POSD13] Anne M. Presanis, David Ohlssen, David J. Spiegelhalter, and Daniela De Angelis. Conflict diagnostics in directed acyclic graphs, with applications in Bayesian evidence synthesis. Statistical Science, 28(3):376–397, August 2013. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1377696942.

Pourahmadi:2011:CEG

[Pou11] Mohsen Pourahmadi. Covariance estimation: The GLM and regularization perspectives. Statistical Science, 26(3):369–387, August 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1320066926.

Peng:2014:TMV

[PQW14] Liang Peng, Yongcheng Qi, and Fang Wang. Test for a mean vector with fixed or divergent dimension. Statistical Science, 29 (1):113–127, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1399645740.

Pratt:1986:SRC

[Pra86] John W. Pratt. [Savage revisited]: Comment. Statistical Science, 1(4):498–499, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177013523.

Pratt:1990:ABE

[Pra90] John W. Pratt. [Applications in business and economic statistics: Some personal views]: Comment. Statistical Science, 5(4): 397–399, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012013.

Press:1987:RMA

[Pre87] S. James Press. [A review of multivariate analysis]: Comment. Statistical Science, 2(4):430–432, November 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013117.

Papaspiliopoulos:2007:GFP

[PRS07] Omiros Papaspiliopoulos, Gareth O. Roberts, and Martin Sköld. A general framework for the parametrization of hierarchical models. Statistical Science, 22(1):59–73, February 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1185975637.

Penttinen:2000:RAP

[PS00] Antti Penttinen and Dietrich Stoyan. Recent applications of point process methods in forestry statistics. Statistical Science, 15(1):61–78, February 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009212674.

Prado:2011:VPR

[PS11] Raquel Prado and Bruno Sansó. The 2004 Venezuelan Presidential Recall Referendum: Discrepancies between two exit polls and official results. Statistical Science, 26(4):517–527, November 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1330437933.

Piegorsch:1998:SAE

[PSES98] Walter W. Piegorsch, Eric P. Smith, Don Edwards, and Richard L. Smith. Statistical advances in environmental science. Statistical Science, 13(2):186–208, May 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1028905935.

Polson:2015:PAS

[PSW15] Nicholas G. Polson, James G. Scott, and Brandon T. Willard. Proximal algorithms in statistics and machine learning. Statistical Science, 30(4):559–581, November 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1449670858.

Patil:2003:GNS

[PT03] G. P. Patil and C. Taillie. Geographic and network surveillance via scan statistics for critical area detection. Statistical Science, 18(4):457–465, November 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1081443229.

Pena:2010:CGC

[PT10] Daniel Peña and Ruey S. Tsay. A conversation with George C. Tiao. Statistical Science, 25(3):408–428, August 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1294167967.

Pericchi:2011:QAD

[PT11] Luis Pericchi and David Torres. Quick anomaly detection by the Newcomb–Benford law, with applications to electoral processes data from the USA, Puerto Rico and Venezuela. Statistical Science, 26(4):502–516, November 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1330437932.

Perlman:1999:ENT

[PW99] Michael D. Perlman and Lang Wu. The Emperor’s new tests. Statistical Science, 14(4):355–369, November 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1009212517.  See

comment [Ber99].

Pfannkuch:2000:STS

[PW00] Maxine Pfannkuch and Chris J. Wild. Statistical thinking an statistical practice: Themes gleaned from professional statisticians. Statistical Science, 15(2):132–152, May 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1009212754.

Puccetti:2015:EDC

[PW15] Giovanni Puccetti and Ruodu Wang. Extremal dependence concepts. Statistical Science, 30(4):485–517, November 2015. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1449670855.

Phillips:2005:CSO

[PY05] Peter C. B. Phillips and Jun Yu. Comment: A selective overview of nonparametric methods in financial econometrics. Statistical Science, 20(4):338–343, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1137076648.  See

[Fan05b].

Pyke:1989:AEP

[Pyk89] Ron Pyke. [Asymptotics via empirical processes]: Comment. Statistical Science, 4(4):357–360, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012397.

Qian:2020:LMM

[QKM20] Tianchen Qian, Predrag Klasnja, and Susan A. Murphy. Linear mixed models with endogenous covariates: Modeling sequential treatment effects with application to a mobile health study. Statistical Science, 35(3):375–390, August 2020. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-35/issue-3/Linear-Mixed-Models-with-
Endogenous-Covariates--Modeling-Sequential-Treatment/
10.1214/19-STS720.full.

Quinn:2002:ICM

[QM02] Kevin M. Quinn and Andrew D. Martin. An integrated computational model of multiparty electoral competition. Statistical Science, 17(4):405–419, November 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1049993200.

Quintana:2022:DDP

[QMJM22] Fernando A. Quintana, Peter Müller, Alejandro Jara, and Steven N. MacEachern. The dependent Dirichlet process and related models. Statistical Science, 37(1):24–41, February 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-1/
The-Dependent-Dirichlet-Process-and-Related-Models/
10.1214/20-STS819.full.

Quintana-Orti:2019:KRP

[QOS19] Gregorio Quintana-Ortí and Amelia Simó. A kernel regression procedure in the 3D shape space with an application to online sales of children’s wear. Statistical Science, 34(2):236–252, May 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1563501640.

Quessy:2024:GCM

[Que24] Jean-François Quessy. A general construction of multivariate dependence structures with nonmonotone mappings and its applications. Statistical Science, 39(3):391–408, August 2024. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-39/issue-3/A-General-Construction-of-
Multivariate-Dependence-Structures-with-Nonmonotone-
Mappings/10.1214/23-STS916.full.

Raade:1997:CHB

[Råd97] Lennart Råde. A conversation with Harald Bergström. Statistical Science, 12(1):53–60, February 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1029963262.

Raftery:1989:EVA

[Raf89] Adrian E. Raftery. Extreme value analysis of environmental time series: An application to trend detection in ground-level ozone: Comment: Are ozone exceedance rates decreasing? Statistical Science, 4(4):378–381, November 1989. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012401.

Raftery:1992:SMD

[Raf92] Adrian E. Raftery. [Statistical methods for data with longrange dependence]: Comment: Computational aspects of fractionally differenced ARIMA modeling for long- memory time series. Statistical Science, 7(4):421–422, November 1992. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011125.

Raftery:2014:BPP

[RAG14] Adrian E. Raftery, Leontine Alkema, and Patrick Gerland. Bayesian population projections for the United Nations. Statistical Science, 29(1):58–68, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1399645729.

Ramsay:1988:MRSa

[Ram88a] J. O. Ramsay. Monotone regression splines in action. Statistical Science, 3(4):425–441, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012761.

Ramsay:1988:MRSb

[Ram88b] J. O. Ramsay. [Monotone regression splines in action]: Rejoinder. Statistical Science, 3(4):459–461, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012766.

Ramachandran:1993:RPA

[Ram93] Vijaya Ramachandran. Randomization in parallel algorithms. Statistical Science, 8(1):65–69, February 1993. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011085.

Randles:2007:CRV

[Ran07] Ronald Herman Randles. A conversation with Robert V. Hogg. Statistical Science, 22(1):137–152, February 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1185975641.

Rao:1987:PFOa

[Rao87a] C. Radhakrishna Rao. Prediction of future observations in growth curve models. Statistical Science, 2(4):434–447, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013119.

Rao:1987:PFOb

[Rao87b] C. Radhakrishna Rao. [Prediction of future observations in growth curve models]: Rejoinder. Statistical Science, 2(4): 467–471, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013126.

Rao:1988:SMF

[Rao88] C. Radhakrishna Rao. [Selection models and the file drawer problem]: Comment. Statistical Science, 3(1):131, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013017.

Rao:1989:GAI

[Rao89] C. R. Rao. [The geometry of asymptotic inference]: Comment. Statistical Science, 4(3):229–231, August 1989. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012484.

Rao:1992:RFF

[Rao92] C. Radhakrishna Rao. R. A. Fisher: The founder of modern statistics. Statistical Science, 7(1):34–48, February 1992. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011442.

Rao:2011:IFB

[Rao11a] J. N. K. Rao. Impact of frequentist and Bayesian methods on survey sampling practice: A selective appraisal. Statistical Science, 26(2):240–256, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:/

/projecteuclid.org/euclid.ss/1312204015. See discussion

[Mee11, Sed11, Slu11].

Rao:2011:R

[Rao11b] J. N. K. Rao. Rejoinder. Statistical Science, 26(2):266–270, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1312204019.

Riani:2014:PFC

[RAP14] Marco Riani, Anthony C. Atkinson, and Domenico Perrotta. A parametric framework for the comparison of methods of very robust regression. Statistical Science, 29(1):128–143, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1399645741.

Roussas:2020:CFJ

[RB20] George G. Roussas and Debasis Bhattacharya. A conversation with Francisco J. Samaniego. Statistical Science, 35(2):321– 333, May 2020. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1591171234.

Rothenhausler:2023:DRG

[RB23] Dominik Rothenhäusler and Peter Bühlmann. Distributionally robust and generalizable inference. Statistical Science, 38(4):527–542, November 2023. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-4/Distributionally-Robust-and-
Generalizable-Inference/10.1214/23-STS902.full.

Ritov:2014:BAC

[RBGK14] Y. Ritov, P. J. Bickel, A. C. Gamst, and B. J. K. Kleijn. The Bayesian analysis of complex, high-dimensional models: Can it be CODA? Statistical Science, 29(4):619–639, November 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1421330550.

Robert:2004:ISI

[RC04] Christian P. Robert and George Casella. Introduction to the special issue: Bayes Then and now. Statistical Science, 19 (1):1–2, February 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1089808270.

Robert:2011:SHM

[RC11] Christian Robert and George Casella. A short history of Markov chain Monte Carlo: Subjective recollections from incomplete data. Statistical Science, 26(1):102–115, February 2011. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1307626568.

Robert:2009:HJT

[RCR09a] Christian P. Robert, Nicolas Chopin, and Judith Rousseau. Harold Jeffreys’s theory of probability revisited. Statistical Science, 24(2):141–172, May 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1263478373.

Robert:2009:RHJ

[RCR09b] Christian P. Robert, Nicolas Chopin, and Judith Rousseau. Rejoinder: Harold Jeffreys’s theory of probability revisited. Statistical Science, 24(2):191–194, May 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1263478380.

Reddy:2006:MLA

[RD06] Srinivas K. Reddy and Mayukh Dass. Modeling on-line art auction dynamics using functional data analysis. Statistical Science, 21(2):179–193, May 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1154979820.

Read:2004:CNJ

[Rea04] Campbell B. Read. A conversation with Norman L. Johnson. Statistical Science, 19(3):544–560, August 2004. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1110999314.

Reid:1988:SMSa

[Rei88a] N. Reid. Saddlepoint methods and statistical inference. Statistical Science, 3(2):213–227, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012906.

Reid:1988:SMSb

[Rei88b] N. Reid. [Saddlepoint methods and statistical inference]: Rejoinder. Statistical Science, 3(2):237–238, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012914.

Reid:1994:CSD

[Rei94] Nancy Reid. A conversation with Sir David Cox. Statistical Science, 9(3):439–455, August 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177010394.

Reid:1995:IBE

[Rei95a] N. Reid. [Inference based on estimating functions in the presence of nuisance parameters]: Rejoinder. Statistical Science, 10(2):193–196, May 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010036.

Reid:1995:RCI

[Rei95b] N. Reid. The roles of conditioning in inference. Statistical Science, 10(2):138–157, May 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177010027.

Reinhart:2018:RRS

[Rei18a] Alex Reinhart. Rejoinder: “A Review of Self-Exciting Spatio-Temporal Point Processes and Their Applications”. Statistical Science, 33(3):330–333, August 2018. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:

//projecteuclid.org/euclid.ss/1534147226. See [Rei18b,

Oga18, Sch18, Zhu18].

Reinhart:2018:RSE

[Rei18b] Alex Reinhart. A review of self-exciting spatio-temporal point processes and their applications. Statistical Science, 33(3):299– 318, August 2018. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/

euclid.ss/1534147221. See [Oga18, Sch18, Zhu18] and rejoinder [Rei18a].

Reid:1989:GAI

[RF89] N. Reid and D. A. S. Fraser. [The geometry of asymptotic inference]: Comment. Statistical Science, 4(3):231–233, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012485.

Rainforth:2024:MBE

[RFIS24] Tom Rainforth, Adam Foster, Desi R. Ivanova, and Freddie Bickford Smith. Modern Bayesian experimental design. Statistical Science, 39(1):100–114, February 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-39/issue-1/Modern-Bayesian-Experimental-
Design/10.1214/23-STS915.full.

Ramdas:2023:GTS

[RGVS23] Aaditya Ramdas, Peter Grünwald, Vladimir Vovk, and Glenn Shafer. Game-theoretic statistics and safe anytime-valid inference. Statistical Science, 38(4):576–601, November 2023. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-4/Game-Theoretic-Statistics-
and-Safe-Anytime-Valid-Inference/10.1214/23-STS894.
full.

Rukhin:1987:SSWa

[RH87a] Andrew L. Rukhin and H. K. Hsieh. Survey of Soviet work in reliability. Statistical Science, 2(4):484–495, November 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013128.

Rukhin:1987:SSWb

[RH87b] Andrew L. Rukhin and H. K. Hsieh. [Survey of Soviet work in reliability]: Rejoinder. Statistical Science, 2(4):503, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013134.

Randles:2004:ISI

[RHC04] Ronald H. Randles, Thomas P. Hettmansperger, and George Casella. Introduction to the special issue: Nonparametric statistics. Statistical Science, 19(4):561, November 2004. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1113832719.

Richardson:2014:NBS

[RHGF14] Amy Richardson, Michael G. Hudgens, Peter B. Gilbert, and Jason P. Fine. Nonparametric bounds and sensitivity analysis of treatment effects. Statistical Science, 29(4):596–618, November 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1421330549.

Rice:1986:SPI

[Ric86] John A. Rice. [A statistical perspective on ill-posed inverse problems]: Comment. Statistical Science, 1(4):522–523, November 1986. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177013528. See [O’S86b].

Rice:2015:CRO

[Ric15] John A. Rice. A conversation with Richard A. Olshen. Statistical Science, 30(1):118–132, February 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1425492443.

Ripley:1994:NNR

[Rip94] B. D. Ripley. [Neural networks: A review from statistical perspective]: Comment. Statistical Science, 9(1):45–48, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010644.

Rissanen:1987:RTF

[Ris87] Jorma Rissanen. [Rational transfer function approximation]: Comment. Statistical Science, 2(2):156–157, May 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013347.

Ritov:2011:RWD

[Rit11] Ya’acov Ritov. A random walk with drift: Interview with Peter J. Bickel. Statistical Science, 26(1):150–159, February 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1307626571.

Raftery:1992:PMC

[RL92] Adrian E. Raftery and Steven M. Lewis. [Practical Markov chain Monte Carlo]: Comment: One long run with diagnostics: Implementation strategies for Markov chain Monte Carlo. Statistical Science, 7(4):493–497, November 1992. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011143.

Robertson:2023:RRA

[RLLKV23a] David S. Robertson, Kim May Lee, Boryana C. López-Kolkovska, and Sofía S. Villar. Rejoinder: Response-adaptive randomization in clinical trials. Statistical Science, 38(2):233– 239, May 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-2/
Rejoinder-Response-Adaptive-Randomization-in-Clinical-
Trials/10.1214/23-STS865REJ.full.

Robertson:2023:RAR

[RLLKV23b] David S. Robertson, Kim May Lee, Boryana C. López-Kolkovska, and Sofía S. Villar. Response-adaptive randomization in clinical trials: From myths to practical considerations. Statistical Science, 38(2):185–208, May 2023. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-2/Response-Adaptive-Randomization-
in-Clinical-Trials--From-Myths-to/10.1214/22-STS865.
full.

Ripamonti:2017:CFV

[RLQ17] Enrico Ripamonti, Chris Lloyd, and Piero Quatto. Contemporary frequentist views of the 2 × 2 binomial trial. Statistical Science, 32(4):600–615, 11 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Ridgeway:2007:CDD

[RM07] Greg Ridgeway and Daniel F. McCaffrey. Comment: Demystifying double robustness: A comparison of alternative strategies for estimating a population mean from incomplete data. Statistical Science, 22(4):540–543, November 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1207580168. See [KS07a].

Reimherr:2013:QDF

[RN13] Matthew Reimherr and Dan L. Nicolae. On quantifying dependence: A framework for developing interpretable measures. Statistical Science, 28(1):116–130, February 2013. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1359468411.

Roberts:1988:EDS

[Rob88] Harry V. Roberts. [Employment discrimination and statistical science]: Comment. Statistical Science, 3(2):167–170, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012897.

Robbins:1989:WSS

[Rob89] Herbert Robbins. [Who solved the Secretary Problem?]: Comment. Statistical Science, 4(3):291, August 1989. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012495.

Roberts:1990:ABEa

[Rob90a] Harry V. Roberts. Applications in business and economic statistics: Some personal views. Statistical Science, 5(4):372– 390, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012008.

Roberts:1990:ABEb

[Rob90b] Harry V. Roberts. [Applications in business and economic statistics: Some personal views]: Rejoinder. Statistical Science, 5(4):399–402, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012014.

Robinson:1991:BGTa

[Rob91a] G. K. Robinson. That BLUP is a good thing: The estimation of random effects. Statistical Science, 6(1):15–32, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011926.

Robinson:1991:BGTb

[Rob91b] G. K. Robinson. [That BLUP is a good thing: The estimation of random effects]: Rejoinder. Statistical Science, 6(1): 48–51, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011933.

Roberts:1993:CST

[Rob93a] Harry V. Roberts. [Can statistics tell us what we do not want to hear? the case of complex salary structures]: Comment. Statistical Science, 8(2):171–176, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011009.

Robinson:1993:ISC

[Rob93b] G. K. Robinson. [Is industrial statistics out of control?]: Comment. Statistical Science, 8(4):395–397, November 1993. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010780.

Robert:1995:CCM

[Rob95] Christian P. Robert. Convergence control methods for Markov chain Monte Carlo algorithms. Statistical Science, 10(3):231– 253, August 1995. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177009937.

Robert:2011:DBP

[Rob11] Christian P. Robert. Discussion of “Is Bayes Posterior just Quick and Dirty Confidence?” by D. A. S. Fraser. Statistical Science, 26(3):317–318, August 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1320066919. See [Fra11a].

Roeder:1994:DFRa

[Roe94a] Kathryn Roeder. DNA fingerprinting: A review of the controversy. Statistical Science, 9(2):222–247, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010488.

Roeder:1994:DFRb

[Roe94b] Kathryn Roeder. [DNA fingerprinting: A review of the controversy]: Rejoinder. Statistical Science, 9(2):267–278, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010496.

Rosenbaum:1987:RSCa

[Ros87a] Paul R. Rosenbaum. The role of a second control group in an observational study. Statistical Science, 2(3):292–306, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013232.

Rosenbaum:1987:RSCb

[Ros87b] Paul R. Rosenbaum. [The role of a second control group in an observational study]: Rejoinder. Statistical Science, 2(3): 313–316, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013237.

Rosen:1988:ASH

[Ros88] L. A. Rosen. Animal studies of human hazards. Statistical Science, 3(3):298–305, August 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012830.

Rosenthal:1992:PMC

[Ros92] Jeffrey S. Rosenthal. [Practical Markov chain Monte Carlo]: Comment. Statistical Science, 7(4):498, November 1992. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011144.

Rosenberger:1996:NDA

[Ros96] William F. Rosenberger. New directions in adaptive designs. Statistical Science, 11(2):137–149, May 1996. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1038425657.

Rosenbaum:1999:CAC

[Ros99] Paul R. Rosenbaum. Choice as an alternative to control in observational studies. Statistical Science, 14(3):259–304, August 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1009212410.

Rosenbaum:2002:CAR

[Ros02] Paul R. Rosenbaum. Covariance adjustment in randomized experiments and observational studies. Statistical Science, 17(3): 286–327, August 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1042727942.

Rosenbaum:2006:CPD

[Ros06] Paul R. Rosenbaum. Comment: The place of death in the quality of life. Statistical Science, 21(3):313–316, August 2006. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1166642432.

Rosenberger:2015:CNF

[Ros15] William F. Rosenberger. A conversation with Nancy Flournoy. Statistical Science, 30(1):133–146, February 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1425492444.

Rosenbaum:2017:GSE

[Ros17] Paul R. Rosenbaum. The general structure of evidence factors in observational studies. Statistical Science, 32(4):514–530, 11 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Rosenbaum:2023:CWR

[Ros23a] Paul R. Rosenbaum. Can we reliably detect biases that matter in observational studies? Statistical Science, 38(3):440– 457, August 2023. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
journals/statistical-science/volume-38/issue-3/Can-
We-Reliably-Detect-Biases-that-Matter-in-Observational-
Studies/10.1214/23-STS882.full.

Rosychuk:2023:CMT

[Ros23b] Rhonda J. Rosychuk. A conversation with Mary E. Thompson. Statistical Science, 38(3):514–524, August 2023. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-3/A-Conversation-with-Mary-
E-Thompson/10.1214/22-STS877.full.

Rinott:2018:CDP

[ROSS18] Yosef Rinott, Christine M. O’Keefe, Natalie Shlomo, and Chris Skinner. Confidentiality and differential privacy in the dissemination of frequency tables. Statistical Science, 33(3):358– 385, August 2018. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1534147228.

Royall:1989:ITP

[Roy89] Richard Royall. [Investigating therapies of potentially great benefit: ECMO]: Comment. Statistical Science, 4(4):318– 319, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012387.

Royall:1991:ESRa

[Roy91a] Richard M. Royall. Ethics and statistics in randomized clinical trials. Statistical Science, 6(1):52–62, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011934.

Royall:1991:ESRb

[Roy91b] Richard M. Royall. [Ethics and statistics in randomized clinical trials]: Rejoinder. Statistical Science, 6(1):83–88, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011942.

Racine-Poon:1992:PMC

[RP92] Amy Racine-Poon. [Practical Markov chain Monte Carlo]: Comment. Statistical Science, 7(4):492–493, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011142.

Robert:2024:EBC

[RP24] Christian P. Robert and Dennis Prangle. Editorial: Bayesian computations in the 21st century. Statistical Science, 39(1):1– 2, February 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-1/
Editorial-Bayesian-Computations-in-the-21st-Century/
10.1214/23-STS920.full.

Rosenthal:1988:SMF

[RR88] Robert Rosenthal and Donald B. Rubin. [Selection models and the file drawer problem]: Comment: Assumptions and procedures in the file drawer problem. Statistical Science, 3(1): 120–125, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013014.

Roberts:2001:OSV

[RR01] Gareth O. Roberts and Jeffrey S. Rosenthal. Optimal scaling for various Metropolis–Hastings algorithms. Statistical Science, 16(4):351–367, November 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1015346320.

Richardson:2014:ABS

[RR14a] Thomas S. Richardson and James M. Robins. ACE bounds; SEMs with equilibrium conditions. Statistical Science, 29(3): 363–366, August 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1411437515.

Richardson:2014:CER

[RR14b] Thomas S. Richardson and Andrea Rotnitzky. Causal etiology of the research of James M. Robins. Statistical Science, 29(4):

459–484, November 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1421330543.

Robert:2016:NBC

[RR16] Christian P. Robert and Judith Rousseau. Nonparametric Bayesian clay for robust decision bricks. Statistical Science, 31(4):506–510, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1484816577.

Robert:2017:HPP

[RR17] Christian P. Robert and Judith Rousseau. How principled and practical are penalised complexity priors? Statistical Science, 32(1):36–40, February 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1491465624.

Rossell:2023:ABV

[RR23] David Rossell and Francisco Javier Rubio. Additive Bayesian variable selection under censoring and misspecification. Statistical Science, 38(1):13–29, February 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-1/Additive-Bayesian-Variable-
Selection-under-Censoring-and-Misspecification/10.1214/
21-STS846.full.

Reshef:2020:EIE

[RRSM20] Yakir A. Reshef, David N. Reshef, Pardis C. Sabeti, and Michael Mitzenmacher. Equitability, interval estimation, and statistical power. Statistical Science, 35(2):202–217, May 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1591171227.

Rabinowitz:1997:REA

[RS97] Daniel Rabinowitz and Steven Shea. Random effects analysis of children’s blood pressure data. Statistical Science, 12(3): 185–194, August 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1030037908.

Rice:2008:CME

[RS08a] Kenneth Rice and David Spiegelhalter. Comment: Microarrays, empirical Bayes and the two-groups model. Statistical Science, 23(1):41–44, February 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1215441280. See [GM05a].

Rosenberger:2008:HCD

[RS08b] William F. Rosenberger and Oleksandr Sverdlov. Handling covariates in the design of clinical trials. Statistical Science, 23 (3):404–419, August 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1233153066.

Roberts:1995:BCS

[RSG95] G. O. Roberts, S. K. Sahu, and W. R. Gilks. [Bayesian computation and stochastic systems]: Comment. Statistical Science, 10(1):49–51, February 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010127.

Rinott:2022:SAQ

[RSK22] Yosef Rinott, Tomer Shoham, and Gil Kalai. Statistical aspects of the quantum supremacy demonstration. Statistical Science, 37(3):322–347, August 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-37/issue-3/Statistical-Aspects-of-the-
Quantum-Supremacy-Demonstration/10.1214/21-STS836.full.

Robins:2007:CPD

[RSLGR07] James Robins, Mariela Sued, Quanhong Lei-Gomez, and Andrea Rotnitzky. Comment: Performance of double-robust estimators when “inverse probability” weights are highly variable. Statistical Science, 22(4):544–559, November 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1207580169.

Rigollet:2012:SEE

[RT12] Philippe Rigollet and Alexandre B. Tsybakov. Sparse estimation by exponential weighting. Statistical Science, 27(4): 558–575, November 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1356098556.

Ramirez:2013:ADA

[RTW13] Jorge M. Ramirez, Enrique A. Thomann, and Edward C. Waymire. Advection–dispersion across interfaces. Statistical Science, 28(4):487–509, November 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1386078875.

Rinaldo:2019:CSI

[RTW19] Alessandro Rinaldo, Ryan J. Tibshirani, and Larry Wasserman. Comment: Statistical inference from a predictive perspective. Statistical Science, 34(4):599–603, November 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1578474025.

Rubin:1990:APT

[Rub90] Donald B. Rubin. [On the application of probability theory to agricultural experiments. essay on principles. section 9.] comment: Neyman (1923) and causal inference in experiments and observational studies. Statistical Science, 5(4):472–480, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012032.

Rubin:2006:CIT

[Rub06a] Donald B. Rubin. Causal inference through potential outcomes and principal stratification: Application to studies with “censoring” due to death. Statistical Science, 21(3):299–309, August 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.

ss/1166642430. See comments [Fie06a, Kor06].

Rubin:2006:R

[Rub06b] Donald B. Rubin. Rejoinder. Statistical Science, 21(3):319–321, August 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1166642434.

Rukhin:1990:DDT

[Ruk90] Andrew L. Rukhin. [Developments in decision-theoretic variance estimation]: Comment. Statistical Science, 5(1):113–116, February 1990. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177012270.

Rubin:2006:ECE

[RW06] Donald B. Rubin and Richard P. Waterman. Estimating the causal effects of marketing interventions using propensity score methodology. Statistical Science, 21(2):206–222, May 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1154979822.

Roeder:2009:GWS

[RW09] Kathryn Roeder and Larry Wasserman. Genome-wide significance levels and weighted hypothesis testing. Statistical Science, 24(4):398–413, November 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1271770339.

Robertson:2023:OMH

[RWR23] David S. Robertson, James M. S. Wason, and Aaditya Ramdas. Online multiple hypothesis testing. Statistical Science, 38(4):557–575, November 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-4/Online-Multiple-Hypothesis-
Testing/10.1214/23-STS901.full.

Ryan:1995:IBE

[Rya95] Louise M. Ryan. [Inference based on estimating functions in the presence of nuisance parameters]: Comment. Statistical Science, 10(2):189–193, May 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177010035.

Ryan:2003:EBE

[Rya03] Louise Ryan. Epidemiologically based environmental risk assessment. Statistical Science, 18(4):466–480, November 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1081443230.

Ryan:2015:CNL

[Rya15] Louise Ryan. A conversation with Nan Laird. Statistical Science, 30(4):582–596, November 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1449670859.

Sunley:1992:RCN

[SAA^{+}92] Judith Sunley, N. Altman, J. F. Angers, D. Banks, D. Duffy, J. Hardwick, C. Leger, M. Martin, D. Nolan, A. Owen, D. Politis, Peter Arzberger, K. Roeder, T. N. Sriram, T. Stukel, Z. Ying, Keith Crank, Nell Sedransk, James R. Maar, Michael R. Chernick, Cindy L. Christiansen, Agnes M. Herzberg, and R. L. Tweedie. Readers’ comments to the new researchers’ committee report. Statistical Science, 7(2):256–266, May 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011365.

Smith:2019:GCL

[SAC19] Anna L. Smith, Dena M. Asta, and Catherine A. Calder. The geometry of continuous latent space models for network data. Statistical Science, 34(3):428–453, August 2019. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1570780978.

Sakaguchi:1989:WSS

[Sak89] Minoru Sakaguchi. [Who solved the Secretary Problem?]: Comment. Statistical Science, 4(3):292–293, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012496.

Staudt:2020:SMC

[SAL^{+}20] Thomas Staudt, Timo Aspelmeier, Oskar Laitenberger, Claudia Geisler, Alexander Egner, and Axel Munk. Statistical molecule counting in super-resolution fluorescence microscopy: Towards quantitative nanoscopy. Statistical Science, 35(1):92– 111, February 2020. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1583226031.

Sampson:1986:SSS

[Sam86] Paul D. Sampson. [Size and shape spaces for Landmark data in two dimensions]: Comment. Statistical Science, 1(2):229–234, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013700.

Samuels:1989:WSS

[Sam89] Stephen M. Samuels. [Who solved the Secretary Problem?]: Comment: Who will solve the Secretary Problem? Statistical Science, 4(3):289–291, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012494.

Samuelson:1991:SFC

[Sam91] Paul A. Samuelson. Statistical flowers caught in amber. Statistical Science, 6(4):330–338, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011570.

Sampson:1999:CRS

[Sam99] Allan R. Sampson. A conversation with I. Richard Savage (with the assistance of Bruce Spencer). Statistical Science, 14(1): 126–148, February 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009211808.

Sampson:2007:CIO

[Sam07] Allan R. Sampson. A conversation with Ingram Olkin. Statistical Science, 22(3):450–475, August 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1199285043.

Samaniego:2008:CMH

[Sam08] Francisco J. Samaniego. A conversation with Myles Hollander. Statistical Science, 23(3):420–438, August 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1233153067.

Samworth:2018:RPL

[Sam18] Richard J. Samworth. Recent progress in log-concave density estimation. Statistical Science, 33(4):493–509, November 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1543482055.

Savitz:1988:HSH

[Sav88] David A. Savitz. Human studies of human health hazards: Comparison of epidemiology and toxicology. Statistical Science, 3(3):306–313, August 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012831.

Sawyer:1991:GSP

[Saw91] Stanley Sawyer. [The growth and stabilization of populations]: Comment: The geographical structure of populations. Statistical Science, 6(3):280–281, August 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011697.

Sanz-Alonso:2022:SAM

[SAY22] Daniel Sanz-Alonso and Ruiyi Yang. The SPDE approach to Matérn fields: Graph representations. Statistical Science, 37(4):519–540, November 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-37/issue-4/The-SPDE-Approach-to-Mat%
c3%a9rn-Fields-Graph-Representations/10.1214/21-STS838.
full.

Schofield:2016:YOC

[SB16] Matthew Schofield and Richard Barker. 50-year-old curiosities: Ancillarity and inference in capture–recapture models. Statistical Science, 31(2):161–174, May 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1464105035.

Shah:2023:DEF

[SB23] Rajen D. Shah and Peter Bühlmann. Double-estimation-friendly inference for high-dimensional misspecified models. Statistical Science, 38(1):68–91, February 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-1/Double-Estimation-Friendly-
Inference-for-High-Dimensional-Misspecified-Models/
10.1214/22-STS850.full.

Strumbelj:2024:PPF

[SBCC^{+}24] Erik Strumbelj, Alexandre Bouchard-Côté, Jukka Corander, Andrew Gelman, Håvard Rue, Lawrence Murray, Henri Pesonen, Martyn Plummer, and Aki Vehtari. Past, present and future of software for Bayesian inference. Statistical Science, 39(1):46–61, February 2024. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-39/issue-1/Past-Present-and-Future-of-
Software-for-Bayesian-Inference/10.1214/23-STS907.full.

Samuel-Cahn:1992:CES

[SC92] Ester Samuel-Cahn. A conversation with Esther Seiden. Statistical Science, 7(3):339–357, August 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011231.

Schaible:1994:SAE

[SC94] Wesley L. Schaible and Robert J. Casady. [Small area estimation: An appraisal]: Comment. Statistical Science, 9(1): 82–84, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010650.

Schonlau:2017:OCW

[SC17] Matthias Schonlau and Mick P. Couper. Options for conducting Web surveys. Statistical Science, 32(2):279–292, May 2017. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1494489816.

Schervish:1987:RMAb

[Sch87a] Mark J. Schervish. [A review of multivariate analysis]: Rejoinder. Statistical Science, 2(4):432–433, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013118.

Schervish:1987:RMAa

[Sch87b] Mark J. Schervish. A review of multivariate analysis. Statistical Science, 2(4):396–413, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013111.

Schmeiser:1992:PMC

[Sch92] Bruce Schmeiser. [Practical Markov chain Monte Carlo]: Comment. Statistical Science, 7(4):498–499, November 1992. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011145.

Schafer:1994:MII

[Sch94a] Joseph L. Schafer. [Multiple-imputation inferences with uncongenial sources of input]: Comment. Statistical Science, 9(4): 560–561, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010271.

Schervish:1994:BMT

[Sch94b] Mark J. Schervish. [Bootstrap: More than a stab in the dark?]: Comment. Statistical Science, 9(3):408–410, August 1994. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010390.

Schucany:2004:KSO

[Sch04] William R. Schucany. Kernel smoothers: An overview of curve estimators for the first graduate course in nonparametric statistics. Statistical Science, 19(4):663–675, November 2004. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1113832731.

Schenker:2011:DCB

[Sch11] Nathaniel Schenker. Discussion of “Calibrated Bayes, for Statistics in General, and Missing Data in Particular” by R. J. A. Little. Statistical Science, 26(2):179–184, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1312204004. See [Lar11].

Schoenberg:2018:CRS

[Sch18] Frederic Paik Schoenberg. Comment on “A Review of Self-Exciting Spatio-Temporal Point Processes and Their Applications” by Alex Reinhart. Statistical Science, 33(3):325–326, August 2018. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1534147224. See [Rei18b, Rei18a].

Sedransk:2010:MRD

[SCN^{+}10] Nell Sedransk, Lawrence H. Cox, Deborah Nolan, Keith Soper, Cliff Spiegelman, Linda J. Young, Katrina L. Kelner, Robert A. Moffitt, Ani Thakar, Jordan Raddick, Edward J. Ungvarsky, Richard W. Carlson, and Rolf Apweiler. Make research data public? — not always so simple: A dialogue for statisticians and science editors. Statistical Science, 25(1):41–50, February 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1280841732.

Scott:1986:CSR

[Sco86] David W. Scott. [Computers in statistical research]: Comment. Statistical Science, 1(4):442–444, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013511.  See

[Edd86a].

Scott:2017:PSE

[Sco17] James G. Scott. Prior specification is engineering, not mathematics. Statistical Science, 32(1):29–32, February 2017. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1491465622.

Su:2009:BMD

[SCtW^{+}09] Zhan Su, Niall Cardin, the Wellcome Trust Case Control Consortium, Peter Donnelly, and Jonathan Marchini. A Bayesian method for detecting and characterizing allelic heterogeneity and boosting signals in genome-wide association studies. Statistical Science, 24(4):430–450, November 2009. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1271770341.

Stewart:2006:OCA

[SDD06] Katherine J. Stewart, David P. Darcy, and Sherae L. Daniel. Opportunities and challenges applying functional data analysis to the study of open source software evolution. Statistical Science, 21(2):167–178, May 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1154979819.

Spiegelhalter:1993:BAEa

[SDLC93a] David J. Spiegelhalter, A. Philip Dawid, Steffen L. Lauritzen, and Robert G. Cowell. Bayesian analysis in expert systems. Statistical Science, 8(3):219–247, August 1993. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010888.

Spiegelhalter:1993:BAEb

[SDLC93b] David J. Spiegelhalter, A. Philip Dawid, Steffen L. Lauritzen, and Robert G. Cowell. [Bayesian analysis in expert systems]: Rejoinder. Statistical Science, 8(3):277–283, August 1993. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010898.

Sedransk:2011:DIF

[Sed11] J. Sedransk. Discussion of “Impact of Frequentist and Bayesian Methods on Survey Sampling Practice: A Selective Appraisal” by J. N. K. Rao. Statistical Science, 26(2):260–261, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1312204017. See [Rao11a].

Segal:2007:CCC

[Seg07] Mark R. Segal. Chess, chance and conspiracy. Statistical Science, 22(1):98–108, February 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1185975639.

Segers:2012:NIM

[Seg12] Johan Segers. Nonparametric inference for max-stable dependence. Statistical Science, 27(2):193–196, May 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1340110867.

Seidenfeld:1986:ASP

[Sei86] Teddy Seidenfeld. [The axioms of subjective probability]: Comment. Statistical Science, 1(3):354–356, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013616.  See

[Fis86a].

Seidenfeld:1992:RFF

[Sei92] Teddy Seidenfeld. R. A. Fisher’s fiducial argument and Bayes’ Theorem. Statistical Science, 7(3):358–368, August 1992. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011232.

Sen:1987:RMA

[Sen87] Pranab Kumar Sen. [A review of multivariate analysis]: Comment. Statistical Science, 2(4):426–428, November 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013115.

Seneta:1993:LCS

[Sen93] Eugene Seneta. Lewis Carroll’s “Pillow Problems”: On the 1993 centenary. Statistical Science, 8(2):180–186, May 1993. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011011.

Senn:2003:CJN

[Sen03] Stephen Senn. A conversation with John Nelder. Statistical Science, 18(1):118–131, February 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1056397489.

Senn:2009:C

[Sen09] Stephen Senn. Comment. Statistical Science, 24(2):185–186, May 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1263478378.

Seneta:2016:MCM

[Sen16] Eugene Seneta. Markov chains as models in statistical mechanics. Statistical Science, 31(3):399–414, August 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1475001235.

Severini:1995:IBE

[Sev95] Thomas A. Severini. [Inference based on estimating functions in the presence of nuisance parameters]: Comment. Statistical Science, 10(2):187–189, May 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177010034.

Spiegelhalter:1990:BBC

[SF90] David J. Spiegelhalter and Laurence S. Freedman. [Biostatistics and Bayes]: Comment. Statistical Science, 5(3):292–294, August 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012096.

Sherlock:2010:RWM

[SFR10] Chris Sherlock, Paul Fearnhead, and Gareth O. Roberts. The random walk Metropolis: Linking theory and practice through a case study. Statistical Science, 25(2):172–190, May 2010. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1290175840.

Spiegelhalter:2009:CCS

[SG09] David Spiegelhalter and Harvey Goldstein. Comment: Citation statistics. Statistical Science, 24(1):21–24, February 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1255009005. See [Hal09].

Seaman:2013:WMM

[SGJC13] Shaun Seaman, John Galati, Dan Jackson, and John Carlin. What is meant by “missing at random”? Statistical Science, 28(2):257–268, May 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1369147915.

Sebastiani:2003:SCF

[SGKR03] Paola Sebastiani, Emanuela Gussoni, Isaac S. Kohane, and Marco F. Ramoni. Statistical challenges in functional genomics. Statistical Science, 18(1):33–70, February 2003. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1056397486.

Small:2006:REV

[SGKR06] Dylan S. Small, Joseph L. Gastwirth, Abba M. Krieger, and Paul R. Rosenbaum. R-estimates vs. GMM: A theoretical case study of validity and efficiency. Statistical Science, 21(3): 363–375, August 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1166642441.

Si:2010:LAB

[SGZW10] Zhangzhang Si, Haifeng Gong, Song-Chun Zhu, and Ying Nian Wu. Learning active basis models by EM-type algorithms. Statistical Science, 25(4):458–475, November 2010. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1300108231.

Swanson:2014:TGA

[SH14] Sonja A. Swanson and Miguel A. Hernán. Think globally, act globally: An epidemiologist’s perspective on instrumental variable estimation. Statistical Science, 29(3):371–374, August 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1411437517.

Stein:2019:CPI

[SH19] Michael L. Stein and Ying Hung. Comment on “Probabilistic Integration: A Role in Statistical Computation?”. Statistical Science, 34(1):34–37, February 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1555056028.

Shafer:1986:CPD

[Sha86a] Glenn Shafer. [Combining probability distributions: A critique and an annotated bibliography]: Comment. Statistical Science, 1(1):135–137, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013826. See [GZ86a].

Shafer:1986:SR

[Sha86b] Glenn Shafer. Savage revisited. Statistical Science, 1(4):463– 485, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013518.

Shafer:1986:SRR

[Sha86c] Glenn Shafer. [Savage revisited]: Rejoinder. Statistical Science, 1(4):499–501, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177013524.

Shafer:1987:PESa

[Sha87a] Glenn Shafer. [Probabilistic expert systems in medicine: Practical issues in handling uncertainty]: Comment. Statistical Science, 2(1):37–38, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177013431. See [Spi87a].

Shafer:1987:PESb

[Sha87b] Glenn Shafer. [Probabilistic expert systems in medicine: Practical issues in handling uncertainty]: Rejoinder. Statistical Science, 2(1):41–42, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177013434. See [Spi87a].

Shafer:1987:PJA

[Sha87c] Glenn Shafer. Probability judgment in artificial intelligence and expert systems. Statistical Science, 2(1):3–16, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013426.

Shafer:1990:UDPb

[Sha90a] Glenn Shafer. [The unity and diversity of probability]: Rejoinder. Statistical Science, 5(4):457–462, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012029.

Shafer:1990:UDPa

[Sha90b] Glenn Shafer. The unity and diversity of probability. Statistical Science, 5(4):435–444, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012021.

Shamir:1993:PAL

[Sha93] Ron Shamir. Probabilistic analysis in linear programming. Statistical Science, 8(1):57–64, February 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011084.

Shao:2003:IBS

[Sha03] Jun Shao. Impact of the bootstrap on sample surveys. Statistical Science, 18(2):191–198, May 2003. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1063994974.

Shafer:2016:RSO

[Sha16] Glenn Shafer. Reconciling the subjective and objective aspects of probability. Statistical Science, 31(4):552–554, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1484816584.

Shafer:2018:MFB

[Sha18] Glenn Shafer. Marie-France Bru and Bernard Bru on dice games and contracts. Statistical Science, 33(2):277–284, May 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1525313146.

Shafer:2020:CIT

[Sha20] Glenn Shafer. Comment: Illusions, then and now. Statistical Science, 35(2):173–174, May 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1591171224. See [MG20a].

Shafer:2021:CHL

[Sha21] Glenn Shafer. Comment: On the history and limitations of probability updating. Statistical Science, 36(2):191–195, May 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-36/issue-2/
Comment-On-the-History-and-Limitations-of-Probability-
Updating/10.1214/21-STS765A.full.

Shepp:1992:CYV

[She92] Larry Shepp. A conversation with Yuri Vasilyevich Prokhorov. Statistical Science, 7(1):123–130, February 1992. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011453.

Sheather:2004:DE

[She04] Simon J. Sheather. Density estimation. Statistical Science, 19(4):588–597, November 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1113832723.

Shibata:1987:RTF

[Shi87] Ritei Shibata. [Rational transfer function approximation]: Comment. Statistical Science, 2(2):157–158, May 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013348.

Shmueli:2010:EP

[Shm10] Galit Shmueli. To explain or to predict? Statistical Science, 25 (3):289–310, August 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1294167961.

Shpitser:2014:CGA

[Shp14] Ilya Shpitser. Causal graphs: Addressing the confounding problem without instruments or ignorability. Statistical Science, 29 (3):367–370, August 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1411437516.

Stangl:2012:CID

[SII12] Dalene Stangl, Lurdes Y. T. Inoue, and Telba Z. Irony. Celebrating 70: An interview with Don Berry. Statistical Science, 27 (1):144–159, February 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1331729988.

Silverman:1991:CKR

[Sil91] B. W. Silverman. [Choosing a kernel regression estimator]: Comment: Should we use kernel methods at all? Statistical Science, 6(4):430–433, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011591.

Silverman:2009:CBC

[Sil09] Bernard W. Silverman. Comment: Bibliometrics in the context of the UK research assessment exercise. Statistical Science, 24 (1):15–16, February 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1255009003.

Simes:1991:ESR

[Sim91] R. John Simes. [Ethics and statistics in randomized clinical trials]: Comment. Statistical Science, 6(1):78–80, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011940.

Simon:1997:CMS

[Sim97] Richard M. Simon. A conversation with Marvin A. Schneiderman. Statistical Science, 12(2):98–102, May 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1029963425.

Singpurwalla:1986:SSA

[Sin86] Nozer D. Singpurwalla. [Some statistical applications of Poisson’s work]: Comment. Statistical Science, 1(2):177–178, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013694. See [Goo86a].

Singpurwalla:1987:SSW

[Sin87] Nozer D. Singpurwalla. [Survey of Soviet work in reliability]: Comment. Statistical Science, 2(4):497–499, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013130.

Singpurwalla:1989:EVA

[Sin89] Nozer D. Singpurwalla. [Extreme value analysis of environmental time series: An application to trend detection in ground-level ozone]: Comment. Statistical Science, 4(4):386– 387, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012405.

Singh:1994:SAE

[Sin94] Avinash C. Singh. [Small area estimation: An appraisal]: Comment. Statistical Science, 9(1):84–87, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010651.

Singpurwalla:1995:SDE

[Sin95] Nozer D. Singpurwalla. Survival in dynamic environments. Statistical Science, 10(1):86–103, February 1995. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010132.

Singpurwalla:2016:FTS

[Sin16a] Nozer D. Singpurwalla. Filtering and tracking survival propensity (reconsidering the foundations of reliability). Statistical Science, 31(4):521–540, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1484816580. See rejoinder

[Sin16b].

Singpurwalla:2016:RCU

[Sin16b] Nozer D. Singpurwalla. Rejoinder: Concert unlikely, “Jugalbandi” perhaps. Statistical Science, 31(4):555–557, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1484816585. See [Sin16a].

Sjolander:2012:AMC

[SJL^{+}12] Arvid Sjölander, Anna L. V. Johansson, Cecilia Lundholm, Daniel Altman, Catarina Almqvist, and Yudi Pawitan. Analysis of 1: 1 matched cohort studies and Twin studies, with binary exposures and binary outcomes. Statistical Science, 27 (3):395–411, August 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1346849946.

Salzman:2011:SMR

[SJW11] Julia Salzman, Hui Jiang, and Wing Hung Wong. Statistical modeling of RNA-seq data. Statistical Science, 26(1): 62–83, February 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1307626566.

Steffey:1991:BGT

[SK91] Duane Steffey and Robert E. Kass. [That BLUP is a good thing: The estimation of random effects]: Comment. Statistical Science, 6(1):45–47, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011931.

Schweinberger:2020:EFM

[SKBS20] Michael Schweinberger, Pavel N. Krivitsky, Carter T. Butts, and Jonathan R. Stewart. Exponential-family models of random graphs: Inference in finite, super and infinite population scenarios. Statistical Science, 35(4):627–662, November 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-35/issue-4/
Exponential-Family-Models-of-Random-Graphs--Inference-
in-Finite/10.1214/19-STS743.full.

Skinner:1994:MII

[Ski94] Chris Skinner. [Multiple-imputation inferences with uncongenial sources of input]: Comment. Statistical Science, 9(4): 561–563, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010272.

Stone:2014:SWA

[SKKS14] Lawrence D. Stone, Colleen M. Keller, Thomas M. Kratzke, and Johan P. Strumpfer. Search for the wreckage of air France flight AF 447. Statistical Science, 29(1):69–80, February 2014. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1399645730.

Simpson:2015:BVC

[SLR15] Daniel Simpson, Finn Lindgren, and Håvard Rue. Beyond the valley of the covariance function. Statistical Science, 30 (2):164–166, May 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1433341472.

Slud:2011:DIF

[Slu11] Eric Slud. Discussion of “Impact of Frequentist and Bayesian Methods on Survey Sampling Practice: A Selective Appraisal” by J. N. K. Rao. Statistical Science, 26(2):262–265, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1312204018. See [Rao11a].

Small:1989:SST

[Sma89] Christopher G. Small. [A survey of the statistical theory of shape]: Comment. Statistical Science, 4(2):105–108, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012584.

Smith:1987:UPA

[Smi87] Adrian F. M. Smith. [Uncertainty, policy analysis and statistics]: Comment. Statistical Science, 2(3):286–287, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013230.

Smith:1989:HGY

[Smi89a] Adrian F. M. Smith. [How to get your first research grant]: Comment. Statistical Science, 4(2):134–136, May 1989. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012592.

Smith:1989:EVAa

[Smi89b] Richard L. Smith. Extreme value analysis of environmental time series: An application to trend detection in ground-level ozone. Statistical Science, 4(4):367–377, November 1989. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012400.

Smith:1989:EVAb

[Smi89c] Richard L. Smith. [Extreme value analysis of environmental time series: An application to trend detection in ground-level ozone]: Rejoinder. Statistical Science, 4(4):389–393, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012407.

Smith:1992:SMD

[Smi92a] Richard L. Smith. [Statistical methods for data with longrange dependence]: Comment. Statistical Science, 7(4):422– 425, November 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011126.

Smith:1992:SPC

[Smi92b] Richard L. Smith. [Statistics, probability and chaos]: Comment: Relation between statistics and chaos. Statistical Science, 7(1):109–113, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011449.

Smith:1995:CDL

[Smi95] Adrian Smith. A conversation with Dennis Lindley. Statistical Science, 10(3):305–319, August 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177009940.

Smith:2002:SAB

[Smi02] Richard L. Smith. A statistical assessment of Buchanan’s vote in Palm beach county. Statistical Science, 17(4):441– 457, November 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1049993203.

Samartsidis:2017:CBM

[SMJN17] Pantelis Samartsidis, Silvia Montagna, Timothy D. Johnson, and Thomas E. Nichols. The coordinate-based meta-analysis of neuroimaging data. Statistical Science, 32(4):580–599, 11 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Splawa-Neyman:1990:APT

[SNDS90] Jerzy Splawa-Neyman, D. M. Dabrowska, and T. P. Speed. On the application of probability theory to agricultural experiments. essay on principles. section 9. Statistical Science, 5(4): 465–472, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012031.

Snell:1997:CJD

[Sne97] J. Laurie Snell. A conversation with Joe Doob. Statistical Science, 12(4):301–311, November 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1030037961.

Straf:1994:CMM

[SO94] Miron Straf and Ingram Olkin. A conversation with Margaret Martin. Statistical Science, 9(1):127–145, February 1994. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010657.

Sobel:1993:BAE

[Sob93] Michael E. Sobel. [Bayesian analysis in expert systems]: Comment. Statistical Science, 8(3):269–273, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010895.

Solomon:1986:SSA

[Sol86] Herbert Solomon. [Some statistical applications of Poisson’s work]: Comment. Statistical Science, 1(2):174–176, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013692.

Solo:1987:RTF

[Sol87] V. Solo. [Rational transfer function approximation]: Comment. Statistical Science, 2(2):158–159, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013349.

Solow:1994:RSP

[Sol94] Andrew R. Solow. [Report on statistics and physical oceanography]: Comment. Statistical Science, 9(2):213–215, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010486.

Solow:2003:SAS

[Sol03] Andrew R. Solow. Statistics in atmospheric science. Statistical Science, 18(4):422–429, November 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1081443226.

Sorensen:2005:CSO

[Sør05] Michael Sørensen. Comment: A selective overview of nonparametric methods in financial econometrics. Statistical Science, 20(4):344–346, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1137076649. See [Fan05b].

Spall:1991:BGT

[Spa91] James C. Spall. [That BLUP is a good thing: The estimation of random effects]: Comment: The Kalman filter and BLUP. Statistical Science, 6(1):39–41, February 1991. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011929.

Speed:1990:IRN

[Spe90] T. P. Speed. Introductory remarks on Neyman (1923). Statistical Science, 5(4):463–464, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012030.

Speed:1991:BGT

[Spe91] Terry Speed. [That BLUP is a good thing: The estimation of random effects]: Comment. Statistical Science, 6(1): 42–44, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011930.

Spiegelhalter:1987:PESa

[Spi87a] David J. Spiegelhalter. Probabilistic expert systems in medicine: Practical issues in handling uncertainty. Statistical Science, 2(1):25–30, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013428. See comments

[Wat87, DK87, Sha87a, Lin87a, Spi87b, Sha87b, Lin87b, Spi87c].

Spiegelhalter:1987:PESb

[Spi87b] David J. Spiegelhalter. [Probabilistic expert systems in medicine: Practical issues in handling uncertainty]: Comment. Statistical Science, 2(1):40–41, February 1987. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013433.  See

[Spi87a].

Spiegelhalter:1987:PESc

[Spi87c] David J. Spiegelhalter. [Probabilistic expert systems in medicine: Practical issues in handling uncertainty]: Rejoinder. Statistical Science, 2(1):43–44, February 1987. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013436.  See

[Spi87a].

Spiegelhalter:2004:IBI

[Spi04] David J. Spiegelhalter. Incorporating Bayesian ideas into health-care evaluation. Statistical Science, 19(1):156–174, February 2004. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1089808280.

Spurrier:1988:ODC

[Spu88] John D. Spurrier. [Optimal designs for comparing test treatments with controls]: Comment. Statistical Science, 3(4):485– 486, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012771.

Sun:2018:MIP

[SQH18] Yifei Sun, Jing Qin, and Chiung-Yu Huang. Missing information principle: A unified approach for general truncated and censored survival data problems. Statistical Science, 33(2):261– 276, May 2018. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1525313145.

Shaby:2012:DSM

[SR12] Benjamin Shaby and Brian J. Reich. Discussion of “Statistical Modeling of Spatial Extremes” by A. C. Davison, S. A. Padoan and M. Ribatet. Statistical Science, 27(2):197–198, May 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1340110868. See [DPR12b].

Sabbaghi:2014:CNF

[SR14] Arman Sabbaghi and Donald B. Rubin. Comments on the Neyman–Fisher controversy and its consequences. Statistical Science, 29(2):267–284, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1408368581.

Sofer:2016:NOC

[SRC^{+}16] Tamar Sofer, David B. Richardson, Elena Colicino, Joel Schwartz, and Eric J. Tchetgen Tchetgen. On negative outcome control of unobserved confounding as a generalization of difference-in-differences. Statistical Science, 31(3):348–361, August 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1475001233.

Simpson:2017:PMC

[SRR^{+}17a] Daniel Simpson, Håvard Rue, Andrea Riebler, Thiago G. Martins, and Sigrunn H. Sørbye. Penalising model component complexity: A principled, practical approach to constructing priors. Statistical Science, 32(1):1–28, February 2017. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1491465621.

Simpson:2017:YJK

[SRR^{+}17b] Daniel Simpson, Håvard Rue, Andrea Riebler, Thiago G. Martins, and Sigrunn H. Sørbye. You just keep on pushing my love over the borderline: A rejoinder. Statistical Science, 32 (1):44–46, February 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1491465626.

Shiryaev:1991:EAK

[SRS91] A. N. Shiryaev, Andrew L. Rukhin, and Paul Shaman. Everything about Kolmogorov was unusual.. Statistical Science, 6 (3):313–318, August 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011701.

Singpurwalla:1992:CBV

[SS92] Nozer D. Singpurwalla and Richard L. Smith. A conversation with Boris Vladimirovich Gnedenko. Statistical Science, 7(2):273–283, May 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011367.

Schwarz:1999:EAA

[SS99] Carl J. Schwarz and George A. F. Seber. Estimating animal abundance: Review III. Statistical Science, 14(4):427– 456, November 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009212521.

Soltis:2003:ABP

[SS03] Pamela S. Soltis and Douglas E. Soltis. Applying the bootstrap in phylogeny reconstruction. Statistical Science, 18(2):256–267, May 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1063994980.

Samworth:2018:ESI

[SS18] Richard J. Samworth and Bodhisattva Sen. Editorial: Special issue on “Nonparametric Inference Under Shape Constraints”. Statistical Science, 33(4):469–472, November 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1543482053.

Saarela:2023:REC

[SSM23] Olli Saarela, David A. Stephens, and Erica E. M. Moodie. The role of exchangeability in causal inference. Statistical Science, 38(3):369–385, August 2023. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-3/The-Role-of-Exchangeability-
in-Causal-Inference/10.1214/22-STS879.full.

Samartsidis:2019:ACE

[SSP^{+}19] Pantelis Samartsidis, Shaun R. Seaman, Anne M. Presanis, Matthew Hickman, and Daniela De Angelis. Assessing the causal effect of binary interventions from observational panel data with few treated units. Statistical Science, 34(3):486– 503, August 2019. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1570780981.

Salamon:2000:CDF

[SSS00] Hugh Salamon, Mark R. Segal, and Peter M. Small. Comparing DNA fingerprints of infectious organisms. Statistical Science, 15 (1):27–45, February 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009212672.

Sarwar:2020:DPC

[SSS20] Owais Sarwar, Benjamin Sauk, and Nikolaos V. Sahinidis. A discussion on practical considerations with sparse regression methodologies. Statistical Science, 35(4):593–601, November 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/journals/
statistical-science/volume-35/issue-4/A-Discussion-
on-Practical-Considerations-with-Sparse-Regression-
Methodologies/10.1214/20-STS806.full.

Shafer:2011:TMB

[SSVV11] Glenn Shafer, Alexander Shen, Nikolai Vereshchagin, and Vladimir Vovk. Test martingales, Bayes factors and p-values. Statistical Science, 26(1):84–101, February 2011. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1307626567.

Straf:2013:CSF

[ST13] Miron L. Straf and Judith M. Tanur. A conversation with Stephen E. Fienberg. Statistical Science, 28(3):447–463, August 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1377696945.

Stasny:1994:SAE

[Sta94] Elizabeth A. Stasny. [Small area estimation: An appraisal]: Comment. Statistical Science, 9(1):87–89, February 1994. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010652.

Stewart:1987:CLSa

[Ste87a] G. W. Stewart. Collinearity and least squares regression. Statistical Science, 2(1):68–84, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013439.

Stewart:1987:CLSb

[Ste87b] G. W. Stewart. [Collinearity and least squares regression]: Rejoinder. Statistical Science, 2(1):98–100, February 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013444.

Stein:1989:DAC

[Ste89] Michael L. Stein. [Design and analysis of computer experiments]: Comment. Statistical Science, 4(4):432–433, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012419.

Steele:1990:PAC

[Ste90] J. Michael Steele. [Poisson approximation and the Chen-Stein method]: Comment. Statistical Science, 5(4):424–425, November 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012016.

Steele:1993:I

[Ste93a] J. Michael Steele. In this issue. Statistical Science, 8(1): 1–2, February 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011075.

Steele:1993:MPD

[Ste93b] J. Michael Steele. Missing pieces, derandomization and concluding remarks. Statistical Science, 8(1):76–77, February 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011087.

Steele:1993:PPE

[Ste93c] J. Michael Steele. Probability and problems in Euclidean combinatorial optimization. Statistical Science, 8(1):48–56, February 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011083.

Steel:1994:CWR

[Ste94] David Steel. [Can we reach consensus on census adjustment?]: Comment. Statistical Science, 9(4):517–519, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010265.

Steele:2005:DHF

[Ste05] J. Michael Steele. Darrell Huff and fifty years of how to lie with statistics. Statistical Science, 20(3):205–209, August 2005. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1124891285.

Stern:2011:DSI

[Ste11] Hal Stern. Discussion of “Statistical Inference: The Big Picture” by R. E. Kass. Statistical Science, 26(1):17–18, February 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1307626558. See [Kas11b].

Stein:2020:SSI

[Ste20] Michael L. Stein. Some statistical issues in climate science. Statistical Science, 35(1):31–41, February 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1583226027.

Schefzik:2013:UQC

[STG13] Roman Schefzik, Thordis L. Thorarinsdottir, and Tilmann Gneiting. Uncertainty quantification in complex simulation models using ensemble copula coupling. Statistical Science, 28(4):616–640, November 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1386078881.

Stigler:1986:LMI

[Sti86] Stephen M. Stigler. Laplace’s 1774 memoir on inverse probability. Statistical Science, 1(3):359–363, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013620.

Stigler:1989:FGA

[Sti89] Stephen M. Stigler. Francis Galton’s account of the invention of correlation. Statistical Science, 4(2):73–79, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012580.

Stigler:1990:NML

[Sti90] Stephen M. Stigler. The 1988 Neyman Memorial Lecture: A Galtonian perspective on shrinkage estimators. Statistical Science, 5(1):147–155, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012274.

Stigler:1991:SSN

[Sti91] Stephen M. Stigler. Stochastic simulation in the nineteenth century. Statistical Science, 6(1):89–97, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011943.

Stigler:1994:CPJ

[Sti94] Stephen M. Stigler. Citation patterns in the journals of statistics and probability. Statistical Science, 9(1):94–108, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010655.

Stigler:1996:HS

[Sti96] Stephen M. Stigler. The history of statistics in 1933. Statistical Science, 11(3):244–252, August 1996. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1032280216.

Stigler:2005:F

[Sti05] Stephen Stigler. Fisher in 1921. Statistical Science, 20(1): 32–49, February 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1118065041.

Stigler:2006:INP

[Sti06a] Stephen M. Stigler. Isaac Newton as a probabilist. Statistical Science, 21(3):400–403, August 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1166642443.

Stine:2006:CCT

[Sti06b] Robert A. Stine. Comment: Classifier technology and the illusion of progress. Statistical Science, 21(1):27–29, February 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1149600843. See comments [Han06a].

Stigler:2007:ESM

[Sti07a] Stephen M. Stigler. The epic story of maximum likelihood. Statistical Science, 22(4):598–620, November 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1207580174.

Stigler:2007:WKR

[Sti07b] Stephen M. Stigler. William Kruskal remembered. Statistical Science, 22(2):275–276, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1190905527.

Stigler:2008:KPT

[Sti08] Stephen M. Stigler. Karl Pearson’s theoretical errors and the advances they inspired. Statistical Science, 23(2):261–271, May 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1219339117.

Stigler:2013:TTB

[Sti13] Stephen M. Stigler. The true title of Bayes’s essay. Statistical Science, 28(3):283–288, August 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1377696937.

Stigler:2018:RPF

[Sti18] Stephen M. Stigler. Richard Price, the first Bayesian. Statistical Science, 33(1):117–125, 02 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Schulte:2014:LME

[STLD14] Phillip J. Schulte, Anastasios A. Tsiatis, Eric B. Laber, and Marie Davidian. Q- and A-learning methods for estimating optimal dynamic treatment regimes. Statistical Science, 29(4): 640–661, November 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1421330551.

Stone:1986:GAM

[Sto86a] Charles J. Stone. [Generalized additive models]: Comment. Statistical Science, 1(3):312–314, August 1986. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013607.

Stone:1986:ASP

[Sto86b] Mervyn Stone. [The axioms of subjective probability]: Comment. Statistical Science, 1(3):356–357, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013617.  See

[Fis86a].

Stoyan:1989:SST

[Sto89] Dietrich Stoyan. [A survey of the statistical theory of shape]: Comment. Statistical Science, 4(2):115–116, May 1989. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012588.

Stone:2014:RDH

[Sto14] Lawrence D. Stone. Response to discussion by A. H. Welsh on the AF 447 paper. Statistical Science, 29(1):104–105, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1399645738. See [Wel14].

Strawderman:1990:DDT

[Str90] William E. Strawderman. [Developments in decision-theoretic variance estimation]: Comment. Statistical Science, 5(1):117– 118, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012271.

Small:2017:IVE

[STR^{+}17] Dylan S. Small, Zhiqiang Tan, Roland R. Ramsahai, Scott A. Lorch, and M. Alan Brookhart. Instrumental variable estimation with a stochastic monotonicity assumption. Statistical Science, 32(4):561–579, 11 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Stuetzle:1990:OOM

[Stu90] Werner Stuetzle. [OMEGA (online multivariate exploratory graphical analysis): Routine searching for structure]: Comment. Statistical Science, 5(2):217–219, May 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012171.

Stuart:2010:MMC

[Stu10] Elizabeth A. Stuart. Matching methods for causal inference: A review and a look forward. Statistical Science, 25(1):1–21, February 2010. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1280841730.

Stoffer:2000:SEA

[STW00] David S. Stoffer, David E. Tyler, and David A. Wendt. The spectral envelope and its applications. Statistical Science, 15 (3):224–253, August 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1009212816.

Sudderth:1986:ASP

[Sud86] William D. Sudderth. [The axioms of subjective probability]: Comment. Statistical Science, 1(3):357–358, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013618. See [Fis86a].

Sudbury:1994:DFR

[Sud94] Aidan Sudbury. [DNA fingerprinting: A review of the controversy]: Comment. Statistical Science, 9(2):262–263, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010493.

Suissa:1992:SEI

[Sui92] Samy Suissa. [A survey of exact inference for contingency tables]: Comment. Statistical Science, 7(1):170–172, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011461.

Sunley:1989:HGY

[Sun89] Judith S. Sunley. [How to get your first research grant]: Comment: Thinking about a research proposal? Statistical Science, 4(2):144–146, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012597.

Suppes:1986:ASP

[Sup86] Patrick Suppes. [The axioms of subjective probability]: Comment. Statistical Science, 1(3):347–350, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013613.  See

[Fis86a].

Suppes:1988:IBS

[Sup88] Patrick Suppes. [The interface between statistics and philosophy of science]: Comment: Causality, complexity and determinism. Statistical Science, 3(4):398–400, November 1988. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012755.

Sutradhar:2003:ORM

[Sut03] Brajendra C. Sutradhar. An overview on regression models for discrete longitudinal responses. Statistical Science, 18(3): 377–393, August 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1076102426.

Shafer:2006:SKG

[SV06] Glenn Shafer and Vladimir Vovk. The sources of Kolmogorov’s Grundbegriffe. Statistical Science, 21(1):70–98, February 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1149600847.

Seaman:2018:IDR

[SV18] Shaun R. Seaman and Stijn Vansteelandt. Introduction to double robust methods for incomplete data. Statistical Science, 33(2):184–197, May 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1525313141.

Savitsky:2011:VSN

[SVS11] Terrance Savitsky, Marina Vannucci, and Naijun Sha. Variable selection for nonparametric Gaussian process priors: Models and computational strategies. Statistical Science, 26(1): 130–149, February 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1307626570.

Solomon:1993:BHI

[SW93] Patricia J. Solomon and Susan R. Wilson. [Backcalculation of HIV infection rates]: Comment. Statistical Science, 8(2): 112–114, May 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011000.

Smith:2016:RCA

[SW16] Theresa R. Smith and Jon Wakefield. A review and comparison of age-period-cohort models for cancer incidence. Statistical Science, 31(4):591–610, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1484816588.

Skinner:2017:IDA

[SW17] Chris Skinner and Jon Wakefield. Introduction to the design and analysis of complex survey data. Statistical Science, 32 (2):165–175, May 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1494489809.

Sweeting:2011:DOP

[Swe11] Trevor Sweeting. Discussion of “Objective Priors: An Introduction for Frequentists” by M. Ghosh. Statistical Science, 26 (2):206–209, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1312204008. See [Gho11a].

Switzer:1992:CHS

[Swi92] Paul Switzer. A conversation with Herbert Solomon. Statistical Science, 7(3):388–401, August 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011234.

Switzer:1997:ENR

[Swi97] Paul Switzer. Editor’s note regarding the interview with Professor David Kendall, August 1996 issue. Statistical Science, 12(3):220, August 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1030037910.

Sacks:1989:DACa

[SWMW89a] Jerome Sacks, William J. Welch, Toby J. Mitchell, and Henry P. Wynn. Design and analysis of computer experiments. Statistical Science, 4(4):409–423, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012413.

Sacks:1989:DACb

[SWMW89b] Jerome Sacks, William J. Welch, Toby J. Mitchell, and Henry P. Wynn. [Design and analysis of computer experiments]: Rejoinder. Statistical Science, 4(4):433–435, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012420.

Small:2000:EMR

[SWY00] Christopher G. Small, Jinfang Wang, and Zejiang Yang. Eliminating multiple root problems in estimation (with comments by John J. Hanfelt, C. C. Heyde and Bing Li, and a rejoinder by the authors). Statistical Science, 15(4):313–341, November 2000. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1009213001.

Singh:2011:DBP

[SX11] Kesar Singh and Minge Xie. Discussion of “Is Bayes Posterior just Quick and Dirty Confidence?” by D. A. S. Fraser. Statistical Science, 26(3):319–321, August 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1320066920. See [Fra11a].

Sacks:2012:AYS

[SY12] Jerome Sacks and Donald Ylvisaker. After 50+ years in statistics, an exchange. Statistical Science, 27(2):308–318, May 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1340110875.

Spirtes:2014:UCE

[SZ14] Peter Spirtes and Jiji Zhang. A uniformly consistent estimator of causal effects under the k-triangle-faithfulness assumption. Statistical Science, 29(4):662–678, November 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1421330552.

Tanur:1990:QPE

[Tan90] Judith M. Tanur. [Quantifying probabilistic expressions]: Comment: On the possible dangers of isolation. Statistical Science, 5(1):21–22, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012247.

Tan:2007:CUP

[Tan07a] Zhiqiang Tan. Comment: Understanding OR, PS and DR. Statistical Science, 22(4):560–568, November 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1207580170.

Tanur:2007:WHK

[Tan07b] Judith M. Tanur. William H. Kruskal, mentor and friend. Statistical Science, 22(2):277–278, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1190905528.

Tan:2023:EDA

[Tan23] Linda S. L. Tan. Efficient data augmentation techniques for some classes of state space models. Statistical Science, 38(2): 240–261, May 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-2/
Efficient-Data-Augmentation-Techniques-for-Some-Classes-
of-State-Space/10.1214/22-STS867.full.

Taqqu:2013:BMF

[Taq13] Murad S. Taqqu. Benoît Mandelbrot and fractional Brownian motion. Statistical Science, 28(1):131–134, February 2013. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1359468412.

Thomas:2009:MIM

[TCC^{+}09] Duncan C. Thomas, Graham Casey, David V. Conti, Robert W. Haile, Juan Pablo Lewinger, and Daniel O. Stram. Methodological issues in multistage genome-wide association studies. Statistical Science, 24(4):414–429, November 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1271770340.

Tsiatis:2007:CDD

[TD07] Anastasios A. Tsiatis and Marie Davidian. Comment: Demystifying double robustness: A comparison of alternative strategies for estimating a population mean from incomplete data. Statistical Science, 22(4):569–573, November 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1207580171. See [KS07a].

Teh:2016:BNM

[Teh16] Yee Whye Teh. Bayesian nonparametric modeling and the ubiquitous Ewens sampling formula. Statistical Science, 31 (1):34–36, February 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1455115911. See [Cra16b, Cra16a].

Teichman:1988:LER

[Tei88] Kevin Yale Teichman. A little exposure to Radon. Statistical Science, 3(3):374–376, August 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012841.

Tuyl:2024:CII

[TGM24] Frank Tuyl, Richard Gerlach, and Kerrie Mengersen. On the certainty of an inductive inference: The binomial case. Statistical Science, 39(2):355–356, May 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-39/issue-2/On-the-Certainty-of-an-Inductive-
Inference--The-Binomial/10.1214/23-STS913.full.

Thall:2010:BMD

[Tha10] Peter F. Thall. Bayesian models and decision algorithms for complex early phase clinical trials. Statistical Science, 25 (2):227–244, May 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1290175844.

Thisted:1986:CEDa

[Thi86a] Ronald A. Thisted. Computing environments for data analysis. Statistical Science, 1(2):259–271, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013704.

Thisted:1986:CEDb

[Thi86b] Ronald A. Thisted. [Computing environments for data analysis]: Rejoinder. Statistical Science, 1(2):274–275, May 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177013707.

Thisted:1987:CLS

[Thi87] Ronald A. Thisted. [Collinearity and least squares regression]: Comment. Statistical Science, 2(1):91–93, February 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013442.

Tibshirani:2003:CPN

[THNC03] Robert Tibshirani, Trevor Hastie, Balasubramanian Narasimhan, and Gilbert Chu. Class prediction by nearest shrunken centroids, with applications to DNA microarrays. Statistical Science, 18(1):104–117, February 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1056397488.

Thompson:1991:BGT

[Tho91] Robin Thompson. [That BLUP is a good thing: The estimation of random effects]: Comment. Statistical Science, 6 (1):47, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011932.

Thompson:1994:MCL

[Tho94a] E. A. Thompson. Monte Carlo likelihood in genetic mapping. Statistical Science, 9(3):355–366, August 1994. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010381.

Thompson:1994:DFR

[Tho94b] William C. Thompson. [DNA fingerprinting: A review of the controversy]: Comment. Statistical Science, 9(2):263–266, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010494.

Thomsen:1994:SAE

[Tho94c] Ib Thomsen. [Small area estimation: An appraisal]: Comment. Statistical Science, 9(1):89–90, February 1994. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010653.

Thompson:2002:CVP

[Tho02] Mary E. Thompson. A conversation with V. P. Godambe. Statistical Science, 17(4):458–466, November 2002. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1049993204.

Thompson:2005:CFR

[Tho05] Elizabeth A. Thompson. Comment: Fuzzy and randomized confidence intervals and P -values. Statistical Science, 20(4):

382–383, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1137076657. See [GM05a].

Tibshirani:1994:NNR

[Tib94] Robert Tibshirani. [Neural networks: A review from statistical perspective]: Comment. Statistical Science, 9(1):48–49, February 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010645.

Tierney:1988:SMS

[Tie88] Luke Tierney. [Saddlepoint methods and statistical inference]: Comment. Statistical Science, 3(2):233–234, May 1988. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012911.

Tierney:1992:PMC

[Tie92] Luke Tierney. [Practical Markov chain Monte Carlo]: Comment. Statistical Science, 7(4):499–501, November 1992. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011146.

Titterington:1986:SPI

[Tit86] D. M. Titterington. [A statistical perspective on ill-posed inverse problems]: Comment. Statistical Science, 1(4):519– 521, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013526. See [O’S86b].

Titterington:2004:BMN

[Tit04] D. M. Titterington. Bayesian methods for neural networks and related models. Statistical Science, 19(1):128–139, February 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1089808278.

Tjostheim:2023:SEB

[TJL23] Dag Tjøstheim, Martin Jullum, and Anders Løland. Statistical embedding: Beyond principal components. Statistical Science, 38(3):411–439, August 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-3/Statistical-Embedding-Beyond-
Principal-Components/10.1214/22-STS881.full.

Tjostheim:2019:DMA

[Tjø19] Dag Tjøstheim. Discussion of models as approximations I & II. Statistical Science, 34(4):575–579, November 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1578474021. See

[BBB^{+}19, BBK^{+}19].

Tan:2013:VIG

[TN13] Linda S. L. Tan and David J. Nott. Variational inference for generalized linear mixed models using partially noncentered parametrizations. Statistical Science, 28(2):168–188, May 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1369147910.

Tian:2022:MCP

[TNM22] Qinglong Tian, Daniel J. Nordman, and William Q. Meeker. Methods to compute prediction intervals: a review and new results. Statistical Science, 37(4):580–597, November 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-4/
Methods-to-Compute-Prediction-Intervals--A-Review-and-
New/10.1214/21-STS842.full.

Tjostheim:2022:SDB

[TOS22] Dag Tjøstheim, Håkon Otneim, and Bård Støve. Statistical dependence: Beyond Pearson’s ρ. Statistical Science, 37(1):90– 109, February 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-37/issue-1/
Statistical-Dependence-Beyond-Pearsons-%cf%81/10.1214/
21-STS823.full.

Tighiouart:2010:DFE

[TR10] Mourad Tighiouart and André Rogatko. Dose finding with escalation with overdose control (EWOC) in cancer clinical trials. Statistical Science, 25(2):217–226, May 2010. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1290175843.

Trotter:1986:CLT

[Tro86] H. F. Trotter. [The central limit theorem around 1935]: Comment. Statistical Science, 1(1):92–93, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013819.  See

[Le 86a].

Trumbo:1989:HGYa

[Tru89a] B. E. Trumbo. How to get your first research Grant. Statistical Science, 4(2):121–131, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012590.

Trumbo:1989:HGYb

[Tru89b] B. E. Trumbo. [How to get your first research grant]: Rejoinder. Statistical Science, 4(2):148–150, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012599.

Tsay:1992:SPC

[Tsa92] Ruey S. Tsay. [Statistics, probability and chaos]: Comment: Simplicity and nonlinearity. Statistical Science, 7(1):113–114, February 1992. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177011450.

Tchetgen:2021:GAR

[TSW21] Eric Tchetgen Tchetgen, BaoLuo Sun, and Stefan Walter. The GENIUS approach to robust Mendelian randomization inference. Statistical Science, 36(3):443–464, August 2021. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
journals/statistical-science/volume-36/issue-3/The-
GENIUS-Approach-to-Robust-Mendelian-Randomization-Inference/
10.1214/20-STS802.full.

Tang:2023:PBD

[TT23] Wenpin Tang and Fengmin Tang. The Poisson binomial distribution — old & new. Statistical Science, 38(1):108–119, February 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-1/

The-Poisson-Binomial-Distribution-Old--New/10.1214/
22-STS852.full.

Tufte:1987:DGD

[Tuf87] Edward R. Tufte. [Dynamic graphics for data analysis]: Comment. Statistical Science, 2(4):389–392, November 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013109.

Tufte:1988:CCD

[Tuf88] Edward R. Tufte. A conversation with Cuthbert Daniel. Statistical Science, 3(4):413–424, November 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012760.

Tukey:1987:CDG

[Tuk87] John W. Tukey. [Dynamic graphics for data analysis]: Comment. Statistical Science, 2(4):383–385, November 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013105;
 http://www.jstor.org/stable/2245524.

Tukey:1990:DBG

[Tuk90] John W. Tukey. Data-based graphics: Visual display in the decades to come. Statistical Science, 5(3):327–339, August 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012101;   http://www.jstor.org/stable/2245820.

Tukey:1991:PMC

[Tuk91] John W. Tukey. The philosophy of multiple comparisons. Statistical Science, 6(1):100–116, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177011945;
 http://www.jstor.org/stable/2245714. 1989 Miller Lec-

ture presented at Stanford University.

Turnbull:1987:SPC

[Tur87] Craig D. Turnbull. [Statistics in psychiatry]: Comment. Statistical Science, 2(2):129–132, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013340. See [Eve87a].

Tanner:2010:EDA

[TW10] Martin A. Tanner and Wing H. Wong. From EM to data augmentation: The emergence of MCMC Bayesian computation in the 1980s. Statistical Science, 25(4):506–516, November 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1300108234.

Tille:2017:PSD

[TW17] Yves Tillé and Matthieu Wilhelm. Probability sampling designs: Principles for choice of design and balancing. Statistical Science, 32(2):176–189, May 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1494489810.

Taylor:2013:WPI

[TWDM13] S. W. Taylor, Douglas G. Woolford, C. B. Dean, and David L. Martell. Wildfire prediction to inform fire management: Statistical science challenges. Statistical Science, 28(4):586–615, November 2013. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1386078880.

Tweedie:1998:CRP

[Twe98] Richard Tweedie. Consulting: real problems, real interactions, real outcomes, with a resources appendix by sue Taylor. Statistical Science, 13(1):1–29, February 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1028905971.

Trippa:2023:CAC

[TX23] Lorenzo Trippa and Yanxun Xu. Comment: Advancing clinical trials with novel designs and implementations. Statistical Science, 38(2):216–218, May 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-2/Comment-Advancing-Clinical-
Trials-with-Novel-Designs-and-Implementations/10.1214/
23-STS865C.full.

Tchetgen:2024:IPC

[TYC^{+}24] Eric J. Tchetgen Tchetgen, Andrew Ying, Yifan Cui, Xu Shi, and Wang Miao. An introduction to proximal causal inference. Statistical Science, 39(3):375–390, August 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-3/
An-Introduction-to-Proximal-Causal-Inference/10.1214/
23-STS911.full.

Utts:1986:CSR

[Utt86] Jessica Utts. [Computers in statistical research]: Comment. Statistical Science, 1(4):437–439, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177013509.  See

[Edd86a].

Utts:1991:RMAa

[Utt91a] Jessica Utts. Replication and meta-analysis in parapsychology. Statistical Science, 6(4):363–378, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011577.

Utts:1991:RMAb

[Utt91b] Jessica Utts. [Replication and meta-analysis in parapsychology]: Rejoinder. Statistical Science, 6(4):396–403, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177011585.

Verbelen:2022:MOE

[VACC22] Roel Verbelen, Katrien Antonio, Gerda Claeskens, and Jonas Crevecoeur. Modeling the occurrence of events subject to a reporting delay via an EM algorithm. Statistical Science, 37(3):394–410, August 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-37/issue-3/Modeling-the-Occurrence-of-
Events-Subject-to-a-Reporting-Delay/10.1214/21-STS831.
full.

Vansteelandt:2011:IVE

[VBBG11] Stijn Vansteelandt, Jack Bowden, Manoochehr Babanezhad, and Els Goetghebeur. On instrumental variables estimation of causal odds ratios. Statistical Science, 26(3):403–422, August 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1320066928.

Villar:2015:MAB

[VBW15] Sofía S. Villar, Jack Bowden, and James Wason. Multi-armed bandit models for the optimal design of clinical trials: Benefits and challenges. Statistical Science, 30(2):199–215, May 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1433341478.

vandeGeer:2019:DMA

[vdG19] Sara van de Geer. Discussion of models as approximations I & II. Statistical Science, 34(4):566–568, November 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

https://projecteuclid.org/euclid.ss/1578474018. See

[BBB^{+}19, BBK^{+}19].

vandeGeer:2012:QLR

[vdGM12] Sara van de Geer and Patric Müller. Quasi-likelihood and/or robust estimation in high dimensions. Statistical Science, 27(4): 469–480, November 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1356098551.

vanDyk:2004:HSM

[vDK04] David A. van Dyk and Hosung Kang. Highly structured models for spectral analysis in high-energy astrophysics. Statistical Science, 19(2):275–293, May 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1105714163.

vanDyk:2010:CFS

[vDM10] David A. van Dyk and Xiao-Li Meng. Cross-fertilizing strategies for better EM mountain climbing and DA field exploration: A graphical guide book. Statistical Science, 25(4): 429–449, November 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1300108229.

vanderVaart:2014:HOT

[vdV14] Aad van der Vaart. Higher order tangent spaces and influence functions. Statistical Science, 29(4):679–686, November 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1421330553.

vanderVaart:2019:CBO

[vdV19] Aad van der Vaart. Comment: Bayes, oracle Bayes and empirical Bayes. Statistical Science, 34(2):214–218, May 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/

1563501635. See [Efr19a].

Velleman:1986:IOH

[Vel86] Paul F. Velleman. [Influential observations, high leverage points, and outliers in linear regression]: Comment. Statistical Science, 1(3):412–413, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013628.

Verducci:1987:SPC

[Ver87] Joseph S. Verducci. [Statistics in psychiatry]: Comment. Statistical Science, 2(2):132–133, May 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013341. See [Eve87a].

Verburgt:2021:KPM

[Ver21] Lukas M. Verburgt. Khinchin’s 1929 paper on von Mises’ frequency theory of probability. Statistical Science, 36(3):339– 343, August 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-36/issue-3/
Khinchins-1929-Paper-on-Von-Mises-Frequency-Theory-
of-Probability/10.1214/20-STS798.full.

Vernon:2014:GFB

[VGB14] Ian Vernon, Michael Goldstein, and Richard Bower. Galaxy formation: Bayesian history matching for the observable universe. Statistical Science, 29(1):81–90, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1399645731.

vonHippel:2021:MLM

[vHB21] Paul T. von Hippel and Jonathan W. Bartlett. Maximum likelihood multiple imputation: Faster imputations and consistent standard errors without posterior draws. Statistical Science, 36(3):400–420, August 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-3/Maximum-Likelihood-Multiple-
Imputation--Faster-Imputations-and-Consistent-Standard/
10.1214/20-STS793.full.

Vansteelandt:2014:SNM

[VJ14] Stijn Vansteelandt and Marshall Joffe. Structural nested models and G-estimation: The partially realized promise. Statistical Science, 29(4):707–731, November 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1421330555.

Vats:2021:RGR

[VK21] Dootika Vats and Christina Knudson. Revisiting the Gelman– Rubin diagnostic. Statistical Science, 36(4):518–529, November 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-36/issue-4/
Revisiting-the-GelmanRubin-Diagnostic/10.1214/20-STS812.
full.

Verbeke:2008:FIM

[VMB08] Geert Verbeke, Geert Molenberghs, and Caroline Beunckens. Formal and informal model selection with incomplete data. Statistical Science, 23(2):201–218, May 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1219339113.

VanderWeele:2020:OWL

[VMC20] Tyler J. VanderWeele, Maya B. Mathur, and Ying Chen. Outcome-wide longitudinal designs for causal inference: A new template for empirical studies. Statistical Science, 35(3):437– 466, August 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-35/issue-3/
Outcome-Wide-Longitudinal-Designs-for-Causal-Inference-
-A-New/10.1214/19-STS728.full.

Vovk:2009:SVP

[Vov09] Vladimir Vovk. Superefficiency from the vantage point of computability. Statistical Science, 24(1):73–86, February 2009. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1255009012.

Vovk:2021:TRO

[Vov21] Vladimir Vovk. Testing randomness online. Statistical Science, 36(4):595–611, November 2021. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-4/Testing-Randomness-Online/
10.1214/20-STS817.full.

vonStorch:1994:RSP

[vS94] Hans von Storch. [Report on statistics and physical oceanography]: Comment. Statistical Science, 9(2):215–221, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010487.

VanderWeele:2014:ISA

[VTH14] Tyler J. VanderWeele, Eric J. Tchetgen Tchetgen, and M. Elizabeth Halloran. Interference and sensitivity analysis. Statistical Science, 29(4):687–706, November 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1421330554.

Vovk:2023:CDV

[VW23] Vladimir Vovk and Ruodu Wang. Confidence and discoveries with E-values. Statistical Science, 38(2):329–354, May 2023. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-38/issue-2/
Confidence-and-Discoveries-with-E-values/10.1214/22-
STS874.full.

vanZwet:1993:RAH

[vZ93] Willem R. van Zwet. [Report of the ad hoc committee on design of an experiment on double-blind refereeing]: Comment. Statistical Science, 8(3):327–330, August 1993. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010909.

Wahba:1986:SPI

[Wah86] Grace Wahba. [A statistical perspective on ill-posed inverse problems]: Comment. Statistical Science, 1(4):521–522, November 1986. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177013527. See [O’S86b].

Wahba:1988:MRS

[Wah88] Grace Wahba. [Monotone regression splines in action]: Comment. Statistical Science, 3(4):456–458, November 1988. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012765.

Wahba:2006:C

[Wah06] Grace Wahba. Comment. Statistical Science, 21(3):347–351, August 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1166642438.

Wainer:1987:DGD

[Wai87] Howard Wainer. [Dynamic graphics for data analysis]: Comment: Deja view. Statistical Science, 2(4):388–389, November 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013108.

Wainer:1990:GVW

[Wai90] Howard Wainer. Graphical visions from William Playfair to John Tukey. Statistical Science, 5(3):340–346, August 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012102;   http://www.jstor.org/stable/2245821.

Wainer:2003:JWT

[Wai03] Howard Wainer. John Wilder Tukey: Statistical inventor, discoverer and revolutionary. Statistical Science, 18(3):285–286, August 2003. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1076102416.

Walker:2004:MBA

[Wal04] Stephen G. Walker. Modern Bayesian asymptotics. Statistical Science, 19(1):111–117, February 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1089808276.

Walther:2009:IML

[Wal09] Guenther Walther. Inference and modeling with log-concave distributions. Statistical Science, 24(3):319–327, August 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1270041258.

Wallis:2014:RFG

[Wal14a] Kenneth F. Wallis. Revisiting Francis Galton’s forecasting competition. Statistical Science, 29(3):420–424, August 2014. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1411437521.

Wallis:2014:TPN

[Wal14b] Kenneth F. Wallis. The two-piece normal, binormal, or double Gaussian distribution: Its origin and rediscoveries. Statistical Science, 29(1):106–112, February 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1399645739.

Waller:2018:CTL

[Wal18] Lance A. Waller. A conversation with Tom Louis. Statistical Science, 33(3):444–457, August 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1534147232.

Wang:2006:CEE

[Wan06] Wenbin Wang. Comment: Expert elicitation for reliable system design. Statistical Science, 21(4):456–459, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177334518. See [BQW06a].

Wang:2012:QCQ

[Wan12] Yazhen Wang. Quantum computation and quantum information. Statistical Science, 27(3):373–394, August 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1346849945.

Ware:1989:ITPa

[War89a] James H. Ware. Investigating therapies of potentially great benefit: ECMO. Statistical Science, 4(4):298–306, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012384.

Ware:1989:ITPb

[War89b] James H. Ware. [Investigating therapies of potentially great benefit: ECMO]: Rejoinder. Statistical Science, 4(4):337– 340, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012393.

Wasserman:2011:FI

[Was11] Larry Wasserman. Frasian inference. Statistical Science, 26 (3):322–325, August 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1320066921.

Watson:1987:PES

[Wat87] Stephen R. Watson. [Probabilistic expert systems in medicine: Practical issues in handling uncertainty]: Comment. Statistical Science, 2(1):30–32, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013429. See [Spi87a].

Watson:1989:SST

[Wat89] Geoffrey S. Watson. [A survey of the statistical theory of shape]: Comment. Statistical Science, 4(2):113–115, May 1989. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012587.

Waterman:1990:PAC

[Wat90] Michael S. Waterman. [Poisson approximation and the Chen-Stein method]: Comment. Statistical Science, 5(4):427–429, November 1990. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177012018.

Wallsten:1990:QPE

[WB90] Thomas S. Wallsten and David V. Budescu. [Quantifying probabilistic expressions]: Comment. Statistical Science, 5(1): 23–26, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012248.

Wang:2022:LPD

[WBR^{+}22] Zitong Wang, Mary Grace Bowring, Antony Rosen, Brian Garibaldi, Scott Zeger, and Akihiko Nishimura. Learning and predicting from dynamic models for COVID-19 patient monitoring. Statistical Science, 37(2):251–265, May 2022. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/journals/
statistical-science/volume-37/issue-2/Learning-and-
Predicting-from-Dynamic-Models-for-COVID-19-Patient/
10.1214/22-STS861.full.

Winter:2024:EDB

[WCL^{+}24] Steven Winter, Trevor Campbell, Lizhen Lin, Sanvesh Srivastava, and David B. Dunson. Emerging directions in Bayesian computation. Statistical Science, 39(1):62–89, February 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-1/
Emerging-Directions-in-Bayesian-Computation/10.1214/
23-STS919.full.

Wegman:1986:CSR

[Weg86a] Edward J. Wegman. [Computers in statistical research]: Comment. Statistical Science, 1(4):449–451, November 1986. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013515.

See [Edd86a].

Wegman:1986:SPR

[Weg86b] Edward J. Wegman. Some personal recollections of Harald Cramér on the development of statistics and probability. Statistical Science, 1(4):528–535, November 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013531.

Wegman:1989:HGY

[Weg89] Edward J. Wegman. [How to get your first research grant]: Comment. Statistical Science, 4(2):146–148, May 1989. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012598.

Weisberg:1986:IOH

[Wei86] Sanford Weisberg. [Influential observations, high leverage points, and outliers in linear regression]: Comment. Statistical Science, 1(3):414–415, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013629.

Weinberg:1987:SSW

[Wei87] Elliot H. Weinberg. [Survey of Soviet work in reliability]: Comment. Statistical Science, 2(4):499–501, November 1987. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177013131.

Weinstein:1988:LS

[Wei88] Jack B. Weinstein. Litigation and statistics. Statistical Science, 3(3):286–297, August 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012829.

Weissman:1989:EVA

[Wei89] Ishay Weissman. [Extreme value analysis of environmental time series: An application to trend detection in ground-level ozone]: Comment. Statistical Science, 4(4):385–386, November 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177012404.

Weihs:1991:C

[Wei91] C. Weihs. Comment. Statistical Science, 6(4):344–348, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177011573.

Weir:1994:DFR

[Wei94] B. S. Weir. [DNA fingerprinting: A review of the controversy]: Comment. Statistical Science, 9(2):266–267, May 1994. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010495.

Welsch:1986:IOH

[Wel86] Roy E. Welsch. [Influential observations, high leverage points, and outliers in linear regression]: Comment. Statistical Science, 1(3):403–405, August 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013625.

Welsh:1988:RBR

[Wel88] A. H. Welsh. [Rank-based robust analysis of linear models. I. Exposition and review]: Comment. Statistical Science, 3 (2):258–259, May 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012916.

Wells:2009:CSR

[Wel09] Martin T. Wells. A conversation with Shayle R. Searle. Statistical Science, 24(2):244–254, May 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1263478385.

Welsh:2014:D

[Wel14] A. H. Welsh. Discussion. Statistical Science, 29(1):101–102, February 2014. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1399645736. See response [Sto14].

Westfall:2011:DMT

[Wes11] Peter H. Westfall. Discussion of “Multiple Testing for Exploratory Research” by J. J. Goeman and A. Solari. Statistical Science, 26(4):604–607, November 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1330437940. See [GS11a].

Wilson:2006:ADC

[WGHR06] Alyson G. Wilson, Todd L. Graves, Michael S. Hamada, and C. Shane Reese. Advances in data combination, analysis and collection for system reliability assessment. Statistical Science, 21(4):514–531, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177334527.

Watson:2016:AMR

[WH16a] James Watson and Chris Holmes. Approximate models and robust decisions. Statistical Science, 31(4):465–489, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.

ss/1484816572. See rejoinder [WH16b].

Watson:2016:RAM

[WH16b] James Watson and Chris Holmes. Rejoinder: Approximate models and robust decisions. Statistical Science, 31(4):516– 520, November 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1484816579. See [WH16a].

Weller:2016:RNH

[WH16c] Zachary D. Weller and Jennifer A. Hoeting. A review of nonparametric hypothesis tests of isotropy properties in spatial data. Statistical Science, 31(3):305–324, August 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1475001231.

Wikle:2019:CNC

[WH19] Christopher K. Wikle and Jay M. Ver Hoef. A conversation with Noel Cressie. Statistical Science, 34(2):349–359, May 2019. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/euclid.ss/1563501646.

Wheeler:2021:CMB

[Whe21] Gregory Wheeler. Comment: Moving beyond sets of probabilities. Statistical Science, 36(2):201–204, May 2021. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-2/Comment-Moving-Beyond-Sets-
of-Probabilities/10.1214/21-STS765C.full.

Whittaker:1993:BAE

[Whi93a] Joe Whittaker. [Bayesian analysis in expert systems]: Comment. Statistical Science, 8(3):273–276, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010896.

Whittle:1993:CHD

[Whi93b] Peter Whittle. A conversation with Henry Daniels. Statistical Science, 8(3):342–353, August 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177010911.

Winkler:1986:CPD

[Win86] Robert L. Winkler. [Combining probability distributions: A critique and an annotated bibliography]: Comment. Statistical Science, 1(1):138–140, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013828. See [GZ86a].

Winkler:1990:QPE

[Win90] Robert L. Winkler. [Quantifying probabilistic expressions]: Comment: Representing and communicating uncertainty. Statistical Science, 5(1):26–30, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177012249.

Wittes:1987:SPM

[Wit87] Janet Wittes. [The statistical precision of medical screening procedures: Application to polygraph and AIDS antibodies test data]: Comment. Statistical Science, 2(3):228–230, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177013218.

Witting:1991:CLS

[Wit91] Hermann Witting. A conversation with Leopold Schmetterer. Statistical Science, 6(4):437–447, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011593.

Wu:2010:MAE

[WL10] Tong Tong Wu and Kenneth Lange. The MM alternative to EM. Statistical Science, 25(4):492–505, November 2010. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1300108233.

Wang:2021:NMB

[WLLH21] Y. X. Rachel Wang, Lexin Li, Jingyi Jessica Li, and Haiyan Huang. Network modeling in biology: Statistical methods for gene and brain networks. Statistical Science, 36(1):89–108, February 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-36/issue-1/
Network-Modeling-in-Biology--Statistical-Methods-for-
Gene-and/10.1214/20-STS792.full.

Wolpert:2004:ALS

[WM04] Robert L. Wolpert and Kerrie L. Mengersen. Adjusted likelihoods for synthesizing empirical evidence from studies that differ in quality and design: Effects of environmental tobacco smoke. Statistical Science, 19(3):450–471, August 2004. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1110999310.

Wegman:2007:CDG

[WM07] Edward J. Wegman and Wendy L. Martinez. A conversation with Dorothy Gilford. Statistical Science, 22(2):291–300, May 2007. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1190905530.

Wang:2019:CVA

[WMB19] Yixin Wang, Andrew C. Miller, and David M. Blei. Comment: Variational autoencoders as empirical Bayes. Statistical Science, 34(2):229–233, May 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1563501638. See [Efr19a].

Wikle:2013:MSM

[WMHL13] Christopher K. Wikle, Ralph F. Milliff, Radu Herbei, and William B. Leeds. Modern statistical methods in oceanography: A hierarchical perspective. Statistical Science, 28(4): 466–486, November 2013. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1386078874.

Wolter:1986:RMA

[Wol86] Kirk M. Wolter. [Regression models for adjusting the 1980 Census]: Comment. Statistical Science, 1(1):24–28, February 1986. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1177013808. See [FN86a].

Wolf:1990:QPE

[Wol90] Charles Wolf. [Quantifying probabilistic expressions]: Comment. Statistical Science, 5(1):31–32, February 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012250.

Wolfe:2004:RSS

[Wol04a] Douglas A. Wolfe. Ranked set sampling: An approach to more efficient data collection. Statistical Science, 19(4):636– 643, November 2004. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1113832728.

Wolpert:2004:CJB

[Wol04b] Robert L. Wolpert. A conversation with James O. Berger. Statistical Science, 19(1):205–218, February 2004. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1089808283.

Wong:1995:BCS

[Won95] Wing Hung Wong. [Bayesian computation and stochastic systems]: Comment. Statistical Science, 10(1):52–53, February 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177010128.

Woodall:1993:ISC

[Woo93] William H. Woodall. [Is industrial statistics out of control?]: Comment. Statistical Science, 8(4):397–399, November 1993. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010781.

Witztum:1994:ELS

[WRR94] Doron Witztum, Eliyahu Rips, and Yoav Rosenberg. Equidistant letter sequences in the book of Genesis. Statistical Science, 9(3):429–438, August 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010393.

Weihs:1990:OOMa

[WS90a] Claus Weihs and Heinz Schmidli. OMEGA (online multivariate exploratory graphical analysis): Routine searching for structure. Statistical Science, 5(2):175–208, May 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012166.

Weihs:1990:OOMb

[WS90b] Claus Weihs and Heinz Schmidli. [OMEGA (online multivariate exploratory graphical analysis): Routine searching for structure]: Rejoinder. Statistical Science, 5(2):222–226, May 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012173.

Wang:2020:QSQ

[WS20] Yazhen Wang and Xinyu Song. Quantum science and quantum technology. Statistical Science, 35(1):51–74, February 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1583226029.

Whitney:2019:CMD

[WSC19] David Whitney, Ali Shojaie, and Marco Carone. Comment: Models as (deliberate) approximations. Statistical Science, 34(4):591–598, November 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://

projecteuclid.org/euclid.ss/1578474024. See [BBB^{+}19,

BBK^{+}19].

Willinger:1995:SSH

[WTLW95] Walter Willinger, Murad S. Taqqu, Will E. Leland, and Daniel V. Wilson. Self-similarity in high-speed packet traffic: Analysis and modeling of Ethernet traffic measurements. Statistical Science, 10(1):67–85, February 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177010131.

Wu:1993:ISC

[Wu93] C. F. J. Wu. [Is industrial statistics out of control?]: Comment. Statistical Science, 8(4):399–400, November 1993. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010782.

Waterman:1994:SCS

[WV94] Michael S. Waterman and Martin Vingron. Sequence comparison significance and Poisson approximation. Statistical Science, 9(3):367–381, August 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177010382.

Wang:2016:QAM

[WWZ16] Yazhen Wang, Shang Wu, and Jian Zou. Quantum annealing with Markov chain Monte Carlo simulations and D-wave quantum computers. Statistical Science, 31(3):362–398, August 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1475001234.

Wynn:1993:ISC

[Wyn93] H. P. Wynn. [Is industrial statistics out of control?]: Comment. Statistical Science, 8(4):400–402, November 1993. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010783.

Woodroofe:2002:C

[WZ02] Michael Woodroofe and Tonglin Zhang. Correction. Statistical Science, 17(3):379, August 2002. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1042727945.

Wellner:2012:ISI

[WZ12] Jon Wellner and Tong Zhang. Introduction to the special issue on sparsity and regularization methods. Statistical Science, 27(4):447–449, November 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1356098549.

Wagenmakers:2024:JBH

[WZG24] Eric-Jan Wagenmakers, Sandy Zabell, and Quentin F. Gronau. J. B. S. Haldane’s rule of succession. Statistical Science, 39(2): 346–354, May 2024. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-39/issue-2/
J-B-S-Haldanes-Rule-of-Succession/10.1214/23-STS912.
full.

Xie:2014:RPI

[XHC14] Yihui Xie, Heike Hofmann, and Xiaoyue Cheng. Reactive programming for interactive graphics. Statistical Science, 29 (2):201–213, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1408368571.

Xiang:2015:API

[XR15] Alice Xiang and Donald B. Rubin. Assessing the potential impact of a nationwide class-based affirmative action system. Statistical Science, 30(3):297–327, August 2015. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1439220715.

Xia:2011:FMT

[XT11a] Yingcun Xia and Howell Tong. Feature matching in time series modeling. Statistical Science, 26(1):21–46, February 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/

1307626560. See discussion [Han11b, Ion11, CT11, Yao11].

Xia:2011:R

[XT11b] Yingcun Xia and Howell Tong. Rejoinder. Statistical Science, 26 (1):59–61, February 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1307626565.

Xiang:2019:OSE

[XYY19] Sijia Xiang, Weixin Yao, and Guangren Yang. An overview of semiparametric extensions of finite mixture models. Statistical Science, 34(3):391–404, August 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https:/

/projecteuclid.org/euclid.ss/1570780976.

Yang:1999:CLC

[Yan99] Grace L. Yang. A conversation with Lucien Le Cam. Statistical Science, 14(2):223–241, May 1999. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1009212249.

Yang:2008:CNL

[Yan08] Grace L. Yang. Comment: The 2005 Neyman lecture: Dynamic indeterminism in science. Statistical Science, 23(1): 69–75, February 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1215441284. See [Bri08a].

Yao:2011:DFM

[Yao11] Qiwei Yao. Discussion of “Feature Matching in Time Series Modeling” by Y. Xia and H. Tong. Statistical Science, 26(1):

57–58, February 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1307626564.

Yan:2017:HSM

[YB17] Xiaohan Yan and Jacob Bien. Hierarchical sparse modeling: A choice of two group lasso formulations. Statistical Science, 32(4):531–560, 11 2017. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Yang:2016:FIS

[YK16] Shu Yang and Jae Kwang Kim. Fractional imputation in survey sampling: A comparative review. Statistical Science, 31(3): 415–432, August 2016. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1475001236.

Yu:2023:VIC

[YNS23] Xuejun Yu, David J. Nott, and Michael Stanley Smith. Variational inference for cutting feedback in misspecified models. Statistical Science, 38(3):490–509, August 2023. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-3/Variational-Inference-for-
Cutting-Feedback-in-Misspecified-Models/10.1214/23-
STS886.full.

Young:1990:OOM

[You90] Forrest W. Young. [OMEGA (online multivariate exploratory graphical analysis): Routine searching for structure]: Comment: Industrial strength VEDA. Statistical Science, 5(2): 219–222, May 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177012172.

Young:1991:C

[You91] Forrest W. Young. Comment. Statistical Science, 6(4):349– 352, November 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011574.

Young:1994:BMTa

[You94a] G. Alastair Young. Bootstrap: More than a stab in the dark? Statistical Science, 9(3):382–395, August 1994. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010383.

Young:1994:BMTb

[You94b] G. Alastair Young. [Bootstrap: More than a stab in the dark?]: Rejoinder. Statistical Science, 9(3):411–415, August 1994. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177010391.

Yu:2022:SPR

[YS22] Bin Yu and Chandan Singh. Seven principles for rapidresponse data science: Lessons learned from COVID-19 forecasting. Statistical Science, 37(2):266–269, May 2022. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-37/issue-2/Seven-Principles-for-Rapid-
Response-Data-Science--Lessons-Learned/10.1214/22-STS855.
full.

Yu:2020:MMO

[YSR20] Ruoqi Yu, Jeffrey H. Silber, and Paul R. Rosenbaum. Matching methods for observational studies derived from large administrative databases. Statistical Science, 35(3):338–355, August 2020. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-35/issue-3/
Matching-Methods-for-Observational-Studies-Derived-
from-Large-Administrative-Databases/10.1214/19-STS699.
full.

Yu:1995:BCS

[Yu95] Bin Yu. [Bayesian computation and stochastic systems]: Comment: Extracting more diagnostic information from a single run using Cusum path plot. Statistical Science, 10(1):54–58, February 1995. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1177010129.

Yu:2006:CMN

[Yu06] Bin Yu. Comment: Monitoring networked applications with incremental quantile estimation. Statistical Science, 21(4):483– 484, November 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177334523. See [CJLV06a].

Ying:2006:CYS

[YZ06] Zhiliang Ying and Cun-Hui Zhang. A conversation with Yuan Shih Chow. Statistical Science, 21(1):99–112, February 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1149600848.

Zabell:1989:RFHa

[Zab89a] Sandy Zabell. R. A. Fisher on the history of inverse probability. Statistical Science, 4(3):247–256, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012488.

Zabell:1989:RFHb

[Zab89b] Sandy Zabell. [R. A. Fisher on the history of inverse probability]: Rejoinder. Statistical Science, 4(3):261–263, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1177012491.

Zabell:1992:RFF

[Zab92] S. L. Zabell. R. A. Fisher and fiducial argument. Statistical Science, 7(3):369–387, August 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011233.

Zabell:1994:CWK

[Zab94] Sandy Zabell. A conversation with William Kruskal. Statistical Science, 9(2):285–303, May 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177010498.

Zabell:2023:SLJ

[Zab23] Sandy Zabell. The secret life of I. J. Good. Statistical Science, 38(2):285–302, May 2023. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-38/issue-2/The-Secret-Life-of-I-J-Good/
10.1214/22-STS870.full.

Zahn:1991:ASP

[Zah91] Douglas A. Zahn. [Avoiding statistical pitfalls]: Comment. Statistical Science, 6(3):264–267, August 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011692.

Zaslavsky:1994:MII

[Zas94] Alan M. Zaslavsky. [Multiple-imputation inferences with uncongenial sources of input]: Comment: Using the full toolkit. Statistical Science, 9(4):563–565, November 1994. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010273.

Zaslavsky:2011:SBM

[Zas11] Alan M. Zaslavsky. Sampling from a Bayesian menu. Statistical Science, 26(2):235–237, May 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1312204013.

Zhang:2015:WDC

[ZC15] Hao Zhang and Wenxiang Cai. When doesn’t cokriging outperform kriging? Statistical Science, 30(2):176–180, May 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1433341475.

Zeitouni:2018:CRV

[Zei18] Ofer Zeitouni. A conversation with S. R. S. Varadhan. Statistical Science, 33(1):126–137, 02 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

Zellner:1987:TPH

[Zel87] Arnold Zellner. [Testing precise hypotheses]: Comment. Statistical Science, 2(3):339–341, August 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013241.

Zelen:1990:BBC

[Zel90] M. Zelen. [Biostatistics and Bayes]: Comment. Statistical Science, 5(3):294–295, August 1990. CODEN STSCEP.

ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012097.

Zelen:1991:ESR

[Zel91] M. Zelen. [Ethics and statistics in randomized clinical trials]: Comment: The ethos of clinical trials. Statistical Science, 6 (1):81–83, February 1991. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011941.

Zellner:2009:C

[Zel09] Arnold Zellner. Comment. Statistical Science, 24(2):187–190, May 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1263478379.

Zhang:2011:SAG

[Zha11a] Heping Zhang. Statistical analysis in genetic studies of mental illnesses. Statistical Science, 26(1):116–129, February 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1307626569.

Zhang:2011:DBP

[Zha11b] Tong Zhang. Discussion of “Is Bayes Posterior just Quick and Dirty Confidence?” by D. A. S. Fraser. Statistical Science, 26 (3):326–328, August 2011. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1320066922. See [Fra11a].

Zhuang:2018:CRS

[Zhu18] Jiancang Zhuang. Comment on “A Review of Self-Exciting Spatio-Temporal Point Process and Their Applications” by Alex Reinhart. Statistical Science, 33(3):323–324, August 2018. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/

1534147223. See [Rei18b, Rei18a].

Zidek:1988:GOC

[Zid88] James V. Zidek. [Golden oldies: Classic articles from the world of statistics and probability]: Comment. Statistical Science, 3 (1):87–90, February 1988. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177013004. See [Hot88c].

Zidek:1989:HGY

[Zid89] James V. Zidek. [How to get your first research grant]: Comment. Statistical Science, 4(2):132–134, May 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012591.

Zhang:2008:MSC

[ZJ08] Zhihua Zhang and Michael I. Jordan. Multiway spectral clustering: A margin-based perspective. Statistical Science, 23(3): 383–403, August 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1233153065.

Zhou:2024:SFO

[ZJ24] Tianjian Zhou and Yuan Ji. Statistical frameworks for oncology dose-finding designs with late-onset toxicities: a review. Statistical Science, 39(2):243–261, May 2024. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-39/issue-2/Statistical-Frameworks-for-
Oncology-Dose-Finding-Designs-with-Late-Onset/10.1214/
23-STS895.full.

Zheng:2009:RTG

[ZJZ^{+}09] Gang Zheng, Jungnam Joo, Dmitri Zaykin, Colin Wu, and Nancy Geller. Robust tests in genome-wide scans under incomplete linkage disequilibrium. Statistical Science, 24(4):503– 516, November 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1271770345.

Zhao:2019:CWC

[ZKS19] Qingyuan Zhao, Luke J. Keele, and Dylan S. Small. Comment: Will competition-winning methods for causal inference also succeed in practice? Statistical Science, 34(1):72–76, February 2019. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/euclid.ss/
1555056032.

Zheng:2008:CQF

[ZL08] Tian Zheng and Shaw-Hwa Lo. Comment: Quantifying the fraction of missing information for hypothesis testing in statistical and genetic studies. Statistical Science, 23(3):321–324, August 2008. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1233153060. See [NMK08a].

Zheng:2015:CRC

[ZLG15] Gang Zheng, Zhaohai Li, and Nancy L. Geller. A conversation with Robert C. Elston. Statistical Science, 30(2):258–267, May 2015. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.
ss/1433341482.

Zeger:1993:RMD

[ZLH93] Scott L. Zeger, Kung-Yee Liang, and Patrick Heagerty. [Regression models for discrete longitudinal responses]: Comment. Statistical Science, 8(3):304–306, August 1993. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177010902.

Zhou:2010:BCM

[ZLK10] Yan Zhou, Roderick J. A. Little, and John D. Kalbfleisch. Block-conditional missing at random models for missing data. Statistical Science, 25(4):517–532, November 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1300108235.

Zhou:2010:GPU

[ZLS10] Hua Zhou, Kenneth Lange, and Marc A. Suchard. Graphics processing units and high-dimensional optimization. Statistical Science, 25(3):311–324, August 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1294167962.

Zheng:2009:ISI

[ZMG09] Gang Zheng, Jonathan Marchini, and Nancy L. Geller. Introduction to the special issue: Genome-wide association studies. Statistical Science, 24(4):387, November 2009. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1271770337.

Zigler:2021:BCI

[ZP21] Corwin M. Zigler and Georgia Papadogeorgou. Bipartite causal inference with interference. Statistical Science, 36(1):109–123, February 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL https://projecteuclid.

org/journals/statistical-science/volume-36/issue-1/
Bipartite-Causal-Inference-with-Interference/10.1214/
19-STS749.full.

Zhang:2009:CER

[ZS09] Kai Zhang and Dylan S. Small. Comment: The essential role of pair matching in cluster-randomized experiments, with application to the Mexican universal health insurance evaluation. Statistical Science, 24(1):59–64, February 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1255009010. See [IKN09a].

Zhao:2006:GDB

[ZSCW06] Y. Zhao, J. Staudenmayer, B. A. Coull, and M. P. Wand. General design Bayesian generalized linear mixed models. Statistical Science, 21(1):35–51, February 2006. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1149600845.

Zollner:2009:UGD

[ZT09] Sebastian Zöllner and Tanya M. Teslovich. Using GWAS data to identify copy number variants contributing to common complex diseases. Statistical Science, 24(4):530–546, November 2009. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL http://projecteuclid.org/euclid.ss/
1271770347.

Zubin:1987:SPC

[Zub87] Joseph Zubin. [Statistics in psychiatry]: Comment: The biometric approach to psychiatry. Statistical Science, 2(2):121– 125, May 1987. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL http://projecteuclid.org/
euclid.ss/1177013337. See [Eve87a].

Zuo:2021:GND

[Zuo21] Yijun Zuo. On general notions of depth for regression. Statistical Science, 36(1):142–157, February 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-1/On-General-Notions-of-Depth-
for-Regression/10.1214/20-STS767.full.

Zhao:2019:TSI

[ZWS^{+}19] Qingyuan Zhao, Jingshu Wang, Wes Spiller, Jack Bowden, and Dylan S. Small. Two-sample instrumental variable analyses using heterogeneous samples. Statistical Science, 34(2):317– 333, May 2019. CODEN STSCEP. ISSN 0883-4237 (print),

2168-8745 (electronic). URL https://projecteuclid.org/
euclid.ss/1563501644.

Zhao:2021:DIV

[ZWS21] Sen Zhao, Daniela Witten, and Ali Shojaie. In defense of the indefensible: a very Naïve approach to high-dimensional inference. Statistical Science, 36(4):562–577, November 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745

(electronic). URL https://projecteuclid.org/journals/
statistical-science/volume-36/issue-4/In-Defense-of-
the-Indefensible--A-Very-Na%c3%afve-Approach/10.1214/
20-STS815.full.

Zhang:2012:GTC

[ZZ12] Cun-Hui Zhang and Tong Zhang. A general theory of concave regularization for high-dimensional sparse estimation problems. Statistical Science, 27(4):576–593, November 2012. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1356098557.

Zhou:2021:CTF

[ZZS21] Jie Zhou, Xiao-Hua Zhou, and Liuquan Sun. Comparison of two frameworks for analyzing longitudinal data. Statistical Science, 36(4):530–541, November 2021. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL https://projecteuclid.org/journals/statistical-
science/volume-36/issue-4/Comparison-of-Two-Frameworks-
for-Analyzing-Longitudinal-Data/10.1214/20-STS813.full.